A Complete Bibliography of Publications in Minerva: A Review of Science, Learning and

Policy

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

08 October 2025 Version 1.33

Title word cross-reference

$22.50 [Mol05]. $29.95 [Son21]. 2 × 2 [DH64]. $49.00 [Mar20]. h [DGR19].

-Index [DGR19].

0-226-35933-6 [Mol05].

1 [Eck04, Son21, Bel98b]. 1-85973-421-9 [Eck04]. 1-89153-72-4 [Gre05]. 1-891853-73-2 [Gre05]. 1620 [Stu98]. 1850 [Bel98b]. 1870 [Mac99a]. 19 [Fre23b, áRH23, SREM^{+}23]. 1915 [Jen01]. 1920s [BB81, Hol81, RLH91, Sol75]. 1930s [AA88, Gen87, Wer07, Zim06]. 1940s [Pet18]. 1950s [Fel19]. 1955 [Bod64]. 1960s [EB22, Lew80, S.88c]. 1970s [Gir79, Kor79, MS12]. 1972 [KHB^{+}72]. 1977 [BEK12]. 1980s [Bun81, CG22, MS12]. 1984 [Kri01]. 1990s [Div98]. 19th [Lub02, Von73, Waa15].

2 [Fle20, Kis24]. 21st [Pra12]. 21st-Century [Pra12]. 241pp [Kis24].

3-16-147557-7 [Bre04].

47/1 [Wei09a]. 47/3 [Wei09b].

5 [Hen21, KHB^{+}72].

’60s [Fel19].

9 [Eck04]. 978 [Fle20, Hen21, Kis24, Son21]. 978-0-262-04855-2 [Til24]. 978-0-262-34232-2 [Kis24]. 978-0-262-53687-5 [Hen21]. 978-0-691-20205-1 [Son21]. 978-0-7456-6578-8 [Bro20]. 978-1-5261-2005-2 [Fle20]. 978-1-74210-520-8 [Pre23]. 978-1-74210-521-5 [Pre23].

= [DH64].

A. [Sta08]. Abby [Bra15]. Abduction [Ano67a]. abeyance [Hoo84]. ABM [Dot72]. Abortion [Woo09]. Above [SV14]. Abraham [Thi04]. Abroad [SJPA14, Xu08, Lak73, MR95, MRC95]. absorptive [Koi95]. Abstracting [HD25]. Abundant [Til24]. abuse [Kem06, Sad91]. Academe [Kog99, Ash96]. Academia [AR20, KBK13, LK22, SV10, Vel18, Pét95]. Academic [AKW70, Ano66a, Ano67b, Ano80a, Ano94, BG23a, Bau20, BGB18, Ben13, Bét81, BGS16, Blo97, BG23b, Boz75, Can11, CSKF85, Con77, DCJ22, ESC^{+}88, EK98, EV12, FD04, Fel09, Fin07, Gei88, GL22, HR15, Har10, Hes23, HvL11, Hoo03, Hor22, HS99, Hun85, Jac09, Joh85, Joh09, Kaa11, Kai13, KP02, KE01, Ker89, KHS24, Kor20, Law81, MT19, Maz05, Mic78, MTE15, Mos08, MK19, MB12, Näs25, NS25, Nov17, Orl86, OCW20, OT20, PJLP16, Pri98, Rey21, Rim98, Rou19, Rus94, S.71a, STH^{+}77, STK^{+}91, Sch05, Sch07, Sco07, Shi95, Smi07, Str64, TH24, THM17, TMP10, Tel22, TP14, TT24, VSR14, Waa15, Whi09, WHF18, ZW25, ZJE10, Ama70, Ano69c, Ano68l, Ano69-32, Ano70a, Ano71c, Ano73c, Ano74b, Ano76d, Ano83d, Ano86d]. academic [Ano90d, Ano90e, AA66, Ash69, Aus80, Bis73, Cra67b, Cra67a, DK74, Etz83, EP91, Etz94, Etz96, Eus95, Gil69, Gil72, Hoo84, Hun81, Hut72, Joh72a, Kie85, LP79, Lob84a, MLS91, Mel66, Mom83, Mon95, Moo96, Muc85, Nol83, O’B66, PP68, Pét96a, Pra67, Puc72, Rie83, Ros72, Rue83, Rüe86, S.70c, S.74b, S.77, SJ77, S.83, S.88a, Sha73, Sha89b, Shi69, Shi75, Shi94b, SH07, Tha80, Ula83, Wan95, Was88, Web77, Wil83b]. Academic- [Hoo03]. Academicians [RYI13]. Academics [AR20, Näs21]. Académie [Cro81, Cro86]. Academy [Ano65b, Cao98, Cro78, FAGH01, For92, Har04, Her21, Küh84, Lan87, NW04, Thi04, Win94, Cha68, Hav66, Kön04, Lew77]. Academy/Industry/State [FAGH01]. accelerator [Gib70]. accelerators [MI81]. Acceptable [EB22]. Access [FBF^{+}21, Tau21, TSB24, Tro00]. Accidents [Tur10]. accommodations [Kor62]. Account [Fis07]. Accountability [SK00, ML83, Nic77, Smi96]. Accounting [RdR15, Sto66, SSG68]. Accounts [Bel68b, Bow67a, Jam68, Ano67x, oC68]. Accreditation [Far01, MBG^{+}92, OS21, Orl92, SV14, SAJ92]. Accumulate [HFSdR19]. Accumulation [BC20]. achievement [Ano71g, Rue83]. Across [GFM21, HJ25, JM19, KS01]. Act [Ano66j, OS21, Ros72, SH18, Ano65p]. acting [Ano67-40, Ber11]. Action [Ano69n, Bun81, CDRH25, CÁCC20, Par02, Woo09, Ano69-46, Ano88a, Ano90c, BCL76, Lit07, MBG^{+}92, S.88b, SAJ92]. actions [Ano69p]. Activism [MKK22]. Activities [VP22]. Activity [BFGS14, CG22, GMY73]. Actorhood [Whi12]. Actors [SDvL19, Sta02b]. Ad [Ano83e, Mor85]. Adair [Fle20]. Adair-Toteff [Fle20]. Adamant [Ano68k]. Adams [Ano67m, Ano67l, Dup90]. adaptability [Phi78, Ron79]. Adaptation [De 97, TC05, Bus63, Ric65]. Added [Tob19]. Address [Ano64a, Ano65a, Ano66b, Ano67c, Ano67-39, Ano65u, Ano67-40]. Administration [Ano62a, Ano62b, WM20]. Administrative [Som92]. admission [Mer79]. Admissions [Orl99]. ado [Lit07]. Adolf [Tru06]. Adoptions [GOR25]. Adrian [Bel97]. Advanced [Nel21]. Advancement [BPM84, Koh78, Shi92c]. Advantage [Tro00]. Advent [End09, Mor82]. adventure [Kol78]. adventurous [Wet85]. Adversity [Cue97]. Advice [Bri06, FM10, McG79, Cla74, Dot72, Ezr80, MA69, Rav78]. adviser [Jon72, Mis73]. Advisers [Two99, Abe65, Bro65, Mor92, Nic72]. Advising [CG02]. Advisor [PK09]. Advisory [GP77, Het90, SL24, KHB^{+}72, Som92]. Advocacy [MS12, Orl76]. Aeronautics [Het90]. affair [Ano65m, Ano69-43, Doe70, SNSB^{+}80]. Affairs [Fox97, Wei12b, Wei15]. Affect [Wei15]. Affected [BG23b]. Affective [vdB19]. Affinities [Tra05]. Affirmative [Ano88a, Bun81, S.88b, Ano90c, MBG^{+}92, SAJ92]. Affordances [FSHdR18]. Africa [Kis24, Ano69c, Ano71d, Ano86c, Ano88b, AA66, Boz75, Cou79, Cra66, Cra67b, Cra67a, ED91, HBW^{+}18, Hys64, Kah16, Lee07, MT67, Mel63, O’B66, PvR16, Pra67, Rot86, S.86, Sub66]. African [Cha68, Aic63, Ano89, Bio63, CS63, GF02, Per71, Ric65]. Africans [FG20]. After [Pet08b, Ano69m, Gei97, Häf97, Hoo66, Mor66b, S.88c, Spa66, Sto65b]. Aftermath [Guo13, Ano67d, Ano68a, Ano68g, Rei07, Sei88]. against [Ano66u, Gro85]. Age [Aro12, BPM84, Pei84, Rad19, Tyf12, Waa15, Koh93]. Agencies [SV14, Gra66, Per71]. Agency [FH13, Nel21, VS20, Ano67h, Ano67i]. Agency-Energy [Nel21]. Agenda [CÁCC20, Fra22, GMD^{+}24, LT16, MBS14, VS20]. Agendas [Lev98]. agents [Sch87b]. ages [SSG68]. aggressive [Ano69p]. agitation [Ano66c, Ano66d, Ano66n, Ano69-29, ESC^{+}88]. agitprop [DF96]. ago [Ano68i]. Agreement [Zim05]. Agreements [SSK21]. Agricultural [Cam81, Cas05, DeJ93, Fin07, Hei06, KBD75, Per71, Sch83, Ves94]. aid [Ano88b]. aims [Bic74]. Air [Het91, SJ77, Web77]. Akademgorodok [Gel00]. Alan [Sch03a, Sol03]. Albert [Hol81]. Alchemy [Abb09]. Alcohol [EB22]. Alexander [Ano66b, Ano67-39]. Alignment [MBS14, For74]. All-Knowing [Mat08]. Allegations [Maz89, GDHH85]. Allen [Bel97]. Allied [Mac99b, Ric88]. Allison [Bel97, Gus94]. allocation [Mon88]. Alternative [Rad19, Wil20]. Alternatives [Roy84, S.75a]. Althoff [Bro91]. Always [IH25]. Am [GCD23]. Ambiguities [Mad18]. Ambitions [Ban14]. America [Ano67b, Ano69r, Ber05, Bra92, Car88, Eng99, Hug06, KL94, Mar20, Orl76, Par02, Sch86, Ves88]. American [Mar06, Van15, Ano67e, Ano68-29, Ano75f, Ano76d, Ano86c, Ass93, BW70, BK70, BBSS03, Bow67b, Bri05, Cas05, Cue97, Etz83, Fel09, Fox97, Gei88, Gro82, Hol81, Hoo87, Kam09, KK81, KE01, KOT14, KV19, Lad70, Lag87, Lak73, Lev98, Mac00a, Maz89, McC76, Mer24, Mor82, Orl92, PP68, POSC07, PJLP16, Rei97, S.75b, S.83, S.84b, Sap71, ST96, Sch03a, Sha92a, Shi73, Shi70, Smi79, SK00, Sör06, Ste74, Tro00, Tsv14, Tur87b, Two99, Wil71, ZC75]. Among [CT13, FG20, HL10, Lau05, RYI13, THM17, Tau21, TSL19, Jor70, Mac99b]. AMP [Woo75]. Analysis [ACGG14, CT13, Fre23b, GFM21, Kaa11, Lem20, LBP10, Pan11, Pei11, Rob21, SBvE25, WR15, vR10, Lin81]. Analyzing [GOR25]. Ancient [BH15]. and/or [MSD21]. Andrea [Mel24]. Angelina [Ban15]. Anglo [Mar06]. animal [LM94]. Ann [Bel97]. Anne [Joh15]. anniversary [Ano69a]. anonymity [Ano75a]. Anti [Ano66c, BWF24, S.71b, SVMN15, Ano68y]. Anti-Nuclear [SVMN15]. anti-parliamentary [Ano68y]. Anti-Racism [BWF24]. Anti-science [S.71b]. Anti-Walloon [Ano66c]. Anticipating [Gus12]. antidote [Hey87]. antigens [Löw87]. anxieties [Lub07]. Anxiety [EL22]. Anxiety-Ridden [EL22]. Aotearoa [MMMS^{+}23]. Aotearoa/ [MMMS^{+}23]. Apart [Roz05, SRTC10]. Apollo [Het97, Het97, Vol21b]. applicable [SH02]. Application [Log09, Gib70, Hod80, Muk64, Nay67, ST93, Sne83]. Applications [HvL11]. Applied [Ort16, Pal90, Ves84, Wei65b]. Applying [Ort16]. Appointment [Con77, STH^{+}77, Ano66z, Ano71c, Ano76d, Ano83d, Ano90d, Ano90e, S.71a, S.77]. Appointments [Ano83f, Hoo66, Mor66b, Spa66, Ano83e, CSKF85, Sto65b]. Appraisal [Näs21]. Approach [CHJZSG20, EL22, FBGGH24, OPBCM15, OT20, TMP10, GP77, Het91]. Approaches [HLE24, Sör06, Wil20]. AQUAMETH [LBP10]. Arab [Nou11, Oky70, Waa70, Zah70, Zah06]. Arabia [Maz05]. Arabian [KV19]. arcadia [Aus75]. Archaeological [Ano89]. Archaeologists [Ano89]. Archive [Ano93a]. Arena [Can11]. Argentina [BGB18, Bek13, Fel19, KL03, NS25]. Argentine [Ano66y, Ano67d]. argument [SG72]. Arguments [Rou24]. Armed [Sel08]. arms [Wei87a]. Aron [Wei64a]. Arons [Ano80a]. Arrangements [GV18]. Arrow [Goo99]. Art [Ano69b, Röd17, Rou19, Ano69-40]. Artefactual [MTE15]. article [Bla67, Ric65, Wei69]. Articulate [BFdR25]. Arts [OT20, Win94]. ARWU [CT13]. ascendancy [Shi78]. ASEAN [Sto11]. Ashby [Ahm89]. Ashok [Ban15]. Asia [Ano63k, Fis63, S.86, Sil64]. Asian [Aic63, Lev98]. Ask [Vel18]. aspect [Cou79]. Aspects [Gro80, Aga92, Ano64r, Aro64, Joh72c, Kid83, Psa71, Psa72, Sco71, Sei81]. Asquith [Ahm89]. Assessing [Chr08, SRTC10]. Assessment [Cop02, Cro86, AA88, Ano75a, Ano90d, Ano90e, Cro81, RR71, Tri71a]. Assessments [Weh15]. assistance [Mor64]. associates [Mac94]. Associations [DRC^{+}14, Lew84, Var81]. Astronautics [Ano65b]. Astronomy [Hei17, Tau21, VGM^{+}24, Bar15]. Astrophysics [Hei11]. at-Risk [ZW25]. Atom [JK09]. Atomic [BPM84, Pei84, Ste74]. atomist [Shi80]. Atoms [PFC18]. atque [Mac07a]. Attempt [Pui05]. Attempts [Bar98, Aue80]. attrition [Ano88b]. audiences [Dub63]. Augustus [Gli05]. Australia [Ano66k, Enc71, Fra91, Kuk11, TIC02, WV17]. Australian [CA06, Enc00, KP02, Lin81, Phi78, Ron79, WC18]. Austrian [HN77, McL99]. Authoritarian [Puc72]. Authoritarianism [KV19]. Authority [CCSM18, Lub04, Rog13, Whi11, WGL18, Dei79]. Authors [Ano00c, Ano01d, Ano02c, Ano08, Tau21, YB16]. Authorship [JM19]. Autobiographical [Sti12]. autobiographies [Arm07]. Autobiography [Bel97, Bra07b]. Automatic [Hes23]. Autonomy [AA66, CCSM18, Gam97, Nic77, O’B66, Pri78, Pri98, SV14, Ano66y, Ano66-27, Ano96k, Dra84, Küh84, Mic78, Muc85, RLH91, S.71a, ST93]. Ave [Mac07a]. Averted [PL19]. Aviation [Dow17, Dow18].

B [Goo99, Hop11]. Bacon [Ans03]. Baker [Cha24]. balance [Mil72, Pry71]. Balancing [Sto72]. Balfour [Mac08]. Banaras [Ano66g]. Bandola [Fra22]. Bandola-Gill [Fra22]. Bangladesh [MJ78]. Bank [Sul91]. Barbara [Smi08]. Barcelona [Sal05]. Barriers [CHJZSG20, Hor22]. Barthe [Ber11]. Based [EL22, FFL^{+}15, Fre23b, GP04, Etz93]. Basic [Ano65b, Cal04, Kay66a, Pie12, Sap17, Sch14, Woo84, Joh65b, Joh71b, Nic77, Rot66, Sha84, Tou67, Wei65b, Wei71]. Battles [Xie21]. Bayh [Val11]. Bayley [Mac99a]. be [Col98, Des09, ED91, ML86, Sch78, SB13]. Beardsley [BB81]. Became [Ben13, áRH23]. Becoming [SBvE25]. Bee [Rod12]. Beek [Fre23a]. Before [CG22]. Begin [GCS00]. Beginning [Kra06, Ano69m, Rei07]. Behavior [MEM20, Pan11]. Behaviour [NM16]. Behind [Kra18, Mel24]. Being [Jun22]. belated [Ano69w]. Belgium [Ano66c, PD17]. beliefs [Orl96]. Bell [Hod80]. Belle [Pan93]. Ben [Fre87, S.87, Sch93, Shi87]. Ben-David [Fre87, S.87, Sch93, Shi87]. Benchmarking [Bru09, ZW25]. Benefits [WR15, Joh71b, Wei71]. Bennett [Pre23]. Benno [Ano68g]. Benoît [Wil20]. Berg [Eck04]. Berkeley [Ada09, S.73b]. Berkeleys [Bre95, Rot04]. Berlin [Ano80a, DF75, Pla99, S.72c, SNSB^{+}80]. Berman [Ben13]. Berne [Rue83]. Bestowing [EL22]. better [Mon95, Wil83a]. Between [BB24, DCJ22, EV12, FG20, HJ16, HBW^{+}18, Hoo08, Jac08, Küh84, MS13, Näs25, NS25, RG21, VV12, Vav98, Vol21b, vR10, Ano77e, Ano94, Ash66, Aue65, Blo97, CMR95, CH85, Cro94, Daa74, HLE24, Hut72, Kid66, Kön07, LP79, Maz73, Mel24, Pan93, Ped77, Rai96, Rob94, Spe85, Wei93b, Wil07, dGBM^{+}21]. Beyond [BA20, DT12, GS05, KHS24, Kra18, Lyn03, SL22, SL24, Val11, WG97, Eis87]. bias [SS79, Wil20]. Bibliometric [AG25]. Bibliometrics [Rey21]. Big [Col00, Ras02, Rül23, Raž94, Doe70]. Bigger [Ras02]. bill [Ano69-31]. binary [Tom96]. Biochemistry [Tru06]. Biodiversity [PKvdW23]. Bioethics [Sel08, Bal99]. Biographical [AA04, vR10]. Biographies [McL99, Arm07]. Bioinformatics [LBA16]. Biological [Bal02, GF02, Lit07]. Biológicas [SM97]. Biologists [Ros02]. Biology [BB24, KL03, Ros02, Jud80, Koh76, Mul72]. Biomedical [ALH09, Bra93, Lai19, MKK22, Pol05, Ras02, RdR15, SM97, Gra66, Löw87, Rei80, Wei65a]. Biomedicine [HSR19]. Biosciences [McG16]. Biosecurity [MMMS^{+}23]. Biotechnologies [FP17]. Biotechnology [NB02, Hen21]. Birth [Bos05, Het97, KL03, NBGT05, Bal99]. Black [And97, CBKN17, Gam97]. Blindness [Wil20]. Blockbuster [Ras06]. Board [Koh85, Kro03, KHB^{+}72, Plo75]. bodies [Lim95]. Body [Mon98, Ano69-32]. Boffins [Rol08]. Bogus [TT24]. Bohm [For08]. Bok [Orl99]. Bologna [GL22]. bolshevisation [Con96]. bomb [Gro96b, Gis04]. Bombay [Alt70, Don71, Par71a]. Book [Abb09, Akm24, AE77, ATIS70, Ano82a, Ano96a, Ano96b, Ano96c, AWZ79, AOT81, AS13, Bal99, Ban15, BBZA73, BM94, Bel97, Bel98a, BDK77, Ben98, Ben13, Ber11, Bro20, BPM84, BMSC87, BWW90, Cap98, Cha24, Coh98, CGTPM72, CSGZ^{+}74, CBJ95, DBL^{+}92, Div98, DBDM77, Dup84, End09, Eng24, Eus97, Fle20, FSH88, FHE94, FSMH75, FGT80, FS81, Fra22, Fre23a, FFML95, FW94, GG73, Goo99, GPM73, Gra12, Gre05, GNNBD84, GT89, GC91, Gro99, HRKZ71, HCRG88, Ham93, HD93, HR90, HC71, Hop11, Jen97, Joh15, Jos90, KKF95, Kel80, KZP^{+}74, Kis24, Kri97, Kro23, KBD72, LWZB83, LZ87, LC72, LJ93, Mar20, MDG95, MP78, MR79, MPM^{+}76, MM88, Mer24, MZP75, MA75, MDH94, MJSB85, MNS75, NGZG74, Orl91, Orl99, PCW79, PSNB80, Pei84, PPF89, PHA83, PK85, Pra12]. Book [Pre23, RBM73, RH89, Rim98, RSR^{+}70, Rob98, Rog13, RC88, RAV76, RH90, SZL81, SCJP71, SGM92, SRM92, Smi08, Son21, SS77, SH78, Stu98, SSS^{+}71, TKMB72, Til24, TRGN74, TGGS83, TSH86, TK86, Vol21a, VMD^{+}91, VLJG93, Wal10, Weh15, WKC^{+}70, WGN80, Wei78c, WG85, Whi09, WS91, Wil20, Wil85, ZWJK78, Zim98, ZJ81, ZLD^{+}76, pHPP^{+}79]. Books [Dan67, Har02, Het97]. Borderless [Ano01b, Dav01, Far01, Fie01, Mid01]. Borders [KS01]. born [Sch78]. bottleneck [Mor89]. Boundaries [FP17, MG24, Ole13, RG21, KL94]. Boundary [ALH09, Chl23, ELSS24, Kle23, NSB25, Röd17, Sap17, Vuo15, vR10]. Boundary-Work [ALH09, ELSS24, NSB25, Sap17]. Bounding [VV12]. Bounds [WG97]. Bourdieu [AK11, Cam11, Hes11, McG16]. Bowen [Orl99]. Box [CBKN17]. Boyd [Smi99]. Bragg [Jen01]. Brain [SM25, Shi18, Fri69, Gru68, Gru76, Joh65a, Joh66, Kid65, Mui69, Psa71, Psa72, Sco71, Smi79, Tho69, WG74]. Branco [Ano65c, Ano67-41]. Brazil [Ano67-41, DV98, PS01, Sá05, Sch78, Sch94]. Brazilian [Ano65c]. Breakdown [Bus63]. Breakers [Cai02]. breeder [Wei78b]. breeding [Pal90]. Bricolage [BG23a]. Bridges [Sol03]. Bridging [BWF24, Elz12, Goñ25]. brief [Ano69x]. Briefen [Bre04]. Bringing [AK11, PKvdW23]. Britain [Rob98, Ano68x, Ano69s, Arn00, AA66, Bra07a, Bra93, Div94, Dra84, Fis78, Gis04, Gun67, Hut70b, Kei81, Kin97, Lee98, Mac70, Mac71, Maz97, Moo99, O’B66, Pal90, Ric88, Ros67b, ST93, Var70, Von73, Wal96, Wal98, Woo09]. British [Ann82, Ano63g, Ano63l, Ano63m, Bal02, Bel67, Bel94, Blu69, BBSS03, CW64, Coo67b, Cox94, Div91, Fis07, Goo02, GP77, Hab73, JS65, Kam09, Ked93a, Ked93b, KMB24, Kol85, Mac00b, Mac94, Mos78, Pau67, Pri78, Shi94a, Var81, Wil97, Wil66, WB65, Zim06]. broadcasters [Rot90]. broadens [Ano67-35]. Broker [Bro08]. Brokerage [dGBM^{+}21]. Brokering [BG23a]. Brown [Hop11]. Bruce [Bel97, KHB^{+}72]. Buckminster [Ank07a]. budgets [BK70, Orl71b]. Buffer [Mon98]. Building [Gel00, LS21, MG24, Sol03, RSS97]. Buildings [CS15, Ano90a, Gol90]. Built [Mol05]. Burden [Fos97]. bureau [Dev80]. Bureaucracy [WS25]. Bureaucratic [HS99, WS25]. Bureaucratisation [GKL98]. Bureaucratization [Coc09]. Bureaucrats [Whi98]. Burgers [Som08b]. Burning [Mos08]. Bus [Ano67f]. Business [IH25, Met10, Ras02, Rya01, Tri94, Ano69-45]. Busy [Rod12]. Butenandt [Tru06]. Buzzwords [CÁCC20].

C [Bal99, Gem03, Het97, Mac99a, Mar20]. Caesar [S.72b]. Calculative [HdRW17]. California [Rai01, Ano67-33, Inn92, Mel79, Som92, SC04]. Call [AR20, Ham08]. Callahan [Ban15]. Calling [Fle20, Ano73c, FV79]. Callon [Ber11]. Cambodian [KMBC24]. Cambridge [Bel98b, Bro20, Hen21, Til24, Ano69-45, Ano75g, Cah05, HJ07, HJ08, Wal98, Wan95, Wil97, Zim63]. Came [Het97]. Campaign [Fre23b, Rot86]. Campbell [Pra12]. Campinas [DV98]. Campus [BWF24, CH24, Ano68-29]. campuses [Ano68b]. Can [Col98, Dot72, Dow17, Dow18, ED91, Pei11, Ros02]. Canada [Bis73, Doe70, FAGH01, Joh72a, Lem20, Ros72, SSK21]. Canadian [Ano73a, Aue80, CG22, FF90, Glo76, Gro71, Hal75, Joh65a, Law81, Por70]. Canal [Maz84]. Cancer [CBKN17]. candidates [Cro86]. Canton [Rue83]. Capabilities [TP16]. Capacity [HBW^{+}18, TP16, Koi95]. Capital [Bau20, BC20, Bro11, BBMR17, MARMSG20, Pan11, Joh67, Sco67, Tho67, Tho68]. Capitalism [Sch07, Gen23]. Captain [Ank07a]. Carayannis [Pra12]. carbuncles [Ano69-32]. Care [MMMS^{+}23, Sch25, Ano70j]. Career [BGP14, Kor20, Lau17, MK19, Näs21, SF25, Waa15, Bét90, Cro75]. Careers [BL16, MK19, Näs25, SL22, TH24, WHF18]. Carl [Bjö01]. Carmichael [Akm24]. Carnation [PFC18]. Carnegie [BPM84, Lag87, Par00]. Carol [Høy24]. Cartesian [Dub63]. Cartography [HH20b]. Case [And97, BBU19, BK18, HJ16, Hei11, Hei17, HH20a, Hes13, Kar24, Léc05, LAN14, MS11, MG24, MB10, Mül14, MB12, NSB25, PKvdW23, Rim98, Tob19, ZW25, Ano65o, Ano67g, Ano68o, AM74, BK70, DK74, Dot72, Enc70, Gib70, Gis69, Gis70, Gru70, Hal75, Hod80, Joh65a, Joh70, Joh73a, Löw87, Mel79, Mic78, Mon95, Mor92, Rey87, Ros67a, ST93, Sko69, Wal96, Woo72]. case-study [Hod80]. Castelo [Ano65c, Ano67-41]. Catalonia [Sal05]. Categories [PW09]. Catherine [Ben98]. Catholic [Ano67b]. cause [Ano69-34, Bak78]. Causes [BPM84, Glä24]. Cautious [HH20c]. CBW [SH02]. CBW-applicable [SH02]. CEEPAT [Sul95]. celebrate [S.92]. Censorship [SREM^{+}23]. Census [LB13]. centenary [S.92]. Centennial [Lin01]. Center [Het97, MK19, OCW20, Ano93a, Het97]. Centers [MEM20]. Central [DA93, Dup63, McG79, Pag02, GP77, Ano67h, Ano67i, Tom98, Wei93b]. Centralisation [HD25]. centralised [Sch83]. Centre [Ash96, Ros68, Von73, Ano70i, BD70, Cla70b, Cro70, Hun80, Rai96, RC67, Ros67a, Shi91]. Centres [EK98, KB99, Cha94, FAGH01]. Centricity [RvDG20]. Centro [SM97]. Century [And07, Bel98b, Cap98, Lub02, Pra12, Waa15, Ano68i, Ash67, Cro78, Cro79, Cro86, Gre87, Gro92, Hah75, LP79, Mor82, Par68, PK80, RSS97, Rüe92, S.92, Sha92a, Sha92b, Shi92b, Shi92c, Szc68, Tur87b, Von73, Wei87b, Fox97, Guo13, Neh06, Sut98]. CERN [Gib70]. certain [Ano64r, Seg85]. cessation [Ano67p]. Ceylon [Ano66j, Ano66-27, Pie64]. Chain [BPM84, Pei84]. Chairman [Pas78]. Chakanetsa [Kis24]. Challenge [DKCL^{+}23, Gar01, HH20a, TP16, Smi08]. Challenges [BDG10, Far01, Kal18a, Kay12, Kru04, Mid01, Häf74, Fra22]. Challenging [FP17, Pei11]. Cham [Akm24]. Chamber [Sal05]. Chancellor [Ano63s, Ano64a, Ano65a, Ano65u, Ano66b, Ano67c, Ano67-39, Ano67-40, Enc07]. Chancellors [Cla95]. chancellorship [Ano67-42]. Change [Aag17, BFGS14, BBU19, BDG10, Glä24, Goo02, Hor22, Kri97, KBK13, MT19, McG16, VS20, Wei99, WC18, KPL^{+}89]. Changer [Rey21]. Changes [Ash00, FAGH01, Kru04, Mül14, Sch25, Ano66x]. Changing [Ban14, Bel98a, Bja01, CÁCC20, End09, Ged95, Gib00, GV18, Kal18b, Lub02, Moo07, Rog13, Whi11, WGL18, Xu08, Car78, ST96, KKK10]. Chapman [Sta08]. chapter [Kol83, Mos78]. Chapters [Wei93a]. Characteristics [BZ21]. Charging [BM18]. Charles [Bak87]. Charmed [DT12]. Charrette [HR09]. Chemical [GF02, Joh02, Mac00b, Tra05, AA88, Cra66, Div94, Joh85, LP79, Maz90, Sub66]. chemicals [Hab73]. Chemistry [HvL11, Log11, MB12, VSR14, LP79, Shi80, Abb09]. Chest [VL17]. Chicago [Mol05, Ano75f, Ano86d, Bul80, Din75, S.92]. Chile [Bar15, BBU19, BK18, KBWA21]. Chilean [UMC^{+}23]. China [Ano66f, Ano67p, Cao14, DM73, Guo13, QZL25, SLL22, Sut97, Sut98, THM17, WBL15, Xu08]. Chinese [Ano67p, Cao98, Cao04, Hay93, Hay94, Kim82, Küh84, SSF15, TH24, XOR21]. Choice [HJ12, Mad64, Wei00, Zim00a, AM74, Dub65, Gil77, Kay66b, Mac76, MI81, Orl72, Roy84, Sne83, Tou64, Tou66b, Tou66a, Wei63, Wei64b, Wei65b, Wei65a, Wei66, Wei69, Wei84, WE72, Zim87]. Choose [CL22]. Christian [Eng24]. Christopher [Fle20]. Chronicle [Ano62d, Ano63a, Ano63b, Ano63c, Ano64b, Ano64c, Ano64d, Ano64e, Ano65d, Ano65e, Ari63]. Chunyan [Cai20]. Churchill [Wal98]. Chutzpah [Wil07]. Chymistry [Abb09]. Chymists [Abb09]. circle [Sch86, DT12]. Circulation [MS13, Joh67, Sco67, Tho67, Tho68]. Citation [KMB24]. Citations [Woe13]. Citizen [AD21, HLE24, HL22, Jas03, S.76]. citizenship [Lob84a, Rue83, Wei78a]. City [Bre95, Bro00, Ano66-28, Ano69d]. Civic [Ano69c, HL22]. Civil [Ano68b, Bun81, MA69, McG72, Cla70a, Hut69]. Civilian [Two99, WBJ^{+}77]. Civility [Mac16]. Claim [God14]. Claiming [HJ25]. claims [Moo96]. Clapperton [Kis24]. Clark [Rot95, SC04]. class [Hut70a, Jon71, Lub07, Var71]. Classic [HJ12]. classical [Eis87]. cleansing [Ano68c]. Cleft [Cam11]. Client [TIC02]. Climate [BBU19, Gra12, Ibr23, KB99, MH16, Mey11, Nol99a, Rad19, Röd17, UMC^{+}23, Wei99]. Clinical [CBKN17, SM25]. Clinician [ALH09, HSR19, dGBM^{+}21]. Closing [Kön07, OT20]. cloth [Gre05]. Cluster [CBB20]. Co [And22, LNST20, Var81, YB16, Hun80]. Co-Authors [YB16]. Co-existing [LNST20]. co-operation [Hun80]. Co-operative [Var81]. Co-production [And22]. Coalitions [Wei99]. coast [Ano69-27]. Code [Cai02, CM76]. Codes [And22, SH02]. Codespeak [LS21]. Coefficients [HL10]. Cognitive [Hei11, MS13]. Cohen [Sta08]. Cold [AS13, Rim98, Ben07a, Aro12, Ber05, Bor21, Elz12, Eng24, Fel19, Gro96b, Lyn03, PP25, Som08a]. Collaboration [Ash66, Ban15, BG23b, BA20, HR09, Hei11, Kid66, PP25, Ros67a]. Collaborations [Cha24, CBB20, GCS00, Ole13, Sku19, Thu10]. Collaborative [MEM20]. Collaborators [YB16]. Collective [Jac08, BCL76]. Collectives [For08]. Collectivisation [EK98, Loc89]. College [Ano67n, Lea98, Orl99, Wal98, Ano66a, Zim63, Par71a, Wal96, Wer07]. Colleges [And97, Car88, KKK10, Alt70, Ano76d, Don71, Lim95, S.83]. Collider [HK00]. colonial [Rai96]. Colonialism [Mel66]. Columbia [Ano69f, Ano68p, McG72, Wal92]. Come [Cas96]. Coming [Gar01, Sch86, Waa15, Hol81]. Commandos [Cai02]. comment [Eus95, Fra78, Fre67, Pas78, Plo75]. Commentary [Zim00a]. Comments [BBD^{+}92, BMR^{+}92, DKM^{+}92, Joh71a, Hun80, WE72]. Commercial [CTLL22, SSF15, Kei81]. Commercialization [Col07, KOT14, Nov17, Ren08, Sjö07]. Commission [Ano62g, Ano64n, Bod64, Bun81, Coh98, Gus94, Ano65i, MTS20, Pas78]. Commissions [Ano68z, Ano65n, Ano66l]. commitment [Etz96]. Committee [Ano65q, Ano83e, Ash68, Bel68b, Bjö01, Bow67a, Bul80, Het90, Hut75, Jam68, Kin97, MA69, Sha87, Ano64r, Ano65p, Ano86d, Mor85, S.90, oC68, Ano65b, Ano83f, Cla95, Rai64]. Committees [ELSS24, Som92]. Commodification [Jac09]. Commonalities [Kog97]. Commonwealth [Lee06]. commotion [Ano69-37]. Communication [EBMP24, Mac99a, VV12, XH22, Mel79, S.84b, Woo75]. communist [Jor70, Pét93, RLH91]. Communists [Hoo87]. Communities [AG25, Hes23, HD22, KBWA21, Sto11, Hun80, Mac76]. Community [Bul80, HLE24, HBW^{+}18, Hol81, Mac16, Pig68, Zim77, Doe70, Dub63, Enc70, Mad64, Pan93, Sch91, Sch78, SB80, Sko69, Ves84, Von73]. Comparability [LB13]. Comparative [SSK21, WBW18, Gra66]. Comparing [FSHdR18, MT19, PD17, SL18, SL24]. Comparison [Ash00, GM03b, HvL11, Kor20, KB99, VSR14, Div94]. compensation [Sch06]. Competencies [SF25]. Competing [OCW20, Ste74]. Competition [Bjö01, Can11, FFM16, Gar01, Vol21b, Gas71, Pol72]. Competitions [Rou24]. Competitive [BGP14]. Competitiveness [Bru09, WO16]. Complementarity [Bon10, Jac08, Tau21]. Complementary [SRTC10]. Completed [KKK10]. Complex [Cao04]. complexity [Kay66b, Tou64, Tou66b, Tou66a, Wei66]. Composite [ZW25]. Computer [Gra12]. Computing [HBW^{+}18, NSB25]. Concentration [BC17, KMB24, NS25]. Concept [Fre23b, HJ12, PJLP16, Ray09, Eam85]. Conception [SGSB23]. Conceptions [BI22, Som08a]. Concepts [Fel19, FP18]. Conceptual [EKB22]. Conceptualizing [FH13, MO17]. concerning [Dub63]. concert [Ano69q]. concession [Ano69w]. Concordia [Mon95]. condemnation [Ano67-27]. Condition [EK01, FV79, S.84a, Sch79]. Conditions [EKB22, Whi12, Ano68l, BD65, PK80]. Condon [Mar20]. Conference [AD21, Ano69n, Ano67n, Hal63, Ano63s]. Conferences [Ano65n, Ano66l]. confidentiality [Ano75a, Ano90d, Ano90e]. Configurational [Pei11]. conflagration [Ano69d]. Conflict [FG20, Hei17, KOT14, Rül23, Sel08]. Conflicting [Hay94, Näs25]. Conflicts [BWF24, Etz96, HV23, YB16]. confrontation [Ano69b]. Congestion [BM18]. Congregation [Ano65a, Ano64a, Ano66b, Ano67c, Ano67-39]. congress [PK80, Ano89, Aro12]. Congressional [MS12]. Congressmen [Gus94]. CONICET [Bek13, NS25]. Conjuration [Vos23a, Vos23b]. Connecting [Met10]. connection [Ben07b]. Conor [Ano64a, Ano65a]. Conquest [Gla08]. Conrad [Ano94, Eus95]. Consensus [Mac16]. Consequences [Fra22, MARMSG20, Orl99, Rog13, SLL22, Ano74b]. Conservation [Bru25, Cam81, Kro03]. conservative [Wet85]. Considerations [PP68]. Considering [Orl99]. Consolations [Edl24]. Conspiracy [Vos23a, Vos23b]. Constituting [RvDG20]. constitution [S.83]. constitutional [Hun81]. Constraints [FSHdR18, SV14]. Constructing [Rai01]. Construction [EL22, Edl24, LB13, SKS13]. Constructions [KB23]. Constructivism [Hoo08]. Consultants [SV10, WM20]. Consultation [AA71]. Consuming [Rou24]. Contagious [Woo09]. Containing [JK09]. Contamination [Col07]. Contemporary [BT19, Bor04, DRC^{+}14, HI06, Mül14, PJLP16, S.88a, Sha89b, Shi92a, Was88]. Contents [Ano01a, Ano02a]. Contest [Guo13]. Context [ACGG14, CHJZSG20, LNST20, Tom98, Cha94, Ezr80, Pap96]. Context-specific [LNST20]. Contexts [Ban14]. Continental [Dav01]. Continued [Ano67j, Ano67l]. continues [Ano67-28, Ano67-30]. Continuing [Ano66d, Ano69y, BC17]. Continuity [BEK12, For92]. contra [McG72]. Contract [Jen97, Vav98]. contracting [Ano62h, Rey87]. Contraction [Bek13, Hay88]. Contradiction [KOT14]. Contradictions [Woo03].

Contributing [Kaa11]. Contribution [God10, Roy84, Sch93]. Contributions [PD17, Ano96i, Bak78]. Contributors [Ano62e, Ano63d, Ano63e, Ano63f, Ano64f, Ano64g, Ano64h, Ano64i, Ano65f, Ano65g, Ano65h, Ano66e, Ano67k, Ano68d, Ano68e, Ano68f, Ano69e, Ano69z, Ano70b, Ano70c, Ano70d, Ano71a, Ano71b, Ano72a, Ano72b, Ano74a, Ano75b, Ano75c, Ano75d, Ano75e, Ano76a, Ano76b, Ano76c, Ano77a, Ano77b, Ano77c, Ano77d, Ano78a, Ano78b, Ano78c, Ano78d, Ano79a, Ano79b, Ano79c, Ano79d, Ano80b, Ano80c, Ano80d, Ano80e, Ano81a, Ano81b, Ano82b, Ano83a, Ano83b, Ano83c, Ano84a, Ano84b, Ano85a, Ano85b, Ano85c, Ano86a, Ano86b, Ano87a, Ano95a, Ano95b, Ano95c, Ano96d, Ano96e, Ano96f, Ano96g, Ano97, Ano98a, Ano99a, Ano99b, Ano99c, Ano00a]. Contributorship [YB16]. Control [Gam97, RR71, Sch83, Tri71a, Wei87a, BPM84]. Controlling [Mac88, Pét96a, Rei80]. Controls [EV12]. Controversies [Fra22]. Controversy [Jac06, Sol75, Vos23a, Vos23b, Vuo15, Xie21, Ano65k, Cla74, Inn92, PK80, Shi80]. Convention [Hoo08]. Conventional [Des09]. Convergence [CDRH25, GMD^{+}24, GFM21, WS25]. Conversational [Win20]. convocation [Ano65u]. Cooperation [Pap04, SSK21, Vol21b, WBL15, Aug63, Hut72]. Cooperative [LAN14, Mar20]. Coordination [Hes13]. Coping [ETMT70, FGR25]. Coproducing [Pfi15]. Coproduction [Win20]. Coral [Nef14]. Córdoba [Ano69a, Ano69r]. Core [OT20]. Cornford [Wil97]. Corona [HV23]. Corporate [Mer24, Rya01, VW95]. Corporation [Kei81, Lag87, LAN14, Par00]. Correction [Vos23b]. correctness [Cox94]. Correspondence [BW71, Ber71, ET67, Mel71, MBJ67, Par71b, Tri71b]. Cosmopolitan [Pan93]. cost [Gau73]. Costs [WO16]. Cottle [Goo99]. Could [Høy24]. Council [Ano63r, Ano65q, Bek13, DeJ93, Elz12, FM10, GP77, Kol83, ST93, Ano66t, Ano67-32, Ano62f, Kin97]. Councils [Sha89a]. counsellors [SM72]. count [Ben07b]. Counter [SREM^{+}23]. Counter-Tactics [SREM^{+}23]. Countering [Tob19]. Countries [Nol99a, Til00, TC05, WW12, Ano71h, AA66, Ded63, Gra66, Jor70, Lip64, Mor64, Mor66a, Osb71, Qua71, Sal66, Sch87a, Zim71]. country [And69, Bas67, Coo67a, Nay67, Nay73, Sch80a]. Counts [Har08]. Course [Enc06]. Courses [BL16]. Court [S.83]. courts [Maz77, Mer79]. Coverage [Lai19]. COVID [Fre23b, áRH23, Chl23, SREM^{+}23, Vos23a, Vos23b]. COVID-19 [Fre23b, áRH23, SREM^{+}23]. Create [YB16]. created [Kel84]. Creating [Bri05, Kin97, Ben13]. Creation [CHJZSG20, FAGH01, Hug06, Lem20, Ass93, Wid07]. Creativity [Lin01]. Credentials [Akm24, Cro81]. Credibility [EB22, HvL11, HFSdR19]. Crisis [Bro20, Enc00, JE97, Mus22, SLL19, SK00, Tyf12, Ano65v, Ano66-28, Ano67-33, Ano70e, Aro64]. Criteria [Con77, Des09, Dub65, FFL^{+}15, Gib70, Orl72, STH^{+}77, Wei63, Wei64b, Wei00, Zim00a, Ano71c, Ano76d, Ano83d, Gil77, MI81, S.77]. criterion [SL89, Sne83, Wei84]. Critical [GM09, Weh15, Bun81, Wil20]. Critique [PL19, Sha17, Lin81]. Croatia [Ano67y]. Crop [Kam09]. Crops [Bra15]. Crosland [Rob98]. Cross [BA20, HH20a, Kle23, McG16]. Cross-Disciplinary [BA20]. Cross-Field [McG16]. Cross-Sector [Kle23]. Cross-Sectoral [HH20a]. Crossdisciplinary [Kle23]. Crossed [RG21]. Crossing [Ole13]. Crossroads [KBWA21, Rou07, Hus77]. Cruise [Ano64a, Ano65a]. Culminations [MEM20]. cultivation [Ano71d]. Cultural [Aro12, BL16, Gem03, Gui06, Mül14, PP25, Rou07, Tyf12, Ano66f]. Culture [Bel98a, FD04, Hay98, Her13, Kog99, LBA16, Log09, Rob98, Tou66b, Zim98, Cum90, Kel85, Kim82, KPL^{+}89, Mol05]. Cultures [Hei17, HFSdR19, Lin01, XOR21, RLL84, Wan95, Wei64b]. Curious [Whi09]. Currents [Lau05]. Curtis [Goo99]. Curzon [Gho88]. cybernetics [Kol78]. Cycle [WV17, BCL76]. cyclic [Woo75]. Cyclical [FBGGH24]. Czech [KHS24, Win94]. Czechia [Chl23]. Czechoslovakia [Fre68, NZ68].

D [BPM84, Het97, Til24, Vol21a]. Dale [O’S01]. Damrosch [Bel98a]. danger [McG72, Ula83]. dangers [Rei80]. Danish [Ano77e, Kam09, KMB24, Ped77, PJLP16, SH18]. Dar [Ano67n]. Darsee [Mor85]. Daston [Sha17]. Data [Bar23, Gra12, áRH23, SDvL19]. Data-Political [áRH23]. Data-Sharing [Bar23]. Database [LBP10]. David [Bal99, Bel98a, Pra12, Cha24, For08, Fre87, S.87, Sch93, Shi87, Son21]. Davis [Som92]. day [Ano67-40]. Dayé [Eng24]. days [Sha87]. DC [Gre05, Gre05]. De-Facto [MN13]. De-scribing [FF10]. deal [Aue65]. dealing [Ano85d]. Death [Bal99, KL03, Ano68g]. debate [Dre71, Etz94, Sho72]. debates [Rav78]. Debray [Ano67g]. decade [S.72a]. Decentralised [Nay73]. Deceptive [JE97]. Decision [KHB^{+}72, Kön07, Log09, Rey87, Gil67, Ros67a, S.83]. Decision-Making [Kön07, Rey87, Gil67, Ros67a]. Decisions [Ano65i, DS16, FGR25, Hau17]. declaration [Ano69a]. declares [Ano90c]. Decline [BH15, Ano96k, BD70, Cla70b, Cro70, Muc85]. declining [Bét90]. Decoupling [HP19, SD08]. Dedisciplining [FB12]. Deep [BFdR25]. Deepening [Ano68g, MARMSG20]. Defence [Bal02, Two99]. Defending [Kri07]. Defense [Wei87a]. Defining [BM24, Sjö07]. Degrees [Akm24, HJ07, HJ08]. Delaware [Ano91]. Deliberation [HI06, MTS20]. Deliberative [Goñ25, HL22, TD10]. Delivery [Näs25]. demand [Gis69, Gis70, Gru70, Joh70, Woo72]. Demands [Fal21, Ano68k, Ano69q, Mic78, Rüe92]. Demise [Elz12, Kol83]. Democracy [Est04, Goñ25, Hop11, Kri22, Lan05, Lyn03, Mac97, MK17, Pra12, Sal00, TD10, Daa74, Hal75, Jon76, Jos94, Mac73, Shi89, Wei90, Fre23a, Ber11]. Democratic [Bro13, Gro99, Häf97, Hav66, Kar97, Pes09, Ano66p]. democratisation [Por70]. Democratizing [Kön07]. demographic [Sto66, SSG68]. demonstrations [Ano67f, Ano69v]. Demos [Tho10]. Denmark [LH19, áRH23, Sch80a, Sch87a]. Dennis [Hen21]. Department [Din75, Kal18b, Wal92, Bar98, Cla95, Eis13]. Departments [Kor20, Mah00, MB12, OT20]. departure [Shi92a, Shi92b]. Dependence [Har02]. Dependencies [Nef14, SL18]. Dependency [Lai20]. Dependent [TMP10]. Derangement [Ano69-27]. Derek [Orl99]. Describing [Str03]. Design [HR09, LB13, Div91]. desire [SH07]. destruction [Ano81c]. determinants [Zim81]. Dethoff [Het97]. deutsch [Bre04]. deutsch-englische [Bre04]. Deutscher [CG02]. Developed [WW12]. Developing [HD22, Til00, TC05, Lin81, Mor66a, Nay73, Osb71, Qua71, Sal66, Zim71]. Development [And97, BGB18, BGS16, Bow02, BPM84, FAGH01, GS05, GL22, MK19, SF25, Tau12, Tur98, Zap22, Ahm89, Ama70, Ano62h, Ano62f, Ano65j, Ano65r, Ano65s, Ano65t, Ano67-32, Ano73b, Bau77, Bie67, Bor67, Cou79, Cro75, Deb69, Dob66, Etz93, Fis78, Fre67, Fre68, FF90, Gra71, Hay94, Hun80, Joh74b, Kei81, Koi95, Lim74, MA69, Mul72, NZ68, Pav76, Per71, RD93, Sal91, Sil64, Tom96, Ano63j, Pra12]. developmental [Col86]. Developments [Tom98, WBL15, Sul95]. Devices [HdRW17, Näs21]. Dialogue [Goñ25, Gui06]. Diamond [TSB24]. Diary [Het97]. Dictating [Gli05]. Dictator [Gli05]. Dictatorship [Bek13]. Did [Tur98]. Diego [Rai01]. Differences [BZ21, Kog97, THM17]. Different [FSHdR18, Moo96]. Differentiation [Waa15]. Diffusion [Bra12]. digging [Ano68m]. digging-in [Ano68m]. Digital [Hes23, dSNC23, SGSB23, Til24, TC05]. dignity [Ano73c]. Dike [Ano67c]. Dilemma [Lag87, Woe15, AM74, S.74f, Wil71]. Dilemmas [FHS09, Gam97, Hay98, KE01, Sel08, Wil97]. dilution [Bis73, Joh72a, Ros72]. Dimension [Joh15]. Dip [MS11]. Dip-Pen [MS11]. Diplomacy [SSK21, Sto11]. Diplomat [Gem03]. direct [Wil86]. Directed [Kim08]. Direction [Bri05, FBH23]. Directions [DT12, Xu08, Cra86]. Disciplinary [BA20, GAL22, God14, GL22, VV12]. Discipline [Bey06, Bru09, Fox06, Hro13, Joh09, SM97, Din75, Mor82, Tha80]. Disciplines [GL22, KMB24, Mah00, Tel22, FV79]. disciplining [MK19]. Disconnect [MBS14]. Discontents [Mac97]. Discontinuity [BEK12]. Discourse [Kal18a, Kle23, PJLP16, Rod12, ZW25]. Discourses [Sto11]. Discovery [O’S01, Gas71, Pol72, Ben98]. Discrimination [Ack08, Tob88]. Discursive [KV19, Lem20]. Discussion [Des09, Eus95, KPL^{+}89]. Discussions [AC23, SGSB23, Ano72c, Gro92]. Disease [Fer23, Woo09]. dishonesty [Maz89]. disintegrating [Ker89]. disintegration [Ano68s]. disjunction [For92]. dismissal [Ano80a]. dismissed [Ano65l]. disorder [Ano67-30, Ano67-31, Ano68r, Ano69y, Ano69-39, Ano69-44, Ano74b, ETMT70, Sch72]. disorganisation [Mel74]. Dispersion [BC17, Jos94]. Displacement [Aag17, Ano69i]. Displaying [BM24]. Disputes [Chl23, HV23, Maz73]. disregarded [Ano69w]. disruption [AKW70, Ano69f]. disruptions [Ano90a, Gol90]. Disruptive [CBKN17]. Dissent [Har06, Pla99]. Dissonance [FG20]. distinctiveness [Elk77]. Distressing [Bal02]. distribution [Car63]. Disturbances [Ano66g, Ano69g, S.88c, Sei88]. diverse [Car78]. Diversification [Kog97]. Diversity [BWF24, FH13, GOR25, PGO24, MBG^{+}92, Orl92, Phi78, Ron79, SAJ92].

Diversity-Related [GOR25]. divestment [Ano86c, Rot86]. Divide [TC05]. Divided [Pos04]. Divides [HJ25]. Diving [BFdR25]. Divining [Col00]. Division [áRH23, Tom96]. Divisions [Sch03a]. DNA [AA04, Col07]. Do [Ano82e, DCJ22, Enc00, Kis24, Kor20, Shi94b, Wei15, ZJE10, Bra12, HSR19, Lau17, Léc05, SM25]. Docility [BRWB19]. Doctoral [BA20, ELSS24, FFM16, Jun22, MBS14, Tor18, Thu10]. Does [Ano90e, AR20, Col98, CBB20, Gel00, Ano90d]. dogmatism [RLH91]. Doing [SM25, SF25]. Dokumenten [Bre04]. Dole [Val11]. Domains [JM19]. dominance [Cam85]. Dominique [Wil20]. Dons [Tuc00]. Don’t [Ham08, UMC^{+}23]. Doomed [SB13]. down [Ano90d, Ano90e]. Dr [Ano66b, Ano67c, Ano67-39, Mor85]. Dr. [Ano64a, Ano65a, Ano67m, Ano67l, Des65, Par65]. Draft [Ano67n]. drain [Fri69, Gru68, Gru76, Joh65a, Joh66, Kid65, Mui69, Psa71, Psa72, Sco71, Smi79, Tho69, WG74]. Drawbacks [FSHdR18]. Drift [Fin07, Har10, Kai13]. Driven [Jun22, SBvE25]. Driving [Jun22]. Drug [EB22, Ras06]. DSIR [Hut70a, Var70]. Dual [DPB^{+}20]. Dueling [YB16]. Duress [Ked93a, Ked93b]. During [Ada09, Bek13, Har04, SBvE25, Bey06, Dev80, McG79, NW04, Par02]. Durkheim [Wei76]. Dutch [Ben98, Ban14, CH85, Daa74, HvL11, Loc89, Sne83, Sta02a]. dva [DH64]. dvazhcy [DH64]. dyestuffs [Hab73, LP79]. Dynamic [CT13]. Dynamics [Bon08, GL22, Har10, WC18]. Dysfunctional [Lai20]. dziady [Ano68a].

e-book [Pre23]. e-Science [Kar24]. Each [Bar23]. Early [Abb09, Ben98, For08, Näs21, RR05, Yeo02, Bra07a, Bul80, DF96, Hut70a, Mis73, Mos78, RLH91, Var70, Wei87b]. Earth [Ank07a]. Earthquake [Mar11]. East [Hys64, Tom98, Hay94, And22, Ano63k, Bor21, Cha68, Cou79, Cra66, Cra67a, De 97, Fis63, Gui06, HBW^{+}18, Hav66, Lev98, MT67, Oky70, Ole13, Pla99, Ree75, SL18, Sil64, Spe85, Sub66, Waa70, Zah70]. Eastern [Wei93b, Jor70, Kol04, TMP10]. Eaton [Akm24]. Eaves [Goo99]. école [Shi84]. Ecological [Bru25]. Ecology [BPM84, CH85, Duf66, Str64, BPM84]. Economic [Bec73, Ben13, Ful00, GS05, Joh74a, MBS14, Mis74, Pla99, Pol00, TMP10, WV17, Zim00b, Ano64o, Ano65s, Etz93, FF90, Glo76, Gro82, Hun80, Joh65b, Joh71b, Joh72c, MA69, Rud68, Wei71, Wil64]. Economically [WW12]. Economics [Ano69-29, MS12, Sco71, Ano67z, Ano69-36, Att73, Cra86, Joh65a, Joh66, Joh73a, Kid65, Pét96a, Psa71, Psa72, Sch80b, Ano67m, Ano67l]. Economies [AGF09, Sá05]. Economists [Cam81, Gre05, Kor20]. Economy [Bor04, Gro99, Met10, Sca03, Sco05, Tyf12, Woe13, Zap22, CW64, Pal90, Rot81]. Ecosystem [SDvL19]. Ecosystems [Bru25]. ed [Abb09, Eck04, Fre23a]. Eddington [Sta08]. edited [Bal99, Bel97, Coh98, Goo99, Gro99, Het97]. Edition [Cai20, Pos04]. Editorial [Hes23, HH20c, Sha00, Wei08, Wei09a, Wei09b, WT12]. eds [Akm24, AS13, End09, Eng24, Fle20, Gre05, Mel24, Mul20, Rog13, Smi08, Wal10, Weh15, Wil20]. Educated [Rou19]. Educating [Ski06]. Education [Ano63j, Ano65j, Ano66j, Ano73e, Ano01b, BBSS03, Bow02, BEK12, Bro11, BZ21, Bul02, Dav01, De 97, Div98, Far01, Fie01, Fin07, FH13, Gar01, GOR25, Har10, JS65, Jun22, KSC^{+}64, KV19, Kog97, Koh85, Kol83, KBK13, LBP10, LB13, Lev98, Met08, Mid01, MSD21, Moo07, Nou11, OB17, Opi04, Pou07, RBB13, Rya01, SC12, SRTC10, SLM22, Sco01, SD08, Str64, Sut98, Til00, Til24, Tom98, Tro00, TIC02, Umu86, Wal98, Wil97, Ahm89, Ann82, Ano66k, Ano69-31, Ano70g, Ano70h, Ano71f, Ano72c, Ano81d, Att73, Bic74, BCL76, Bow67b, Bro91, Cou79, DA93, Das64, Div94, Duf66, Gaj68, Gua88, Hus83, Kin63, Lay73, Lev73, Lew80, Lin81, Mat69, Mil72, Orl92, PH80, Pry71, Psa88, PS65b, PS65a, RLL84, Rat65, Rüe75, S.71c, S.74c]. education [S74e, S.82, S.84a, Sad91, Sal91, ST96, Sha89a, Sul95, Tom96, Tri94, Tro76, Wal96, Wei79, Wil86, Woo72, WB65, dGdMS^{+}69, Akm24, Cla95, Goo99, Mus22]. Educational [Bla67, Dou03, Hus83, SH02, Wan95, Ano66u, Ano69-38, Ano69-45, Ano84c, Hun80, Mot64, RG73, Sto65a, Sto66, Wil71]. Edward [Ano96l, Ben98, Fle20, Ano96h, Ano96i, Bul96, Hal00, Mac16, Mar20, Orl96, Sha95, Wei96]. Edwards [Gra12]. Effect [CTLL22, MKK22, Nov17, SH18, Ben07a]. Effective [NM16, Woe13, Cum90]. Effectiveness [CTLL22]. Effects [Glä24, Hor22, McG16, Tal09, Vol21b]. Efficiency [Psa88]. effort [Car63]. egalitarianism [RG73]. ego [Aus75]. Eight [BC20]. eighteenth [Hah75]. eighteenth-century [Hah75]. Einstein [GR06, Hol81, Low01]. Elaine [Akm24]. Election [Cao98, Cro81, Cro86]. electrical [Div94]. electron [MI81]. Electronics [Léc05]. eleven [Lob84b]. Elias [Pra12]. Elite [Cao98, HL10, Hei11, Lau05, Wei87b, Con96, Tro76]. Elites [Kar97, KMB24, Kor20, Tri79]. Elitism [S74a]. Elizabeth [Ben13]. Ellen [BPM84, Fra22]. Elusive [HL10]. Emergence [Lau17, Tom96, Vel18]. Emergency [Rab25]. Emergent [FG92, Smi79]. Emerging [CHJZSG20, Gus12, Rob21, Sá05, SMaR21, SGSB23, Sla11, Bra92]. Emigration [Gis69, Gis70, Gru70, Joh70, Ano63g, Ano63h, Wil66]. Émile [Wei76]. Eminent [HdRW17, Kor20]. Emory [Mor85]. Emotions [BM24]. Emphasis [CH24]. Empire [Far05, SJ05, Mac99a]. Empirical [CL22, GFM21, LBP10, RYI13]. Empiricism [Nye07, LM94, Ric07]. Employees [KHS24]. Employers [Div91, RBB13, Div94]. Employment [Can11, Kin63]. Empowerment [HL22]. Emulation [POSC07]. Enacting [PvR16]. Encyclopedic [Dol05]. End [SH18, Bak87, S.88c]. endangered [Tro76]. Endless [Ano98b]. endowment [Mac71]. Ends [HP19, Ano69o]. enemies [Ano94, Rob94, S.74b]. Energy [Hei17, Nel21, PFC18, PP25, Pui05, SVMN15, FG92, Häf74, HN77, MI81, Ste74, WBJ^{+}77, Eis13]. Engagement [AD21, FF10, Goñ25, HH20b, Sta08]. Engaging [BM18, HI06]. Engine [TMP10, Ben13]. Engineer [Mac00b]. Engineering [Ada09, BL16, Eis13, Ibr23, Joh09, Kal18b, Kön04, Kri01, MBS14, PD17, ST93, bre06, Div91, Gua88, RLL84, Bal99]. Engineers [Mac00b, Mac94, THM17, Div94, Koh93]. England [Div98, Gre87, KKK10, Lim95, Lub02]. englische [Bre04]. English [Ano65k, Ano67o, Ano68h, Ano69i, AA66, Gar01, ML86]. English-speaking [AA66]. Engwall [Rog13]. Enhance [CBB20]. Enhancing [QZL25, WW19]. Enigma [Eis13]. Enlightened [bre06]. Enlightenment [Som08a, DF96, Rob98]. Ennobling [Eis13]. Enoch [Ano69-35]. Entanglements [Eng24]. enterprise [Fox73, Woo84]. Enterprises [Etz93, Mar11]. enthusiasm [Ano69m, Hey87]. Entraide [Gui06]. Entrepreneur [Gem03, Shi18]. Entrepreneurial [CTLL22, DCJ22, Etz83, SB13]. Entrepreneurship [Kah16, Pra12, Cai20]. entry [Hod80]. Environment [BPM84, KV25, Neh06, Sör06, Coo72, PD71, ST96]. Environment-Focused [KV25]. Environmental [Ank07b, Chr08, Gre05, HR09, Low05, Maz97, Mel72, MK19, Rae06, Rit71, Roz04, Sch25]. Environments [BA20]. Epidemic [Kah16]. Epigenetic [SM25]. episode [Aue65, Wei79]. Epistemic [BFdR25, Dav20, Fal21, FSHdR18, Glä24, HLE24, HV23, Hei17, HL22, HFSdR19, Kai13, Pfi15, WM20]. Epistemological [BG23a, Mad18, MTE15]. Epistemologies [BB24]. époque [Pan93]. EPSRC’s [MPW16]. Equality [Dha84, KKK10, Lev73, Bét81]. Equity [HK11, NS25, Sla11, Psa88]. Era [BH15, BZ21, Kri00, Mer24]. Eric [Mer24]. Erika [Mul20]. Ernst [Het97]. Erratum [Ano05a, Dow18]. Erwin [McL99]. eSec [Kar24]. Esha [vdB19]. especially [ML81]. Essay [Ber11, Chr08]. essays [Ked93a, Ked93b]. Establish [Bar98]. Establishing [Mel63]. Establishment [Pet08b, SM97, Wal98, Ano66p, BD65, Bul80, Léc87, Mor82]. Esteve [RR05]. Estimating [Pou07]. Estonian [Tom96]. ETH [Kru08]. Ethic [Ker89, Hoo84, Kie85, Lob84a, Nol83, Rie83, Ula83, Wil83b]. Ethical [SH02, Seg85]. Ethics [Bal99, NB02, SDvL19, SD08, Tal09, Mon95]. Ethos [BT19, Wei87a, Rüe86, Shi75]. Etzkowitz [Cai20]. EU [Kön07, Tal09, TD10, ZW25]. EUR [Bre04]. Europe [Aug63, BD65, Ber05, Bro00, Bux03, Cra86, DA93, Dav01, HI06, Kra06, Kru04, Orl76, Pag02, Rai96, SLL19, TMP10, Wei93b, Yeo02]. European [Ano68i, Ban14, CÁCC20, Fox06, Geu98, GP04, Hau17, Jor70, Kol04, LB13, MTS20, MG24, Nol99a, Nol99b, Pap04, Pfi15, PGO24, Pig68, RvDG20, Sör06, Spe85, Sta03, Tom98]. EUropeanisation [CR25]. Européenne [Gui06]. Euros [Edl24]. evaluate [Mor85]. Evaluating [SRTC10]. Evaluation [ACGG14, BGS16, BS11, BM24, Dem17, DS16, ELSS24, End09, GM03b, Hau17, Her21, HH20c, OS21, SLL22, Tob19, VP22, MI81, Mor71a, Sto72, ZM71]. Evaluations [EL22]. Evaluative [FFM16]. Evangelism [RG73]. Evaporation [Ano69m]. Evelyn [Gro99]. Events [PvR16, Wal98]. Everything [HLE24, ML86]. Evidence [BI22, Joh19, FG07]. Evolution [Aag17, DR12, HLE24, Har02, Kim08, Kit10, TT24, Vol21b, EP91, Tur87b, PL19]. Evolutionary [BB24, Kim08]. Examination [DGR19, ELSS24]. Examinations [Ano69-28]. Examining [OB17]. example [CG02, Jon66, SSG68]. Excellence [FP18, HJ16, HK11, Kit10, LH19, Mul20, NS25, Rot95, FAGH01, WO16]. Excellent [HdRW17]. Exchange [Pan11, Ano67p, MR95, MRC95]. exchanges [Rai96]. Exclusion [WO16, ZW25]. Executive [Gre05, Alb63]. Exhibition [Lin01]. existing [LNST20]. Exit [FGR25]. Exit-Decisions [FGR25]. Expansion [BGB18, Bek13, BBSS03, Hei06, Zap22, Gau73, Pri78, Wei64a]. expect [Wil64, Wil86]. Expectations [Lem20, Rob21]. Expected [HSR19]. Expenditures [Bor67, Pou07, Fre67]. Experience [BL16, CA06, Koh76, Mis73, Mon88, Pap96, Sin62]. Experiences [KMBC24, AR20]. experiment [Wil74]. Experimental [Kim08, Hey87]. Experimentation [Kro23]. Experiments [GCS00, Sti12]. Expert [Cla74, Tob19, Mor73, Mor92, Rav78]. Expertise [CBKN17, Dem17, Gru17, Hop11, LS21, MMMS^{+}23, SL22, Tur10, Arn07, Pap96, Bro20]. Experts [BM18, LS21, vR10, Maz73, Rot90]. Exploiting [Low92]. Exploration [Kim08, Gas07]. Exploring [DS16, Edl24, Joh15]. Export [EV12]. Exporting [Kru05]. Expression [BWF24, Ano90a, Gol90]. Expressions [HJ25]. Expulsion [Ano67p, Sch06]. external [Ano94, Rob94, Sha92b]. Eyal [Bro20]. Eyewitness [Fis07].

F [God10, Gre05, Pra12]. Fabrikant [Mon95]. Face [AC23, Rüe92]. faced [Bra07a]. facilities [Ano63k]. factions [Ano68t]. Facto [MN13]. Factor [RdR15]. Factors [BC20, CCSM18]. Factory [Col07]. Facts [Rad19]. Faculty [Ano90b, BWF24, CBB20, Ano85d, Nye75, S.90]. Failed [Vav98]. Failures [Mey11]. Fair [EG06]. Faith [Lub04, Ano71e]. Fake [Akm24]. Fall [PK09]. Familiengeschichte [Bre04]. Family [Bre04]. Famine [PL19]. Far [WG97, Ros67b, And22, Spe85]. fares [Ano67f, Ano69t]. Fashioning [Fox06]. Fate [Bjö01, vdW01]. Fathoming [Gre07b]. Fear [Rad19]. Fecundity [Rei97]. Federal [Ano92, Nol82, SB92, Het90, Klo71, S.71a, Ano65i, BC85, Bri06, FG92, Fin77, Joh65b, MLS91, Mer79, Nol83]. Federation [Ano63l, Ano63m, Pet08b]. Feeling [BM24]. fees [Ano67r, Ano67s, Ano67t]. fellowships [Ass93, Cob76]. Female [BL16, TH24]. Feminism [Gro99]. Fermi [Hol74]. Fictions [Rad19]. Fiducial [HS99]. Field [ALH09, Bau20, BGB18, Bek13, BB24, HJ16, HBW^{+}18, Mad18, McG16, MB12, Pan11, Pfi15, SGSB23, WV17, Pét96a, RD93]. Fields [HvL11, KV25, Ort16, Tor18, Wil71]. fifteen [Gra66]. Fifty [Ban15, God02]. figure [Bro91]. financed [Kei81]. Financial [WR15, Ano64s, Ano70j, For74, Mon88]. Financialization [Hen21]. Finding [Gel00, HK88, IH25]. Findings [FFL^{+}15, SGSB23]. finds [Ano69-34]. Finland [Kaa11]. Finnish [MS23, Vuo15]. Fiore [Mel24]. Fire [LS21, Mos08]. First [Ano69-45, Bar98, KL03, LBP10, WG97, Gro92, Rüe92, S.92, Sha92a, Sha92b, Shi92b, Shi92c, Wan95, Pan93]. Fish [SV10]. FIT [Yan20]. Five [Ano65l]. fixer [Wei93a]. Flash [Rae08]. Fleck [For08, Pei11]. Flesh [Lub02]. flicker [Ano68j]. Flood [Ano69h]. Flow [MKK22, Mac88]. Flows [AG25]. Focused [KV25]. Foetal [Ted95]. Follow [Ada09]. Following [PJLP16]. Fondation [Gui06]. Food [Hut72, Kön07]. For-Profit [Met08]. Fordist [Tho10]. Foreign [Par02, Smi79, Ves94, Ano67p, Sin62]. forerunner [Het90]. Forests [KV25]. Foreword [Ano96j]. Forgotten [Thi04]. Form [Ack08, Est04, Rei97, Ano69-40, Mel72, Rit71, Tur87b]. Formalization [Woe15, NM16]. Formation [FBH23, GCS00, GCD23, GP77, Hol81, Mac73, SJ90]. Formed [Kai13]. Former [Häf97, Raž94]. Forms [Bon10, Pes09]. formula [Mon88]. formulation [WE72]. Forschungsrat [CG02]. fortunes [Car78]. Forty [BPM84, Pei84, Pei84]. Forum [AC23]. Forward [SR12]. Foundation [BPM84, Cob76, KK81, MS12, Sch03a, Lag87, Pfe70, AA88, Bux03, Cra86, Fis78, Gli05, Koh76, Koh78, KS01, Mac89, Pag02, Par02, Ric90b, Sch03a, Sch83, Sta95, Ves94, Wei93b, Wei97, Xie21]. Foundations [BPM84, Bul02, Fox97, Gei88, MTE15, BC85, Coi81, Gro82, SS79, Sei81]. Founding [BZ21, Kin97]. Four [Har02, HFSdR19, Nol99a, WE72]. Fourth [Ano63o, Ano63p, Ano63q, Jun22]. Fox [Gro99]. Fractured [GL22]. fragile [SB80]. Framework [GMD^{+}24, Hes13, LNST20, RR71]. frameworks [Tri71a]. Framing [PGO24]. France [Rob98, Stu98, Ano63o, Ano63p, Ano63q, Ano66h, Ano66u, Ano67-36, Ano70f, BD70, BBV11, CMR95, Cas05, Cla70b, Cro70, Cro75, Cro79, Cro86, DGR19, Fox73, Gir79, Kor79, Kra18, Lab08, Léc87, LP79, Nye75, Pap73, Pap88, Pav76, PD17, Sch03b, SJ90, Tri79, Von73, Wei87b]. Franco [RR05]. Frank [Son21]. Franklin [KHB^{+}72]. fraud [Ano85d, Ano86d, GDHH85]. Fraudulent [Akm24, MO17]. Frederic [Gem03]. Frederick [Gus12]. Free [Ano66i, Ano68k, BWF24, DF75, Ham05, Ano66r, Sch87b, SNSB^{+}80]. Freedom [Ano90a, Ano94, Blo97, EV12, Gol90, KP02, Pri98, Rus94, Smi07, AKW70, Ano66a, Ano67b, Ano69c, Ano68l, Ano78f, Ano80a, AA66, Bak78, Boz75, CSKF85, Cox94, Cra67b, Cra67a, ESC^{+}88, Eus95, Hun81, McG78, Mel66, Moo96, O’B66, Pra67, S.70c, S.73a, S.74b, S.84b, STK^{+}91, Shi94b, Shi95, Aro12]. French [AC23, Aro64, Bri77, Cro78, Cro81, Gau73, Pap96, Sco01, Shi80, Wei76, Wei79]. fresh [CH85]. fresh-water [CH85]. Friedrich [Bro91]. friend [Ano96l]. Friends [Mac99b]. Friendships [Gui06]. Frontier [FP18, HK00, Kro03]. Frontières. [Sco01]. Frontiers [Ano01b, DR12, Hab73]. fruitful [Hut72]. Frustrated [GR06]. Fukushima [SLL19]. full [Sch86]. Fuller [Ank07a]. fumbling [Ano69j]. Function [EBMP24, Rei97, Woo03]. Functional [BB24, Sei81]. functions [Sha92a, ZM71]. fund [DK74, Lan87]. Fundamental [Div91, Sch80a, Hod80, Mor64]. Funded [Rou19]. Funders [Kru04, WBW18]. Funding [Aag17, BM24, CCSM18, EL22, FSHdR18, GM03b, GV18, HBW^{+}18, HH20a, HK11, Kin97, Lai20, Rou24, Tor18, Tur98, VS20, WGL18, Xie21]. Funds [Sá05]. Further [Ano69i, Ano69-29, Ron79, Ano68m, Fra78]. Future [Bri05, Gel00, Kru08, Mos08, Pap04, Pre23, Rob03, Rou24, WW12, Ano70f, Ash67, Par68, Szc68, Gro99]. Futures [Dou03].

G [Ben98, BPM84, Orl99, Pei84, Pra12]. G7 [Joh19]. gains [Sto72]. Galison [Sha17, Zim98]. Game [Kam09, Rey21]. gap [Bro72, Nel71]. Gaps [Kön07, vR13]. Garland [Het97]. Gatekeeping [Can11]. Gathering [Mac99a]. Gaza [Sul91]. Gender [EK01, FvSA01, Kar24, KKK10, PS01, THM17, VL17]. Gene [Guo13, vdB19]. general [Plo75]. Generals [Ano68n]. generation [Hoc87]. Generations [MT19]. genesis [Gho88]. Genetic [Bal99, SSF15]. Genetics [Die05, Pan11]. genome [HK88, Lem20]. Genovese [Ano65m]. geographies [KL94]. Geohistorical [THE08]. Geology [Tur87b]. German [Ano92, Eck04, Hav66, Klo71, MLS91, Mer79, Pas78, SB92, Aic63, Ano74b, Ano78e, Bus63, De 97, DPB^{+}20, Fin07, Häf97, Hav66, Hen77, Kal18b, KBK13, Lew80, Lob84b, Mac89, MT19, MB12, Nol82, OS21, Rai64, Rei07, Rüe75, S.82, Sch72, Sch05, Sei81, SV14, Sta95, TSB24, Tsv14, Tur87a, VSR14, Vos23a, Vos23b, Waa15]. Germany [Ano62f, Ano63r, Ano66t, Ano67a, Ano67-32, Ano68y, Ano70l, Ano73c, AR20, Ash00, BD69, Bes68, BC85, Bra07a, Bri06, Bro91, CG02, Dei79, FG92, For74, FvSA01, Gil78, Hei06, Joh85, Klo68, Neh06, Nol83, Pfe70, Pos04, PD17, Pri98, Ric90a, Ric90b, Rie69, SG72, SL18, Von73]. Getting [Rou19, vR13, vR14, Wal92]. GeV [Gib70]. Ghana [Ano64a, Ano65a, Ano66b, Ano67-39, Ano69-43, Aus75, Wil74]. Giant [HH20c]. Gift [Kön04]. Gil [Bro20]. gilded [Koh93]. Gill [Fra22]. Gillispie [Bak87]. Gini [HL10]. given [Ano67-40]. Giving [KS01]. Gläser [End09, Rog13]. Glennan [Het97]. Global [Bra15, BZ21, CT13, D’S06, Gra12, Mah00, Mar06, Mul20, Nol99a, Nou11, Rab25, SLL22, Son21, Zap22, Sch94, Cha24]. globalisation [Sch91]. Globalism [Par00, Par02]. Globality [Sch91]. Go [FHS09, Gre05, Ros67b]. goals [Ano65b, Kay66a]. Godin [Wil20]. Goes [Guo13]. gold [Aus80, Hol74]. Good [BPM84, LNST20, Yan20]. Google [DGR19]. got [SM25]. Governance [BEK12, End09, Fel09, GM09, Glä24, HI06, Jac09, Kön07, Kri22, PW09, Rab25, Rim98, SDvL19, SSF15, TP14, Whi11, WW19, Woe13, Woe15, Woo09, Bro07a]. Governing [Jas03, SDvL19, Tal09, Woe15, Lim95]. Government [Ano98b, Ash68, Bel68b, Bru09, Cai20, CW64, Cra67a, DV98, Doe70, Dup63, Hab73, Jam68, Kay12, Kro03, Lai20, Lew77, MT67, Mon98, Pes09, Sap71, Shi79, Thu10, Ano62h, Ano63n, Ano64o, Ano65i, Ano69t, Ano70i, Ash66, Aue65, BW70, BK70, Bow67a, Cro94, De 78, Ged95, Gil78, Hal63, Hun85, Hut70b, Hut72, Kid66, Mac70, Puc72, Rot81, Rue83, S.71a, S.83, SG72, Var70, Wal88, oC68]. Governmental [Pat85, Pav76, Het90, Het91, Mic78, PK80, Rit66, RS77]. Governmentalisation [Wei96]. Governmentally [Phi78, Kei81]. Governments [SS79, Vav98, Pay80]. Gown [Bro00]. gradSERU [BI22]. Graduate [WHF18, Joh73a]. graduates [Smi79]. graduation [Ano67-40]. Grand [Kal18a]. Grant [Rob21, Rou24, Vel18]. Grants [Ano64r, Ash68, BGP14, Edl24, Hut75, Sha87, Ano93a, Cro79]. Great [Joh02, Bro91, S.76, Sho72, Tur87a, Ano68x, Ano69s, Bra07a, Bra93, Dra84, Fis78, Hut70b, Kei81, Mac70, Mac71, Pal90, Ric88, ST93, Var70, Von73]. greatness [Gre87]. Greece [Ano68n, Ano84c, BH15, Han65, Han68, Hol68, Sio69]. Greek [Ano68c, Ano70k, Psa88]. Green [Gre05, Log11, Tau21, PL19]. Gregg [Sch03a, Sol03]. Gribbe [AS13]. Grip [LK22, vR14]. Groping [Gen87]. Grosart [Joh71a]. Gross [Sel08]. ground [Ano69h]. Group [Ano90b, Cao98, SGSB23, Hol74, Mul72, S.90]. Groups [HFSdR19, Etz92, Ted95]. Growing [LM04]. Growth [Bar15, FFM16, Ano64o, Bec73, BD64, BD69, CW64, Joh74a, Klo71, Kol85, Lan87, Mis74, Pol67, Rie69, Rot81, Rud68, Wei67, Wil64]. Grudging [Bor21]. guarantee [Ano90d, Ano90e]. Guenther [Bre04]. Guide [Lim95, SF25]. Guidebooks [Kea13]. Gulbenkian [Coh98]. Gunnar [Lag87].

H [BPM84, Gis04, KHB^{+}72, Kri97]. H-Bomb [Gis04]. Habitus [MT19]. Hacettepe [Ree75]. Halfway [Wil07]. Hall [Ano67-40]. Halting [Ano69-30]. Hamburg [Gil78]. handbook [Deb69]. Handelshochschule [Tri94]. Handful [Het97]. Handling [MO17]. Hands [Bel68a, Elv68]. Hans [Kro23]. Hans-Jörg [Kro23]. Hansen [Het97]. Happened [Gei97]. Hard [Tal09]. hardcover [Mar20, Til24]. Hardest [VP22]. Harding [Gro99]. Harness [GS05]. Harris [Ano67-40]. Harrison [BPM84]. Harvard [Ano70e, Ano83f, Sha84]. Havemann [Hav66]. Haven [Vol21a]. Hazard [Col07]. hazards [Maz90]. Heads [ZJE10]. Health [AC23, ALH09, Fox97, MKK22, PD17, Rab25, Wei97, Wei93b]. Hebrew [Bar98, Mac08]. heels [Ano68m]. Heidelberg [Car78]. Helen [Akm24, Gro99]. Helix [Ano98b, Cai20, CL22, HP19, MS11, Pra12, Thu10]. Heller [Stu98]. Help [Ros02]. Henry [Het97, Stu98, Cai20, Jon66, O’S01]. Heretical [Wei64a]. heritage [Lew80]. Hermann [Het97]. Hesitancy [Fre23b]. Heterodoxy [SREM^{+}23]. Heterogeneity [BGB18, Nov17]. Heterogeneous [Kri01]. Heyday [Rol08]. Heyworth [Kin97]. Hidden [LBA16]. HidroAysén [BK18]. Hierarchy [Lew84, Lom79]. High [Hei17, Kri97, Mar06, PP25, Ano69h, Mac88, MI81]. high-energy [MI81]. high-technology [Mac88]. Higher [Akm24, Ano63j, Ano66k, Ano70g, Ano70h, Ano73e, BBSS03, Bow02, BEK12, BZ21, Bul02, Cla95, Dav01, De 97, Far01, FH13, Gar01, GOR25, Gua88, Har10, KSC^{+}64, Kin63, KV19, Kog97, Kol83, KBK13, LBP10, LB13, Lev98, Met08, Mid01, MSD21, Moo07, Mus22, Nou11, Pou07, PS65b, PS65a, RYI13, Rat65, RBB13, Rya01, Sad91, SC12, SRTC10, SLM22, Sco01, Sha89a, Str64, Sut98, Til00, Til24, Tom98, Tro00, TIC02, Umu86, Wal96, Wil97, Ahm89, Ann82, Ano66u, Ano69-31, Ano71f, Ano72c, Ano84c, Att73, BCL76, Bow67b, Bro91, Cou79, DA93, Das64, Duf66, Gaj68, Hay94, Lay73, Lew80, Mat69, Orl92, PH80, Psa88, Rüe75, S.71c, S.74c, S74e, S.82, ST96, Sul95, Tom96, Tro76, Wil86, Wil71, Woo72, dGdMS^{+}69, Ano66j]. Hinder [Lau17]. Hindi [Ano65k, Ano67o, Ano68h]. Hindu [Ano66g]. Hinterland [SF25]. Hintz [Mer24]. Hippocrates [Rei80]. Hippocratic [Ash69]. Hiring [CBB20]. Historical [Dou03, Eck04, Kai13, Kim08, Kri80, Mac97, PP25, Rot99, Sch14].

Historically [And97]. Historiographic [Kay97]. historiography [Jud80]. History [Abb09, Arn00, Bel98b, BEK12, BPM84, Cap98, FGR25, For08, Fox06, Het97, Kar06, Kor06, Kuk11, Lea98, Mon98, Ort16, PFC18, Pet08a, Pre23, Som08a, VV12, Vol21a, WV17, vdB19, Coi81, HK88, Kim82, Kol83, Mom83, Mos78, Tha80, Var70]. HIV [Kah16]. Hlávka [Win94]. Hoc [Ano83e, Mor85]. Holders [TSL19]. hole [Ano70i]. holidays [Ano69-42]. home [Lak73]. Honest [Bro08, Her21]. honesty [Hoo89]. Hong [SJPA14]. Honorary [HJ08, HJ07]. hope [Ano68j, Ano88b]. Hopkins [FD04, Rim98]. Hopping [WHF18]. Horizon [Dou03]. horses [S.73c]. Horton [Elk77]. Hospitals [Gam97]. Hot [DR12]. House [Ano65b, Gre05, Ren08]. housing [Ano68l]. Hughes [Mol05]. Hull [Hay88]. human [Ano78f, Gau73, HK88, Joh67, Löw87, Pat85, Sco67, Tho67, Tho68, RvDG20, Mol05]. Human- [RvDG20]. Humanism [Kri78]. humanistic [FV79, S.84a]. Humanities [Bux03, CA06, Har04, HJ25, XOR21, Thi04]. humanity [Dub63]. Humboldtian [Kru05]. Humility [Jas03]. hundred [RLL84]. Hungarian [Pét93]. Hungary [Pét96a]. Hunley [Het97]. Hurricane [Mar11]. husbandry [LM94]. Hybrid [Kai13]. hybridisation [Cha94]. Hygiene [Sch03b]. Hyper [FFM16]. Hyper-Competition [FFM16]. Hypothesis [BH15]. Hypotheticality [Häf74].

Ian [Lo 25]. Ibadan [Ano67c, Ano67-40]. ICSPS [Elz12]. Idea [Cal04, Ash67, Col86, Cou79, Kol83, Par68, Row64, Shi92a, Szc68]. Ideal [AGF09]. ideals [Ano69q, Pan93]. Ideas [Har07, Hay94, Mul20, Cha94, Hoc87, Wal10]. Identities [Deu07]. Identity [FBH23, GCD23, Kal18a, KB23, KOT14, MSD21, PvR16, S.74c, dGBM^{+}21]. Idle [Ank07b]. if [CH24]. II [Ano66m, Ano67i, Ano67s, Ano67u, Ano68h, Ano68v, Ano69s, Ano69u, Ano70h, Ano82e, Ano85d, Ano90b, Ano90e, Ano93c, DDS^{+}87, Dub65, Eps86, GMY73, Jon72, MR74, MT67, Maz73, Mor73, PH80, Pet08b, Plo75, Rob94, S.82, Sch83, Sei88, Tou66b, Ula83, Wei64b, dGdMS^{+}69]. III [Bel98b, Ano67t, Ano67v]. I’ll [Win20]. Illiberal [SLM22]. Illicit [EB22]. Image [Zim98]. Imaginaries [Hes23, JK09, SL18]. Imagined [KL94, Ray09]. Imagining [Jun22]. Immanuel [Coh98]. Immersion [HH20a]. Immiseration [VW95]. Immobile [KHS24]. Impact [BC20, BGP14, CBB20, DS16, EKB22, Fra22, Gen23, Kal18b, MTE15, Par02, RdR15, SBvE25, WGL18, Law81, Löw87, Sin62]. Impact-Driven [SBvE25]. Impacts [KW11, XOR21]. Impediment [Mar11]. Imperative [SH02]. imperial [Ano73c, Joh85, Pfe70, Mac94, Pfe70]. implantation [Eis79]. Implanting [Joh09]. Implementing [HP19, VS20]. Implications [DR12, TIC02, CMR95, Mom83]. Impressions [Hal63, KP02]. Imprinting [PGO24]. improve [Dot72]. Improved [Col98, ED91, Mon95]. improvement [Kar65, Mor66a, Umu86]. Improving [Bal99]. in-and-between [dGBM^{+}21]. In-between [HLE24]. In-house [Ren08]. Inauguration [Mac08]. Inbred [KHS24]. Inbreeding [Hor22, KHS24, TSL19]. Incentive [QZL25]. Incentives [XOR21]. Incentivized [SMaR21]. Incidence [Joh19]. Inclusivity [WW19]. Incommensurability [Pei11]. Increased [Ano67r, Ano67s, Ano67t]. Increasing [BC17, HL10]. Indefinite [Bru09, Mor92]. Independence [BA20, GAL22]. Independent [Mer24]. Index [Ano97, Ano00b, Ano01c, Ano02b, Ano03, Ano04, Ano05b, DGR19, Til24, Bre04]. India [Mac99a, Ano66n, Ano67-31, AM74, Bha79, Das64, Gil69, Kar63, KSC^{+}64, Rai96, SSK21, Shi69, SB80]. Indian [AR92, Ano64j, Ano64k, Ano64l, Ano64m, Ano67u, Ano67v, Ano67w, Ano73b, Bét81, Des65, Dha84, Edu66, Gaj68, Gil72, Kar65, Mah64, Muk64, Par65, Sha73, Sin62, Sin92, Woo72]. Indicator [ZW25]. indicators [FG07]. Indies [Pay80]. Indirect [Ack08, Wil86]. indiscipline [Ano67z]. Individual [Etz92, Jac08, WO16, MLS91]. individuals [Wil71]. Indonesia [Ano66d, Ano66o]. Indonesian [Ano66o, Ano66x]. Inducing [EL22]. Industrial [Ano63l, Ano65q, Bon08, Jun22, Kra18, Mat08, Ren08, SVMN15, DK74, Hod80, Koi95, Kor62, Nay67, Pap73, Pav76, Teu83, VW95, Wil86, Woo84]. industrialisation [Ama70]. Industries [Ano63l, Ano63m]. Industry [Ano63m, Ano98b, Cai20, CA06, DV98, Far05, FAGH01, Gib00, Léc05, MS23, Mar20, MB10, Ren08, Sal05, Thu10, WBW18, Ano63l, LP79, Lew84, Shi70, Tou66b, Var81]. Inequalities [NS25]. Inequality [BC17, EK01, HL10, Kar24, WW19]. inferior [Hut70a, Jon71, Var71]. Influence [BBSS03, GCD23, HJ08, RG21, Bha79, HJ07, ML83]. Influences [BL16, Ste74]. Influential [Gui06]. Informal [Sto11, Woo75]. Informatics [PKvdW23]. Information [Ano08, HD25, HN77, Mac99a, Mac99b, Shi18, TD10, Ano63n, Dan67, Mac88, Mel79, Ric88, Vic63]. Information-Transmission [TD10]. Infrastructure [CR25]. Infrastructures [Pap04, PKvdW23, Pfi15]. ING [Doe70]. Ingram [Goo99]. Initiative [AS13, BBV11, MG24, Wei87a]. initiatives [Ano92, SB92]. Injustice [HL22]. Innovation [Cai20, CL22, Cob76, De 97, DT12, FBGGH24, FSHdR18, God08, God10, Har02, Hen21, Kah16, Kay12, Kis24, LM04, LAN14, Mar20, MKK22, MPW16, Pei11, Pra12, QZL25, SS22, Tyf12, Val11, VS20, WW12, WW19, Wil20, vR14, Hun80, KBD75, Lew84, Maz75, RD93, S.82, Shi80, Wil86, Cai20, God12, God14, Rog13]. Innovations [WGL18, EP91, Rav78]. Innovativeness [Fal21]. Innovaton [POSC07]. Innovator [HH20c]. Input [Sto66]. Input-output [Sto66]. Inquiry [RYI13]. Ins [SB13, TC05]. Inscribing [FF10]. Insights [Sch14]. Inspecting [Ano67x]. Instalment [Ans03]. Institute [And91, DPB^{+}20, Etz94, HD25, Log09, Lom79, Pfe70, Sap17, Sch03b, Ano66h, GR06, Smi99]. Institutes [Kra18, PD17]. Instituting [Aro12]. Institution [Häf97, Mac00b, MG24, RBB13, Dev80, MT67, RSS97, Wei72a, Rai01]. institution-building [RSS97]. Institutional [AA04, BGB18, BZ21, CHJZSG20, CT13, Fel19, FSHdR18, FH13, Har10, MT19, MTE15, MB12, PGO24, RBB13, SV14, SBvE25, SB13, WO16, Bul80, Léc87, Mel72, Rit71].

Institutionalisation [BFGS14, ZM71]. Institutionalising [MH16]. Institutionalization [Vel18]. Institutionalized [Bul02]. Institutions [Ano62f, Coc09, GP04, Hop11, Jos94, Kay97, LBP10, LB13, PJLP16, RYI13, Ray09, Rei97, Str64, SJ05, Wei74, Ano67-32, Ano84c, DM73]. Instructed [Mac94]. instruction [Ano67o, Ano68h, Gaj68]. Instructions [Ano00c, Ano01d, Ano02c]. Instrument [BGS16, Kay12, Shi70]. Instrumental [Log11]. Integrating [CDRH25, Häf97]. Integration [De 97, Pfi15, Pay80, Per71, Sch87a]. Integrative [HR09]. Integrity [PF21]. intellect [S.72c]. Intellectual [Ber05, Far01, Fox06, Hay98, Isr08, Lan05, Rog13, Mor73, Rüe75]. Intellectuals [Mor92, Lub07]. Intellectuelle [Gui06]. Intelligence [Ano67h, Ano67i, Mac99a, Mac99b, Jon76]. Inter [Kol83, Smi99]. Inter-War [Smi99]. Interaction [GL22]. Interactional [MTE15]. Interactional-Institutional [MTE15]. Interactions [BM18]. Interdisciplinarity [BH15, LH19, Mad18, OB17, SMaR21, SGSB23, Shi18]. Interdisciplinary [CS15, DKCL^{+}23, GAL22, HR09, KV25, LH19, MEM20, MK19, Pan11, Pfi15, UMC^{+}23, WV17, AA88]. Interest [Etz96, Roz04, Ted95, Ves88]. Interests [Bar03, BS07, DeJ93]. Interfaces [BG23a]. Intermediaries [KW11]. Intermediary [Hes13]. Internal [Con96, MI81, Ano94, Rob94, Sha92b]. Internally [SMaR21]. International [And91, Ano01e, Ban15, Bau20, Can11, Dav20, DCJ22, GM03b, Kal18b, Koh85, Nol99a, Ole13, Roz05, Sal64, Wei12b, Wei15, Xu08, XOR21, Zim77, CH85, Hus83, Joh67, Kid83, MR95, MRC95, Ros67a, Sch87a, Sco67, Tho67, Tho68, Von73, Elz12]. Internationalisation [Ack08, Geu98, Jos94]. Internationalised [Lev98]. Internationalism [Roz04, Wid95]. Internationality [Som08a, Ric90b]. Internationalization [TIC02]. Internationalizing [KBWA21]. Internet [MK17]. Interpersonal [DCJ22]. Interplay [Kah16, KBK13, Vol21b]. Interpretation [SGSB23]. Interpretive [Sto11]. Interprofessional [OB17]. Interrogating [UMC^{+}23]. intervention [Hab73]. Interview [FHS09, Ano82e]. Interwar [Eng99, Kra18]. Intra [CT13]. Intra-Institutional [CT13]. Introducing [Pet18]. Introduction [BDG10, BM18, Cas05, Dem17, GM03a, GV18, GP04, HV23, HNW04, NSG03, Pes09, Pét96b, Pet08a, Wid01, Wid07, Zan02, Het97, Mom83]. invent [Fal21]. Invented [Die05]. Invention [Ben98, God12, SL18, Kel85]. Inventions [Kei81, SH18]. Inventors [Mer24]. Investigaciones [SM97]. Investigating [AG25]. investigation [Pat85]. investigations [Dot72]. investigators [Etz92]. Investment [DR12, NS25]. Invisible [Ben98]. invitation [S.72b]. involvement [Ano70j]. Ir [Bru25]. Irish [Gis69, Gis70, Gru70, Joh70]. Irreconcilables [Ano69j]. irreconciliables [Ano69-28]. ISBN [Bre04, Bro20, Eck04, Fle20, Gre05, Hen21, Kis24, Mol05, Pre23, Son21, Til24]. ISIS [Sto11]. Islamic [Deu07]. Island [Maz84]. Isolation [KE01, Sal66, Ves84]. Isomorphic [WC18]. Israel [Ano67-38, KBD75, MSD21, Sch87a, Teu83]. Israeli [Fre23b, MS23, Shi65]. Issue [Dem17, GV18, HV23, Wei09a, Wei09b, Chu85, Orl92]. Issues [AD21, LB13, Maz97, Joh65b, ST96]. Italian [Ano68t, Cam85, Pan93]. Italy [Ano64p, Ano64q, BFGS14, Gua88, Hol74, Woo03]. Itches [Ano69-32]. itself [Ano90c]. IV [Cap98, Ano67w]. Ivan [RLH91]. Ivory [Aus81, CG22].

J [Akm24, Bal99, Cha24, Goo99, Het97, Pra12]. J. [Lan87]. James [Het97]. Jamie [Akm24]. Jan [Som08b]. January [KHB^{+}72]. Janze [Eck04]. Japan [Ano70m, Cum90, ETMT70, Hay98, KKK10, Koi95, Low01, Low05, Suz01]. Japanese [Ano69y, Ano73e, Cum73, Kak69, Kuw01, Lon69, Mat08]. Jason [Gre05]. Jennifer [Het97]. Jensen [Tur06]. Jerusalem [Mac08]. Jetliners [Dow17, Dow18]. Jochen [End09, Rog13]. Johan [AS13]. John [Mor85, Son21, Ano67-40, Ban15, Bux03, Dup90, Smi99]. Johns [Rim98, FD04]. Johnson [Het97]. Joint [Pet08b]. Jones [BH15]. Jörg [Kro23]. Josef [Win94]. Joseph [Fre87, S.87, Sch93, Shi87, Sta03]. Journal [HH20c, RdR15, TSB24, Ano66q]. Journalist [Lai19]. Journalists [Rot90, Maz84]. Journals [TT24]. Jr [Bro08]. judgement [Mar88, S.75a, Wil83b, Whi09]. Junior [Rou19, Ano69-45]. Just [Ham08]. justifying [Moo96]. Justin [Cha24]. Justyna [Fra22].

Kaiser [Bey06, Sch06]. Kaiserreich [Pos04]. Kantian [Sha17]. Karachi [Ano65u]. Karve [Des65, Par65]. Katherine [Fra22]. Kauri [MMMS^{+}23]. Kedourie [Ked93a, Ked93b]. Keith [Het97]. Keller [Gro99]. Kenneth [Goo99]. Kenya [Eis79]. Kerala [Ano66n]. Kerr [Rot95, SC04]. Key [FFL^{+}15, NM16]. Killing [Bal02]. Kinchy [Bra15]. Kinds [Rou07]. Kingdom [Ano63h, Ano67m, Ano67l, Hut72, Lee06, Mah00, McG79, Pri98, WG97]. Kitcher [Bro13]. Klaus [Ano82e]. Knopfelmacher [Ano65o, Hoo66, Mor66b, Spa66, Sto65b]. Know [Dow17, Dow18]. Knowing [Mat08, Ray09]. Knowledge [AG25, And22, BG23a, BDG10, BK18, Bru25, Bul02, CA06, CTLL22, Cha24, DCJ22, EKB22, Etz94, FP17, Fer23, Gen23, GM09, GV18, GP04, GCD23, Gru17, GL04, Hau17, HS99, Jon76, Kal18b, Kay97, Kri22, MEM20, Met10, MK17, NSG03, OB17, Par02, Pes03, POSC07, Pra12, RYI13, Ray09, Rog13, Rou07, Sca03, Sco05, Sco07, Ski06, Son21, SL24, Woe15, WM20, Yeo02, Zah06, Zap22, Zim05, vR13, vR14, DF96, Eam85, Eis87, EP91, Fre87, KL94, Kid83, Kim82, Lan87, Rav78, Ric90a, Sch93, SH07, dlM03, Gen23]. Knowledge-Based [GP04]. Knowledge-Nexus [Kri22]. known [Lub07]. Kong [SJPA14]. Korea [JK09, SJPA14, Shi18]. Korean [Ano67a, KHCJ19]. Kraemer [Mul20]. Kraemer-Mbula [Mul20]. Krige [Ban15]. Krippendorff [Ano68o]. Kurt [Bar98]. Kushner [Bal99]. Kwapong [Ano66b, Ano67-39]. Kyoto [Nol99b].

L [Bel97, Mul20, Sel08]. Labor [OCW20]. Laboratories [Hod80, KV19, vR13, Ash66, Dei79, Kid66, Rit66, VW95]. Laboratory [CS15, Hut69, KL03, Mos78, And70, Cla70a, PK80, Swa70, TL77]. Labour [Tho10, Mac88, Stu98]. Lack [NM16, Ano82e]. Laden [Som08b]. Lagemann [BPM84]. Lagos [Ano65v]. Lambright [Het97]. Lamont [Whi09]. Lancaster [Rey87]. Landscape [BDG10, BZ21, Fer23, Kle23, Rai01, TSB24]. Lane [Gem03]. Langley [Het97]. Language [Cal04]. Lanka [Aus81, De 78]. Large [Häf97, Rav78, SB80, TL77]. Lars [Rog13]. Lascoumes [Ber11]. Lassman [Mar20]. Last [Ans03, Sha87]. Lasting [Gen23]. Late [Ano68r, Gir79, Kor79, Mor82]. Later [BPM84, Pei84]. Latest [Bek13]. Latin [Ano69r, Cue97, Lev98, Sch86, Ves88]. Lattimore [Rim98]. Launius [Het97, Vol21a]. Laura [BB81]. Law [Bou08, BEK12, BBMR17, FP17, WC18, Ano67d, Ano84c, dGdMS^{+}69]. Lawrence [Jen01, Lan87, Abb09]. Laxen [BPM84]. Layered [Edq03]. Layering [Aag17]. Lead [BPM84]. Leaders [Lee07]. leadership [Bro72, Nel71, Gem03]. Leading [Wal98]. League [Wei97]. Learn [Pei11]. Learning [BL16, Bja01, Bor04, BA20, D’S06, Har02, Joh75, Lub02, Lub04, RYI13, Ric90b, RG21, Rot99, SL22, Zim06, Ano66m, Ano66s, Ano70j, Ano71d, Aus75, Coi81, Hay94, ML89, Shi78, Shi92c, Wil74]. Legacy [Coe06, Dup90, PGO24, Cro79, Pét95, Sch06]. Legal [Kra18, RR71, Tri71a, Sei81]. Legendary [Hoo03]. legislation [Muc85]. Legislative [Tom98]. legislator [Hen77]. Legitimacy [Ole13, SV14]. Leibniz [Pos04]. Lending [Wil74]. Lens [OB17]. Leo [Ano80a]. Leone [Ano70j]. Less [WW12]. Lessons [HK88, Inn92, Mar11, Ort16, POSC07, Rya01, TP14, Tru06, Mon95, Pap96]. Letter [Arn00]. leucocyte [Löw87]. Level [BA20, HP19, Näs25, Sku19, Yli14, Ano63k]. Lewin [Bar98]. Lewis [Rim98]. Liberal [OT20, Pes09, Shi89]. Liberalism [KV19]. Libraries [Dan67, Joh75]. Library [Mos08, Mos08]. lie [Ula83]. Life [FFM16, Mül14, Rou19, SD08, Sut97, Tru06, WV17, Kol78, MJ78, SJ77, Shi91, Web77, Wei93a]. life-time [Kol78, MJ78]. Lifelong [SL22]. Light [Wil20, Das64]. Like [Høy24]. Limitations [RS77, Nol82]. Limits [NSB25]. Line [Two99]. Linear [SL24]. Lines [Lau17]. Linguistic [Ano67y]. Links [Sku19, CH85]. Lionel [Rim98]. Lise [McL99]. Lists [Kea13]. Literature [Bar06, GM09, Rou07]. little [Ano71e, Lit07]. lived [BBV11]. Lives [FFM16, Ras06]. Livestock [Woo09]. Livin [Mir08]. Living [BFdR25, Dav20, Wil83a]. Lobbies [SVMN15]. Local [Bul80, Dav20, Rae06, Hun80]. Locally [MSD21]. location [Str64]. Lock [SB13]. Lock-Ins [SB13]. loco [Ano67q]. Logic [Fer23, Zim98]. Logical [Ric07]. Logics [BZ21, MBS14, WS25]. London [Vol21a, Ano67m, Ano67l, Ano67z, Ano69-29, Ano69-36]. Long [Ban15, Kaa11, Orl99]. Long-Term [Kaa11, Orl99]. Longino [Gro99]. Look [SR12, Win20, Att73]. Looks [Xu08]. Lorraine [Sha17]. losses [Sto72]. lottery} [Rou24]. Louvain [Ano66c, Ano68s]. Love [Maz84]. Loyalty [Lan05, Inn92]. Lubricants [Win20]. Luck [Näs25]. Ludwik [For08, Pei11].

Luigi [Weh15]. Lund [Blo97]. Lundin [AS13]. Lutter [Gre05]. Luxembourg [PD17].

M [Abb09, BPM84]. MA [Hen21, Til24]. Maasen [Smi08]. Machinations [Het97]. Machine [Gra12, RvDG20]. Machine-Centricity [RvDG20]. Machineries [FF10]. Machinery [Het97]. Macintyre [Kuk11]. Mackenzie [Bel97]. Maclaurin [God08]. Macro [Sku19]. Macro-Meso-Micro-Level [Sku19]. Made [RG21]. Madrid [GR06]. Magnets [Mah00]. Maharaj [Ban15]. Mainstreaming [WW19]. maintain [Pri78]. Maintaining [Bey06]. Maintenance [Kar97, Dev80]. major [Ano69-33]. Make [OPBCM15, RvDG20, ZJE10]. Makers [God02]. Making [Ben07b, Bro08, Chl23, FF10, Kle25, Kön07, Log09, Lub02, MS23, MN13, PC02, PK80, RdR15, SL24, Tsv14, UMC^{+}23, Van15, WG97, Gil67, Gus94, Hus83, Rey87, Ros67a]. Malaria [BH15]. Malawi [Mic78]. Malaysia [Lim74, SJPA14]. Male [TH24]. Malign [LJ84]. Malta [Ano81c]. Management [GL22, Sch05, SL18, Smi99, WS25, Ano70j, Daa74, Div94, Koh76, Lak68, Mon95, Nay73, SH07, SV10]. Managerial [KBK13]. Managerialism [SC12]. Managing [Ano96k, Bja01, Bro07a, BM24, Chr08, SC04, Suz01, Yeo02]. Manchester [Fle20, Wal96]. mandatory [Mor92]. manifestation [Hey87]. Manifesto [Ano66p, HR15, Ano69r]. Mannheim [Tri94]. manpower [Ano63i, Gis69, Gis70, Gru70, Joh70, Kin63]. manufacturing [Ano63l]. manuscript [Cro86]. many [Fin77, Pay80, S.70a]. Mao [Sut97]. Map [Rot99]. Mapping [BS11, Bru25, BBMR17, HLE24, Sla11, TSB24, THE08, HK88, Bru25]. Maps [Bru25]. March [Kor20]. Marcovich [Joh15]. Margins [Bro06]. Margit [Eck04]. Marine [Kor06, Roz04, Sch25]. Marja [Weh15]. MARK [Wal06, Eng24, Hen21, Hop11, Xie21]. Market [Ben13, Dou03, Hen21, Sjö07]. Markets [Bou08, Fie01, Kaa11]. Marshall [Bux03]. Marxism [Sta08]. Masked [AC23]. Masks [AC23]. Mass [Mus22, Tro00]. Massachusetts [Etz94, Gil77]. Masses [Var12]. Material [Zim98]. Materials [BBV11, CS15, Eis13]. Mathematics [Kel85, Kid07, PD17, Sch80a, Sch87a, Tau21]. Matter [Enc00, Tel22, Tur98, KHB^{+}72]. Matters [Kar06, KV25, MN13, PF21]. Matthew [Mul20]. Maurice [Rob98]. Mavhunga [Kis24]. Max [Bre04, Kem08]. Maxwell [Pre23]. May [S.90]. Mbula [Mul20]. McCarthy [Hoo87]. McCarthyism [Bad00]. McCurdy [Kri97]. McLean [Mul20]. McNamara [Two99]. Mean [Kis24]. Meaning [Kle25]. Meaning-Making [Kle25]. Meanings [DR12]. Means [ACGG14, HP19, Mil07, SH18]. Measurement [Fel09, ML89]. Measuring [VFMGEE13]. mechanical [Div94, Kel85]. Mechanics [For08, Kel85]. Media [SDvL19, Lay73]. Mediate [DCJ22]. Mediating [TD10]. Medical [Bro11, Cue97, Sch03a, Gil77, Gis69, Gis70, Gru70, Joh70, Löw87, Mil72, Pry71, Rei07, Smi79, Wei87b]. Medicine [Hen21, Mat08, Pic07, Rou07, Sel08, Fis78, LM94, Bjö01]. Medicines [Hau17]. Medieval [Geu98]. Mediocrity [Her13]. Medium [Vos23a, Vos23b, Ano67o, Ano68h, Gaj68]. Mediumistic [Vos23a, Vos23b]. Meer [Fra22]. Meeting [Ano63i, Röd17]. meetings [Ano90a, Gol90]. Mega [Cha24]. Megalomania [Wil07]. Meitner [McL99]. Mellanby [BPM84]. Member [Alb63]. Members [Ano90b, Ano69-45, Lim95, S.90]. Membership [Mac00b]. memoir [Shi87]. Memorandum [Ano63j]. memorial [BB81]. Memory [Woe13]. Men [Die05, HdRW17, Het97, Ras02, Ano71e, Mac94]. Merely [Het97]. Merit [HJ12, Sne83]. Merton [Lea98]. Meso [Sku19]. Mess [Vos23a, Vos23b]. messenger [Woo75]. Metaphor [Col07]. Metascience [PP23]. Methodological [HJ25, LB13]. Methodology [McG16]. Methods [AG25, Rot99, Tur98, Lak68]. Metrics [Ack08]. Mexican [Lom79, Sch79]. Mexico [Ano66-28, Ano67f, Ano69d, FFL^{+}15, LM94, PvR16]. Meyer [Son21]. Michael [Bal99, Div98, Bak78, Gus12, S.76, Sel08, Til24]. Michel [Ber11]. Michele [Whi09]. Micro [MB10, Sku19, Yli14]. Micro- [MB10]. Micro-Level [Yli14]. Micronesia [Kro03]. Microphysics [Zim98]. Microscope [Ben98, Ben98]. Middle [LBA16, Oky70, Ole13, Waa70, Zah70, Bor21, Ree75]. Migrate [Bra12]. Migration [Hoc87, Lau05, Wal10]. Mihajlov [Ano66q, Ano67-27]. Milan [Gua88]. Mile [Maz84]. Militants [Ano69k]. Military [Bek13, NB02, Ano70k, WBJ^{+}77]. millennium [Cas96]. millstones [Ano77e, Ped77]. Mind [Gem03, Hen21]. minds [Cla08]. Minerva [Ano62c, Ano96i, Aro12, Mac16, S.72a, Tau12, Wei00, Wei12a]. Mines [Mac94]. Mining [Lab08, Mac94]. Minister [Ano63j]. minor [Gla74]. Minority [Ano90b, S.90]. miscellany [Ano69l]. Misconduct [Ano85e, Ano91, Chu85]. Missing [Div98, Var12]. Mission [ACGG14, KKK10, Mac08, MSD21]. Missions [Het97, Wei65b]. Mississippi [And97]. Mitigating [Kah16]. Mix [LBP10]. Mobile [PC02, Shi70]. Mobilisation [Par00]. Mobilising [SM25]. Mobility [Ack08, Bau20, Can11, Dav20, DCJ22, HH20a, Bri77, CMR95]. Mobilization [BK18]. Mode [GL04, HJ16, NSG03, Pra12]. Model [Ano65p, Bra12, CL22, HP19, Sch03a, SL24, Sto65a]. Modeling [Tob19]. Modelling [KB99, MH16, Mel24]. Models [Fel19, Gra12, Mel24, MO17, Ged95]. moderates [Ano69k]. Moderating [MARMSG20]. Modern [Abb09, Ben98, Bos05, Mar20, Pol05, Rot95, THE08, Yeo02, Shi89]. Modernity [Bra07a, Bro06, Pui05]. Modernizer [Bor21]. Modes [Bru25]. modest [Mor71b]. Mohr [Bre04]. Molecular [KL03, Jud80, Koh76, Mul72]. Money [Ada09, FHS09]. Monica [Mel24]. Montreal [RLL84]. Montyon [Cro79]. Moon [Vol21a]. Moral [AGF09, Cla08, Sel08]. Morality [EB22, Ano87b, Ano87c, DDS^{+}87, Eps86, Gro96a, S.85]. morning [Ano69m]. morphology [AA88]. Most [Kor20]. Motifs [Lai19]. Motivation [OS21]. Motor [Het97]. Movement [PP23, Ano66r, Ano68l, Ano69-34, Kid83, Mac71, Ric90a, Sch93]. Movements [Pfi15, SVMN15, Ano67-36, Ano69-30]. Moving [WR15]. Mr. [Ano69-35]. MTA [Mir08]. Much [Lit07]. muckrakers [Wei69]. Multi [Kit10]. Multi-Scalar [Kit10]. Multidisciplinarity [MARMSG20]. Multidisciplinary [BA20]. Multiple [WC18]. Municipal [Wal96]. Muting [S.73b]. My [Ano96l]. Myrdal [Lag87]. Mysore [Woo72]. Mysticism [Sta08]. Myth [Chl23, PL19, Pla99, Gro80]. Myth-Making [Chl23].

N [Gra12]. Name [GOR25]. names [Fin77]. Nanobiotechnology [Kim08]. Nanolithography [MS11]. Nanomedicine [Sla11]. Nanoscale [Joh15]. Nanotechnology [Eis13, FFL^{+}15, Kea13, MB10]. Narratives [MS23]. Narrowing [FFM16]. NASA [Het97, Kri97, Ban15, Het97]. Nasar [Fra22]. Nascent [HBW^{+}18]. Nation [Kri07, MH16]. National [Aag17, Ano69n, Bek13, BBV11, BPM84, CBB20, DT12, EV12, FM10, Gra66, Gre87, Hut69, Kei81, Lau17, MBS14, MS12, Mos78, Nol99a, Nou11, Ric90a, Roz04, Sch03b, Whi12, Woo72, Xie21, Ano65b, Ano69-39, Ano69-42, Bro72, Cla70a, CH85, DA93, DK74, Dei79, Han65, Kay66a, Mac89, Nel71, Por70, Ric90b, Sha84, TL77, Ves88, Ano65b, Ano65i, Bey06, Car78, Het90, Lan87, Log09]. Nationalism [Ano73a, SKS13, Glo76]. nationality [Pan93]. Nations [CP99, GFM21, Wei97]. NATO [Kri00]. Natural [Cam81, Kim82, Pri88]. Nature [Bal99, Bja01, Bru25, Gla08, GL04, Eam85]. Naukovedenie [Rab76]. Navigating [Näs21, Lo 25]. Nazi [Gil78, Hei06]. Necessity [Roz04]. need [Por70, Shi94b]. needs [Ano63i, Hun80]. neglect [LJ84]. Neglected [Fer23, S.70b]. Negotiated [LT16]. Negro [Ano67-34, Ano69g, Ano68v]. Nejedlý [Win94]. Neo [Sha17]. Neo-Kantian [Sha17]. Neoliberal [Bru09, Smi08]. Neoliberalism [Weh15]. Netherlands [BD67, Jul67, RdR15, Som08b]. Network [Gui06, Hei11, Kit10, MS12, SDvL19, Sto11]. Networks [CBKN17, DCJ22, FAGH01, GCD23, KBWA21, RG21, Rül23, THE08, VV12, VFMGEE13]. Neuberg [Bjö01]. Neuromorphic [NSB25]. Neuroscience [Hen21]. Neuroscientific [Smi08]. Neutrality [Teu83, Wid95]. New-Public-Management [GL22]. Newton [Far05]. next [S.72a]. Nexus [Kri22]. Nico [Gen23]. Nigeria [Ahm89, Ano65i, Ano65t, Ano65v, Aus80, Eis79]. Nigerian [Kol85]. Niklas [AS13]. Nine [Mos08]. Nineteenth [Bel98b, Ash67, Cro78, Cro79, Cro86, LP79, Mor82, Par68, Szc68, Tur87b, Wei87b]. Nineteenth-Century [Bel98b, Cro86, LP79, Tur87b]. Nixon [Orl86]. No [S.71c, Ano67d]. Nobel [Jen01, Kri01, Bjö01, Cao04, Cao14, Lin01, Low01, McL99, O’S01, Wid95]. Nomos [Pet18]. Non [Met08, Var12, WHF18]. Non-Academic [WHF18]. Non-Publics [Var12]. Non-Traditional [Met08]. Nonprofit [WBW18]. Noordung [Het97]. Normal [KB23, Tur10]. Normalizing [SLL19]. Normative [SH02]. Norms [Mül14, EP91]. North [Aue80, Sku19]. Northern [Ano65r]. Northwestern [ESC^{+}88]. Norway [Ben07b, GL04, SSK21]. Norwegian [Ano69-41, Enc70, Sko69]. Note [Sti12].

Notes [Sal91]. Notices [XH22]. Notions [LNST20, Näs25]. November [dGdMS^{+}69]. Nuclear [Arn00, BPM84, Bul01, FG92, JK09, Joh09, Maz90, Pei84, PFC18, Pui05, Sal05, SVMN15, Cam85, Häf74, HN77, WBJ^{+}77]. nullity [Raž94]. Numbers [God02, Mel24, PW09]. Nursing [Vuo15, VL17]. Nutrition [FHS09].

O [Ano67c]. oath [Ash69, Inn92]. Obama [WM20]. object [SH07]. Objective [Yan20]. Objectivity [Sha17]. Objects [Röd17]. obligation [Hoo89]. Obligations [Seg85, Ano82c, Ano82d, SM72, Wei78a]. O’Brien [Ano64a, Ano65a]. obscure [Ano69d, SH07]. observation [Kem06]. Observations [Bie67, Joh74b, PKvdW23, Sei88]. Observatory [Dev80]. Obstacles [Shi92a]. occupation [Ano68p, Ano69-36, Hah75]. Occupational [KBK13, Nov17]. Occupied [HR15]. Ocean [Gre07b]. Oceanography [Ben07a]. Oceans [Roz05]. O’clock [Mos08]. Odds [Ano69o]. OECD [Ben07b, DT12, FFL^{+}15, Hal63]. Off [CHJZSG20, Enc06, Bel68a, Elv68, NB02, WHF18]. Office [Gre05, McG72]. Officer [Sch03a]. officers [Ano67-37]. Offices [GOR25]. Official [Des09, God02, Sal05]. officials [Cob76]. Offs [TMP10, ZJE10]. Ogburn [God10]. Ohnesorg [Ano68g]. Old [Ano92, SB92, Vav98, Bak87]. Oligarchy [Hor22]. Olkin [Goo99]. One [Pay80, RLL84, Vav98, Ano66w]. one-party [Ano66w]. Ontario [MR74, Mon88, Mon98]. Ontological [OB17]. Open [Tau21, TSB24, Coh98]. Opening [CBKN17, Ano68r]. Opennes [OPBCM15]. openness [Eam85, S.74c]. operation [Hun80]. Operational [Lak68]. operations [Mis73]. operative [Var81]. Opinion [Par00, Es77, Pie64, Rot90]. Opinions [Wer07]. Opportunities [Kay12, Shi92a]. opportunity [Low92]. Opposition [Maz75, Ano68y, HN77, Lad70, McC76]. optical [Dev80]. Optimal [Yan20]. optimum [Sut73]. options [Zim81]. Oracle [Col00]. ordeal [Ano67-29]. order [Shi78]. organic [Hab73]. Organisation [Dup63, Lom79, McG79, Wei97, And70, Ano62a, Ano62b, Div94, Hun80, Muk64, Pfe70, Swa70, Wil72]. Organisational [BFGS14, EP91, TP16, Whi12]. Organisations [CCSM18, Kit10, Kor62]. organised [Hut70b, Mac70, Var70]. Organising [Gun67, Ros67b, LH19]. Organization [Elz12, Hes13]. Organizational [BZ21, CT13, CTLL22, FBH23, Kra18, Kri97, KBK13, Ole13, SMaR21, Sap17, TP14, WS25, MSD21]. Organizations [CHJZSG20, FvSA01, Kai13, MSD21, SL24]. Organizing [NM16]. Orientation [CTLL22]. Oriented [ACGG14]. origin [Ano69d]. Origins [Mac16, And91, DeJ93, Eam85, Etz93, Gil72, Hut75, Kei81, Kel85, Mos78, Mul72, Ves94]. Orr [Smi99]. Orthodoxy [Shi80]. Osmania [Ano67-42]. Other [Bar23, Roz04, Xie21, Fin77]. Otto [Dev80]. Outcome [Nel21]. Outline [GAL22]. outlook [Hun80]. Outpost [HK00]. Output [BC17, CBB20, Sto66]. Overcoming [KE01]. overcrowding [Ano68r]. Overeducation [Joh19]. overheads [Tou66b]. Overinterpreted [Mel24]. Overlapping [Lev98]. Overseas [Kol83, Lee98, Ano67r, Ano67s, Ano67t]. owe [Léc05]. Own [FD04]. Oxford [Bel98b, Cap98, Eck04, And07, Ano72c, Arn00, HJ07, HJ08, Lea98, Zim63, Cap98]. Oxford/New [Eck04].

P [BPM84, Cha24, Mol05, Lub07]. p-p-p-prigs [Lub07]. Pace [Val11]. Pacific [Kro03]. Painting [Gre05]. Pakistan [Ano62g, Ano69x, Ano69-38]. Palestine [Mac08]. Palestinian [Sul91, Sul95]. Pandemic [HV23, Lai19]. Panels [DS16]. paper [Gre05, Joh71a]. paperback [Pre23, Son21]. Paradigms [Lyn03]. Paradox [Dow17, Dow18, Woe13]. Paragraphs [Rot07]. parentis [Ano67q]. Paris [Hah75]. Park [MMMS^{+}23]. Parliament [Alb63, Boy66]. Parliamentary [Bha79, Ano68y]. Parliaments [Cop02]. Parsons [Van15]. Part [Bel98b, Ano70g, Ano70h, Ano93c]. Participation [Hag04, HL22, Jas03, MTS20, Nov17, Tho10, Ano69-45]. Participative [Est04]. Participatory [CDRH25, Kri22, Hal75]. Particle [NBGT05, HK88]. particularism [Ano67y]. Partisanship [Wil83b]. Partnership [Jen01, RvDG20, Woo09]. Partnerships [CA06, Fer23, LAN14]. party [Ano66w]. Passionate [Nye07]. Past [Kru08, Pap04, Rob03, S.72a]. Pasteur [Ano66h]. patent [Etz94]. Patenting [BFGS14, CG22, Kaa11, SH18]. patents [Kei81]. Path [Col00, Har02, Lai20, Nef14, SL18, TMP10]. Path-Dependent [TMP10]. pathfinder [Häf74]. Paths [SF25]. Pathways [Isr08]. Patronage [Cra86, MS12, Cob76, Fin77, Fox73, Lan87, Low92, Sta95, Tur87b]. Pattern [Ano63m, Ano63r, Lin81]. Patterning [MEM20]. Patterns [Bux03, Gib00, KMB24, LBP10, Mor71a, ZM71, Tur87a, Zim71]. Paul [Gra12, Thi04]. Pavlov [RLH91]. pbk [Bro20]. Peace [BWF24]. Peer [Bar23, BM24, Col98, DS16, FB12, HH20c, OS21, VP22]. Peer-to-Peer [Bar23]. Peers [Bel97, Ano90d, Ano90e, Cro86, Roy84, S.75a]. Pellizzoni [Weh15]. Pen [MS11]. Peng [Bra07b]. Peninsula [KV19]. Pennsylvania [CS15]. People [Kor06, Bro07b, Rue83]. Perceived [BC20]. Perception [JM19]. Perceptions [ALH09]. Perfect [Rot07]. Performance [BGP14, Fel09, SRTC10]. Performing [LS21, Yan20]. Period [Smi99, Bra93, DF96]. Periodical [Bro05]. Peripherality [Lom79]. peripheries [Cha94]. Periphery [BGB18, DV98, OT20, Pui05, SM97, Ash96, Nye75, Shi91, Von73]. permanent [Ano66z, CSKF85, Shi95]. Perpetual [Ano69-37]. Persistence [Kar24, Kak69]. persisting [Lew80]. Personal [Ray09, SDvL19, VFMGEE13, Ann82, Ash96, Rie83, Sei88, Zim81]. Personnel [Zap22]. persons [ML81]. Perspective [Dou03, For08, GP04, HD22, MPW16, PP25, RBB13, SSK21, SS22, VL17, Yli14, Hus83]. Perspectives [AA04, Ano01e, BL16, Eck04, GAL22, Ibr23, Kai13, Nou11, TH24]. Peter [Zim98, Mus22, Sha17]. Pethrick [Akm24]. Ph.D. [Rud68]. phage [Mul72]. phantasm [Ano68q]. Pharmaceuticals [Sjö07]. phase [Ano67-28]. PhD [FBF^{+}21, GCD23, TSL19]. Phenomenology [Kro23]. Philanthropic [Bul02, KK81, Lag87, BPM84]. Philanthropy [And97, BB81, Cue97, Gam97, Kar97, Kay97, Ric90b, RSS97, Sch03b, Wei97, Koh85]. Philip [Bro13]. Philosophers [Ros02]. Philosophic [Hro13]. Philosophical [Aga92, PP25, Kri80]. Philosophy [Ben98, Bro04, Har08, HH20b, Pet08a, Ros02, Hey87, Ric07]. Physical [Hut69, Mos78, Cla70a]. physicist [ML83]. Physicists [McL99, For74, Gro96b]. physico [AA88]. physico-chemical [AA88]. Physics [And07, Cah05, FGR25, Gas71, Gli05, Hei17, Kri01, NBGT05, PP25, Pfe70, Cam85, HK88, Hol74, Hol81, Mac89, MI81, Pol72, Jen01]. PIAAC [Joh19]. Pielke [Bro08]. Pierre [Ber11, AK11]. Pillars [DPB^{+}20]. pioneering [Ano82e]. Pioneers [KL03]. Pipeline [SL22, vdW01]. Pitfall [Nef14]. pitfalls [BCL76]. Pittsburgh [Mar20]. Place [CS15, EBMP24, IH25, VGM^{+}24, BC85, Hol74]. plan [Ano63o, Ano63p, Ano63q]. Planning [Mat69, And69, Bla67, GP77, Joh74b, Küh84, Lim74, McG78, Sto66, Wei74, Woo72, BPM84]. Plans [CÁCC20]. Plant [Kam09, Pal90]. plants [Zim81]. Platform [dSNC23]. Platformization [dSNC23]. Platforms [MB10, Pol05]. Play [Kam09]. plea [Ded66, Sch83]. Pliable [Tsv14]. Plugs [vR13]. Point [Two99, Tom96]. Points [Röd17, Shi92a, Shi92b]. Poland [BT19, Bie67, Bor67, Con96, Fre67, Muc85]. Polanyi [Bak78, Gus12, Ray09, S.76]. poles [Spe85]. polices [Fri69]. Policies [Edq03, GS05, GDHH85, Lai20, MTS20, Nou11, Pla99, Ano69t, Che71, Gru68, Mui69, Orl71a, Rat65, Tho69]. Policy [ACGG14, Ano64s, Ano85d, BG23a, Bod64, BS11, Bri06, BM18, Bro04, Bro08, CG02, CR25, DR12, EKB22, Elz12, EB22, Fel19, Fox97, Gre05, GP77, HJ16, Ham05, HM21, HS99, Kak69, Kal18a, Kay12, Kra06, Kru04, Lem20, Lon69, McG16, MN13, MKK22, Muk64, MPW16, Nol99a, Nol99b, Par02, RR05, Rou24, Sá05, SSK21, Shi18, Sjö07, SD08, Sti12, Tau12, Tho10, Tom98, TC05, Two99, Vos23a, Vos23b, WR15, WG97, WM20, dlM03, vR10, Alb63, Ano64j, Ano64k, Ano64l, Ano64m, Ano64o, Ano64p, Ano64q, Ano66v, Ano69-38, Ano69-41, Ano71g, Ano86c, Ano91, BD67, BS07, Bha79, Boy66, Bra92, Bri77, Bro91, Bro72, Cam85, CW64, Chu85, DH64, Don72, Dup90, Enc71, Etz94, FG92, Gee74, Gen87, Gil67, Gil77, Glo76, Gro71]. policy [Gus94, Han65, HK88, Hei06, HN77, Hus83, Jos88, Jul67, Kap66a, Kap66b, Koh78, Lak73, Lis67, Mac73, Mah64, Mos78, Nel71, Nic65, Pap88, Pfe70, Pig68, PK80, Rah65, Rak65, Ron79, S.70b, Sal64, Sap71, Sch94, Shi65, Sho72, Smi73, Sne83, Ste74, Teu83, Tra67, Wil71, Mel24]. Policy-Relevant [HJ16]. Policymaking [EKB22, FP18]. Polish [Mat69]. Political [Ano70j, Bro04, Chl23, CH24, Ful00, Het97, LB13, Mad18, Pes03, Pie12, PL19, Pol00, áRH23, Sta08, Sut97, Tyf12, Zim00b, Ano66x, DF96, Ezr80, For74, Mac76, MT67, MJ78, Pal90, Rad83, Rei07, S.70c, Wei93b, Cox94, Xie21]. politicians [Ano67-37, Ste74]. Politicised [Lee98]. Politicization [Zim06]. politicl [Tob88]. Politics [Ank07b, Ano66m, Ano66s, Aus75, Bar03, Bra15, Fel19, Gra12, HJ08, Ibr23, KW11, Kor06, KB99, Lob84a, Mac00a, Mel24, Moo99, Neh06, Pet08a, Rae06, Ras03, Smi08, Som08b, Tho10, Wei99, Hal75, HJ07, Kak69, Lip64, Lon69, Pay80, Pie64, Pri88, PS65b, PS65a, S.72c, S.88a, Sch80b, Sha89b, Was88, Bro08]. Politics/Policy [Fel19]. Polity [Bro20]. Pollution [BPM84, Sto72]. Polyphasia [FG20]. polytechnique [Shi84]. Polytheistic [KMB24]. Pondering [Ank07b]. Pooling [Kit10]. Poor [Kuk11]. Pop [Ano69q]. Popp [Ben13]. Popular [Bro05]. Popularization [Sta08]. Portfolio [WR15]. Portrait [Bre04]. Portugal [PFC18, TSL19]. Portuguese [SC12]. posed [ST96]. Position [CT13]. Positioning [FH13]. Positions [Waa15]. Positive [Ano88b, Psa71, Psa72, Sco71]. possibilities [Nol82]. possibility [Ano68q]. Post [CG02, Hoo03, KB23, Kri00, Kro03, SLL19, Sut97, Tho10]. Post-Academic [Hoo03]. Post-Fordist [Tho10]. Post-Fukushima [SLL19]. Post-Mao [Sut97]. Post-Normal [KB23]. Post-Sputnik [Kri00]. Post-War [CG02, Kro03]. Postcolonial [Lee07]. Postdoc [Näs25]. Postdoctoral [FFM16, Mül14, Ass93]. Postgraduate [Wal98]. Postwar [Rai01, KL94]. potent [Gro80]. Potential [AG25, FG20, Nef14, NB02, WBJ^{+}77]. Pour [Gui06]. Poverty [ZW25]. Powell [Cha24, Ano69-35]. Power [BB24, BBMR17, BPM84, Har07, JK09, Kay97, OB17, Sal05, Vuo15, Woo03, Ano67z, Ano69r, Ano69x, Ano73c, Ano73d, AA71, BS07, Bis73, Het91, Joh72a, Jon76, Klo71, Maz90, Ros72]. Powering [Het97]. powers [BD72, Zim73]. pp [Bre04, Bro20, BPM84, Eck04, Fle20, Gre05, Hen21, Mar20, Mol05, Pei84, Pre23, Son21, Til24]. pp. [Akm24, Vol21a]. Practical [HvL11, OT20, Smi99, DeJ93, Mor66a, Nic72, Row64]. Practice [Cal04, Dav20, EL22, Hes23, Kön07, MK19, PC02, SGSB23, SH18, SD08, THE08, dGBM^{+}21, Chu85, Löw87, Pav76]. Practice-Based [EL22]. Practices [BG23a, ELSS24, Fal21, Kal18b, KOT14, Kri22, Näs21, SF25, Vel18]. Practising [KV25]. Practitioner [KB23]. Pragmatic [AGF09, Ezr80]. Prague [Ano68l]. praxis [Ano67-29]. Pre [BEK12, Aue65, Mos78, Tha80]. Pre-History [BEK12, Mos78, Tha80]. pre-war [Aue65]. Predatory [TT24]. Predicting [Dob66]. Prediction [MH16, Mar11]. Preface [Mac01, Mac02, Mac03, Mac04a, Mac04b, Mac05, Mac06a, Mac06b, Mac07b, Mac07c, McL03, PW09]. Prehistory [Tob88]. Preliminaries [Bal02]. Prelude [Ano68r, Pét93]. Preparation [Mor82]. Preparedness [Mar11]. Presence [Lee98]. Present [Kru08, Mac97, PP25, Rob03, Waa15, Wil97, Gru76]. preserving [Ano96k]. presidency [Orl86]. president [Ano62h, Lob84b, Abe65, Ano90b, Bro65, Gre05, PK09]. Presidents [BW70]. Press [Bro20, Fle20, Gre05, Hen21, Kis24, Mar20, Mol05, Pre23, Son21, Til24, Vol21a]. Pressure [ELSS24, Fre23a]. pressures [Ama70]. Prestige [DGR19, MS12]. Price [Til24]. pride [Ano71e]. prigs [Lub07]. Primary [BRWB19]. Princeton [Son21, KL94]. Principals [Cla95]. Principe [Abb09]. Principles [Kön07, Ano69-31]. Prioritization [Nef14]. priority [Gas71, Pol72]. privatdozent [Bus63]. Private [AD21, Cue97, Cum73, Fer23, Fra91, Gar01, KK81, Kra18, Kru04, Lan87, POSC07, RvDG20, BC85, Nol82, S.82, Shi73, BPM84]. Privilege [Pet18]. Prize [Bjö01, Jen01, Cao04, Cao14, Kri01, Lin01, Low01, O’S01]. Prizes [Edl24, Kay12, Cro79, Wid95]. Pro [Wil20]. Problem [BBSS03, Gru17, Het97, Kay97, Kle25, Ano69-35, Bri77, S.70b, Zim87]. Problem-Solving [Kle25]. Problems [Fri69, Gil67, Gru68, Kim82, Lai20, Mui69, NM16, Tho69, Vic63, vR13, Enc70, Nic72, Rei80, Seg85, Sko69, WE72]. Procedural [FM10]. Procedures [Bjö01, OS21]. Process [Nol99b, Ano83f]. Processes [SSK21]. processing [Raž94]. Production [BK18, DPB^{+}20, Fer23, GV18, GFM21, GL04, Kal18b, KV19, MTS20, NSG03, Pes03, PD17, Pra12, Rog13, TP14, VFMGEE13, vR14, And22, Lew84, Cha24]. Productive [RG21]. Productivity [CP99, JS65, SJPA14, THM17, TSL19, WB65, Ano68-27, Bla68, Sch87a, Sch80b]. Prof [Hav66]. Profession [Maz05, Nov17, Sch05, Sco07, Ano70a, Ash69, BD72, Bét90, Gil69, Gil72, Gua88, Joh85, Ker89, S.88a, Sha73, Sha89b, Shi69, Was88, Zim73]. Professional [Div94, Har10, SV14, Cro75, Kin63]. Professionalism [BI22, RR05]. Professionalizable [HSR19]. Professionalization [Waa15]. Professionals [GF02]. Professions [Lub02, Gla74, Wei79]. Professor [OS21, Pet18, Ano82e, Ano94, Bak87, Elk77, Ked93a, Ked93b, KHB^{+}72, Sco07, Tsv14, Wei64a]. Professoriate [Maz05]. Professors [BBMR17, CG22, Rue83, Whi09, Ano65l, Ano67-37, Ano68n, BW70]. Profiles [Sco07]. Profiling [Mac00b]. Profit [Met08]. Profiting [EP91]. Program [Ban14, KPL^{+}89, Kri97]. Programme [GF02, LH19, Sta03, Ano73b, Ded66, Koh76, Sul95, Bal02]. Programmes [Eng99, Joh73a]. Programs [OT20]. Progress [Ano68s, Ano90c, MBG^{+}92, SAJ92, Ted95]. Progressive [BWF24]. Prohibitions [SH02]. Project [Edl24, FSHdR18, Vol21b]. Projectification [Näs25, Tor18]. Projects [Nel21, Rül23]. Proliferating [FBF^{+}21]. Promise [HD25, JE97, Sut98]. Promises [Nel21]. Promote [Lau17]. Promoting [KKK10, Nay67]. promotion [Bri77, Coi81, FF90, ML81, Osb71, Teu83]. Proof [Fos97]. proper [Bic74, Dub63, Rot81]. Property [Far01, Isr08, Etz94]. prophetic [Ano71g]. Proposal [Nol99b]. Proposals [Ano66h, Ano69-38, Rob21, KPL^{+}89, Mor71b]. prospect [Ano75g]. Prospects [GAL22, Pap04, Nol83, Vic63]. Protecting [Sap17]. Protection [BPM84, Zim06]. Prussian [Har04, NW04, Thi04]. pseudoscience [Gro85]. Psychiatric [SM25]. psychological [GMY73]. Psychology [Bar98]. Public [AD21, And97, Bek13, Bel68b, Bow67a, BS11, BPM84, Coc09, CCSM18, FF10, Fer23, Goñ25, GL22, Hag04, HI06, HR15, Ham05, Jam68, Kra18, Kru04, Lan05, LM04, MTS20, MMMS^{+}23, Mey11, Nel21, NS25, PvR16, POSC07, RvDG20, SVMN15, Sch05, Sjö07, Sla11, TC05, Val11, Wei93b, WBW18, Whi11, Woe15, Xie21, Yan20, vR10, Arn07, Bel67, Coo67b, Dub63, Eam85, Es77, KK81, Kar63, Koh93, Mel79, Pau67, Pie64, Rot90, S.74d, SG72, Woo72, oC68, AC23]. Public-Private [Fer23]. Publication [THM17]. Publications [Bel97, Woe13, XOR21, Cro81]. Publics [FF10, Var12]. published [ML86]. Publishing [Kal18b, VSR14]. Pumpkin [Gus12]. Pure [DeJ93, Sap17, Rou24]. Purity [Sap17]. purse [Bel67, Coo67b, Pau67]. Pursuit [Cao14, LT16, Bét81, Cha84, Rot95]. Pushing [MG24].

Quadruple [CL22, MS11, Pra12]. qualifications [Cro81]. Quality [Des09, LNST20, RBB13, SRTC10, TP14, Ano90d, Ano90e, ED91, ML89]. Quantification [DKCL^{+}23]. quantity [Maz90]. Quantum [For08, Rob21, Hod80]. Quasi [Var12]. Quasi-Publics [Var12]. Quest [Aro12, SV14, Gas07]. questions [Gun67]. Quincy [Dup90].

R [Bel97, BPM84, Het97, Mor85]. R&D [Nay73, Pou07, Mer24]. Race [Orl99, Vol21a]. Racial [Gro99]. Racism [BWF24]. Radiation [Raž94]. Radical [Lo 25]. Radicalism [Enc03, Ano68g]. Radiotherapy [Gil77]. Radium [Ren08]. Raffaella [Wal10]. Ramanujan [Shi91]. ramifications [Ano69a]. Rampant [Ano67-30, Ano67-31]. Rancorous [Ano69-39]. Randall [Gre05]. Rangers [MMMS^{+}23]. Rank [MSD21]. Rankings [HdRW17]. Rapid [BB24, Ama70]. rashes [Ano69-32]. rate [Rud68]. Rather [Rey21]. rationalism [Ezr80]. Rationality [Aga92, Est04, FM10, Gro85, FG92]. Rationalization [KHCJ19]. Rationalizing [Gre05, WBW18]. ray [Ano88b]. Re [Fal21, Kem08, MK19, Str03, dlM03, Smi79]. Re-Describing [Str03]. Re-disciplining [MK19]. re-emergent [Smi79]. Re-invent [Fal21]. Re-Reading [Kem08]. Re-Thinking [dlM03]. reach [Ano69h]. Reaching [Bar15, Vol21a]. Reaction [BPM84, Pei84]. Reactionary [Shi84, Shi16]. Reactions [SLM22]. reactor [Wei78b]. Reactors [Dow17, Dow18]. Readers [Tau21]. Reading [DKCL^{+}23, Kem08, Pes03]. Readjustment [KMBC24]. reaffirmed [Ano88a, S.88b]. Real [Des09]. Reality [Kra18, Rad19]. Really [Léc05]. reappraisal [Sch86]. rearguard [Ano69-46]. Rebellion [Ano69-40]. Rebels [Ano68t]. recapture [Hol74]. Recapturing [Sta02a]. Reception [Pui05]. recherche [SJ90]. Recognition [MB12, Arn07]. Recollections [Hoo84]. Recombinant [Col07]. Recommendations [Ano62f, Ano66t, Ano67-32, Ano69-41, QZL25, Ano67n, ML81]. Reconceiving [HS99]. Reconciliation [BB24]. Reconfiguration [SBvE25]. Reconfigure [MB10]. Reconfiguring [Rai96, Rog13]. Reconsideration [Pag02, McC76]. reconsidered [Kol78, MJ78]. Record [Cro81]. recovery [Dra84]. Recruitment [Ano90b, OCW20, Rey21, Bus63, S.90]. Red [Wer07]. Redbrick [Bel97]. redemptive [Ano73d]. redirection [TL77]. Rediscovering [Fle20]. redistribution [Dei79]. reduction [Fri69, Gru68, Mui69, Tho69]. Redux [End06]. Reef [Nef14]. referee [ZM71]. Refiguring [Kle23]. Reflecting [Wei12a]. Reflections [Bak87, Bau77, Das64, Dre71, Dub63, Elk77, FM10, Gru76, Hoo89, Jud80, Kin63, Lob84b, Mac97, Sch05, Shi91, Wil97, Ash96, S.88c]. Reflexive [Hes11]. Reflux [WG74]. Reform [Ano67w, GL22, JE97, Lee07, Lev98, Sca03, SLL22, Tom98, Woo03, Yli14, Ahm89, Ano66h, Ano66u, Ano67u, Ano67v, Ano69m, Ano68q, Ano69-30, Ano70l, Ano70m, Ano73e, Bes68, Edu66, Gho88, Klo68, RG73, Wei79]. reformers [Ano69j, Ano68u]. reforms [Ano69-28, PH80, Wan95]. Reframing [Kle25]. refusals [Ano68k]. Regal [Tuc00]. Regime [Bey06, Kin97, RR05, Tor18, Ano70k, Bak87, Ano65c]. Regimes [AD21, Bon08, Pes03]. Region [Nou11, Sto11]. Regional [Nou11, Pay80, Etz93, Hun80]. Régis [Ano67g]. Regulated [Hag04]. Regulating [SDvL19]. Regulation [Far01, FP17, LS21, WM20, Aue80, Joh85, Pat85]. Regulatory [CBKN17, Dem17, Hau17]. Reich [Eck04, NW04]. Reimagining [MKK22]. Reiss [Bal99]. rejection [RC67, Ros67a, Ros68]. Relate [Par02, Sch25]. Related [GOR25, Ano69v]. Relation [Kuk11]. Relational [McG16]. Relations [Ano98b, CÁCC20, DV98, FAGH01, Gib00, Mat08, MB10, Met10, Mon98, Roz05, Whi11, Woo03, Aue65, Jor70, LP79]. Relationship [DCJ22, HBW^{+}18, Mac99b, VFMGEE13]. Relationships [Bar23, Rog13, WGL18]. Relentless [Ano68u]. Relevance [BFdR25, HJ25, Kit10, Kle25, KV25, Mul20, SM25, SF25, SBvE25, Bru25, Mis73, SL89]. Relevant [HJ16, IH25, MSD21, SF25, SL24, Gun67]. reliable [FG07]. Religion [FG20, Lub04, Top02]. Reluctant [Ano68u]. Remaking [Til24]. Remarks [Coi81, ML81, Dub65, Wei64a]. remedy [Gru76]. Renaissance [Bar06]. Renewal [De 97, Ano83e]. Renewed [Ano67-33]. Rentier [Maz05]. reorganisation [Ano66t, S74e, Umu86]. Repertoires [BG23a, FFM16]. reply [Rus94]. Report [Ano62h, Ano62g, Ano63k, Ano64n, Ano64r, Ano65p, Ano65r, Ano65q, Ano78e, Ano83e, Ano83f, Ano86d, Coh98, Gro92, Mor85, S.90, Ano65b, Ano65i, Ano69-45, MLS91, Plo75, Rai64, Sch83, WE72, Das64]. Reporting [Maz84, Maz97, Nel21, Maz90]. Reports [Ano63l, Ano63m, Cro86, Lin81]. Representation [Hop11]. Representations [BBU19]. Representatives [Ano65b]. Republic [Ano65i, Ben98, BC85, Bri06, FG92, Nol83, dlM03, S.76, Wei72a, Wei78a, Zim78, Ano92, Bar03, Ful00, Häf97, Har04, Hav66, KHS24, MLS91, Mer79, Nol82, Pol62, Pol00, SB92, SJ05, Zim00b]. Republicanism [Col00]. reputation [ML89]. Research [Aag17, ALH09, ACGG14, And91, Ano63o, Ano63p, Ano63q, Ano65q, Bas67, Bek13, BBV11, BRWB19, BBU19, BGP14, Blu69, BG23b, Bri06, Bri05, BC17, BI22, BM24, Cal04, CBKN17, Cas05, CDRH25, Cha24, CS15, Coc09, Coo67a, CR25, CCSM18, Cue97, CBB20, DeJ93, DR12, Don72, DPB^{+}20, EL22, End09, EB22, EK98, Fer23, FM10, FP18, FFM16, FSHdR18, FBH23, Fre68, GMD^{+}24, Gei97, GS05, GM03b, GAL22, Goo99, HJ16, HR09, HBW^{+}18, HFSdR19, Het97, HK11, HD22, Jac09, KW11, Kin97, Kle23, Kle25, Kra18, Kra06, Kru04, LNST20, LM04, Lau17, Lem20, LH19, LAN14, Lom79, LT16, MA69, Mac99b, MEM20, MS13, Mar11, MARMSG20, MTE15, MKK22, MK19, Mul20, Nef14, NZ68, Nel21, NM16, NB02, Nol99a, Nol99b, Nov17, Ole13]. Research [OPBCM15, Opi04, OCW20, Pap04, Pie12, PD17, Ras02, Ren08, Rit66, Ros68, Rou24, RdR15, Sá05, SMaR21, ST93, SM25, SDvL19, Sap17, Sch14, SGSB23, Sch07, SH18, Sha89a, SJPA14, Shi18, SLL22, SF25, SH02, Sku19, Smi99, SK00, SC04, TH24, TSL19, TP14, Tor18, Tur98, Tyf12, UMC^{+}23, VS20, VL17, WR15, WBW18, Wil64, Woe13, WV17, XOR21, Zap22, dGBM^{+}21, AA88, Ama70, Ano62a, Ano62b, Ano62h, Ano63l, Ano63m, Ano65b, Ano65s, Ano66z, Ano69-41, Ano70f, Ano82e, Ano85d, Ano85e, Ano91, Ano93a, Ass93, Aue80, BS07, Ben07b, BK70, BC85, Bor67, Bra93, Bul80, Che71, Chu85, CH85, Cra86, Cra66, Cro79, DA93, Dan67, DK74, Deb69, ED91, Etz92, Fox73, Fre67, Gir79, Gra71, Gra66, GDHH85, Gro82, Hah75, Hei06, Het90]. research [Het91, Hun80, Hus77, Hus83, Hut70b, Hut72, Joh65b, Joh71b, KBD75, Kay66a, Kei81, Kor79, Kor62, Lak68, Low92, Löw87, Mac70, Mac71, MLS91, MR95, MRC95, MI81, Maz89, Mel74, Mil72, Mis73, Mor85, Mor64, Nay67, Nic77, Orl71b, Orl71a, Orl72, Pal90, Pat85, Pav76, Per71, Pry71, Rab76, RC67, Ros67a, Rot66, S.73a, SS79, Sch80a, ST96, Sch80b, Sch83, Sch86, Sei81, SJ90, SL89, Sub66, Ted95, Tou67, Var70, Var81, Ves84, Ves88, Wei71, Wei74, Wet85, Woo75, Woo84, WE72, Zim71, Zim81, BEK12]. Researcher [GCD23, LT16]. Researchers [BFdR25, Kaa11, KMBC24, OPBCM15]. Reshape [Fal21]. Reshaping [Häf97, Rot99]. Resilience [Fre23a]. Resistance [Ano66u, Ano67j, Ano67-41]. Resolution [Mus22]. Resolving [FG20]. Resources [Cam81, Kit10, BS07, Sha92a]. Respiration [MS13]. Respond [SD08, Tor18]. responding [GDHH85]. response [Cla95]. Responses [Ano69p, Hay88, ST96]. responsibilities [Ver81]. Responsibility [Bad04, Ano69c, Orl80, Rad83, NSB25]. Responsibilization [NSB25]. Responsible [MPW16]. restless [Per68]. restoration [Cro94]. Restructuring [Coh98, SMaR21, WS25]. Results [CBB20]. Retention [Ano90b, S.90]. Rethinking [Kay97, Lo 25, Ray09]. retirement [Mor92]. Retraction [XH22]. Retractions [MO17]. Retreat [Fre23a, Mus22]. retrospect [Ann82]. Retrospective [SS22]. Return [Geu98]. Returnees [SJPA14]. Returning [KMBC24]. Returns [Joh19, WR15]. Reunification [Häf97]. Revelle [Rai01]. Review [Abb09, Akm24, AS13, Bal99, Ban15, Bel97, Bel98a, Ben98, Ben13, Ber11, Bla67, Bro08, Bro20, BM24, BPM84, Cai20, Cap98, Cha24, Chr08, Coh98, Col98, DS16, Div98, End09, Eng24, Eus97, Fle20, Fra22, Fre23a, FB12, Gen23, GM09, GAL22, Goo99, Gra12, Gre05, Gro99, Har02, Hen21, Hop11, HH20c, Jen97, Joh15, Kem08, Kis24, Kri97, Kro23, Kru08, Lo 25, Mar20, Mel24, Mer24, Mul20, Orl99, Pra12, Pre23, QZL25, Ric65, Rim98, Rob98, Rog13, Smi08, Son21, Stu98, Til24, VP22, Vol21a, Wal10, Weh15, Wei69, Whi09, Wil20, Xie21, Zim98, Bun81, Ham93, LJ93, Orl91, Roy84, Wil85, Cla95]. Reviewers [OS21]. Reviewing [Ano65q]. Reviews [Ano96c, KZP^{+}74, TGGS83, AE77, ATIS70, Ano82a, Ano96a, Ano96b, AWZ79, AOT81, BBZA73, BM94, BDK77, BPM84, BMSC87, BWW90, CGTPM72, CSGZ^{+}74, CBJ95, DBL^{+}92, DBDM77, Dup84, FSH88, FHE94, FSMH75, FGT80, FS81, FFML95, FW94, GG73, GPM73, GNNBD84, GT89, GC91, HRKZ71, HCRG88, HD93, HR90, HC71, Jos90, KKF95, Kel80, KBD72, LWZB83, LZ87, LC72, MDG95, MP78, MR79, MPM^{+}76, MM88, MZP75, MA75, MDH94, MJSB85, MNS75, NGZG74, PCW79, PSNB80, Pei84, PPF89, PHA83, PK85, RBM73, RH89, RSR^{+}70, RC88, RAV76, RH90, SZL81, SCJP71, SGM92, SRM92, SS77, SH78, SSS^{+}71, TKMB72, TRGN74, TSH86, TK86, VMD^{+}91, VLJG93, WKC^{+}70, WGN80, Wei78c, WG85, WS91, ZWJK78, ZJ81, ZLD^{+}76, pHPP^{+}79]. Revised [HM21]. Revisionism [Bul01]. Revisited [Bel97, NSG03, Wer07, Gro96a]. Revisiting [BH15, HJ12, Mac00a, Zap22]. Revolution [Het97, Jun22, Ort16, Ano66f, Ano68q, PL19]. revolutionaries [Ano68u]. Reward [LBA16]. RFF [Gre05]. Rheinberger [Kro23]. Rhetorical [Lai19, Tob19]. Rhetorics [SC12]. Rhodesia [Ano65r, Ano66a]. Rice [And91]. Richard [Fra22, Goo99, End09, Rog13]. Ridden [EL22]. Riddle [Gei08]. right [Ano75f, Mer79, S.75b]. Rights [Bun81, Far01, POSC07, Ano68b, Ano70k, Ano78f, McG72]. Riots [Ano67-34, Ano68v]. Rise [Ada09, Elz12, HL10, HBW^{+}18, Par00, Par02, PK09, Rab25, THE08, BD70, Cla70b, Cro70, Lew77]. Rising [SV14]. Risk [Dem17, DR12, Fos97, Lai19, Mar88, ZW25, Gro85]. Risks [Chr08, Jon72]. Rituals [PvR16]. River [Orl99]. Robbins [Ano64r, Das64, Rai64]. Robert [BPM84, Mul20, Pei84]. Roberto [Wal10]. Roberts [BPM84]. Robin [Elk77]. Robinson [Hen21]. Robots [RvDG20]. Rockefeller [AA88, Ano93a, BB81, Bux03, Cra86, Fis78, Gli05, Koh76, Koh78, KS01, Mac89, Pag02, Par02, Ric90b, Sch03a, Sch83, Sta95, Sta03, Ves94, Wei93b, Wei97]. Rocket [Het97]. Roger [Bal99, Het97, Bro08, Rai01, Vol21a]. Role [BBMR17, EKB22, EK98, FHS09, GF02, GCD23, LM04, Lem20, MARMSG20, PGO24, Sal05, SH02, Ano66x, Ano67n, BD65, Cam81, Coi81, Häf74, Hei06, Joh74b, Léc87, Lim74, RD93, Rot81, Ted95, Wil86]. Roles [DRC^{+}14, KKK10, Lub02, Cha94]. Roman [Ano69-42]. Romania [Sad91]. Rome [Ano66-29]. Roots [Geu98]. rose [Fin77, Koh85]. Roth [Bre04]. Rothschild [WE72]. Routines [Tau21]. Rowett [Smi99]. Royal [Kön04, Mac94]. Ruestow [Ben98]. Rules [BM24]. Ruml [BB81]. Rupert [God08]. rural [Sol75]. Russell [Ano94, Eus95]. Russia [KS01, Sol03]. Russian [For92, Jos94, Tom98, TT24].

S [Bro08, Mer24, Rim98, Sta08]. S. [Shi91]. Sabine [Smi08]. saccharin [PK80]. Sachs [BPM84, Pei84]. safeguards [Rav74]. Safety [Kön07, LS21, Rav74]. Sake [Bru25, FD04, Ano69-40]. Salaam [Ano67n]. Saltelli [Mel24]. salvation [S.71c]. San [Rai01]. Sanderson [Div98]. Sandra [Gro99]. Sanitarians [Koh93]. sans [Sco01]. Sarah [Akm24]. Saudi [Maz05]. Saving [Gre07a]. Scalar [Kit10]. Scale [DS16, Ano69-39, Ano69-42]. Scanning [Dou03]. scarcity [Bra93]. Scazzieri [Wal10]. Scholar [Thi04, DGR19]. Scholarly [Ano78f, BA20, KBWA21]. Scholars [Bel98a, HD22, KMBC24, MT19, MN13, Rou19, MR95, MRC95, Pét93, Sch87b]. Scholarship [Eng99, Mat08, Kri80]. School [Ano67m, Ano67l, Ano69-29, Div98, IH25, Mac94, Ano67z, Ano69-36, SSG68]. Schools [Smi79, WC18, Gla74]. Schrödinger [McL99]. Schumpeter [God08]. Science [AR92, ALH09, AK11, Ano63n, Ano63r, Ano64o, Ano65b, Ano71h, Ano73b, Ano01e, Aro12, AGF09, Bad00, Bad04, Bal99, BG23a, Bar06, Bar03, Ben13, BT19, BBSS03, Bon08, Bon10, Bor21, Bos05, Bou08, BS11, Bra15, Bro05, BM18, Bro04, Bro08, Bro13, Cam11, Cao04, Cao14, CG02, Cha24, CS15, Col00, CÁCC20, Cue97, DF96, DRC^{+}14, Dem17, DR12, Dup90, DPB^{+}20, Eck04, Edq03, Eis13, Elz12, Eng24, EBMP24, EK98, EK01, EG06, FG20, FBGGH24, Fel19, FG07, Fel09, FGR25, FM10, FP18, FFM16, Fos97, Fox06, FvSA01, Ful00, Gei88, GR06, God02, Goñ25, Gre07a, Gre87, Gro80, Gro85, Gro99, HLE24, Hag04, HI06, Hal63, Ham05, Ham08, Har08, Har06, HJ25, HL22, Hes11, Hes13, HM21, Hoo03, HS99, Ibr23, Jas03]. Science [Jen97, Joh72b, Jos88, JE97, Kal18a, Kar24, KB23, KE01, Kin97, Kis24, KMB24, Kle23, KV25, KKK10, Koe07, Koh85, Kor06, Kra06, Kri00, Kro03, KB99, Kuw01, Lab08, Lai20, Lem20, Log09, Low05, Lyn03, Mac00a, Mac94, Mac97, Mac08, MTS20, Mah00, MS23, MBS14, Mar20, MS12, Mat08, Maz77, MMMS^{+}23, McG16, Mel24, MB10, Mey11, MN13, Mül14, Nef14, dSNC23, NS25, Nou11, Oky70, Ort16, Pan11, Pap88, PW09, Pet08a, PK09, PS01, Pol00, Pol05, PD17, Pui05, Qua71, Rad83, Rai01, Ras02, Ras03, Ren08, Rob98, RR05, Röd17, Rou24, Rou07, Roz04, Rül23, SSK21, Sal00, Sap71, Sch94, Shi16, SLL22, Sin92, Som08a, Som08b, Sta02a, Sta02b, Sta08, Sti12, Stu98, SL24, Sut97, Suz01, SJ05, Tau12, THE08, Tho10]. Science [TD10, Top02, UMC^{+}23, VV12, Vuo15, Waa70, WBL15, WR15, Wei72b, Wei99, Wei12a, WBW18, Wei12b, Wei15, WG97, Wid95, Wid07, Xie21, Xu08, Zah70, Zim00b, Zim06, vR10, vdW01, Abe65, Ano63i, Ano66m, Ano66s, Ano66t, Ano67-32, Ano71g, Ano73a, Ano73d, Arm07, Aue65, Bak78, Bar87, BD69, BD72, Bha79, Bie67, Bri77, Bro65, Bro72, Bul80, BB81, Cam85, Cha84, Che71, Chu85, Cob76, Con96, Cou79, Cro75, Cum90, DK74, DeJ93, Ded63, Ded66, Div91, Dob66, Doe70, Dub63, Eam85, Eis79, Elk77, Enc71, Es77, Etz83, Etz93, Etz96, Gas07, Gee74, Gen87, Gil67, Glo76, Gro71, GP77, Gun67, Gus94, Hab73, Hod80, Hol74, Hut72, Joh72c, Kak69, Kel84, Kim82, Koh76, Koh78, Koh93, Kol78, Lak73, Lew84, LM94, Lon69]. science [Mac71, Mac73, McG78, Mor66a, Mor71a, MJ78, Mos78, Mot64, Muk64, Nel71, Orl71a, Orl72, Orl76, Osb71, OO64, Pét93, Pfe70, Phi78, Pol62, Pol67, Raž94, Rei80, Ric07, Rie69, Ron79, Ros67b, RLH91, Rot81, RS77, S.70b, S.71b, S.74d, S.76, S.84b, Sal91, ST93, Sch79, Sch93, SG72, Sha84, Shi91, Shi80, Sho72, Smi73, Smi96, Sne83, Sol75, Teu83, Tha80, Tur87a, Ves94, Wan95, Wei65b, Wei65a, Wei67, Wei72a, Wei78a, Wei84, Wei96, Wil71, Zim73, Zim78, ZM71, ZC75, Ano62f, AS13, Eck04, Hop11]. science-based [Etz93]. Science-Policy [BG23a]. science-production [Lew84]. Science-Public [MMMS^{+}23]. Sciences [Coh98, Eis13, End09, IH25, Kön04, Kuk11, Mil07, MK19, Rog13, Roz04, Sta03, Thi04, Whi11, Win94, Woe15, Xie21, XOR21, Hay93, Lew77, Nye75, Pri88, S.73b, Sho72, SL89, Ano65b, Cao98, Cro78, For92, Har04, Küh84, Lan87, Lew77, NW04, Cro81, Cro86, Coh98]. Scientific [AG25, Ano62g, Ano62f, Ano63o, Ano63p, Ano63q, Ano64p, Ano64q, Ano65q, Ano65s, Ano66v, Ash00, Aue80, Aug63, BC20, BGB18, Bek13, BD64, BD67, BB24, BRWB19, BBU19, BC17, Bro11, BBMR17, BPM84, CP99, Cro81, Dav20, DM73, DRC^{+}14, Dup63, EKB22, Fos97, Fox73, FvSA01, GCS00, GV18, GP04, GFM21, GP77, Häf97, Hah75, HD25, HJ12, JM19, Jor70, Jos94, Jul67, Kap66a, Kap66b, KBD75, Lau05, Lis67, MA69, Mac99b, MS13, McG79, dSNC23, Nic65, Ort16, Pan11, Pap04, PF21, PP23, Pet08b, Pfe70, Pig68, Pla99, QZL25, Rah65, Rak65, Rav78, Rod12, SSK21, SM97, Sch87a, Som08b, THE08, Ted95, Tra67, VV12, Vic63, VFMGEE13, Wei65b, Wei65a, Wei69, Wei00, WGL18, Zim77, Zim00a, Alb63, Ano64j, Ano64k]. scientific [Ano64l, Ano64m, Ano67-32, Ano70f, Ass93, BD65, BD70, CMR95, Car63, CW64, Cha94, Cla70b, CH85, Cro70, Cro79, DH64, Dev80, Doe70, Dot72, Dub63, Dub65, ED91, Enc70, Ezr80, Fis78, Fre87, Gir79, GMY73, Han65, Het90, Het91, Hoc87, Jon72, Kay66b, Kid83, Kor79, Lak68, Low92, Mad64, Mah64, Mar88, MI81, Mel74, Mel79, Mul72, Nye75, Pan93, Pap96, PK80, Rab76, Rai96, RD93, Ric88, Ric90a, Ros67a, Rot90, Row64, Roy84, Sal64, ST93, Sch91, Sch93, Sch78, Shi65, Sko69, Sne83, Tou64, Tou66b, Tou66a, Ves84, Ves88, Von73, Wei63, Wei64b, Wei66, Wei84, Wet85, Wid95, BEK12]. Scientism [Bak87]. Scientist [ALH09, Bal02, FFM16, Nov17, Sha17, Two99, vdB19, CM76, Mis73, Nic72, Rei80, Sal66, Shi70]. scientist-advisers [Nic72]. Scientists [Aue65, Bra12, Cao98, Cla70a, DGR19, FHS09, HSR19, Hut69, Hut70b, Hut70a, Jon66, Jon71, Kam09, KP02, Mac70, MS23, MN13, Mül14, Sch25, SD08, Ste74, THM17, Var70, Var71, Wer07, dGBM^{+}21, Ano63g, Ano63h, Ano87b, Ano87c, Bri77, DDS^{+}87, Eps86, Etz83, Gro96a, Gus94, Kid83, MLS91, MR95, MRC95, Orl80, Orl86, Pap73, RD93, S.70a, SM72, S.85, Sch87b, SB80, VW95, Ver81, Ves94, Wil66, Pét93]. Scoones [Lo 25]. Scopus [DGR19]. Scott [Mus22]. Screening [Hes23, Rey21]. scribing [FF10]. Scripps [Rai01]. Script [HM21]. Scrutiny [MS12]. Scuole [Gua88]. SDS [Ano69-44]. Sea [BFdR25]. Search [Bon08, GL04, Pre23, Han65, Mel72, Rit71, Umu86, Ves84]. Searching [Sku19]. Seas [KV25]. Second [Cai20, Hug06, Wer07, Woo75, Dev80, Het91, McG79, Ras02, Rei07]. Secrecy [S.84b, Sha84]. Secretiveness [Gas71, Pol72]. Secrets [Mac99b, Eam85]. Sector [BZ21, Kle23, LM04, Pou07, Sá05]. Sectoral [HH20a]. Security [EV12, Sha84]. Seeds [Bra15]. Segmentation [CT13]. seismology [Mel79]. self [Hun85, Joh85, S.83, SG72]. self-government [Hun85, S.83, SG72]. self-regulation [Joh85]. Selling [POSC07]. Selves [Smi08]. Semantics [Sch14]. Semiconductor [LAN14]. Senator [Joh71a]. Sense [Bro08, MS23]. Sensemaking [MS23]. sequencing [HK88]. servants [Cla70a, Hut69]. Service [Mac99b, Ano71d, Mil72, Pry71, Shi92c]. Services [Bru25]. Serving [Lab08]. setting [Sch72]. Settings [TT24]. Seventeenth [Cap98, Gre87]. Seventeenth-Century [Cap98, Gre87]. Shadow [God08]. Shah [Ano68g, vdB19]. Shake [EG06]. Shape [MN13, Orl99]. Shapes [dGBM^{+}21]. Shaping [BS07, Ben98, Hen21, Lem20, Nol99a, Rob21, Shi18, Gei97]. Share [Bar23, Kol85]. Sharing [Bar23, RYI13]. Shifting [Bux03, CH24, Fer23, Hei11, Rod12]. Shifts [PP25]. Shils [Ano96h, Ano96i, Bul96, Fle20, Hal00, Mac16, Orl96, Sha95, Wei96]. Shinn [Joh15]. Shirakawa [Low01]. Shogren [Gre05]. Short [BBV11, BCL76, Vol21a]. Short-cycle [BCL76]. Short-lived [BBV11]. Should [ML86, MN13, Wil64, Wil86]. Side [VL17]. Siebeck [Bre04]. siege [Ano88b, Ash62]. Sierra [Ano70j]. Signature [Mil07]. Significance [Bro11, Shi91]. Silicon [Ada09]. Similarities [THM17]. Simili [Wal10]. Since [Rob98]. Singapore [Puc72, PS65b, PS65a, Rat65]. Siren [Ano69-43]. Sisyphus [Mac88, Pos04]. Site [MKK22]. siting [Ass93]. Situated [Kal18b]. Situation [Sch05, Ano66o, Han68, Hol68, Ric65, Rüe75, Shi92b, Sio69]. size [Sut73]. Sketch [Rou24]. Skills [HD22, PJLP16]. Skirmishes [Ano68w]. Slack [CTLL22]. Sleepy [HH20c]. Small [Ras02, AM74, Enc70, Sch80a, Sch87a, Sko69]. Smith [Fra22, Til24, Lan87]. Snowbird [HR09]. Social [ALH09, Ano63j, Bad04, Bau20, BBU19, Bor21, BBMR17, Che71, Coh98, DGR19, Eng24, Es77, Fle20, Gee74, Gei88, Hei11, HJ25, Ibr23, IH25, Jen97, Kah16, KP02, Kin97, KV25, Kuk11, Mac99a, MARMSG20, MS12, MTE15, Orl71a, Pes03, PP23, Pic07, Rob21, SDvL19, SS22, Sne83, Sta03, SJ05, Vav98, VGM^{+}24, Xie21, XOR21, ZW25, dGBM^{+}21, Aga92, And70, Ano73a, Bul80, BB81, Con96, Cou79, DF96, Hay93, Joh73b, Orl72, Orl76, Orl86, Pat85, Pri88, RS77, S.73b, Sho72, SL89, Sol75, Swa70, Tur87a, Zim81, PJLP16, Xie21]. Socialism [Ric90a, Mac89, Pét95, Car78]. socialist [Ano67n, Ano69q, Ric90b, Bey06]. Socialization [MBS14]. Socialize [MN13]. Societal [BFdR25, DS16, HV23, Kle25, áRH23, Tel22, WR15]. Societies [Bor04, Gru17, Pes09, Zah06, Ano90a, Gol90, Shi92a]. Society [Bro13, BPM84, CA06, CÁCC20, Eng99, Gen23, GP04, HH20b, McG78, Pes03, Shi18, Ski06, Str03, SSF15, Tel22, TD10, Vav98, Wei12a, Zim06, Ano71d, Léc87, Orl96, Pol67, Shi92c, Wei67, Bey06, Sch06, Son21]. Socio [BL16, LB13, SBvE25, WW19]. Socio-Political [LB13]. Socio-Technical [SBvE25, WW19]. Sociocultural [HD22]. Sociological [KBWA21, BD64, GMY73]. Sociologist [Kor20, Bul96]. Sociology [Cam11, Din75, End09, Fis07, Hal00, Hes11, Sta95, TSL19, Tur98, VSR14, Wal92, Eis87, Fre87, Mor82, Orl96]. Sociotechnical [JK09]. Soddy [Gus12]. Soft [Tal09]. soil [SJ77, Web77]. Solid [Léc05]. Solidarity [Zim78]. solids [Hod80]. Solovey [Eng24, Xie21]. Solve [YB16]. Solving [Kle25]. Some [Aro64, Ash96, Dub65, Gun67, Joh65b, Joh72c, Kid83, Mon95, Mor66a, Mor71b, Nic72, Rie83, S.88c, Sei88, Psa71, Psa72, Sco71, Pap96]. Sometimes [YB16]. sophisticated [Teu83]. Soul [Gre07a]. Source [HvL11, áRH23]. South [Ano86c, Ano88b, Ano89, Boz75, GF02, Kah16, Rot86, Sal91, Ano63k, Ano67a, FG20, Fis63, JK09, PvR16, Sil64, Sku19, Mul20]. South-East [Ano63k, Fis63, Sil64]. Soviet [Ama70, Ano67p, DF96, Eng99, For92, Gil67, Gra71, Gro96b, HD25, Jos88, Kol78, Lew77, Lew84, Mac88, MJ78, Rab76, RLH91, Sol75, Sol03, Tsv14].

Sovietization [Kol04]. Soyuz [Vol21b]. Space [Ban14, HSR19, Het97, Kri97, LT16, SJ05, VGM^{+}24, Vol21a, WG97, Ban15]. Spaceflights [Het97]. Spaces [AD21, BFdR25, Dav20, Fer23]. Spaceship [Ank07a]. Spain [Ano65s, Ano66i, Ano66p, Ano66r, Ano67j, Ano67-35, Ano70g, Ano70h, DGR19, Gli05, NBGT05, SVMN15, SM97]. Spanish [Pui05]. speak [Ano75f, S.75b]. speaking [AA66]. Special [Dem17, GV18, HV23, Mac99b, Ano83f, Ted95]. Specialty [God12, Mul72]. Species [PKvdW23, Tro76]. specific [LNST20]. Specs [Het97]. Spectacle [áRH23]. Spectre [Sch07]. Speech [CH24, Cox94]. Spelman [BB81]. Sphere [SVMN15, TT24, KK81]. Spheres [MS13, S.72c]. Spin [CHJZSG20, NB02, TMP10, ZJE10]. Spin-Off [CHJZSG20, NB02]. Spin-Offs [TMP10, ZJE10]. Splice [Kro23]. Split [Kro23]. Sponsored [BK70, Orl71b]. Sponsorship [Nol99b]. Springer [Akm24]. Sputnik [Het97, Gei97, Kri00, Pla99]. Sri [Aus81, De 78]. SSH [HH20a]. stabilisation [Wei93b]. Stable [BC17]. staff [Tom96]. Staffing [CS63]. Staffs [Ano66t]. Stakes [Val11]. Staking [God14]. Stalin [Gro96b]. Stalinist [Con96]. standard [Aus80]. Standardisation [And22]. Standardization [BK18]. Standards [Log09, Bét81, Bic74]. Stanford [Ada09, KL94, KHB^{+}72, KPL^{+}89, Léc05, Low92, S.72c]. Stars [Bar15]. Starting [Wal92]. State [And97, AS13, Blo97, Car88, FAGH01, GS05, HP19, Lab08, Léc05, Mat08, Maz05, MKK22, S.83, Sap71, TT24, Ano66w, Ano69-44, Ano73c, Ano75g, Bel94, Dub63, Gee74, Gir79, Gru76, Kin63, Kor79, Nol83, Pét95]. Statecraft [Yan20]. Statement [Ano64r, Ano69-31, Ano90b, Bun81]. Statements [MSD21, PGO24]. States [BPM84, McG72, Rad19, S.86, Ano62h, Ano67-33, Ano69-27, Ano70i, AR20, Aue65, BD69, Bra93, Bun81, CMR95, Che71, Cla74, Cum90, Dup63, Dup90, Etz96, Fin77, Gei97, Gen87, Hun81, Hun85, JK09, Kid83, Kor62, Lai20, Mac88, Maz97, Moo99, Orl71a, Par00, PK09, Ras02, Rie69, Rit66, RS77, S.83, S.84a, Shi78, Shi79, Smi73, Smi07, Spe85, Tha80, Tri79, VW95]. statesmen [Jon66, Ste74]. statistician [Léc87]. Statistics [Des09, God02, Deb69, Léc87]. Stature [McL99]. Status [CT13, EL22, Lee06, Hun81, MLS91]. Status-Bestowing [EL22]. steady [Ano69-44, Ano75g]. Steering [FBH23]. Stehr [Gen23]. STEM [SL22, BWF24, FBF^{+}21]. Stenlås [AS13]. Stephen [Fle20]. Steps [Col00, Ano69i]. Stewart [Fra22]. Sticky [Lai20]. Stigma [XH22]. Still [AR20, Gas07, Ros67b, Shi94b]. Stimulates [OPBCM15]. Stockholm [BM18, STK^{+}91]. stocktaking [Tou64]. Stories [Bos05, Lai19]. Story [PKvdW23]. STP [MN13]. strain [Shi75]. Strains [Kor62, Ano92, SB92]. Strategic [FH13, TP16, Wei87a, WS25, BS07]. Strategically [LH19]. Strategies [FG20, Tob19, Wei74]. Strategy [Mac08]. Stratification [CT13, KMB24]. Stratum [Div98]. Straughan [Bal99]. Strengthening [Kri00]. Stress [EL22, SLL19]. Stress-Inducing [EL22]. Stress-Testing [SLL19]. Stretching [FP17]. strike [Ano67-38]. Striking [Hol74]. Strip [Sul91]. Structural [BGB18, BB24, Yli14]. Structure [CCSM18, Hei11, Woo03, Rai96, ZM71]. Structures [Hei11]. Structuring [NM16]. Struggle [Bra15, Ano67-35, Cla70a, Hut69, Pri78, Shi84]. Struggles [EB22, MB12]. Struggling [Sch78]. Struve [Dev80]. STS [Goñ25, MN13, Rad19, Roz05, WHF18]. Stuart [Kuk11]. Student [Ano66x, Ano67z, Ano67-36, Ano67-38, Ano69q, Ano69r, Ano68z, Ano70k, BL16, Enc03, GCD23, Lee06, MBS14, Moo99, AKW70, Ano66d, Ano66i, Ano66n, Ano66r, Ano67f, Ano67j, Ano67p, Ano68k, Ano69p, Ano69x, Ano69-35, Ano69-33, Ano90a, ETMT70, Fis63, Gol90, Per68]. Students [Ano66w, Ano67-37, Ano69n, Ano68y, Ano68x, Ano69s, Lee98, Lee06, Thu10, WHF18, Aic63, Ano64s, Ano67a, Ano67e, Ano67p, Ano67r, Ano67s, Ano67t, Ano69b, Ano69g, Ano69t, Ano81d, Lip64]. Studiengruppe [Bri06, Coe06, Kra06]. Studies [Abb09, AK11, Aro12, CL22, DT12, Elz12, Eng99, Gen23, God12, God14, Hes11, Kin97, Mad18, MO17, Ort16, Pfi15, Sku19, Tau12, VV12, VGM^{+}24, WW12, Wil20, Bio63, Div91, Löw87, Mel72, Rit71, Sol75, Rob98]. Study [Ano83f, BBU19, CBB20, FHS09, God08, God10, Hei11, Hes13, LNST20, LAN14, Mül14, NS25, Pic07, Rou19, SJPA14, THE08, VS20, WBW18, BK70, DK74, Gib70, Gra66, GMY73, Hal75, Hod80, Kim82, MJ78, Rab76, Ros67a, ST93, Sin62, Tha80]. Stuhlinger [Het97]. styles [Wet85]. Subject [LBP10, Sha17, vdB19, Mom83]. Subjects [OB17, Sta02b, Pat85]. subordination [Gil72]. subservience [Puc72]. such [Lub07]. Suddenly [Het97]. suggested [Gru76]. Suggestions [Ano63r, Mor66a]. Super [HK00]. Superconducting [HK00]. supersonic [Cla74]. Supervision [Mül14]. supply [Gis69, Gis70, Gru70, Joh70]. Support [Bux03, Gli05, Ano64s, BC85, Blu69, Cro79, DK74, Enc71, For74, Gro82, Het90, Het91, Joh65b, Mac71, Mon88, Pav76, SG72, Sei81, SL89, Wil71]. supported [Phi78]. Supporting [TH24]. Suppression [SREM^{+}23]. Supreme [S.83]. Survey [BI22, Hro13, Tur87b]. surveys [MLS91]. Surviving [FGR25, Goo02]. Suspension [Ano66y]. Sustainability [Log11]. Sutter [Smi08]. Sweden [Blo97, BEK12, HJ16, Hal63, PKvdW23, Pet18, S74e, Sjö07, Wid95, AS13]. Swedish [Bod64, Ano64s, BS07, Bod64, EB22, Hus77, MK19, PH80]. Sydney [Pre23, Enc03]. Symbiosis [FBF^{+}21]. Symbol [Pie12]. Symbolic [Rou19]. symposium [S.92]. Synergy [MS13]. Synthetic [LP79, Hab73]. System [Aag17, Cas05, FH13, GMD^{+}24, Hes13, Kah16, Kim08, NM16, Pla99, QZL25, Ama70, Gra71, Klo71, PP68, Por70, Sto65a, Zim63, ZM71]. System-of-Systems [GMD^{+}24]. Systemforschung [Bri06, Coe06, Kra06]. Systemic [BC20]. Systems [Bri06, DT12, End09, GMD^{+}24, HL22, Lai20, LM04, Lau17, Pra12, Ano96k, DA93]. Szöllösi [Eck04]. Szöllösi-Janze [Eck04].

T [BPM84, Het97]. Tacit [Ray09]. Tackling [Mus22]. Tactics [SREM^{+}23]. takeover [Pét93]. Taking [KS01]. Tale [Hes23, Rot04]. talented [Ano81d, ML81]. Talking [Ort16]. Tanzania [Ano67n]. Targeting [WW19]. Task [VP22]. Taxonomy [dSNC23]. Teacher [Tob19, Bus63]. teachers [Ano67a, Ano82c, Ano82d, Kay84, Ker89, Lad70, McC76, Mor92, Seg85]. Teaching [Ano66t, Ano66z, Bja01, Wei12b, Ano93b, Ano93c, Cra66, Mil72, Mom83, Pry71, S.73a, S.93, Sub66, BPM84]. Team [Smi99]. Technical [Ber11, Div98, Har10, IH25, Mac76, McG79, Mor64, Ree75, SBvE25, Ano70f, Jor70, WW19]. technician [Ano63k]. Technicization [ZW25]. Techno [SKS13]. Techno-Nationalism [SKS13]. Technocracy [RR05]. Technological [Ano63o, Ano63p, Ano63q, Bek13, God08, God10, GP04, MB10, SJ90, Tri79, VP22, Wal98, AM74, Bro72, Dan67, FG92, Lew77, Lew84, Maz75, Nel71, Rav78, Wal96, Wei93a]. Technologies [CBKN17, Gus12, Jas03, PW09, Rob21]. technologists [S.70a, Shi84]. Technology [AK11, AS13, BBSS03, BG23b, Bro07b, Bru09, Cop02, Dem17, Eis13, FD04, GS05, God02, Goñ25, HI06, Jun22, Kaa11, Kis24, Kor06, Kri97, Kuw01, MTS20, Mah00, Mat08, Mol05, PS01, Pol05, PD17, RLL84, SKS13, Sal00, Stu98, Suz01, TC05, VV12, Wal96, Wei90, Wei12b, Wei15, Ano63i, Ano73b, Bau77, Bra07a, Bra92, Gro80, Kel84, Koi95, LM94, Mac88, Maz84, Pap88, Rad83, RR71, Rud68, Sal91, Sch94, Spe85, Teu83, Tri71a, Wil83a, Etz94, Log09, Pfe70]. Technoscience [Var12, Weh15]. Technoscientific [Deu07]. Telephone [Hod80]. Temptations [Jon72]. Tendencies [Pap04]. Tension [HJ16, RG21]. Tensions [BDG10, EV12, HJ12, NS25, SMaR21]. Tenure [WHF18, Ano83e, CSKF85, Mor92, Shi95]. Term [Kaa11, Orl99]. Terradas [RR05]. Territorial [NS25]. Territory [HLE24, Tel22]. Terry [Joh15]. Tertiary [Opi04, Lin81, Tou66b]. Test [Vol21b]. Testimony [Fos97, MLS91]. Testing [SLL19, SSF15]. Their [BGP14, DGR19, KKK10, Lee98, OPBCM15, Rog13, VFMGEE13, Whi12, Ano94, BFdR25, Coi81, Etz92, Maz89, Pri78, Rob94]. Them [UMC^{+}23]. Theme [Kol04, Bak87, Dub63]. themselves [Ula83]. Theoretical [Hes13, VL17]. Theory [CHJZSG20, Ful00, Pav76, Pol00, PC02, Zim00b, Hod80, Roy84]. Therapeutic [Her24, Vuo15]. There [EBMP24, SK00, Kie85, Ros67b, Sut73]. Think [Mol05, Whi09]. Thinking [Par02, Rot07, Top02, dlM03]. Third [Ano65q, Ano67-27, Eck04, NW04, Vol21b, Bau77, Eck04]. Thomas [Mar20, Mol05]. Thomasine [Bal99]. Thomasma [Bal99]. Thorn [VL17]. Those [Tuc00]. Thought [Fle20, For08, Kri80]. Thought-Collectives [For08]. Thoughts [Ano94, Cra66, Hoo66, Jon76, Mor66b, Nol83, Spa66, Sto65b, Sub66, Rie83, Ron79]. threats [Sha92b]. Three [Aue80, HvL11, Zim71, Ano96k, Lin81, Maz84]. tickets [Ano69q]. Tide [Mar06]. Tiger [Gus12]. Tijssen [Mul20]. Time [CT13, Rou24, Kol78, MJ78]. Time-Consuming [Rou24]. Time-Dynamic [CT13]. Timely [BG23b]. times [DH64]. Timing [Nol99b]. tissue [Ted95]. Tizard [Jon66]. TNO [vR13]. Tobit [CT13]. Today [Mid01, Pes09]. Together [PKvdW23]. Tomorrow [Het97, Mid01]. Tool [Rey21, Tal09, Sto66, BGS16]. Tools [Hes23, KW11, PW09]. Top [TP14].

Toronto [Ros72]. Toteff [Fle20]. Touch [Vuo15]. Toulouse [Nye75]. Tower [CG22]. Towers [Aus81]. Town [Bro00]. Track [WHF18]. tractor [AM74]. Trade [KW11, Law81, Kay84, S.83]. Trading [MPW16]. Tradition [Fox06, Ahm89, Ama70, Gil72, Kak69, Mat69, Shi91]. Traditional [Met08, Kim82]. Traditionalist [BWF24]. traditions [Rei07, Rüe92]. Training [FBF^{+}21, Thu10, Tor18, Ano63k, Con96, ED91, MR95, MRC95, Pap73, Rit66, SJ90]. Trajectories [Kor20]. Trajectory [Joh09]. trans [Joh72b, Wei72b]. trans-science [Joh72b, Wei72b]. Transatlantic [Ban14]. Transcendentalism [Hoo08]. Transfer [CTLL22, FD04, GS05, GM09, SKS13]. Transformation [Cas05, DA93, De 97, SGSB23, SBvE25, SB13, Sut98]. Transformations [SJ05, Tsv14]. Transformative [Aag17, DR12, RG21]. Transforming [Kal18a, Whi12, Mul20]. Transgenic [Bra15]. Transition [Ano98b, PvR16, TMP10, Küh84, Tur87a]. Translational [HSR19, Hen21]. Transmission [TD10]. Transnational [And22, CBKN17, Can11, Eng24, MG24]. transplantation [Ted95]. Transport [Ano69t, Cla74]. Trash [Rae08]. Travel [Het97, Ano93a]. Travelling [FP18, PJLP16]. Treasure [VL17]. treatment [Maz89]. Trees [MMMS^{+}23]. Trenchard [Ano67-40]. Trends [JS65, Tom96, WB65]. Trial [Fos97, Vos23a, Vos23b, Ano67-27]. Tribe [God14]. tribulation [Shi94a]. tribulations [Ano69l, DF75]. Trip [Bor21]. Triple [Ano98b, Cai20, CL22, HP19, MS11, Thu10]. triumph [Ano69x]. Trojan [S.73c]. Tropical [Fer23, AA66]. Trouble [Rül23]. Trowbridge [Gli05]. Truscot [Bel97]. Trust [Bar87, Fre23b, Cro94]. Truth [AC23, FD04, HV23, Pre23]. Tübingen [Bre04]. turbulence [Ano69o]. Turbulent [Lo 25]. Turin [Gua88]. Turkey [Coh68, Ged95, Mar69, Oky68, Ozd68, Ree75, Umu86]. Turmoil [Sut98]. Turn [KBK13]. turned [Ano90d, Ano90e]. Turner [Fle20]. Turning [Two99]. twentieth [RSS97, Fox97, Neh06, Sut98]. Twentieth-Century [Fox97, Neh06, Sut98]. twenty [Gro92, Rüe92, S.92, Sha92a, Sha92b, Shi92b, Shi92c]. twenty-first [Gro92, Rüe92, S.92, Sha92a, Sha92b, Shi92b, Shi92c]. Two [Die05, Gir79, Hes23, Ked93a, Ked93b, Kor79, MT19, McL99, Rot04, Rou07, SL24, TP14, VSR14, Bre95, DH64, MLS91, RLL84, Spe85, Wan95, Wei64b, Wei93b]. Types [FSHdR18, Rou07]. Typhoon [And22]. Typology [Kam09, SF25, DF96].

U.S [Ano65b, Kri97]. U.S. [BC17, BI22, CT13, CBB20, FBF^{+}21, GOR25, Ren08, THM17, TP14, WBL15]. U.S.S.R. [Ano66v, Kap66a, Kap66b, Kin63]. UC [Ada09]. UK [Bro20, Bro11, Joh09, KW11, MH16, SSK21, TD10]. Ukrainian [JE97]. ultimate [Mor89]. Umbrella [KMB24]. Uncertain [Ber11]. Uncertainty [Lo 25, Näs21, Tob19, Mar88, PK80]. Unclear [LH19]. Undefined [HSR19]. Underdeveloped [Ded63, And69, Ano71h, Bas67, Coo67a, Lip64, Mor64, Nay67]. Underexplored [Mel24]. undergraduates [Ano93b, Ano93c, S.93]. Understand [BWF24, McG16, Ros02, UMC^{+}23]. Understanding [EKB22, Hag04, Har10, Pes09, Rou07, S.74d, Sch93, Rül23]. Understandings [LNST20]. Uneasy [BWF24, Joh73a]. Unemployed [Rod12]. UNESCO [Pet08b]. Uneven [KOT14]. ’Unfashionable [Kem06]. Union [Ano69n, Nol99a, Jos88, Kol78, Lew77, Mac88, Nol99b, Rab76, RLH91]. Unions [Ano68z, Law81, Ano66i, Ano66r, Ano68k, Kay84, S.83]. Unique [Jen01]. United [BPM84, Ano62h, Ano63h, Ano67m, Ano67l, Ano67-33, Ano69-27, Ano70i, AR20, Aue65, BD69, Bra93, Bun81, CMR95, Che71, Cla74, Cum90, Dup63, Dup90, Etz96, Fin77, Gei97, Gen87, Hun81, Hun85, Hut72, JK09, Kid83, Kor62, Lai20, Lee06, Mac88, Mah00, Maz97, McG72, McG79, Moo99, Orl71a, Par00, PK09, Pri98, Ras02, Rie69, Rit66, RS77, S.83, S.84a, Shi78, Shi79, Smi73, Smi07, Spe85, Tha80, Tri79, VW95, WG97]. Unity [Eis13, Wei84, DH64]. Universal [BGS16, Cao14, EK01, Tro00, Ano71f, Hal75, Joh73a, Mor73, Shi91]. Universalism [Som08a]. universality [Elk77]. Universities [AR92, Aic63, Ano65i, Ano68v, Ano68-27, Ano69y, Ano74b, Ano75f, Ano75g, Ano78e, Ano86c, Ano88b, Ano92, Arn07, Aro64, Ash62, Aus80, Bek13, Bel68b, Bel94, Blo97, BDG10, Bri05, BC17, BI22, CT13, Cha24, Coh68, Cox94, CBB20, Des65, Div91, Enc00, EBMP24, Fra91, FF90, Geu98, Goo02, GKL98, HL10, Hal00, HdRW17, HJ08, Hun80, Ked93a, Ked93b, KHCJ19, Kit10, KOT14, KHS24, Kol04, Léc05, Mar69, MG24, Maz89, McG72, Oky68, Ozd68, Par65, Pas78, PGO24, Pie64, POSC07, PD17, Ree75, Rie69, Rob03, S.75b, S.86, SB92, Sch07, Sco05, Shi94a, SB13, Sul91, TP14, TP16, TT24, Tsv14, Vav98, Whi12, WS25, Ano63r, Ano64n, Ano65c, Ano65p, Ano67h, Ano67i, Ano67-34, Ano68c, Ano68i, Ano69k, Ano68t, Ano68x, Ano69s]. universities [Ano68-28, Ano69u, Ano76d, Ano94, Ash66, Ash68, Bel67, Bel68a, BD69, Bla68, Blu69, Bow67a, CS63, Coo67b, Cra67a, Cro94, Daa74, Dan67, De 78, Elv68, Etz83, Fin77, Gir79, Han68, Hay93, HJ07, Hol68, Hoo87, Jam68, Kar63, Kar65, Kid66, Kol85, Kor79, Lim95, Loc89, MR74, MT67, Mer79, Mon95, Mor82, MJ78, Nol83, Pau67, Por70, Pri78, Ric65, Rob94, Rot86, S.83, S.88c, Sch72, ST96, Shi78, Shi79, Shi92b, Sil64, Sin92, Sio69, Ula83, Ves88, Wei76, Wei79, Wil71, oC68, Aus81, Bod64, Pre23]. University [And69, Ano64a, Ano65r, Ano65t, Ano66t, Ano66-27, Ano67b, Ano67w, Ano67-41, Ano70l, Ano70m, Ano98b, Bel97, Bel98b, Ben13, Bes68, Bét81, Blo97, Bre95, Cai20, CHJZSG20, Cap98, DV98, Dha84, DPB^{+}20, Fle20, Gei97, GM03b, Gib00, Gil78, HdRW17, Hay94, Her24, JS65, Kaa11, Kay84, Klo68, Law81, Lip64, Lob84b, Mac08, MS23, Mar20, Met10, Mid01, Mol05, Muc85, Nov17, Orl99, PvR16, Pet18, Pre23, Rey87, Rot95, SNSB^{+}80, Sca03, Son21, SC04, TH24, Thu10, Til24, Vol21a, Wil72, Yli14, vR14, And70, Ano66a, Ano66o, Ano66p, Ano66y, Ano66z, Ano66-28, Ano67e, Ano67x, Ano67u, Ano67v, Ano67-33, Ano68b, Ano69t, Ano69-30, Ano69-45, Ano77e, Ano82c, Ano82d, Ano90a, Ano90d, Ano90e, Ano96k, Ash67, BW70, BK70, Bic74, Bio63]. university [Boy66, Bus63, Cas96, Col86, Coo72, Dra84, Edu66, Fis63, Ged95, Gho88, Gol90, Gro92, Hen77, Hoo66, Hoo89, Hus77, Joh73b, Ker89, Kol83, Lad70, Lim74, MT67, McC76, Mel63, Mor66b, Nol82, Orl71b, Par68, Pay80, Ped77, Rüe92, S.70c, S.71a, S.82, S.92, Sad91, Sch87b, Sch86, Sha92b, Shi73, Shi89, Shi92a, Spa66, Sto65b, Sut73, Swa70, Wei64a, Wei79, WB65, And07, Ano64r, Ano64s, Ano65a, Ano65u, Ano65v, Ano66b, Ano66g, Ano66-29, Ano67c, Ano67n, Ano67-39, Ano67-40, Ano67-42, Ano69f, Ano68l, Ano68p, Ano68s, Ano69-43, Ano69-45, Ano70i, Ano75f, Ano80a, Ano81c, Ano83e, Ano83f, Ano86d, Ano91, Ash68, BFGS14, Bar98, BBSS03, Bod64, Bro00, Bul80, Car78, CS15, CG22, Cum73, DV98, Din75, DF75, Enc03, Enc06, ESC^{+}88, FD04, Gau73, GS05]. University [GM09, HR15, Hay88, Hoo84, Hut75, Hys64, Inn92, IH25, Joh74b, KHB^{+}72, KPL^{+}89, Klo71, Kol83, Kru05, Lee07, Low92, Lub02, Lub04, LT16, Mar06, McG72, Mon95, Mon98, Mor85, Mül14, MB12, NM16, NB02, PD71, PC02, Puc72, Rai01, Rey87, Ros72, S.75b, STK^{+}91, S.92, SKS13, Seg85, Sha92a, Sha87, Sha84, SBvE25, SK00, Som92, Son21, SC04, Szc68, Wal92, Wal88, Wan95, Wei87a, Woo03, Bel98a, Goo99, Van15]. University-Government [Mon98]. Unlikely [BB24]. Unpacking [Mad18]. unpersuasive [Sch83]. unresolved [S.74f]. Unsustainable [FFM16]. Unwelcome [Kön04]. Uppsala [Blo97]. upside [Ano90d, Ano90e]. Urban [PD71, THE08, Coo72]. Ursula [Fre23a]. US$35.00 [Til24]. USA [Joh09]. Usable [OPBCM15]. Use [Col07, EB22, Kay12, SSK21, SDvL19, Smi99, Kem06, Osb71, Row64, RS77, Sad91, Wei78b]. useful [KL94, Wei72a]. Using [ACGG14, DGR19, LBP10, McG16, MI81]. USSR [Lis67, Tra67]. Utility [HJ16]. Utopia [LK22]. Utopian [Ezr80].

Vaccination [Chl23, Fre23b]. Vaccine [Fre23b]. Vale [Mac07a]. valiant [Ano69-46]. validity [Shi91]. Valley [Gei08, Ada09]. Valparaíso [KBWA21]. Valuation [HFSdR19, Log11]. Value [BS11, LBA16, Nel21, Sla11, Som08b, Tob19, Mot64, Wid07]. Value-Laden [Som08b]. Values [Cao14, DGR19, Mey11, Tal09, Val11, Wei84, Hay94, Wil72]. Vanderbilt [Ano83e]. Variables [Nov17]. Variation [Dav20, HFSdR19]. Variations [Kol04]. Varied [Whi12]. Varieties [FP17]. Vast [Gra12]. Venezuela [Ves84, Ves94]. versus [Bro07b, Div91, Gam97, Koi95, Mor73, Pét93, Wil71]. Veterinary [LM94]. VI [Bel98b]. via [HD22, XH22]. Vice [Cla95, Ano67-42, Ano63s, Ano64a, Ano65a, Ano65u, Ano66b, Ano67c, Ano67-39, Ano67-40, Enc07]. Vice-Chancellor [Ano63s, Ano64a, Ano65a, Ano65u, Ano66b, Ano67c, Ano67-39, Ano67-40, Enc07]. Vice-Chancellors [Cla95]. vice-chancellorship [Ano67-42]. Vicissitudes [Enc07, Bus63, Lak73]. Victims [KL03]. Victorian [Bro05, Mac71]. victory [Ano69-33]. Vietnam [Ano67e, Ano68-28, Ano69u, Ano69v, HD22, Lad70, McC76, NM16]. Vietnamese [Ano68-29]. view [BD64, Phi78, Rai64, Tom96]. Views [SRTC10, Gir79, Kor79]. Vinck [Wil20]. VINITI [HD25]. Violence [Ano66-29, Rou19]. Virology [vR10]. Virtual [PC02]. Virtue [BRWB19, Ham05]. Virtue-Free [Ham05]. Viruses [MO17]. vis [Smi08]. vis-à-vis [Smi08]. Visibility [KKK10]. Visible [Wer07]. Vision [Mac16, Mar20]. Visions [Dol05, NSB25]. visit [Ano68g]. Visualising [WV17]. Vito [Pan93]. Vol [Bel97]. Volterra [Pan93]. Volume [Ano97, Ano00b, Ano01a, Ano01c, Ano02a, Ano02b, Ano03, Ano05b, Bel98b, Cap98, Ano04]. voting [AA71]. vs [DPB^{+}20, SVMN15]. Vulnerability [Fox97].

W [Cha24, Goo99, Het97, Son21]. W. [God08]. Wales [Lim95]. WALKER [Wal06]. Wallace [Mul20, End06]. Wallerstein [Coh98]. Walloon [Ano66c]. Walpurgisnacht [Hoo84]. Wanted [Deb69]. War [AS13, CG02, Eng24, Guo13, Joh02, Kro03, Pet08b, Rim98, Sch03b, Sel08, Smi99, Wan95, Ano67e, Ano68-28, Ano69u, Ano68-29, Aue65, Ben07a, Lad70, McC76, Aro12, Bor21, Dev80, Elz12, Fel19, Gro96b, Het91, Lyn03, McG79, Pan93, PP25, Ras02, Rei07, Som08a]. Warfare [AS13, Bal02, GF02, Joh02]. Warming [Gra12]. Warning [And22]. warrants [Rot66, Tou67]. Warren [Koh76]. Wars [HV23, Koe07, Wei93b, Ber05]. Wartime [Tru06]. was [Lee98]. Washington [Gre05, Gre05]. Water [Jac06, SV10, CH85]. Watermeyer [Fra22]. waters [Ano69h]. Way [RvDG20, Vel18, Wal92]. weapons [Lit07]. Weaver [Koh76]. Webb [Het97]. Weber [Bre04, Kem06, Kem08]. Webers [Bre04]. Weimar [For74, Har04, SG72, SJ05]. Weinberg [Gib70, HJ12]. Weiss [Høy24]. Weissermel [Ano82e]. Welfare [AS13, Lee06, PJLP16, Bec73, Joh73b, Joh74a, Mis74]. Well [Mos08]. west [Hay94, Aic63, Ano67a, Ano68y, Ano71d, Bes68, CG02, Gui06, Klo68, Klo71, Lew80, Met08, Pay80, SL18, Sul91]. Western [Ano62f, Ano63r, Ano66t, Ano67-32, Aug63, Bau77, Dei79, KPL^{+}89, Kri00]. Where [FHS09, Var12, Cas96]. Which [CL22]. White [Gam97, Gre05]. Whitley [End09, Rog13]. Who [Die05, GCD23, Sha17, vdB19, RC67, Ros67a, Ros68]. Wickliffe [Koh85]. Wider [Gas07, Mel79, Sch72]. Wild [Met08, Met08]. Wilhelm [Bey06, Sch06]. Wilhelmine [SJ05]. Will [Koi95]. William [Orl99, God10, Jen01]. Willing [Smi08]. Willits [Sta03]. Wilson [Ben98, Orl86]. Winning [O’S01]. Wissenschaft [SJ05]. Wissenschaftsrat [ML81]. Within [Tom98, TT24, BFdR25, Ula83, Zim78]. Without [Bal02, God10, Rou24, Enc71]. Wolf [LK22]. Women [Ano01e, EG06, Kid07, Kuw01, Sta02b, TH24, ZC75, vdW01, KE01, KKK10]. wonderful [Low92]. Word [God10]. words [Lub02]. Work [ALH09, Chl23, Col98, ELSS24, Fle20, FFM16, Gel00, HSR19, Kal18a, Mül14, NSB25, PC02, Pos04, Röd17, Rod12, Sap17, Smi99, Vuo15, Ano69-44, Elk77, Lag87]. Worker [Ano68z, Sco07]. Workers [Pet08b, Thu10]. Working [PC02]. World [Ano89, Ban15, Ben98, Ber11, Dev80, Het91, Lee06, McG79, MSD21, Mol05, Pan93, Pet08b, Ras02, Rei07, Rya01, Wan95, Wei97, Whi09, Ano66m, Ano66s, Ano90d, Ano90e, Bau77, Deb69, ML89, Sch91, Sch94, Wei93b, Bro72, Nel71, Til24, Lo 25]. World-Rank [MSD21]. Worlds [Hag04, SRTC10, VSR14]. Worldwide [Zap22]. Worth [FFM16, Näs25]. Write [Høy24]. Writing [Arm07, Vel18]. Wrong [Ort16].

XII [Eck04]. XL [Ano02b]. XLI [Ano03]. XLII [Ano04]. xxi [Pre23]. xxviii [Til24]. XXXIX [Ano01c]. XXXV [Ano97]. XXXVII [Ano00b].

Yale [Ano90b, Vol21a, Ano66z, Ano90c, S.75b, S.90]. Yannick [Ber11]. Yearbook [End09]. Years [Ban15, BPM84, God02, Pei84, RR05, Wei12a, Bra07a, Bre95, Hut70a, Lob84b, Mos78, RLL84]. Yerkes [Dev80]. Ylönen [Weh15]. Yoke [Bra07b]. York [Eck04, Hoo84]. young [MR95, MRC95, RD93]. younger [Kid83]. Yourself [Fal21]. Youth [RD93]. Yugoslavia [Ano66q, Ano69q, Nic65, Rah65, Rak65, Raž94].

Zambia [Ano65r]. Zdenek [Win94]. Zealand [MMMS^{+}23, Opi04]. Zhou [Cai20]. Zipfel [SNSB^{+}80]. Zone [vR10]. Zones [MPW16]. Zoomification [BG23b]. Zurich [Kru08].

References

Ashby:1966:AAF

[AA66] Eric Ashby and Mary Anderson. Autonomy and academic freedom in Britain and in English-speaking countries of Tropical Africa. Minerva, 4(3):317–364, Spring 1966. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

             http://link.springer.com/accesspage/article/10.1007/
             BF01584631.

Ashby:1971:CVP

[AA71] Eric Ashby and Mary Anderson. Consultation or voting power. Minerva, 9(3):400–406, July 1971. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

             springer.com/accesspage/article/10.1007/BF01553363.

Abir-Am:1988:AIR

[AA88] Pnina G. Abir-Am. The assessment of interdisciplinary research in the 1930s: The Rockefeller Foundation and physico-chemical morphology. Minerva, 26(2):153–176, Summer 1988. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

             URL http://link.springer.com/accesspage/article/10.
             1007/BF01096694.

Abir-Am:2004:DIB

[AA04] Pnina G. Abir-Am. DNA at 50: Institutional and biographical perspectives. Minerva, 42(2):191–213, ???? 2004. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1023/B%3AMINE.0000030022.66674.ac.

Aagaard:2017:ENR

[Aag17] Kaare Aagaard. The evolution of a national research funding system: Transformative change through layering and displacement. Minerva, 55(3):279–297, September 2017. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/article/10.1007/s11024-017-9317-1.

Abbri:2009:BRL

[Abb09] Ferdinando Abbri. Book review: Lawrence M. Principe (ed.), Chymists and Chymistry. Studies in the History of Alchemy and Early Modern Chemistry. Minerva, 47(1):115–118, March 2009. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-009-9116-4.

Abelson:1965:PSA

[Abe65] Philip H. Abelson. The President’s science advisers. Minerva, 3 (2):149–158, Winter 1965. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096331.

Akrich:2023:MTP

[AC23] Madeleine Akrich and Franck Cochoy. A masked truth? Public discussions about face masks on a French health forum. Minerva, 61(3):315–334, September 2023. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL https://link.

springer.com/article/10.1007/s11024-023-09493-0.

Amanatidou:2014:UER

[ACGG14] Effie Amanatidou, Paul Cunningham, Abdullah Gök, and Ioanna Garefi. Using evaluation research as a means for policy analysis in a ‘new’ mission–oriented policy context. Minerva, 52(4):419–438, December 2014. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
014-9258-x.

Ackers:2008:IMM

[Ack08] Louise Ackers. Internationalisation, mobility and metrics: a new form of indirect discrimination? Minerva, 46(4):411– 435, December 2008. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-008-9110-2.

Aceros:2021:PIP

[AD21] Juan C. Aceros and Miquel Domènech. Private issues in public spaces: Regimes of engagement at a citizen conference. Minerva, 59(2):195–215, June 2021. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL https://link.springer.

com/article/10.1007/s11024-020-09423-4.

Adams:2009:FME

[Ada09] Stephen B. Adams. Follow the money: Engineering at Stanford and UC Berkeley during the rise of Silicon Valley. Minerva, 47(4):367–390, December 2009. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
009-9138-y.

Albu:1976:BR

[AE77] Austen Albu and Yaron Ezrahi. Book reviews. Minerva, 14(4): 575–585, Winter 1976–1977. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01097640.

Aman:2025:IKF

[AG25] Valeria Aman and Jochen Gläser. Investigating knowledge flows in scientific communities: The potential of bibliometric methods. Minerva, 63(1):155–182, March 2025. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL https://link.

springer.com/article/10.1007/s11024-024-09542-2.

Agassi:1992:RPS

[Aga92] Joseph Agassi. Rationality: Philosophical and social aspects. Minerva, 30(3):366–390, Autumn 1992. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01097644.

Atkinson-Grosjean:2009:MES

[AGF09] Janet Atkinson-Grosjean and Cory Fairley. Moral economies in science: From ideal to pragmatic. Minerva, 47(2):147–170, June 2009. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-009-9121-7.

Ahmed:1989:ATA

[Ahm89] Ajuji Ahmed. The Asquith tradition, the Ashby reform, and the development of higher education in Nigeria. Minerva, 27 (1):1–20, Spring 1989. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096712.

Aich:1963:AAS

[Aic63] Prodosh Aich. Asian and African students in West German universities. Minerva, 1(4):439–452, Summer 1963. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01107187.

Albert:2011:BPB

[AK11] Mathieu Albert and Daniel Lee Kleinman. Bringing Pierre Bourdieu to science and technology studies. Minerva, 49(3): ??, September 2011. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-011-9174-2.

Akmalia:2024:BRS

[Akm24] Frida Akmalia. Book review: Sarah Elaine Eaton, Jamie J. Carmichael, Helen Pethrick, eds., Fake Degrees and Fraudulent Credentials in Higher Education, Springer, Cham, 2023, 291 pp. Minerva, 62(4):639–642, December 2024. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL https://

link.springer.com/article/10.1007/s11024-024-09546-y.

Ackner:1970:AFS

[AKW70] Q. C. Desmond Ackner, G. W. Keeton, and H. W. R. Wade. Academic freedom and student disruption. Minerva, 8(1–4): 100–110, January 1970. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01553201.

Albu:1963:MPE

[Alb63] Austen Albu. The member of Parliament, the executive and scientific policy. Minerva, 2(1):1–20, Autumn 1963. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01100777.

Albert:2009:BWH

[ALH09] Mathieu Albert, Suzanne Laberge, and Brian D. Hodges. Boundary-work in the health research field: Biomedical and clinician scientists’ perceptions of social science research. Minerva, 47(2):171–194, June 2009. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-009-9120-8.

Altbach:1970:BC

[Alt70] Philip G. Altbach. Bombay colleges. Minerva, 8(1–4):520– 541, January 1970. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01553237;   http://
link.springer.com/article/10.1007/BF01553237.

Aurora:1974:DTC

[AM74] G. S. Aurora and Ward Morehouse. The dilemma of technological choice in India: The case of the small tractor. Minerva, 12 (4):433–458, October 1974. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01557639.

Amann:1970:SRD

[Ama70] Ronald Amann. The Soviet research and development system: The pressures of academic tradition and rapid industrialisation. Minerva, 8(1–4):217–241, January 1970. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01553213.

Anderson:1968:UPU

[And69] C. Arnold Anderson. University planning in an underdeveloped country. Minerva, 7(1–2):36–51, Autumn–Winter 1968– 1969. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096807.

Anderson:1970:SOU

[And70] Robert S. Anderson. The social organisation of a university laboratory. Minerva, 8(1–4):297–299, January 1970. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01553218;   http://link.springer.com/article/
10.1007/BF01553218.

Anderson:1991:OIR

[And91] Robert S. Anderson. The origins of the international Rice research institute. Minerva, 29(1):61–89, Spring 1991. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01096298.

Anderson:1997:PSD

[And97] James D. Anderson. Philanthropy, the state and the development of historically black public colleges: The case of Mississippi. Minerva, 35(3):295–309, September 1997. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1023/
A%3A1004277916909.

Anderson:2007:OUC

[And07] R. G. W. Anderson. Oxford University: a century of physics. Minerva, 45(2):217–221, June 2007. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
007-9032-4.

Anduaga:2022:TCP

[And22] Aitor Anduaga. Transnational co-production of knowledge: The standardisation of typhoon warning codes in the Far East, 1900– 1939. Minerva, 60(2):301–323, June 2022. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL https://

link.springer.com/article/10.1007/s11024-021-09457-2.

Anker:2007:BFC

[Ank07a] Peder Anker. Buckminster Fuller as Captain of spaceship Earth. Minerva, 45(4):417–434, December 2007. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
007-9066-7.

Anker:2007:IPA

[Ank07b] Peder Anker. Idle pondering about environmental politics. Minerva, 45(1):93–95, March 2007. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
006-9024-9.

Annan:1982:BHE

[Ann82] Lord Annan. British higher education, 1960–80: a personal retrospect. Minerva, 20(1–2):1–24, Spring–Summer 1982. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01098188.

Anonymous:1962:AORa

[Ano62a] Anonymous. Administration and organisation of research. Minerva, 1(1):117–123, Autumn 1962. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01101457.

Anonymous:1962:AORb

[Ano62b] Anonymous. Administration and organisation of research. Minerva, 1(1):123–131, Autumn 1962. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01101458.

Anonymous:1962:M

[Ano62c] Anonymous. Minerva. Minerva, 1(1):5–17, Autumn 1962. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01101449.

Anonymous:1962:Ca

[Ano62d] Anonymous. Chronicle. Minerva, 1(1):132–138, Autumn 1962. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01101459.

Anonymous:1962:Cb

[Ano62e] Anonymous. Contributors. Minerva, 1(1):138, Autumn 1962. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01101460.

Anonymous:1962:RSCb

[Ano62f] Anonymous. Recommendations of the Science Council for the development of scientific institutions in Western Germany. Minerva, 1(1):87–105, Autumn 1962. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01101455.

Anonymous:1962:RSCa

[Ano62g] Anonymous. Report of the scientific commission of Pakistan. Minerva, 1(1):74–86, Autumn 1962. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01101454.

Anonymous:1962:RPU

[Ano62h] Anonymous. Report to the president of the United States on government contracting for research and development. Minerva, 1 (1):106–116, Autumn 1962. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01101456.

Anonymous:1963:Ca

[Ano63a] Anonymous. Chronicle. Minerva, 1(2):246–254, Winter 1963. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096257.

Anonymous:1963:Cc

[Ano63b] Anonymous. Chronicle. Minerva, 1(3):392–397, Spring 1963. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02251991.

Anonymous:1963:Cd

[Ano63c] Anonymous. Chronicle. Minerva, 1(4):522–529, Summer 1963. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01107192.

Anonymous:1963:Cb

[Ano63d] Anonymous. Contributors. Minerva, 1(2):255, Winter 1963. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01096258.

Anonymous:1963:Ce

[Ano63e] Anonymous. Contributors. Minerva, 1(4):530, Summer 1963. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01107193.

Anonymous:1963:Cf

[Ano63f] Anonymous. Contributors. Minerva, 2(1):130–131, Autumn 1963. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01100785.

Anonymous:1963:EBS

[Ano63g] Anonymous. The emigration of British scientists. Minerva, 1 (3):342–357, Spring 1963. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF02251988.

Anonymous:1963:ESU

[Ano63h] Anonymous. The emigration of scientists from the United Kingdom. Minerva, 1(3):358–380, Spring 1963. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF02251989.

Anonymous:1963:MMN

[Ano63i] Anonymous. Meeting manpower needs in science and technology. Minerva, 1(3):381–391, Spring 1963. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF02251990.

Anonymous:1963:MME

[Ano63j] Anonymous. Memorandum by the Minister of Education and Social Development on higher education. Minerva, 2(1): 82–90, Autumn 1963. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01100782.

Anonymous:1963:RTF

[Ano63k] Anonymous. Report on training facilities at the technician level in South and South-East Asia. Minerva, 1(2):241–245, Winter 1963. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096256.

Anonymous:1963:RFBa

[Ano63l] Anonymous. Reports of the federation of British industries on industrial research in manufacturing industry 1959–60. Minerva, 1(2):226–231, Winter 1963. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096254.

Anonymous:1963:RFBb

[Ano63m] Anonymous. Reports of the federation of British industries on pattern of research in British industry. Minerva, 1(2): 232–240, Winter 1963. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096255.

Anonymous:1963:SGI

[Ano63n] Anonymous. Science, government and information. Minerva, 2 (1):91–117, Autumn 1963. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01100783.

Anonymous:1963:STRa

[Ano63o] Anonymous. Scientific and technological research in France: Fourth plan 1962–1965. Minerva, 1(4):470–492, Summer 1963. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01107189.

Anonymous:1963:STRb

[Ano63p] Anonymous. Scientific and technological research in France: Fourth plan 1962–1965. Minerva, 1(4):493–507, Summer 1963. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01107190.

Anonymous:1963:STRc

[Ano63q] Anonymous. Scientific and technological research in France: Fourth plan 1962–1965. Minerva, 1(4):508–521, Summer 1963. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01107191.

Anonymous:1963:SSC

[Ano63r] Anonymous. Suggestions of the science council on the pattern of new universities in Western Germany. Minerva, 1(2): 217–225, Winter 1963. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096253.

Anonymous:1963:VCC

[Ano63s] Anonymous. Vice-Chancellors’ Conference, 1961. Minerva, 1 (2):209–216, Winter 1963. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096252.

Anonymous:1964:AVC

[Ano64a] Anonymous. Address by the Vice-Chancellor, Dr. Conor Cruise O’Brien, to congregation of The University of Ghana. Minerva, 2 (4):484–491, Summer 1964. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01097540.

Anonymous:1964:Ca

[Ano64b] Anonymous. Chronicle. Minerva, 2(2):263–273, Winter 1964. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096602.

Anonymous:1964:Cd

[Ano64c] Anonymous. Chronicle. Minerva, 2(3):389–402, Spring 1964. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01097326.

Anonymous:1964:Cf

[Ano64d] Anonymous. Chronicle. Minerva, 2(4):531–547, Summer 1964. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01097543.

Anonymous:1964:Ch

[Ano64e] Anonymous. Chronicle. Minerva, 3(1):131–145, Autumn 1964. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01630156.

Anonymous:1964:Cb

[Ano64f] Anonymous. Contributors. Minerva, 2(2):274, Winter 1964. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01096603.

Anonymous:1964:Ce

[Ano64g] Anonymous. Contributors. Minerva, 2(3):403–404, Spring 1964. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01097327.

Anonymous:1964:Cg

[Ano64h] Anonymous. Contributors. Minerva, 2(4):548, Summer 1964. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01097544.

Anonymous:1964:Ci

[Ano64i] Anonymous. Contributors. Minerva, 3(1):146, Autumn 1964. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01630157.

Anonymous:1964:ISPa

[Ano64j] Anonymous. Indian scientific policy. Minerva, 2(2):243–244, Winter 1964. CODEN MINEFY. ISSN 0026-4695 (print),

1573-1871 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF01096598.

Anonymous:1964:ISPb

[Ano64k] Anonymous. Indian scientific policy. Minerva, 2(2):245–246, Winter 1964. CODEN MINEFY. ISSN 0026-4695 (print),

1573-1871 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF01096599.

Anonymous:1964:ISPc

[Ano64l] Anonymous. Indian scientific policy. Minerva, 2(2):247–256, Winter 1964. CODEN MINEFY. ISSN 0026-4695 (print),

1573-1871 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF01096600.

Anonymous:1964:ISPd

[Ano64m] Anonymous. Indian scientific policy. Minerva, 2(4):519– 530, Summer 1964. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01097542.

Anonymous:1964:RUC

[Ano64n] Anonymous. Report of the universities commission 1962. Minerva, 2(4):492–518, Summer 1964. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01097541.

Anonymous:1964:SEG

[Ano64o] Anonymous. Science, economic growth and government policy. Minerva, 2(2):232–242, Winter 1964. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01096597.

Anonymous:1964:SPIa

[Ano64p] Anonymous. Scientific policy in Italy. Minerva, 2(2):210– 224, Winter 1964. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096595.

Anonymous:1964:SPIb

[Ano64q] Anonymous. Scientific policy in Italy. Minerva, 2(2):225– 231, Winter 1964. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096596.

Anonymous:1964:SUG

[Ano64r] Anonymous. Statement to the University grants committee on certain aspects of the Robbins report. Minerva, 2(2): 257–262, Winter 1964. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096601.

Anonymous:1964:SUP

[Ano64s] Anonymous. Swedish University policy: The financial support of University students. Minerva, 3(1):91–98, Autumn 1964. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01630152.

Anonymous:1965:AVC

[Ano65a] Anonymous. Address by the Vice-Chancellor, Dr. Conor Cruise O’Brien, to congregation of the University of Ghana. Minerva, 3(3):343–355, Spring 1965. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01099960.

Anonymous:1965:BRN

[Ano65b] Anonymous. Basic research and national goals: a report to the Committee on Science and Astronautics, U.S. House of Representatives, by the National Academy of Sciences. Minerva, 3(4):499, Summer 1965. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01099413.

Anonymous:1965:BUU

[Ano65c] Anonymous. The Brazilian universities under the Castelo Branco Regime. Minerva, 3(4):555–560, Summer 1965. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01099419.

Anonymous:1965:Ca

[Ano65d] Anonymous. Chronicle. Minerva, 3(2):265–291, Winter 1965. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096340.

Anonymous:1965:Cc

[Ano65e] Anonymous. Chronicle. Minerva, 3(3):408–431, Spring 1965. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01099967.

Anonymous:1965:Cb

[Ano65f] Anonymous. Contributors. Minerva, 3(2):292, Winter 1965. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01096341.

Anonymous:1965:Cd

[Ano65g] Anonymous. Contributors. Minerva, 3(3):432–433, Spring 1965. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01099968.

Anonymous:1965:Ce

[Ano65h] Anonymous. Contributors. Minerva, 3(4):610, Summer 1965. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01099422.

Anonymous:1965:DGF

[Ano65i] Anonymous. Decisions of the government of the Federal Republic of Nigeria on the report of the National Universities Commission. Minerva, 3(2):229–232, Winter 1965. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01096336.

Anonymous:1965:ED

[Ano65j] Anonymous. Education for development. Minerva, 3(2): 233–244, Winter 1965. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096337.

Anonymous:1965:EHC

[Ano65k] Anonymous. The English–Hindi controversy. Minerva, 3(4): 560–591, Summer 1965. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01099420.

Anonymous:1965:FPD

[Ano65l] Anonymous. Five professors dismissed. Minerva, 4(1):135– 145, Autumn 1965. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01585991.

Anonymous:1965:GA

[Ano65m] Anonymous. The genovese affair. Minerva, 4(1):145–149, Autumn 1965. CODEN MINEFY. ISSN 0026-4695 (print),

1573-1871 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF01585992.

Anonymous:1965:CCa

[Ano65n] Anonymous. I conferences and commissions. Minerva, 4(1): 111–135, Autumn 1965. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01585990.

Anonymous:1965:KC

[Ano65o] Anonymous. The Knopfelmacher case. Minerva, 3(4):530– 555, Summer 1965. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01099418.

Anonymous:1965:RCM

[Ano65p] Anonymous. Report of the committee on a “model act for universities”. Minerva, 4(1):80–94, Autumn 1965. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01585986.

Anonymous:1965:RTR

[Ano65q] Anonymous. Report of the third reviewing committee of the council of scientific and industrial research. Minerva, 3(3): 356–384, Spring 1965. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01099961.

Anonymous:1965:RDU

[Ano65r] Anonymous. Report on the development of a university in Northern Rhodesia [Zambia] 1963. Minerva, 3(2):245–260, Winter 1965. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096338.

Anonymous:1965:SRE

[Ano65s] Anonymous. Scientific research and economic development in Spain. Minerva, 3(3):385–391, Spring 1965. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01099962.

Anonymous:1965:UDN

[Ano65t] Anonymous. University development in Nigeria. Minerva, 3 (2):210–228, Winter 1965. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096335.

Anonymous:1965:UKC

[Ano65u] Anonymous. University of Karachi convocation 1965 address by the Vice-Chancellor. Minerva, 4(1):73–79, Autumn 1965. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01585985.

Anonymous:1965:ULC

[Ano65v] Anonymous. The University of Lagos crisis, Nigeria. Minerva, 3 (4):592–609, Summer 1965. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01099421.

Anonymous:1966:AFU

[Ano66a] Anonymous. Academic freedom at university college, Rhodesia. Minerva, 5(1):123–148, Autumn 1966. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF02389801.

Anonymous:1966:AVC

[Ano66b] Anonymous. Address by the Vice-Chancellor, Dr. Alexander Kwapong, to congregation of the University of Ghana. Minerva, 4(4):542–554, Summer 1966. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF02207982.

Anonymous:1966:AWA

[Ano66c] Anonymous. Anti-Walloon agitation at Louvain, Belgium. Minerva, 5(1):100–102, Autumn 1966. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF02389797.

Anonymous:1966:CSA

[Ano66d] Anonymous. Continuing student agitation in Indonesia. Minerva, 5(1):116–123, Autumn 1966. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF02389800.

Anonymous:1966:C

[Ano66e] Anonymous. Contributors. Minerva, 5(1):159, Autumn 1966. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02389804.

Anonymous:1966:CRC

[Ano66f] Anonymous. The cultural revolution, China. Minerva, 5(1): 112–116, Autumn 1966. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF02389799.

Anonymous:1966:DBH

[Ano66g] Anonymous. Disturbances at the Banaras Hindu University. Minerva, 4(3):429–433, Spring 1966. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01584642.

Anonymous:1966:FPI

[Ano66h] Anonymous. France: The Pasteur Institute: Proposals for reform. Minerva, 4(3):408–419, Spring 1966. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01584639.

Anonymous:1966:FSU

[Ano66i] Anonymous. Free students’ unions in Spain. Minerva, 4(4): 597–604, Summer 1966. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF02207994.

Anonymous:1966:HEAb

[Ano66j] Anonymous. The Higher Education Act, Ceylon. Minerva, 5 (1):102–111, Autumn 1966. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF02389798.

Anonymous:1966:HEAa

[Ano66k] Anonymous. Higher education in Australia. Minerva, 4(4): 505–541, Summer 1966. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF02207981.

Anonymous:1966:CC

[Ano66l] Anonymous. I conferences and commissions. Minerva, 4(2): 286–287, Winter 1966. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01584865.

Anonymous:1966:IPW

[Ano66m] Anonymous. II politics in the world of science and learning. Minerva, 4(2):287–311, Winter 1966. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01584866.

Anonymous:1966:ISA

[Ano66n] Anonymous. India students’ agitation in Kerala. Minerva, 4 (3):424–429, Spring 1966. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01584641.

Anonymous:1966:IIU

[Ano66o] Anonymous. Indonesia the Indonesian university situation. Minerva, 4(3):433–451, Spring 1966. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01584643.

Anonymous:1966:MED

[Ano66p] Anonymous. Manifesto for the establishment of a democratic university in Spain. Minerva, 5(1):82–88, Autumn 1966. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02389794.

Anonymous:1966:MJY

[Ano66q] Anonymous. Mihajlov’s journal, Yugoslavia. Minerva, 5(1): 154–158, Autumn 1966. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF02389803.

Anonymous:1966:MFS

[Ano66r] Anonymous. The movement for free student unions, Spain. Minerva, 5(1):148–154, Autumn 1966. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF02389802.

Anonymous:1966:PWS

[Ano66s] Anonymous. Politics in the world of science and learning. Minerva, 4(4):566–568, Summer 1966. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF02207987.

Anonymous:1966:RSC

[Ano66t] Anonymous. Recommendations of the science council on the reorganisation of university teaching staffs in Western Germany. Minerva, 4(2):246–253, Winter 1966. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01584859.

Anonymous:1966:RAH

[Ano66u] Anonymous. Resistance against higher educational reform in France. Minerva, 4(4):579–584, Summer 1966. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF02207989.

Anonymous:1966:SPU

[Ano66v] Anonymous. Scientific policy in the U.S.S.R. Minerva, 4(2): 215–218, Winter 1966. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01584856.

Anonymous:1966:SOP

[Ano66w] Anonymous. Students and the one-party state. Minerva, 4(4): 595–597, Summer 1966. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF02207993.

Anonymous:1966:SRI

[Ano66x] Anonymous. Students’ role in Indonesian political changes. Minerva, 4(4):584–585, Summer 1966. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF02207990.

Anonymous:1966:SUA

[Ano66y] Anonymous. Suspension of university autonomy, Argentine. Minerva, 5(1):93–100, Autumn 1966. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF02389796.

Anonymous:1966:TRP

[Ano66z] Anonymous. Teaching, research and permanent appointment at Yale university. Minerva, 4(2):231–245, Winter 1966. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01584858.

Anonymous:1966:UAC

[Ano66-27] Anonymous. University autonomy in Ceylon. Minerva, 4(4): 568–579, Summer 1966. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF02207988.

Anonymous:1966:UCM

[Ano66-28] Anonymous. The university crisis in Mexico City. Minerva, 4 (4):588–595, Summer 1966. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF02207992.

Anonymous:1966:VUR

[Ano66-29] Anonymous. Violence at the University of Rome. Minerva, 4 (4):585–588, Summer 1966. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF02207991.

Anonymous:1967:ASK

[Ano67a] Anonymous. Abduction of South Korean students and teachers from West Germany. Minerva, 6(1):144–147, Autumn 1967. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01101619.

Anonymous:1967:AFC

[Ano67b] Anonymous. Academic freedom in the Catholic University of America. Minerva, 5(4):621–623, Summer 1967. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01096801;   http://link.springer.com/article/
10.1007/BF01096801.

Anonymous:1967:AVC

[Ano67c] Anonymous. Address by the Vice-Chancellor, Dr. K. O. Dike, to congregation of the University of Ibadan. Minerva, 5(4):

553–557, Summer 1967. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096784.

Anonymous:1967:ALN

[Ano67d] Anonymous. The aftermath of law no. 16912, Argentine. Minerva, 5(2):274–278, Winter 1967. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF02388529.

Anonymous:1967:AUS

[Ano67e] Anonymous. American university students and the war in Vietnam. Minerva, 5(3):465–469, Spring 1967. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF02388587.

Anonymous:1967:BFS

[Ano67f] Anonymous. Bus fares and student demonstrations, Mexico. Minerva, 5(2):301–307, Winter 1967. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF02388533.

Anonymous:1967:CRD

[Ano67g] Anonymous. The case of Régis Debray. Minerva, 5(4):588– 590, Summer 1967. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096792.

Anonymous:1967:CIAa

[Ano67h] Anonymous. The Central Intelligence Agency and the universities. Minerva, 5(3):469–471, Spring 1967. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF02388588.

Anonymous:1967:CIAb

[Ano67i] Anonymous. The Central Intelligence Agency and the universities II 23. Minerva, 5(4):614–620, Summer 1967. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01096799;   http://link.springer.com/article/
10.1007/BF01096799.

Anonymous:1967:CSR

[Ano67j] Anonymous. Continued student resistance in Spain. Minerva, 5 (2):308–312, Winter 1967. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF02388534.

Anonymous:1967:C

[Ano67k] Anonymous. Contributors. Minerva, 5(4):628–629, Summer 1967. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096805;   http://link.springer.com/
article/10.1007/BF01096805.

Anonymous:1967:DALb

[Ano67l] Anonymous. Dr. Adams and the London School of Economics, continued, United Kingdom. Minerva, 5(3):455–458, Spring 1967. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02388584.

Anonymous:1967:DALa

[Ano67m] Anonymous. Dr. Adams and the London School of Economics, United Kingdom. Minerva, 5(2):312–315, Winter 1967. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02388535.

Anonymous:1967:DRC

[Ano67n] Anonymous. Draft recommendations of the conference on the role of the University College, Dar Es Salaam, in a socialist Tanzania. Minerva, 5(4):558–570, Summer 1967. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01096785.

Anonymous:1967:EHM

[Ano67o] Anonymous. English, Hindi and the medium of instruction. Minerva, 6(1):123–131, Autumn 1967. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01101616.

Anonymous:1967:EFS

[Ano67p] Anonymous. Expulsion of foreign students from China and cessation of Chinese–Soviet student exchange, China. Minerva, 5 (2):282–284, Winter 1967. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF02388531.

Anonymous:1967:LP

[Ano67q] Anonymous. In loco parentis? Minerva, 5(4):593–595, Summer 1967. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096794.

Anonymous:1967:IFOa

[Ano67r] Anonymous. Increased fees for overseas students. Minerva, 5 (3):458–462, Spring 1967. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF02388585.

Anonymous:1967:IFOb

[Ano67s] Anonymous. Increased fees for overseas students II. Minerva, 5 (4):604–610, Summer 1967. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096797.

Anonymous:1967:IFOc

[Ano67t] Anonymous. Increased fees for overseas students III. Minerva, 6 (1):147–149, Autumn 1967. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01101620.

Anonymous:1967:IURa

[Ano67u] Anonymous. Indian university reform II. Minerva, 5(2): 242–264, Winter 1967. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF02388527.

Anonymous:1967:IURb

[Ano67v] Anonymous. Indian university reform III. Minerva, 5(3): 391–412, Spring 1967. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF02388579.

Anonymous:1967:IURc

[Ano67w] Anonymous. Indian university reform IV. Minerva, 6(1): 48–80, Autumn 1967. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01101606.

Anonymous:1967:IUA

[Ano67x] Anonymous. Inspecting university accounts. Minerva, 6(1): 149–152, Autumn 1967. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01101621.

Anonymous:1967:LPC

[Ano67y] Anonymous. Linguistic particularism in Croatia. Minerva, 5(4): 623–624, Summer 1967. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096802;   http://
link.springer.com/article/10.1007/BF01096802.

Anonymous:1967:LSE

[Ano67z] Anonymous. The London school of economics: Student power and indiscipline. Minerva, 5(4):610–614, Summer 1967. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096798.

Anonymous:1967:MTT

[Ano67-27] Anonymous. Mihajlov: Third trial and condemnation. Minerva, 5(4):624, Summer 1967. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096803;   http://
link.springer.com/article/10.1007/BF01096803.

Anonymous:1967:NPC

[Ano67-28] Anonymous. The new phase continues. Minerva, 5(4):601– 604, Summer 1967. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096796.

Anonymous:1967:OP

[Ano67-29] Anonymous. The ordeal of praxis. Minerva, 5(4):625– 627, Summer 1967. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096804;   http://
link.springer.com/article/10.1007/BF01096804.

Anonymous:1967:RDC

[Ano67-30] Anonymous. Rampant disorder continues. Minerva, 6(1): 131–140, Autumn 1967. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01101617.

Anonymous:1967:RDI

[Ano67-31] Anonymous. Rampant disorder, India. Minerva, 5(2):284– 301, Winter 1967. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF02388532.

Anonymous:1967:RSC

[Ano67-32] Anonymous. Recommendations of the science council for the development of scientific institutions in Western Germany. Minerva, 5(3):413–428, Spring 1967. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF02388580.

Anonymous:1967:RCU

[Ano67-33] Anonymous. Renewed crisis at the university of California, United States. Minerva, 5(3):462–465, Spring 1967. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF02388586.

Anonymous:1967:RNU

[Ano67-34] Anonymous. Riots at Negro universities. Minerva, 5(4): 620–621, Summer 1967. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096800;   http://
link.springer.com/article/10.1007/BF01096800.

Anonymous:1967:SBS

[Ano67-35] Anonymous. The struggle broadens, Spain. Minerva, 5(3): 447–455, Spring 1967. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF02388583.

Anonymous:1967:SMF

[Ano67-36] Anonymous. Student movements in France. Minerva, 5(4): 590–593, Summer 1967. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096793.

Anonymous:1967:SPP

[Ano67-37] Anonymous. Students, politicians, professors and officers. Minerva, 5(4):595–601, Summer 1967. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01096795.

Anonymous:1967:SSI

[Ano67-38] Anonymous. Students’ strike, Israel. Minerva, 5(3):435–447, Spring 1967. CODEN MINEFY. ISSN 0026-4695 (print),

1573-1871 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF02388582.

Anonymous:1967:UGA

[Ano67-39] Anonymous. University of Ghana: Address by the Vice-Chancellor, Dr. Alexander Kwapong, to congregation. Minerva, 6(1):87–98, Autumn 1967. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01101608.

Anonymous:1967:UIA

[Ano67-40] Anonymous. University of Ibadan: an address by John Harris, acting Vice-Chancellor, given in Trenchard Hall on graduation day. Minerva, 6(1):81–86, Autumn 1967. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01101607.

Anonymous:1967:URC

[Ano67-41] Anonymous. University resistance to Castelo Branco, Brazil. Minerva, 5(2):278–282, Winter 1967. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF02388530.

Anonymous:1967:VCO

[Ano67-42] Anonymous. The vice-chancellorship of the osmania University. Minerva, 6(1):140–144, Autumn 1967. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01101618.

Anonymous:1968:AD

[Ano68a] Anonymous. The aftermath of dziady. Minerva, 6(4):759– 772, Summer 1968. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096562.

Anonymous:1968:CRU

[Ano68b] Anonymous. Civil rights on university campuses. Minerva, 6(3):474, Spring 1968. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096437.

Anonymous:1968:CGU

[Ano68c] Anonymous. The “cleansing” of the Greek universities. Minerva, 6(4):577–597, Summer 1968. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01096550.

Anonymous:1968:Ca

[Ano68d] Anonymous. Contributors. Minerva, 6(2):317, Winter 1968. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01097273.

Anonymous:1968:Cb

[Ano68e] Anonymous. Contributors. Minerva, 6(3):479, Spring 1968. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01096439.

Anonymous:1968:Cc

[Ano68f] Anonymous. Contributors. Minerva, 6(4):794, Summer 1968. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096565;   http://link.springer.com/
article/10.1007/BF01096565.

Anonymous:1968:DRA

[Ano68g] Anonymous. Deepening radicalism: The aftermath of the Shah’s visit and the death of Benno Ohnesorg. Minerva, 6 (2):283–287, Winter 1968. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01097268.

Anonymous:1968:EHM

[Ano68h] Anonymous. English, Hindi and the medium of instruction II. Minerva, 6(2):287–305, Winter 1968. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01097269.

Anonymous:1968:EUC

[Ano68i] Anonymous. European universities a century ago. Minerva, 6 (4):561–576, Summer 1968. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096549.

Anonymous:1968:FH

[Ano68j] Anonymous. A flicker of hope. Minerva, 6(4):618–629, Summer 1968. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096558.

Anonymous:1968:FSU

[Ano68k] Anonymous. Free student unions: Adamant demands, adamant refusals. Minerva, 6(3):463–474, Spring 1968. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01096436.

Anonymous:1968:HCA

[Ano68l] Anonymous. From housing conditions to academic freedom: The movement at Prague University. Minerva, 6(3):440–441, Spring 1968. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096432.

Anonymous:1968:FDH

[Ano68m] Anonymous. A further digging-in of heels. Minerva, 6(4): 772–788, Summer 1968. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096563.

Anonymous:1968:GPG

[Ano68n] Anonymous. Generals and professors in Greece. Minerva, 6 (3):457–463, Spring 1968. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096435.

Anonymous:1968:KC

[Ano68o] Anonymous. The Krippendorff case. Minerva, 6(2):272–281, Winter 1968. CODEN MINEFY. ISSN 0026-4695 (print),

1573-1871 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF01097266.

Anonymous:1968:OCU

[Ano68p] Anonymous. The occupation of Columbia University. Minerva, 6 (4):788–793, Summer 1968. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096564;   http://
link.springer.com/article/10.1007/BF01096564.

Anonymous:1968:PRP

[Ano68q] Anonymous. A phantasm of revolution and the possibility of reform. Minerva, 6(4):630–689, Summer 1968. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01096559.

Anonymous:1968:PDLb

[Ano68r] Anonymous. Prelude to disorder: Late opening and overcrowding. Minerva, 6(3):441–448, Spring 1968. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01096433.

Anonymous:1968:PDLa

[Ano68s] Anonymous. Progress in the disintegration of Louvain University. Minerva, 6(3):436–440, Spring 1968. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01096431.

Anonymous:1968:RFI

[Ano68t] Anonymous. Rebels and factions in Italian universities. Minerva, 6(4):740–759, Summer 1968. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01096561.

Anonymous:1968:RRR

[Ano68u] Anonymous. Relentless revolutionaries: Reluctant reformers. Minerva, 6(4):690–740, Summer 1968. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01096560.

Anonymous:1968:RNU

[Ano68v] Anonymous. Riots at Negro universities II. Minerva, 6(2): 314–316, Winter 1968. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01097272.

Anonymous:1968:S

[Ano68w] Anonymous. Skirmishes. Minerva, 6(2):281–283, Winter 1968. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01097267.

Anonymous:1968:SUGa

[Ano68x] Anonymous. Students and universities in Great Britain. Minerva, 6(4):559–560, Summer 1968. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01096548.

Anonymous:1968:SAP

[Ano68y] Anonymous. Students as an anti-parliamentary opposition in West Germany. Minerva, 6(3):448–457, Spring 1968. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01096434.

Anonymous:1968:SUW

[Ano68z] Anonymous. Students’ unions and workers’ commissions. Minerva, 6(2):305–309, Winter 1968. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01097270.

Anonymous:1968:UP

[Ano68-27] Anonymous. Universities and productivity. Minerva, 6(3): 432–436, Spring 1968. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096430.

Anonymous:1968:UWVa

[Ano68-28] Anonymous. The universities and the war in Vietnam. Minerva, 6(3):474–478, Spring 1968. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096438.

Anonymous:1968:VWA

[Ano68-29] Anonymous. The Vietnamese war on the American campus. Minerva, 6(2):309–314, Winter 1968. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01097271.

Anonymous:1968:ACD

[Ano69a] Anonymous. The anniversary of the Córdoba declaration and its ramifications. Minerva, 7(1–2):95–112, Autumn–Winter 1968– 1969. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096816.

Anonymous:1968:ASC

[Ano69b] Anonymous. Art students in confrontation. Minerva, 7(1–2): 300–305, Autumn–Winter 1968–1969. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01096826;
http://link.springer.com/article/10.1007/BF01096826.

Anonymous:1968:CRA

[Ano69c] Anonymous. Civic responsibility and academic freedom in Africa. Minerva, 7(1–2):73–81, Autumn–Winter 1968–1969. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01096811.

Anonymous:1968:COO

[Ano69d] Anonymous. A conflagration of obscure origin in Mexico City. Minerva, 7(1–2):256–268, Autumn–Winter 1968– 1969. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096821.

Anonymous:1968:Cd

[Ano69e] Anonymous. Contributors. Minerva, 7(1–2):342, Autumn– Winter 1968–1969. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096834;   http://
link.springer.com/article/10.1007/BF01096834.

Anonymous:1968:DCU

[Ano69f] Anonymous. The disruption of Columbia University. Minerva, 7(1–2):310–319, Autumn–Winter 1968–1969. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01096828;   http://link.springer.com/article/
10.1007/BF01096828.

Anonymous:1968:DAN

[Ano69g] Anonymous. Disturbances by and about Negro students. Minerva, 7(1–2):319–325, Autumn–Winter 1968–1969. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01096829;   http://link.springer.com/article/
10.1007/BF01096829.

Anonymous:1968:FWR

[Ano69h] Anonymous. Flood waters reach high ground. Minerva, 7(1–2): 268–279, Autumn–Winter 1968–1969. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01096822.

Anonymous:1968:FST

[Ano69i] Anonymous. Further steps towards the displacement of English. Minerva, 7(1–2):178–234, Autumn–Winter 1968– 1969. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096819.

Anonymous:1968:IFR

[Ano69j] Anonymous. Irreconcilables and fumbling reformers. Minerva, 7(1–2):153–178, Autumn–Winter 1968–1969. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01096818.

Anonymous:1968:MMU

[Ano69k] Anonymous. Militants and moderates in the universities. Minerva, 7(1–2):284–300, Autumn–Winter 1968–1969. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01096825.

Anonymous:1968:MT

[Ano69l] Anonymous. A miscellany of tribulations. Minerva, 7(1–2): 235–256, Autumn–Winter 1968–1969. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01096820.

Anonymous:1968:MAE

[Ano69m] Anonymous. The morning after: Evaporation of enthusiasm and the beginning of reform. Minerva, 7(1–2):113–153, Autumn– Winter 1968–1969. CODEN MINEFY. ISSN 0026-4695 (print),

1573-1871 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF01096817.

Anonymous:1968:NUS

[Ano69n] Anonymous. National union of students in action and in conference. Minerva, 7(1–2):279–282, Autumn–Winter 1968– 1969. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096823.

Anonymous:1968:OET

[Ano69o] Anonymous. Odds and ends of turbulence. Minerva, 7(1–2): 328–333, Autumn–Winter 1968–1969. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01096831;
http://link.springer.com/article/10.1007/BF01096831.

Anonymous:1968:RAS

[Ano69p] Anonymous. Responses to aggressive student actions. Minerva, 7(1–2):305–310, Autumn–Winter 1968–1969. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01096827;   http://link.springer.com/article/
10.1007/BF01096827.

Anonymous:1968:SDY

[Ano69q] Anonymous. Student demands in Yugoslavia: Pop concert tickets and socialist ideals. Minerva, 7(1–2):336–341, Autumn– Winter 1968–1969. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096833;   http://
link.springer.com/article/10.1007/BF01096833.

Anonymous:1968:SPL

[Ano69r] Anonymous. Student power in Latin America: The Córdoba Manifesto. Minerva, 7(1–2):82–87, Autumn–Winter 1968– 1969. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096812.

Anonymous:1968:SUGb

[Ano69s] Anonymous. Students and universities in Great Britain. II. Minerva, 7(1–2):67–72, Autumn–Winter 1968–1969. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01096810.

Anonymous:1968:TFU

[Ano69t] Anonymous. Transport fares, university students and government policies. Minerva, 7(1–2):333–336, Autumn–Winter 1968– 1969. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096832;   http://link.springer.com/
article/10.1007/BF01096832.

Anonymous:1968:UWVb

[Ano69u] Anonymous. The universities and the war in Vietnam. II. Minerva, 7(1–2):325–328, Autumn–Winter 1968–1969. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01096830;   http://link.springer.com/article/
10.1007/BF01096830.

Anonymous:1968:VRD

[Ano69v] Anonymous. Vietnam and related demonstrations. Minerva, 7(1–2):282–284, Autumn–Winter 1968–1969. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01096824.

Anonymous:1969:BDC

[Ano69w] Anonymous. A belated and disregarded concession. Minerva, 7 (3):579–584, Spring 1969. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01103282.

Anonymous:1969:BTS

[Ano69x] Anonymous. A brief triumph for student power in Pakistan. Minerva, 7(4):783–812, Summer 1969. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01099551.

Anonymous:1969:CDJ

[Ano69y] Anonymous. Continuing disorder in Japanese universities. Minerva, 7(4):778–783, Summer 1969. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01099550.

Anonymous:1969:C

[Ano69z] Anonymous. Contributors. Minerva, 7(3):611, Spring 1969. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01103288;   http://link.springer.com/article/
10.1007/BF01103288.

Anonymous:1969:DCC

[Ano69-27] Anonymous. Derangement from coast to coast in the United States. Minerva, 7(4):822–863, Summer 1969. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01099553.

Anonymous:1969:ERI

[Ano69-28] Anonymous. Examinations, reforms and irreconciliables. Minerva, 7(3):507–527, Spring 1969. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01103275.

Anonymous:1969:FAL

[Ano69-29] Anonymous. Further agitation at the London school of economics. Minerva, 7(4):812–822, Summer 1969. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01099552.

Anonymous:1969:HMT

[Ano69-30] Anonymous. Halting movements towards university reform. Minerva, 7(3):527–533, Spring 1969. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01103276.

Anonymous:1969:SPH

[Ano69-31] Anonymous. I statement of principles of the higher education bill. Minerva, 7(4):706–712, Summer 1969. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01099542.

Anonymous:1969:IRC

[Ano69-32] Anonymous. Itches, rashes and carbuncles on the academic body. Minerva, 7(3):587–590, Spring 1969. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01103284.

Anonymous:1969:MSV

[Ano69-33] Anonymous. A major student victory. Minerva, 7(3):569– 579, Spring 1969. CODEN MINEFY. ISSN 0026-4695 (print),

1573-1871 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF01103281.

Anonymous:1969:MFC

[Ano69-34] Anonymous. The movement finds a cause. Minerva, 7(3): 563–569, Spring 1969. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01103280.

Anonymous:1969:MEP

[Ano69-35] Anonymous. Mr. Enoch Powell as a student problem. Minerva, 7(3):590–591, Spring 1969. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01103285.

Anonymous:1969:OLS

[Ano69-36] Anonymous. The occupation of the London school of economics. Minerva, 7(3):584–587, Spring 1969. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01103283.

Anonymous:1969:PC

[Ano69-37] Anonymous. Perpetual commotion. Minerva, 7(3):545–561, Spring 1969. CODEN MINEFY. ISSN 0026-4695 (print),

1573-1871 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF01103278.

Anonymous:1969:PNE

[Ano69-38] Anonymous. Proposals for a new educational policy in Pakistan. Minerva, 7(4):728–750, Summer 1969. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01099544.

Anonymous:1969:RDN

[Ano69-39] Anonymous. Rancorous disorder on a national scale. Minerva, 7(3):595–610, Spring 1969. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01103287;   http://
link.springer.com/article/10.1007/BF01103287.

Anonymous:1969:RFA

[Ano69-40] Anonymous. Rebellion as a form of art for art’s sake. Minerva, 7(3):591–595, Spring 1969. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01103286;   http://
link.springer.com/article/10.1007/BF01103286.

Anonymous:1969:RNR

[Ano69-41] Anonymous. Recommendations for a Norwegian research policy. Minerva, 7(3):465–489, Spring 1969. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01103271.

Anonymous:1969:RHN

[Ano69-42] Anonymous. Roman holidays on a national scale. Minerva, 7 (3):561–563, Spring 1969. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01103279.

Anonymous:1969:SAU

[Ano69-43] Anonymous. The Siren affair and the University of Ghana. Minerva, 7(4):762–778, Summer 1969. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01099549.

Anonymous:1969:SSD

[Ano69-44] Anonymous. A steady state of disorder: The SDS at work. Minerva, 7(3):533–545, Spring 1969. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01103277.

Anonymous:1969:UCF

[Ano69-45] Anonymous. University of Cambridge: First report on participation by junior members in the educational business of the university. Minerva, 7(3):454–464, Spring 1969. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01103270.

Anonymous:1969:VRA

[Ano69-46] Anonymous. A valiant rearguard action. Minerva, 7(3): 499–507, Spring 1969. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01103274.

Anonymous:1970:AP

[Ano70a] Anonymous. The academic profession. Minerva, 8(1–4): 90–99, January 1970. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01553200.

Anonymous:1970:Ca

[Ano70b] Anonymous. Contributors. Minerva, 8(1–4):151, January 1970. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01553208.

Anonymous:1970:Cb

[Ano70c] Anonymous. Contributors. Minerva, 8(1–4):316, January 1970. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01553220;   http://link.springer.com/
article/10.1007/BF01553220.

Anonymous:1970:Cc

[Ano70d] Anonymous. Contributors. Minerva, 8(1–4):474, January 1970. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01553234;   http://link.springer.com/
article/10.1007/BF01553234.

Anonymous:1970:CH

[Ano70e] Anonymous. The crisis at Harvard. Minerva, 8(1–4):136– 143, January 1970. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01553204.

Anonymous:1970:FST

[Ano70f] Anonymous. The future of scientific and technical research in France. Minerva, 8(1–4):412–427, January 1970. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01553226;   http://link.springer.com/article/
10.1007/BF01553226.

Anonymous:1970:HESa

[Ano70g] Anonymous. Higher education in Spain: Part I. Minerva, 8(1– 4):268–283, January 1970. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01553215.

Anonymous:1970:HESb

[Ano70h] Anonymous. Higher education in Spain: Part II. Minerva, 8(1– 4):428–439, January 1970. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01553227;   http://
link.springer.com/article/10.1007/BF01553227.

Anonymous:1970:HCU

[Ano70i] Anonymous. The hole in the centre: University government in the United States. Minerva, 8(1–4):1–7, January 1970. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01553196.

Anonymous:1970:PIF

[Ano70j] Anonymous. Political involvement, financial management and the care of learning in Sierra Leone. Minerva, 8(1–4):284– 294, January 1970. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01553216;   http://
link.springer.com/article/10.1007/BF01553216.

Anonymous:1970:SRU

[Ano70k]  Anonymous.  Student rights under the Greek military

regime. Minerva, 8(1–4):111–115, January 1970. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01553202.

Anonymous:1970:URG

[Ano70l] Anonymous. University reform in Germany. Minerva, 8(1–4): 242–267, January 1970. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01553214.

Anonymous:1970:URJ

[Ano70m] Anonymous. University reform in Japan. Minerva, 8(1–4): 581–593, January 1970. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01553240;   http://
link.springer.com/article/10.1007/BF01553240.

Anonymous:1971:Ca

[Ano71a] Anonymous. Contributors. Minerva, 9(1):159–160, January 1971. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01553195.

Anonymous:1971:Cb

[Ano71b] Anonymous. Contributors. Minerva, 9(3):438, July 1971. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01553370.

Anonymous:1971:CAA

[Ano71c] Anonymous. The criteria of academic appointment. Minerva, 9(2):272–290, April 1971. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01553160.

Anonymous:1971:CLS

[Ano71d] Anonymous. The cultivation of learning and the service of society in West Africa. Minerva, 9(1):122–128, January 1971. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01553190.

Anonymous:1971:PML

[Ano71e] Anonymous. Of pride and men of little faith. Minerva, 9 (1):1–6, January 1971. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01553182.

Anonymous:1971:UHE

[Ano71f] Anonymous. On universal higher education. Minerva, 9(2): 256–271, April 1971. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01553159.

Anonymous:1971:PAS

[Ano71g] Anonymous. A prophetic achievement in science policy. Minerva, 9(4):528–538, October 1971. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01558023.

Anonymous:1971:SUC

[Ano71h] Anonymous. Science in underdeveloped countries. Minerva, 9 (1):101–121, January 1971. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01553189.

Anonymous:1972:Ca

[Ano72a] Anonymous. Contributors. Minerva, 10(2):348, April 1972. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01682428.

Anonymous:1972:Cb

[Ano72b] Anonymous. Contributors. Minerva, 10(4):639, October 1972. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01695910.

Anonymous:1972:ODH

[Ano72c] Anonymous. Oxford discussions on higher education. Minerva, 10(2):295–318, April 1972. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01682423.

Anonymous:1973:NCS

[Ano73a] Anonymous. I nationalism in Canadian social science. Minerva, 11(1):113–120, January 1973. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01544981.

Anonymous:1973:NIP

[Ano73b] Anonymous. A new Indian programme: Science and technology for development. Minerva, 11(4):537–570, October 1973. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01557801.

Anonymous:1973:PSD

[Ano73c] Anonymous. The power of the state and the dignity of the academic calling in imperial Germany. Minerva, 11(4):571– 632, October 1973. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01557802.

Anonymous:1973:RPS

[Ano73d] Anonymous. The redemptive power of science. Minerva, 11 (1):1–5, January 1973. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01544974.

Anonymous:1973:RJH

[Ano73e] Anonymous. The reform of Japanese higher education. Minerva, 11(3):387–414, July 1973. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01557691.

Anonymous:1974:C

[Ano74a] Anonymous. Contributors. Minerva, 12(3):401, July 1974. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01102532.

Anonymous:1974:ACD

[Ano74b] Anonymous. I the academic consequences of disorder in the German universities. Minerva, 12(4):510–514, October 1974. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01557642.

Anonymous:1975:CAA

[Ano75a] Anonymous. The confidentiality and anonymity of assessment. Minerva, 13(2):135–151, Summer 1975. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01097792.

Anonymous:1975:Ca

[Ano75b] Anonymous. Contributors. Minerva, 13(1):131, Spring 1975. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096246.

Anonymous:1975:Cb

[Ano75c] Anonymous. Contributors. Minerva, 13(2):337, Summer 1975. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01097802.

Anonymous:1975:Cc

[Ano75d] Anonymous. Contributors. Minerva, 13(3):497, Autumn 1975. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01097851.

Anonymous:1975:Cd

[Ano75e] Anonymous. Contributors. Minerva, 13(4):645, Winter 1975. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096181.

Anonymous:1975:RSA

[Ano75f] Anonymous. The right to speak in American universities: The University of Chicago. Minerva, 13(2):303–305, Summer 1975. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01097799.

Anonymous:1975:USS

[Ano75g] Anonymous. The universities in a steady state: The prospect from Cambridge. Minerva, 13(2):270–296, Summer 1975. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01097797.

Anonymous:1976:Ca

[Ano76a] Anonymous. Contributors. Minerva, 14(1):166, Spring 1976. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096218.

Anonymous:1976:Cb

[Ano76b] Anonymous. Contributors. Minerva, 14(2):275, Summer 1976. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01107890.

Anonymous:1976:Cc

[Ano76c] Anonymous. Contributors. Minerva, 14(3):404, Autumn 1976. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096280.

Anonymous:1976:CAA

[Ano76d] Anonymous. The criteria of academic appointment in American universities and colleges. Minerva, 14(1):97–117, Spring 1976. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096216.

Anonymous:1976:Cd

[Ano77a] Anonymous. Contributors. Minerva, 14(4):586, Winter 1976– 1977. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01097641.

Anonymous:1977:Ca

[Ano77b] Anonymous. Contributors. Minerva, 15(1):115, Spring 1977. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096323.

Anonymous:1977:Cb

[Ano77c] Anonymous. Contributors. Minerva, 15(2):270, Summer 1977. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096680.

Anonymous:1977:Cc

[Ano77d] Anonymous. Contributors. Minerva, 15(3–4):562, Autumn– Winter 1977. CODEN MINEFY. ISSN 0026-4695 (print),

1573-1871 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF01096732.

Anonymous:1977:DUB

[Ano77e] Anonymous. The Danish university between the millstones. Minerva, 15(3–4):377–386, Autumn–Winter 1977. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01096729.

Anonymous:1978:Ca

[Ano78a] Anonymous. Contributors. Minerva, 16(1):156, Spring 1978. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01102187.

Anonymous:1978:Cb

[Ano78b] Anonymous. Contributors. Minerva, 16(2):354, Summer 1978. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096020.

Anonymous:1978:Cc

[Ano78c] Anonymous. Contributors. Minerva, 16(3):461, Autumn 1978. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096672.

Anonymous:1978:Cd

[Ano78d] Anonymous. Contributors. Minerva, 16(4):606, Winter 1978. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01100336.

Anonymous:1978:RGU

[Ano78e] Anonymous. Report on the German universities. Minerva, 16 (1):103–138, Spring 1978. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01102183.

Anonymous:1978:SFH

[Ano78f] Anonymous. Scholarly freedom and human rights. Minerva, 16 (2):283–326, Summer 1978. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096018.

Anonymous:1979:Ca

[Ano79a] Anonymous. Contributors. Minerva, 17(1):193, Spring 1979. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096265.

Anonymous:1979:Cb

[Ano79b] Anonymous. Contributors. Minerva, 17(2):328, Summer 1979. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096287.

Anonymous:1979:Cc

[Ano79c] Anonymous. Contributors. Minerva, 17(3):480, Autumn 1979. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096226.

Anonymous:1979:Cd

[Ano79d] Anonymous. Contributors. Minerva, 17(4):574, Winter 1979. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01101905.

Anonymous:1980:AFT

[Ano80a] Anonymous. Academic freedom then and now: The dismissal of Leo Arons from the University of Berlin. Minerva, 18(3): 499–520, Autumn 1980. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096954.

Anonymous:1980:Ca

[Ano80b] Anonymous. Contributors. Minerva, 18(1):196, Spring 1980. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096665.

Anonymous:1980:Cb

[Ano80c] Anonymous. Contributors. Minerva, 18(2):364, Summer 1980. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01100254.

Anonymous:1980:Cc

[Ano80d] Anonymous. Contributors. Minerva, 18(3):538, Autumn 1980. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096957.

Anonymous:1980:Cd

[Ano80e] Anonymous. Contributors. Minerva, 18(4):672, Winter 1980. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096130.

Anonymous:1981:Ca

[Ano81a] Anonymous. Contributors. Minerva, 19(2):340, Summer 1981. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096572.

Anonymous:1981:Cb

[Ano81b] Anonymous. Contributors. Minerva, 19(4):652, Winter 1981. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096198.

Anonymous:1981:DUM

[Ano81c] Anonymous. The destruction of the University of Malta. Minerva, 19(1):123–164, Spring 1981. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF02192551.

Anonymous:1981:ETS

[Ano81d] Anonymous. The education of talented students. Minerva, 19 (3):480–497, Autumn 1981. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF02192826.

Anonymous:1982:BR

[Ano82a] Anonymous. Book reviews. Minerva, 20(1–2):230–250, Spring– Summer 1982. CODEN MINEFY. ISSN 0026-4695 (print),

1573-1871 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF01098197.

Anonymous:1982:C

[Ano82b] Anonymous. Contributors. Minerva, 20(1–2):251, Spring– Summer 1982. CODEN MINEFY. ISSN 0026-4695 (print),

1573-1871 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF01098198.

Anonymous:1982:OUTa

[Ano82c] Anonymous. I the obligations of university teachers. Minerva, 20 (1–2):105–208, Spring–Summer 1982. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01098192.

Anonymous:1982:OUTb

[Ano82d] Anonymous. I the obligations of university teachers. Minerva, 20 (1–2):209–212, Spring–Summer 1982. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01098193.

Anonymous:1982:IDW

[Ano82e] Anonymous. II. do we lack pioneering research? An interview with Professor Klaus Weissermel. Minerva, 20(1–2):226– 229, Spring–Summer 1982. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01098196.

Anonymous:1983:Ca

[Ano83a] Anonymous. Contributors. Minerva, 21(1):155, Spring 1983. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01099508.

Anonymous:1983:Cb

[Ano83b] Anonymous. Contributors. Minerva, 21(2–3):348, Summer– Autumn 1983. CODEN MINEFY. ISSN 0026-4695 (print),

1573-1871 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF01097971.

Anonymous:1983:Cc

[Ano83c] Anonymous. Contributors. Minerva, 21(4):477, Winter 1983. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096529.

Anonymous:1983:CAA

[Ano83d] Anonymous. On the criteria of academic appointment. Minerva, 21(4):410–414, Winter 1983. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096525.

Anonymous:1983:RAH

[Ano83e] Anonymous. Report of the ad hoc University committee on appointments, renewal and tenure at Vanderbilt University. Minerva, 21(4):415–443, Winter 1983. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01096526.

Anonymous:1983:RSC

[Ano83f] Anonymous. Report of the special Committee to Study the Appointments Process at Harvard University. Minerva, 21(4): 444–455, Winter 1983. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096527.

Anonymous:1984:Ca

[Ano84a] Anonymous. Contributors. Minerva, 22(1):128, Spring 1984. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02207560.

Anonymous:1984:Cb

[Ano84b] Anonymous. Contributors. Minerva, 22(2):295, Summer 1984. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02207953.

Anonymous:1984:NLH

[Ano84c] Anonymous. A new law for higher educational institutions in Greece. Minerva, 22(2):245–275, Summer 1984. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF02207951.

Anonymous:1985:Ca

[Ano85a] Anonymous. Contributors. Minerva, 23(1):173, Spring 1985. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01097844.

Anonymous:1985:Cb

[Ano85b] Anonymous. Contributors. Minerva, 23(2):319, Summer 1985. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01099949.

Anonymous:1985:Cc

[Ano85c] Anonymous. Contributors. Minerva, 23(3):441, Autumn 1985. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096448.

Anonymous:1985:IPD

[Ano85d] Anonymous. II policy for dealing with faculty fraud in research. Minerva, 23(2):305–308, Summer 1985. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01099947.

Anonymous:1985:MR

[Ano85e] Anonymous. Misconduct in research. Minerva, 23(3):423– 432, Autumn 1985. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096446.

Anonymous:1986:Ca

[Ano86a] Anonymous. Contributors. Minerva, 24(2–3):390, Summer– Autumn 1986. CODEN MINEFY. ISSN 0026-4695 (print),

1573-1871 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF01096618.

Anonymous:1986:Cb

[Ano86b] Anonymous. Contributors. Minerva, 24(4):510, Winter 1986. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096711.

Anonymous:1986:PAU

[Ano86c] Anonymous. I the policy of American universities towards divestment in south Africa. Minerva, 24(2–3):246–343, Summer– Autumn 1986. CODEN MINEFY. ISSN 0026-4695 (print),

1573-1871 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF01096614.

Anonymous:1986:RCA

[Ano86d] Anonymous. Report of the committee on academic fraud the University of Chicago. Minerva, 24(2–3):347–358, Summer– Autumn 1986. CODEN MINEFY. ISSN 0026-4695 (print),

1573-1871 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF01096616.

Anonymous:1987:C

[Ano87a] Anonymous. Contributors. Minerva, 25(3):208, Autumn 1987. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01097791.

Anonymous:1987:MSa

[Ano87b] Anonymous. The morality of scientists. Minerva, 25(4): 502–503, Winter 1987. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01099445.

Anonymous:1987:MSb

[Ano87c] Anonymous. The morality of scientists. Minerva, 25(4): 504–512, Winter 1987. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01099446.

Anonymous:1988:AAR

[Ano88a] Anonymous. Affirmative action reaffirmed. Minerva, 26(4): 600–610, Winter 1988. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096502.

Anonymous:1988:ASA

[Ano88b] Anonymous. The attrition of south Africa’s universities under siege: Positive aid a ray of hope. Minerva, 26(4):579–587, Winter 1988. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096499.

Anonymous:1989:WAC

[Ano89] Anonymous. The world archaeological Congress and the south African archaeologists. Minerva, 27(1):87–125, Spring 1989. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096716.

Anonymous:1990:FED

[Ano90a] Anonymous. Freedom of expression and disruptions at meetings of student societies in university buildings. Minerva, 28 (1):91–96, Spring 1990. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096328.

Anonymous:1990:IPS

[Ano90b] Anonymous. II: President’s statement on the “Recruitment and Retention of Minority Group Members on the Faculty at Yale”. Minerva, 28(2):242–247, Summer 1990. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF02219660.

Anonymous:1990:PAA

[Ano90c] Anonymous. The progress of “affirmative action”: Yale declares itself. Minerva, 28(2):217–220, Summer 1990. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF02219658.

Anonymous:1990:UWTa

[Ano90d] Anonymous. The university world turned upside down: does confidentiality of assessment by peers guarantee the quality of academic appointment? Minerva, 28(3):324–385, Autumn 1990. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096294.

Anonymous:1990:UWTb

[Ano90e] Anonymous. The university world turned upside down: Does confidentiality of assessment by peers guarantee the quality of academic appointment? II. Minerva, 28(4):469–539, Winter 1990. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096652.

Anonymous:1991:MRP

[Ano91] Anonymous. Misconduct in research: The policy of the University of Delaware. Minerva, 29(2):226–230, Summer 1991. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096409.

Anonymous:1992:OSN

[Ano92] Anonymous. Old strains and new initiatives in the universities of The Federal German Republic. Minerva, 30(3):430– 431, Autumn 1992. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01097647.

Anonymous:1993:RAC

[Ano93a] Anonymous. Rockefeller Archive Center grants for travel and research. Minerva, 31(2):260, Summer 1993. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01125438.

Anonymous:1993:TU

[Ano93b] Anonymous. The teaching of undergraduates. Minerva, 31(2): 228–241, Summer 1993. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01125436.

Anonymous:1993:TUP

[Ano93c] Anonymous. The teaching of undergraduates: Part II. Minerva, 31(3):361–374, Autumn 1993. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01098627.

Anonymous:1994:UBT

[Ano94] Anonymous. The universities between their internal and external enemies: Thoughts on Professor Conrad Russell’s academic freedom. Minerva, 32(2):186–187, Summer 1994. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01103357.

Anonymous:1995:Ca

[Ano95a] Anonymous. Contributors. Minerva, 33(2):209, Summer 1995. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01097913.

Anonymous:1995:Cb

[Ano95b] Anonymous. Contributors. Minerva, 33(3):303, Autumn 1995. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01117506.

Anonymous:1995:Cc

[Ano95c] Anonymous. Contributors. Minerva, 33(4):406, Winter 1995. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096521.

Anonymous:1996:BRa

[Ano96a] Anonymous. Book reviews. Minerva, 34(2):189–217, ???? 1996. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00122901.

Anonymous:1996:BRb

[Ano96b] Anonymous. Book reviews. Minerva, 34(3):309–318, ???? 1996. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00120330.

Anonymous:1996:BRc

[Ano96c] Anonymous. Book reviews. Minerva, 34(4):393–401, ???? 1996. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00127074.

Anonymous:1996:Ca

[Ano96d] Anonymous. Contributors. Minerva, 34(1):129, ???? 1996. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF00124207.

Anonymous:1996:Cb

[Ano96e] Anonymous. Contributors. Minerva, 34(2):218, ???? 1996. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF00122902.

Anonymous:1996:Cc

[Ano96f] Anonymous. Contributors. Minerva, 34(3):319, ???? 1996. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF00120331.

Anonymous:1996:Cd

[Ano96g] Anonymous. Contributors. Minerva, 34(4):402, ???? 1996. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF00127075.

Anonymous:1996:ES

[Ano96h] Anonymous. Edward Shils (1910–1995). Minerva, 34(1): 85–93, ???? 1996. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00124203.

Anonymous:1996:ESC

[Ano96i] Anonymous. Edward Shils’ contributions to Minerva. Minerva, 34(1):125–127, ???? 1996. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00124206.

Anonymous:1996:F

[Ano96j] Anonymous. Foreword. Minerva, 34(1):1–5, ???? 1996. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF00124196.

Anonymous:1996:MDP

[Ano96k] Anonymous. Managing decline and preserving autonomy in three university systems. Minerva, 34(2):177–187, ???? 1996. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00122900.

Anonymous:1996:MFE

[Ano96l] Anonymous. My friend edward. Minerva, 34(1):103–123, ???? 1996. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00124205.

Anonymous:1997:CIV

[Ano97] Anonymous. Contributors & index to volume XXXV. Minerva, 35(4):400–404, December 1997. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1023/A%3A1017128500736.

Anonymous:1998:C

[Ano98a] Anonymous. Contributors. Minerva, 36(3):1–2, September 1998. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1017177614249.

Anonymous:1998:ETT

[Ano98b] Anonymous. The endless transition: a “triple helix” of university–industry–government relations. Minerva, 36(3):203– 208, September 1998. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1017159001649.

Anonymous:1999:Ca

[Ano99a] Anonymous. Contributors. Minerva, 37(1):102, March 1999. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1017259215813.

Anonymous:1999:Cb

[Ano99b] Anonymous. Contributors. Minerva, 37(2):199, June 1999. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1023/A%3A1017163232101.

Anonymous:1999:Cc

[Ano99c] Anonymous. Contributors. Minerva, 37(3):300, September 1999. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1017223424274.

Anonymous:2000:C

[Ano00a] Anonymous. Contributors. Minerva, 37(4):416, December 2000. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1017148201150.

Anonymous:2000:IVX

[Ano00b] Anonymous. Index to volume XXXVII. Minerva, 37(4):417– 420, December 2000. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1017156517988.

Anonymous:2000:IA

[Ano00c] Anonymous. Instructions for authors. Minerva, 38(1):121– 127, ???? 2000. CODEN MINEFY. ISSN 0026-4695 (print),

1573-1871 (electronic). URL http://link.springer.com/
accesspage/article/10.1023/A%3A1026440025364.

Anonymous:2001:CV

[Ano01a] Anonymous. Contents of volume 39. Minerva, 39(4):466– 467, ???? 2001. CODEN MINEFY. ISSN 0026-4695 (print),

1573-1871 (electronic). URL http://link.springer.com/
accesspage/article/10.1023/A%3A1017380631757.

Anonymous:2001:FBE

[Ano01b] Anonymous. The frontiers of borderless education. Minerva, 39(1):1, ???? 2001. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1017369916964.

Anonymous:2001:IVX

[Ano01c] Anonymous. Index to volume XXXIX. Minerva, 39(4): 465–466, ???? 2001. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1017450020900.

Anonymous:2001:IA

[Ano01d] Anonymous. Instructions for authors. Minerva, 39(1):143– 149, ???? 2001. CODEN MINEFY. ISSN 0026-4695 (print),

1573-1871 (electronic). URL http://link.springer.com/
accesspage/article/10.1023/A%3A1017374327296.

Anonymous:2001:WSI

[Ano01e] Anonymous. Women in science: International perspectives. Minerva, 39(2):151, ???? 2001. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1017357724049.

Anonymous:2002:CV

[Ano02a] Anonymous. Contents of volume 40. Minerva, 40(4):427– 429, ???? 2002. CODEN MINEFY. ISSN 0026-4695 (print),

1573-1871 (electronic). URL http://link.springer.com/
accesspage/article/10.1023/A%3A1020950028344.

Anonymous:2002:IVX

[Ano02b] Anonymous. Index to volume XL. Minerva, 40(4):425–426, ???? 2002. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1020916608778.

Anonymous:2002:IA

[Ano02c] Anonymous. Instructions for authors. Minerva, 40(1):107– 113, ???? 2002. CODEN MINEFY. ISSN 0026-4695 (print),

1573-1871 (electronic). URL http://link.springer.com/
accesspage/article/10.1023/A%3A1017425806054.

Anonymous:2003:IVX

[Ano03] Anonymous. Index to volume XLI. Minerva, 41(4):421–422, ???? 2003. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/B%3AMINE.0000005279.73279.9b.

Anonymous:2004:IVX

[Ano04] Anonymous. Index to volume XLII. Minerva, 42(4):455– 456, ???? 2004. CODEN MINEFY. ISSN 0026-4695 (print),

1573-1871 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11024-004-3451-2.

Anonymous:2005:E

[Ano05a] Anonymous. Erratum. Minerva, 43(2):219, June 2005. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/s11024-005-7351-x.

Anonymous:2005:IV

[Ano05b] Anonymous. Index to volume 43, 2005. Minerva, 43(4):449– 450, December 2005. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-005-5042-2.

Anonymous:2008:IA

[Ano08] Anonymous. Information for authors. Minerva, 46(4):495– 499, December 2008. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-008-9112-0.

Anstey:2003:BLI

[Ans03] Peter Anstey. Bacon’s last instalment. Minerva, 41(1):89– 92, ???? 2003. CODEN MINEFY. ISSN 0026-4695 (print),

1573-1871 (electronic). URL http://link.springer.com/
accesspage/article/10.1023/A%3A1022240203736.

Ashby:1981:BR

[AOT81] Lord Ashby, Harold Orlans, and Wilhelm Treue. Book reviews. Minerva, 19(4):640–651, Winter 1981. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01096197.

Ahmed:1992:SIU

[AR92] Rais Ahmed and Madhulika Rakesh. Science in the Indian universities. Minerva, 30(1):62–100, Spring 1992. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01096397.

Antal:2020:DAS

[AR20] Ariane Berthoin Antal and Jan-Christoph Rogge. Does academia still call? Experiences of academics in Germany and the United States. Minerva, 58(2):187–210, June 2020. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/article/10.1007/s11024-019-09391-4.

Rogvi:2023:DPS

[áRH23] Sofie á Rogvi and Klaus Hoeyer. A data-political spectacle: How COVID-19 became a source of societal division in Denmark. Minerva, 61(3):335–355, September 2023. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL https://

link.springer.com/article/10.1007/s11024-022-09486-5.

Arikpo:1963:C

[Ari63] Okoi Arikpo. Chronicle. Minerva, 2(1):118–129, Autumn 1963. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01100784.

Armon:2007:WBA

[Arm07] Rony Armon. Writing biographies and autobiographies of science. Minerva, 45(3):295–304, September 2007. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
s11024-007-9043-1.

Arnold:2000:LOH

[Arn00] Lorna Arnold. A letter from Oxford: The history of nuclear history in Britain. Minerva, 38(2):201–219, ???? 2000. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1023/A%3A1026553128414.

Arnoldi:2007:UPR

[Arn07] Jakob Arnoldi. Universities and the public recognition of expertise. Minerva, 45(1):49–61, March 2007. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
006-9028-5.

Aron:1964:SAC

[Aro64] Raymond Aron. Some aspects of the crisis in the French universities. Minerva, 2(3):279–285, Spring 1964. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01097317.

Aronova:2012:CCF

[Aro12] Elena Aronova. The Congress for Cultural Freedom, Minerva, and the quest for instituting “science studies” in the age of Cold War. Minerva, 50(3):307–337, September 2012. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/s11024-012-9206-6.

Autio-Sarasmo:2013:BRP

[AS13] Sari Autio-Sarasmo. Book review: Per Lundin, Niklas Stenlås and Johan Gribbe (eds.), Science for Welfare and Warfare: Technology and State Initiative in Cold War Sweden. Minerva, 51(1):123–126, March 2013. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-013-9223-0.

Ashby:1962:UUS

[Ash62] Eric Ashby. Universities under siege. Minerva, 1(1):18–29, Autumn 1962. CODEN MINEFY. ISSN 0026-4695 (print),

1573-1871 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF01101450.

Ashby:1966:CBU

[Ash66] Eric Ashby. Collaboration between universities and government laboratories. Minerva, 4(3):406–407, Spring 1966. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01584638.

Ashby:1967:FNC

[Ash67] Eric Ashby. The future of the nineteenth century idea of a university. Minerva, 6(1):3–17, Autumn 1967. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01101602.

Ashby:1968:GUG

[Ash68] Eric Ashby. Government, the University grants committee and the universities. Minerva, 6(2):244–256, Winter 1968. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01097260.

Ashby:1968:HOA

[Ash69] Eric Ashby. A Hippocratic oath for the academic profession. Minerva, 7(1–2):64–66, Autumn–Winter 1968–1969. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01096809.

Ashby:1996:CPA

[Ash96] Lord Ashby. Centre and periphery in academe: Some personal reflections. Minerva, 34(1):95–101, ???? 1996. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF00124204.

Ash:2000:SCG

[Ash00] Mitchell G. Ash. Scientific changes in Germany 1933, 1945, 1990: Towards a comparison. Minerva, 37(4):329–354, December 2000. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1004745805047.

Assmus:1993:CPF

[Ass93] Alexi Assmus. The creation of postdoctoral fellowships and the siting of American scientific research. Minerva, 31(2): 151–183, Summer 1993. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01125433.

AlbuMP:1970:BR

[ATIS70] Austen Albu, M.P., Hugh Thomas, Edmund Ions, and Edward Shils. Book reviews. Minerva, 8(1–4):602–623, January 1970. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01553244;   http://link.springer.com/
article/10.1007/BF01553244.

Attiyeh:1973:NLE

[Att73] Richard Attiyeh. A new look at the economics of higher education. Minerva, 11(3):335–347, July 1973. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01557688.

Auerbach:1965:SND

[Aue65] Lewis E. Auerbach. Scientists in the new deal: a prewar episode in the relations between science and government in the United States. Minerva, 3(4):457–482, Summer 1965. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01099411.

Auerbach:1980:SRC

[Aue80] Lewis Auerbach. Scientific research in the Canadian North: Three recent attempts at regulation. Minerva, 18(2):284– 292, Summer 1980. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01100249.

Auger:1963:SCW

[Aug63] Pierre Auger. Scientific cooperation in Western Europe. Minerva, 1(4):428–438, Summer 1963. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01107186.

Austin:1975:AEP

[Aus75] Dennis Austin. Et in arcadia ego: Politics and learning in Ghana. Minerva, 13(2):236–269, Summer 1975. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01097796.

Austin:1980:UAG

[Aus80] Dennis Austin. Universities and the academic gold standard in Nigeria. Minerva, 18(2):201–242, Summer 1980. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01100247.

Austin:1981:ITU

[Aus81] Dennis Austin. Ivory towers? Universities in Sri Lanka. Minerva, 19(2):203–235, Summer 1981. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01096566.

Ashby:1979:BR

[AWZ79] Eric Ashby, J. W. N. Watkins, and John Ziman. Book reviews. Minerva, 17(2):313–327, Summer 1979. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01096286.

Brodin:2020:CDC

[BA20] Eva M. Brodin and Helen Avery. Cross-disciplinary collaboration and scholarly independence in multidisciplinary learning environments at doctoral level and beyond. Minerva, 58(3):409–433, September 2020. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http:/

/link.springer.com/article/10.1007/s11024-020-09397-
3;   http://link.springer.com/content/pdf/10.1007/s11024-
020-09397-3.pdf.

Badash:2000:SM

[Bad00] Lawrence Badash. Science and McCarthyism. Minerva, 38 (1):53–80, ???? 2000. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1026516820931.

Badash:2004:SSR

[Bad04] Lawrence Badash. Science and social responsibility. Minerva, 42(3):285–298, ???? 2004. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/B%3AMINE.0000038308.46951.
ed.

Baker:1978:MPC

[Bak78] John R. Baker. Michael Polanyi’s contributions to the cause of freedom in science. Minerva, 16(3):382–396, Autumn 1978. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096667.

Baker:1987:SEO

[Bak87] Keith Michael Baker. Scientism at the end of the old regime: Reflections on a theme of Professor Charles Gillispie. Minerva, 25(1–2):21–34, Spring–Summer 1987. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01096854.

Balmer:1999:BRI

[Bal99] Brian Balmer. Book review: Improving Nature? The Science and Ethics of Genetic Engineering, by Michael J. Reiss and Roger Straughan; Birth to Death: Science and Bioethics, edited by David C. Thomasma and Thomasine Kushner. Minerva, 37 (1):95–97, March 1999. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1004544424534.

Balmer:2002:KDP

[Bal02] Brian Balmer. Killing ‘without the distressing preliminaries’: Scientists’ defence of the British Biological Warfare Programme. Minerva, 40(1):57–75, ???? 2002. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1023/
A%3A1015009613250.

Baneke:2014:SAD

[Ban14] David M. Baneke. Space for ambitions: The Dutch space program in changing European and transatlantic contexts. Minerva, 52(1):119–140, March 2014. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-014-9244-3.

Baneke:2015:BRJ

[Ban15] David Baneke. Book review: John Krige, Angelina Long Callahan and Ashok Maharaj, NASA in the World: Fifty Years of International Collaboration in Space. Minerva, 53(3):303– 305, September 2015. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-015-9277-2.

Barber:1987:TS

[Bar87] Bernard Barber. Trust in science. Minerva, 25(1–2):123–134, Spring–Summer 1987. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096860.

Bargal:1998:KLF

[Bar98] David Bargal. Kurt Lewin and the first attempts to establish a Department of Psychology at the Hebrew University. Minerva, 36(1):49–68, March 1998. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1004334520382.

Barke:2003:PIR

[Bar03] Richard P. Barke. Politics and interests in the Republic of science. Minerva, 41(4):305–325, ???? 2003. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1023/
B%3AMINE.0000005153.03357.48.

Barbour:2006:RSL

[Bar06] Reid Barbour. Renaissance science and literature. Minerva, 44 (1):113–117, March 2006. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-005-5398-3.

Barandiaran:2015:RSA

[Bar15] Javiera Barandiaran. Reaching for the stars? Astronomy and growth in Chile. Minerva, 53(2):141–164, June 2015. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-015-9272-7.

Barlosius:2023:WSA

[Bar23] Eva Barlösius. “We share all data with each other”: Datasharing in peer-to-peer relationships. Minerva, 61(2):243–263, June 2023. CODEN MINEFY. ISSN 0026-4695 (print),

1573-1871 (electronic). URL https://link.springer.com/
article/10.1007/s11024-023-09487-y.

Bass:1967:RUC

[Bas67] Lawrence W. Bass. Research in an underdeveloped country. Minerva, 5(4):581–585, Summer 1967. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01096790.

Bauer:1977:RWT

[Bau77] Peter Bauer. Reflections on Western technology and “third world” development. Minerva, 15(2):144–154, Summer 1977. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096674.

Bauder:2020:IMS

[Bau20] Harald Bauder. International mobility and social capital in the academic field. Minerva, 58(3):367–387, September 2020. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/article/10.1007/s11024-
020-09401-w.

Bulmer:1981:PSS

[BB81] Martin Bulmer and Joan Bulmer. Philanthropy and social science in the 1920s: Beardsley Ruml and the Laura Spelman Rockefeller memorial, 1922–29. Minerva, 19(3):347–407, Autumn 1981. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02192822.

Benz:2024:SPE

[BB24] Pierre Benz and Felix Bühlmann. Structural power and epistemologies in the scientific field: Why a rapid reconciliation between functional and evolutionary biology is unlikely. Minerva, 62(2):229–251, June 2024. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL https://link.springer.

com/article/10.1007/s11024-023-09520-0.

Beteille:1992:C

[BBD^{+}92] André Beteille, Asa Briggs, Hans Daalder, Michèle Gendreau-Massaloux, Patricia Albjerg Graham, Heinz Maier-Leibnitz, Amrik Singh, Wang Gungwu, and Anthony C. Yu. Comments. Minerva, 30(2):206–241, Summer 1992. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01096883.

Buhlmann:2017:MPL

[BBMR17] Felix Bühlmann, Pierre Benz, André Mach, and Thierry Rossier. Mapping the power of law professors: The role of scientific and social capital. Minerva, 55(4):509–531, December 2017. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/article/10.1007/s11024-
017-9333-1.

Bocock:2003:AIB

[BBSS03] Jean Bocock, Lewis Baston, Peter Scott, and David Smith. American influence on British higher education: Science, technology, and the problem of University expansion, 1945– 1963. Minerva, 41(4):327–346, ???? 2003. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1023/
B%3AMINE.0000005154.25610.b2.

Billi:2019:WLS

[BBU19] Marco Billi, Gustavo Blanco, and Anahí Urquiza. What is the ‘social’ in climate change research? A case study on scientific representations from Chile. Minerva, 57(3):293–315, September 2019. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/article/10.
1007/s11024-019-09369-2;   http://link.springer.com/
content/pdf/10.1007/s11024-019-09369-2.pdf.

Bertrand:2011:MRF

[BBV11] Emanuel Bertrand and Bernadette Bensaude-Vincent. Materials research in France: a short-lived national initiative (1982– 1994). Minerva, 49(2):191–214, June 2011. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
011-9168-0.

Barber:1973:BR

[BBZA73] Bernard Barber, Harvey Brooks, John Ziman, and Mary Anderson. Book reviews. Minerva, 11(3):415–429, July 1973. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01557692.

Birrenbach:1985:PPF

[BC85] Kurt Birrenbach and Helmut T. Coing. The place of private foundations in the support of research in the Federal Republic of Germany. Minerva, 23(3):383–422, Autumn 1985. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096445.

Brint:2017:SRO

[BC17] Steven Brint and Cynthia E. Carr. The scientific research output of U.S. research universities, 1980–2010: Continuing dispersion, increasing concentration, or stable inequality? Minerva, 55(4): 435–457, December 2017. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/article/10.1007/s11024-017-9330-4.

Begin-Caouette:2020:PIE

[BC20] Olivier Bégin-Caouette. The perceived impact of eight systemic factors on scientific capital accumulation. Minerva, 58 (2):163–185, June 2020. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/article/10.1007/s11024-019-09390-5.

Boudon:1976:SCH

[BCL76] Raymond Boudon, Philippe Cibois, and Janina Lagneau. Shortcycle higher education and the pitfalls of collective action. Minerva, 14(1):33–60, Spring 1976. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01096213.

Ben-David:1964:SGS

[BD64] Joseph Ben-David. Scientific growth: a sociological view. Minerva, 2(4):455–476, Summer 1964. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01097538.

Ben-David:1965:SRC

[BD65] Joseph Ben-David. The scientific role: The conditions of its establishment in Europe. Minerva, 4(1):15–54, Autumn 1965. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01585983.

Ben-David:1967:SPN

[BD67] Joseph Ben-David. Scientific policy in the Netherlands. Minerva, 6(1):118–121, Autumn 1967. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01101614.

Ben-David:1968:UGS

[BD69] Joseph Ben-David. The universities and the growth of science in Germany and the United States. Minerva, 7(1–2): 1–35, Autumn–Winter 1968–1969. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01096806.

Ben-David:1970:RDF

[BD70] Joseph Ben-David. The rise and decline of France as a scientific centre. Minerva, 8(1–4):160–179, January 1970. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01553210.

Ben-David:1972:PSP

[BD72] Joseph Ben-David. The profession of science and its powers. Minerva, 10(3):362–383, July 1972. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01556920.

Bonaccorsi:2010:UNK

[BDG10] Andrea Bonaccorsi, Cinzia Daraio, and Aldo Geuna. Universities in the new knowledge landscape: Tensions, challenges, change — an introduction. Minerva, 48(1):1–4, March 2010. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-010-9144-0.

Ben-David:1977:BR

[BDK77] Joseph Ben-David and C. P. Kindleberger. Book reviews. Minerva, 15(3–4):539–561, Autumn–Winter 1977. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01096731.

Beckerman:1973:EGW

[Bec73] Wilfred Beckerman. Economic growth and welfare. Minerva, 11 (4):495–515, October 1973. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01557799.

Bragesjo:2012:CDS

[BEK12] Fredrik Bragesjö, Aant Elzinga, and Dick Kasperowski. Continuity or discontinuity? Scientific governance in the pre-history of the 1977 Law of Higher Education and Research in Sweden. Minerva, 50(1):65–96, March 2012. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
012-9188-4.

Bekerman:2013:SFD

[Bek13] Fabiana Bekerman. The scientific field during Argentina’s latest military dictatorship (1976–1983): Contraction of public universities and expansion of the national council for scientific and technological research (CONICET). Minerva, 51(2): 253–269, June 2013. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-013-9227-9.

Beloff:1967:BUP

[Bel67] Max Beloff. British universities and the public purse. Minerva, 5 (4):520–532, Summer 1967. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096781.

Beloff:1968:HU

[Bel68a] Max Beloff. Hands off the universities? Minerva, 6(4): 601–603, Summer 1968. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096552.

Beloff:1968:UGP

[Bel68b] Max Beloff. The universities, the government and the public accounts committee. Minerva, 6(2):264–265, Winter 1968. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01097263.

Beloff:1994:BUS

[Bel94] Max Beloff. The British universities and the state. Minerva, 32 (2):188–193, Summer 1994. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01103358.

Beloff:1997:BRR

[Bel97] Max Beloff. Book review: Redbrick University Revisited: The Autobiography of “Bruce Truscot” (E. Allison Peers Publications, Vol. I), edited by Ann L. Mackenzie and Adrian R. Allen. Minerva, 35(4):385–386, December 1997. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1023/
A%3A1004383405587.

Beloff:1998:BRW

[Bel98a] Max Beloff. Book review: We Scholars: Changing the Culture of the University, by David Damrosch. Minerva, 36(1): 90–91, March 1998. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1004350924016.

Beloff:1998:HUO

[Bel98b] Max Beloff. The History of the University of Oxford. Volume VI: Nineteenth-Century Oxford, Part 1; A History of the University of Cambridge. Volume III: 1750–1850. Minerva, 36(4): 387–391, December 1998. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1004479032065.

Bennett:1998:BRI

[Ben98] J. A. Bennett. Book review: The Invisible World: Early Modern Philosophy and the Invention of the Microscope, by Catherine Wilson; The Microscope in the Dutch Republic: The Shaping of Discovery, by Edward G. Ruestow. Minerva, 36(1):84–85, March 1998. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1004394723108.

Benson:2007:OCW

[Ben07a] Keith R. Benson. Oceanography and the cold war effect. Minerva, 45(2):223–224, June 2007. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
007-9033-3.

Benum:2007:MRC

[Ben07b] Edgeir Benum. Making research count: Norway and the OECD connection 1965–1980. Minerva, 45(4):365–387, December 2007. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-007-9064-9.

Benneworth:2013:BRE

[Ben13] Paul Benneworth. Book review: Elizabeth Popp Berman, Creating the Market University: How Academic Science Became an Economic Engine. Minerva, 51(4):521–527, December 2013. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-013-9236-8.

Berry:1971:C

[Ber71] R. Stephen Berry. Correspondence. Minerva, 9(4):565–567, October 1971. CODEN MINEFY. ISSN 0026-4695 (print),

1573-1871 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF01558033.

Berghahn:2005:AIC

[Ber05] Volker R. Berghahn. America and the intellectual Cold Wars in Europe. Minerva, 43(1):103–107, ???? 2005. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
004-6642-y.

Berker:2011:BRM

[Ber11] Thomas Berker. Book review: Michel Callon, Pierre Lascoumes and Yannick Barthe, Acting in an Uncertain World: an Essay on Technical Democracy. Minerva, 49(4):509–511, December 2011. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-011-9186-y.

Besson:1968:URW

[Bes68] Waldemar Besson. University reform in West Germany. Minerva, 6(4):614–617, Summer 1968. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01096557.

Beteille:1981:IUA

[Bét81] André Béteille. The Indian university: Academic standards and the pursuit of equality. Minerva, 19(2):282–310, Summer 1981. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096569.

Beteille:1990:CDP

[Bét90] André Béteille. A career in a declining profession. Minerva, 28(1):1–20, Spring 1990. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096324.

Beyler:2006:MDK

[Bey06] Richard H. Beyler. Maintaining discipline in the Kaiser Wilhelm Society during the National Socialist regime. Minerva, 44 (3):251–266, June 2006. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-006-9001-3.

Bieszczad:2025:DRH

[BFdR25] Sarah Rose Bieszczad, Maximilian Fochler, and Sarah de Rijcke. Diving into relevance: How deep sea researchers articulate societal relevance within their epistemic living spaces. Minerva, 63(2):205–229, June 2025. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL https://link.springer.

com/article/10.1007/s11024-025-09577-z.

Baldini:2014:OCI

[BFGS14] Nicola Baldini, Riccardo Fini, Rosa Grimaldi, and Maurizio Sobrero. Organisational change and the institutionalisation of University patenting activity in Italy. Minerva, 52(1):27–53, March 2014. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-013-9243-9.

Bandola-Gill:2023:KBR

[BG23a] Justyna Bandola-Gill. Knowledge brokering repertoires: Academic practices at science-policy interfaces as an epistemological bricolage. Minerva, 61(1):71–92, March 2023. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (elec-

tronic).  URL https://link.springer.com/article/10.
1007/s11024-022-09478-5.

Bozeman:2023:ZCH

[BG23b] Barry Bozeman and Monica Gaughan. The ‘Zoomification’ of collaboration: How timely technology has affected academic research. Minerva, 61(4):467–493, December 2023. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (elec-

tronic).  URL https://link.springer.com/article/10.
1007/s11024-023-09500-4.

Beigel:2018:IES

[BGB18] Fernanda Beigel, Osvaldo Gallardo, and Fabiana Bekerman. Institutional expansion and scientific development in the periphery: The structural heterogeneity of Argentina’s academic field. Minerva, 56(3):305–331, September 2018. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/article/10.1007/s11024-017-9340-2.

Bloch:2014:CRG

[BGP14] Carter Bloch, Ebbe Krogh Graversen, and Heidi Skovgaard Pedersen. Competitive research grants and their impact on career performance. Minerva, 52(1):77–96, March 2014. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/s11024-014-9247-0.

Bianco:2016:AEU

[BGS16] Mariela Bianco, Natalia Gras, and Judith Sutz. Academic evaluation: Universal instrument? Tool for development? Minerva, 54(4):399–421, December 2016. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
016-9306-9.

Baron:2015:MDA

[BH15] Christopher Baron and Christopher Hamlin. Malaria and the decline of ancient Greece: Revisiting the Jones hypothesis in an era of interdisciplinarity. Minerva, 53(4):327–358, December 2015. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-015-9280-7.

Bhaneja:1979:PIS

[Bha79] Balwant Bhaneja. Parliamentary influence on science policy in India. Minerva, 17(1):70–97, Spring 1979. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01096261.

Brint:2022:CPU

[BI22] Steven Brint and Ali O. Ilhan. Conceptions of professionalism in U.S. research universities: Evidence from the gradSERU survey. Minerva, 60(4):535–565, December 2022. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL https://

link.springer.com/article/10.1007/s11024-022-09476-7.

Bickel:1974:AEP

[Bic74] Alexander M. Bickel. The aims of education and the proper standards of the university. Minerva, 12(2):199–206, April 1974. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01553177.

Biernacki:1967:ODS

[Bie67] Andrzej Biernacki. Observations on the development of science in Poland. Minerva, 6(1):18–27, Autumn 1967. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01101603.

Biobaku:1963:ASA

[Bio63] Saburi Biobaku. African studies in an African university. Minerva, 1(3):285–301, Spring 1963. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF02251985.

Bissell:1973:DAP

[Bis73] Claude T. Bissell. The dilution of academic power in Canada. Minerva, 11(1):130–133, January 1973. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01544983.

Bjarnason:2001:MCN

[Bja01] Svava Bjarnason. Managing the changing nature of teaching and learning. Minerva, 39(1):85–98, ???? 2001. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1023/A%3A1010378403760.

Bjork:2001:INC

[Bjö01] Ragnar Björk. Inside the Nobel Committee on Medicine: Prize competition procedures 1901–1950 and the fate of Carl Neuberg. Minerva, 39(4):393–408, ???? 2001. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1023/
A%3A1012767418228.

Betz:1970:SRU

[BK70] Frederick Betz and Carlos Kruytbosch. Sponsored research and university budgets: a case study in American university government. Minerva, 8(1–4):492–519, January 1970. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01553236;   http://link.springer.com/article/
10.1007/BF01553236.

Broitman:2018:KPM

[BK18] Claudio Broitman and Pablo Kreimer. Knowledge production, mobilization and standardization in Chile’s HidroAysén case. Minerva, 56(2):209–229, June 2018. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/article/10.1007/s11024-017-9335-z.

Balakrishnan:2016:LES

[BL16] Balamuralithara Balakrishnan and Foon Siang Low. Learning experience and socio–cultural influences on female engineering students’ perspectives on engineering courses and careers. Minerva, 54(2):219–239, June 2016. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
016-9295-8.

Blaug:1967:EPR

[Bla67] Mark Blaug. Educational planning: Review article. Minerva, 6 (1):43–47, Autumn 1967. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01101605.

Blaug:1968:PU

[Bla68] Mark Blaug. The productivity of universities. Minerva, 6(3): 398–407, Spring 1968. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096425.

Blomqvist:1997:SUA

[Blo97] Göran Blomqvist. State, university and academic freedom in Sweden: The universities of Uppsala and Lund between 1820 and 1920. Minerva, 35(2):171–194, June 1997. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1023/A%3A1004246928195.

Blume:1969:RSB

[Blu69] Stuart Blume. Research support in British universities. Minerva, 7(4):649–667, Summer 1969. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01099539.

Beloff:1994:BR

[BM94] Max Beloff and David B. McLay. Book reviews. Minerva, 32 (3):344–356, Autumn 1994. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01098668.

Brostrom:2018:EES

[BM18] Anders Broström and Maureen McKelvey. Engaging experts: Science–policy interactions and the introduction of congestion charging in Stockholm. Minerva, 56(2):183–207, June 2018. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/article/10.1007/s11024-
017-9331-3;   http://link.springer.com/content/pdf/10.
1007/s11024-017-9331-3.pdf.

Brunet:2024:FRP

[BM24] Lucas Brunet and Ruth Müller. The feeling rules of peer review: Defining, displaying, and managing emotions in evaluation for research funding. Minerva, 62(2):167–192, June 2024. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL https://link.springer.com/article/10.
1007/s11024-023-09518-8.

Bowen:1992:C

[BMR^{+}92] William G. Bowen, William Massy, William C. Richardson, Henry Rosovsky, and George Stigler. Comments. Minerva, 30 (2):175–188, Summer 1992. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096881.

Bulmer:1987:BR

[BMSC87] Martin Bulmer, Michael J. Moravcsik, Barnett Singer, and Arye Carmon. Book reviews. Minerva, 25(3):362–394, Autumn 1987. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01097790.

Bodstrom:1964:SUP

[Bod64] L. Bodström. Swedish University policy: The 1955 Commission on the Swedish Universities. Minerva, 3(1):83–90, Autumn 1964. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01630151.

Bonaccorsi:2008:SRI

[Bon08] Andrea Bonaccorsi. Search regimes and the industrial dynamics of science. Minerva, 46(3):??, September 2008. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/s11024-008-9101-3.

Bonaccorsi:2010:NFC

[Bon10] Andrea Bonaccorsi. New forms of complementarity in science. Minerva, 48(4):355–387, December 2010. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
010-9159-6.

Borowy:1967:ERD

[Bor67] Michael Borowy. Expenditures on research and development in Poland: 1961 to 1965. Minerva, 5(3):357–371, Spring 1967. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02388574.

Borras:2004:LEC

[Bor04] Susana Borrás. The ‘learning’ economy in contemporary societies. Minerva, 42(1):97–100, ???? 2004. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1023/
B%3AMINE.0000017836.97412.45.

Bortolini:2021:GMT

[Bor21] Matteo Bortolini. The grudging modernizer: a trip to the Middle East and Cold War social science. Minerva, 59(2):261–284, June 2021. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL https://link.springer.com/article/10.
1007/s11024-020-09413-6.

Boschiero:2005:SAB

[Bos05] Luciano Boschiero. Stories about the birth of modern science. Minerva, 43(3):311–318, ???? 2005. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
005-6511-3.

Bouchard:2008:SML

[Bou08] Julie Bouchard. Science, markets, and the law. Minerva, 46 (1):143–146, March 2008. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-007-9079-2.

Bowden:1967:UGP

[Bow67a] Lord Bowden. The universities, the government and the public accounts committee. Minerva, 6(1):28–42, Autumn 1967. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01101604.

Bowles:1967:AHE

[Bow67b] Frank Bowles. American higher education in 1990. Minerva, 5 (2):227–241, Winter 1967. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF02388526.

Bown:2002:HED

[Bow02] Lalage Bown. Higher education and development. Minerva, 40 (4):417–420, ???? 2002. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1020947219635.

Boyle:1966:PUP

[Boy66] Edward Boyle. Parliament and university policy. Minerva, 5 (1):3–19, Autumn 1966. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF02389789.

Bozzli:1975:AFS

[Boz75] G. R. Bozzli. Academic freedom in south Africa. Minerva, 13 (3):428–465, Autumn 1975. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01097849.

Bulmer:1984:BRP

[BPM84] Martin Bulmer, Rudolf Peierls, and Kenneth Mellanby. Book reviews: Philanthropic Foundations and the Development of a National Society in the United States (pp. 276–285). Private Power for the Public Good: a History of the Carnegie Foundation for the Advancement of Teaching by Ellen Lagemann. Review by: Martin Bulmer. Forty Years into the Atomic Age (pp. 285–291). The Nuclear Chain Reaction Forty Years Later by Robert G. Sachs. Review by: Rudolf Peierls. Scientific Ecology and the Protection of the Environment (pp. 291– 293). Planning and Ecology by R. D. Roberts, T. M. Roberts; Lead Pollution: Causes and Control by R. M. Harrison, D. P. H. Laxen. Review by: Kenneth Mellanby. Minerva, 22(2): 276–293, Summer 1984. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF02207952;   http://
www.jstor.org/stable/41820566; http://www.jstor.org/
stable/41820567;   http://www.jstor.org/stable/41820568.

Branscomb:1992:AET

[Bra92] Lewis M. Branscomb. America’s emerging technology policy. Minerva, 30(3):317–336, Autumn 1992. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01097642.

Braun:1993:BRP

[Bra93] Dietmar Braun. Biomedical research in a period of scarcity: The United States and Great Britain. Minerva, 31(3):268– 290, Autumn 1993. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01098624.

Brain:2007:MHG

[Bra07a] Robert M. Brain. Modernity: How Germany and Great Britain faced the early years of technology. Minerva, 45(3): 331–335, September 2007. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-007-9051-1.

Bray:2007:PYA

[Bra07b] Francesca Bray. Ho Peng Yoke: an autobiography. Minerva, 45(3):337–340, September 2007. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
007-9050-2.

Braun:2012:WDS

[Bra12] Dietmar Braun. Why do scientists migrate? A diffusion model. Minerva, 50(4):471–491, December 2012. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
012-9214-6.

Brandl:2015:AKS

[Bra15] Barbara Brandl. Abby Kinchy, seeds, science, and struggle: The global politics of transgenic crops. Minerva, 53(2):189–192, June 2015. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-015-9267-4.

Brentano:1995:TBC

[Bre95] Carroll Brentano. The two Berkeleys: City and university through 125 years. Minerva, 33(4):361–371, Winter 1995. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096518.

Brechtken:2004:MWF

[Bre04] Magnus Brechtken. Max Weber: a Family Portrait: Guenther Roth, Max Webers deutsch-englische Familiengeschichte 1800– 1950, mit Briefen und Dokumenten (Tübingen: Mohr Siebeck, 2001), 721 pp. + index, ISBN: 3-16-147557-7, 84 EUR. Minerva, 42(4):445–450, ???? 2004. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-004-2998-2.

bret:2006:EE

[bre06] Patrice bret. Enlightened engineering. Minerva, 44(4):439– 446, December 2006. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-006-9010-2.

Brickman:1977:PMS

[Bri77] Ronald Brickman. The promotion of mobility of scientists: a problem of French science policy. Minerva, 15(1):62–82, Spring 1977. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096320.

Brint:2005:CFN

[Bri05] Steven Brint. Creating the future: ‘new directions’ in American research universities. Minerva, 43(1):23–50, ???? 2005. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-004-6620-4.

Brinckmann:2006:SSS

[Bri06] Andrea Brinckmann. The studiengruppe Für systemforschung: Systems research and policy advice in the Federal Republic of Germany, 1958–1975. Minerva, 44(2):149–166, June 2006. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-006-0004-x.

Brooks:1965:PSA

[Bro65] Harvey Brooks. The President’s science advisers. Minerva, 3 (3):392–398, Spring 1965. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01099963.

Brooks:1972:WLT

[Bro72] Harvey Brooks. “World leadership”, the “technological gap” and national science policy. Minerva, 10(2):323–329, April 1972. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01682425.

Brocke:1991:FAG

[Bro91] Bernhard Vom Brocke. Friedrich Althoff: a great figure in higher education policy in Germany. Minerva, 29(3):269–293, Autumn 1991. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01097879.

Brockliss:2000:GTU

[Bro00] Laurence Brockliss. Gown and Town: The University and the city in Europe, 1200–2000. Minerva, 38(2):147–170, ???? 2000. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1026596910667.

Brown:2004:PPS

[Bro04] Mark B. Brown. The political philosophy of science policy. Minerva, 42(1):77–95, ???? 2004. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1023/B%3AMINE.0000017701.
73799.42.

Brock:2005:PSV

[Bro05] W. H. Brock. Popular science in the Victorian periodical. Minerva, 43(3):319–323, ???? 2005. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
005-6512-2.

Brooke:2006:MM

[Bro06] John Hedley Brooke. Modernity at the margins. Minerva, 44(4): 463–467, December 2006. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-006-9009-8.

Brown:2007:MG

[Bro07a] Gavin Brown. Managing governance. Minerva, 45(4):499– 503, December 2007. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-007-9057-8.

Brown:2007:TVP

[Bro07b] Gavin Brown. Technology versus the people. Minerva, 45 (2):225–229, June 2007. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-007-9035-1.

Brown:2008:RRP

[Bro08] Mark B. Brown. Review of Roger S. Pielke, Jr., The Honest Broker: Making Sense of Science in Policy and Politics. Minerva, 46(4):485–489, December 2008. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
008-9106-y.

Brosnan:2011:SSC

[Bro11] Caragh Brosnan. The significance of scientific capital in UK medical education. Minerva, 49(3):??, September 2011. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/s11024-011-9177-z.

Brown:2013:PKS

[Bro13] Mark B. Brown. Philip Kitcher, science in a democratic society. Minerva, 51(3):389–397, September 2013. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
013-9233-y.

Brown:2020:BRG

[Bro20] Mark B. Brown. Book review: Gil Eyal, The Crisis of Expertise. Polity Press, Cambridge, UK, 2019, 190 pp, ISBN: 978-0-7456-6578-8 (pbk). Minerva, 58(4):657–660, December 2020. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/article/10.1007/s11024-
020-09417-2;   http://link.springer.com/content/pdf/
10.1007/s11024-020-09417-2.pdf.

Bruno:2009:IDC

[Bru09] Isabelle Bruno. The “indefinite discipline” of competitiveness benchmarking as a neoliberal technology of government. Minerva, 47(3):??, September 2009. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
009-9128-0.

Brunet:2025:MMS

[Bru25] Lucas Brunet. Mapping for Mapping’s sake? ecosystems services maps and the modes of (ir)relevance of ecological knowledge for nature conservation. Minerva, 63(2):231–255, June 2025. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL https://link.springer.com/article/10.
1007/s11024-024-09552-0.

Bezuidenhout:2019:DPV

[BRWB19] Louise Bezuidenhout, Emanuele Ratti, Nathaniel Warne, and Dori Beeler. Docility as a primary virtue in scientific research. Minerva, 57(1):67–84, March 2019. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/article/10.1007/s11024-018-9356-2;
 http://link.springer.com/content/pdf/10.1007/s11024-
018-9356-2.pdf.

Benner:2007:SSR

[BS07] Mats Benner and Sverker Sörlin. Shaping strategic research: power, resources, and interests in Swedish research policy. Minerva, 45(1):31–48, March 2007. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
006-9019-6.

Bozeman:2011:PVM

[BS11] Barry Bozeman and Daniel Sarewitz. Public value mapping and science policy evaluation. Minerva, 49(1):1–23, March 2011. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-011-9161-7.

Bielinski:2019:ESC

[BT19] Jacek Bieliński and Aldona Tomczyńska. The ethos of science in contemporary Poland. Minerva, 57(2):151–173, June 2019. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/article/10.1007/s11024-
018-9365-1;   http://link.springer.com/content/pdf/10.
1007/s11024-018-9365-1.pdf.

Bulmer:1980:EIE

[Bul80] Martin Bulmer. The early institutional establishment of social science research: The local community research committee at the University of Chicago, 1923–30. Minerva, 18(1):51–110, Spring 1980. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096661.

Bulmer:1996:ESS

[Bul96] Martin Bulmer. Edward Shils as a sociologist. Minerva, 34 (1):7–21, ???? 1996. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00124197.

Bulmer:2001:NR

[Bul01] Sue Rabbitt Bulmer. Nuclear revisionism. Minerva, 39(2): 259–264, ???? 2001. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1010396913647.

Bulmer:2002:KIH

[Bul02] Martin Bulmer. Knowledge institutionalized: Higher education and philanthropic foundations. Minerva, 40(2):189–201, ???? 2002. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1015790930231.

Bunzel:1981:CRS

[Bun81] John H. Bunzel. A critical review of the statement of affirmative action in the 1980s of the United States commission on civil rights. Minerva, 19(2):311–328, Summer 1981. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096570.

Busch:1963:VPB

[Bus63] Alexander Busch. The vicissitudes of the privatdozent: Breakdown and adaptation in the recruitment of the German university teacher. Minerva, 1(3):319–341, Spring 1963. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF02251987.

Buxton:2003:JMH

[Bux03] William J. Buxton. John Marshall and the humanities in Europe: Shifting patterns of Rockefeller Foundation support. Minerva, 41(2):133–153, ???? 2003. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1023606814200.

Bell:1970:PPA

[BW70] Mr. Laird Bell and Leonard D. White. Presidents and professors in American university government. Minerva, 8(1–4): 440–448, January 1970. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01553228;   http://
link.springer.com/article/10.1007/BF01553228.

Balderston:1971:C

[BW71] F. E. Balderston and Thomas Walsh. Correspondence. Minerva, 9(3):414–417, July 1971. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01553366.

Brint:2024:UPH

[BWF24] Steven Brint, Megan Webb, and Benjamin Fields. An uneasy peace: How STEM progressive, traditionalist, and bridging faculty understand campus conflicts over diversity, antiracism, and free expression. Minerva, 62(3):339–372, September 2024. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL https://link.springer.com/article/10.
1007/s11024-023-09517-9.

Bulmer:1990:BR

[BWW90] Martin Bulmer, Alvin M. Weinberg, and George Weisz. Book reviews. Minerva, 28(1):98–113, Spring 1990. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01096330.

Buckner:2021:ILG

[BZ21] Elizabeth Buckner and Mike Zapp. Institutional logics in the global higher education landscape: Differences in organizational characteristics by sector and founding era. Minerva, 59(1):27–51, March 2021. CODEN MINEFY. ISSN 0026-4695 (print), 1573-

1871 (electronic). URL http://link.springer.com/article/
10.1007/s11024-020-09416-3.

Cassity:2006:HIP

[CA06] Elizabeth Cassity and Ien Ang. Humanities–industry partnerships and the ‘knowledge society’: The Australian experience. Minerva, 44(1):47–63, March 2006. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
005-5412-9.

Conceicao:2020:EAP

[CÁCC20] Cristina Palma Conceição, Patrícia Ávila, Ana Rita Coelho, and António Firmino Costa. European action plans for science– society relations: Changing buzzwords, changing the agenda. Minerva, 58(1):1–24, March 2020. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/article/10.1007/s11024-019-09380-7.

Cahn:2005:CP

[Cah05] Robert Cahn. Cambridge physics. Minerva, 43(3):325–329, ???? 2005. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-005-6513-1.

Cain:2002:CBC

[Cai02] Frank Cain. Code breakers and commandos. Minerva, 40 (3):315–325, ???? 2002. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1019514607410.

Cai:2020:IIR

[Cai20] Yuzhuo Cai. ‘Innovation in innovation’: a review of Henry Etzkowitz and Chunyan Zhou, The Triple Helix: University– Industry–Government Innovation and Entrepreneurship (Second Edition). Minerva, 58(4):651–656, December 2020. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/article/10.1007/s11024-
020-09418-1.

Calvert:2004:IBR

[Cal04] Jane Calvert. The idea of ‘basic research’ in language and practice. Minerva, 42(3):251–268, ???? 2004. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1023/
B%3AMINE.0000038307.58765.b4.

Campbell:1981:RAE

[Cam81] Keith O. Campbell. The role of agricultural economists in the conservation of natural resources. Minerva, 19(4):632–639, Winter 1981. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096196.

Camrosio:1985:DNP

[Cam85] Alberto Camrosio. The dominance of nuclear physics in Italian science policy. Minerva, 23(4):464–484, Winter 1985. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01098701.

Camic:2011:BCS

[Cam11] Charles Camic. Bourdieu’s cleft sociology of science. Minerva, 49(3):??, September 2011. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-011-9176-0.

Cantwell:2011:TMI

[Can11] Brendan Cantwell. Transnational mobility and international academic employment: Gatekeeping in an academic competition arena. Minerva, 49(4):425–445, December 2011. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/s11024-011-9181-3.

Cao:1998:CAS

[Cao98] Cong Cao. The Chinese Academy of Sciences: The election of scientists into the elite group. Minerva, 36(4):323–346, December 1998. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1004421406141.

Cao:2004:CSN

[Cao04] Cong Cao. Chinese science and the ‘Nobel Prize complex’. Minerva, 42(2):151–172, ???? 2004. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1023/
B%3AMINE.0000030020.28625.7e.

Cao:2014:UVS

[Cao14] Cong Cao. The universal values of science and China’s Nobel Prize pursuit. Minerva, 52(2):141–160, June 2014. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/s11024-014-9249-y.

Capp:1998:BRH

[Cap98] Bernard Capp. Book review: The History of the University of Oxford. Volume IV: Seventeenth-Century Oxford. Minerva, 36(4):381–386, December 1998. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1023/A%3A1004462513409.

Carter:1963:DSE

[Car63] C. F. Carter. The distribution of scientific effort. Minerva, 1 (2):172–181, Winter 1963. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096249.

Carmon:1978:DCF

[Car78] Arye Carmon. The diverse and changing fortunes of the University of Heidelberg under National Socialism. Minerva, 16 (4):516–544, Winter 1978. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01100331.

Cardozier:1988:ASC

[Car88] V. R. Cardozier. America’s state colleges. Minerva, 26(4): 549–574, Winter 1988. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096497.

Casper:1996:CMW

[Cas96] Gerhard Casper. Come the millennium, where the university? Minerva, 34(1):69–83, ???? 1996. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00124202.

Castonguay:2005:TAR

[Cas05] Stéphane Castonguay. The transformation of agricultural research in France: The introduction of the American system. Minerva, 43(3):265–287, ???? 2005. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
005-6508-y.

Curran:2020:DCH

[CBB20] Michaela Curran, Quinn Bloom, and Steven Brint. Does cluster hiring enhance faculty research output, collaborations, and impact? Results from a national study of U.S. research universities. Minerva, 58(4):585–605, December 2020. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/article/10.1007/s11024-020-09408-3.

Cumpsty:1995:BR

[CBJ95] N. A. Cumpsty, Harvey Brooks, and Paul Josephson. Book reviews. Minerva, 33(2):193–208, Summer 1995. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01097912.

Cambrosio:2017:ORB

[CBKN17] Alberto Cambrosio, Pascale Bourret, Peter Keating, and Nicole Nelson. Opening the regulatory black box of clinical cancer research: Transnational expertise networks and “disruptive” technologies. Minerva, 55(2):161–185, June 2017. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

Cruz-Castro:2018:AAP

[CCSM18] Laura Cruz-Castro and Luis Sanz-Menéndez. Autonomy and authority in public research organisations: Structure and funding factors. Minerva, 56(2):135–160, June 2018. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/article/10.1007/s11024-018-9349-1;
 http://link.springer.com/content/pdf/10.1007/s11024-
018-9349-1.pdf.

Castro-Diaz:2025:PCI

[CDRH25] Laura Castro-Diaz, Anais Roque, and Michael Hanemann. Participatory convergence: Integrating convergence and participatory action research. Minerva, 63(3):535–555, September 2025. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL https://link.springer.com/article/10.
1007/s11024-024-09547-x.

Carson:2002:SAS

[CG02] Cathryn Carson and Michael Gubser. Science advising and science policy in post-war West Germany: The example of the Deutscher Forschungsrat. Minerva, 40(2):147–179, ???? 2002. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1015708903590.

Colleret:2022:ITP

[CG22] Maxime Colleret and Yves Gingras. Out of the Ivory Tower: The patenting activity of Canadian University professors before the 1980s. Minerva, 60(2):281–300, June 2022. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL https://

link.springer.com/article/10.1007/s11024-021-09458-1.

Cooper:1972:BR

[CGTPM72] William Mansfield Cooper, June Goodfield-Toulmin, H. C. Perejra, and Peter Mathias. Book reviews. Minerva, 10(2): 332–347, April 1972. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01682427.

Cramer:1985:DFW

[CH85] Jacqueline Cramer and Rob Hagendijk. Dutch fresh-water ecology: The links between national and international scientific research. Minerva, 23(4):485–503, Winter 1985. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01098702.

Clemens:2024:PSC

[CH24] Mario Clemens and Christian Hochmuth. Political speech on campus: Shifting the emphasis from “if” to “how”. Minerva, 62(3):373–396, September 2024. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL https://link.

springer.com/article/10.1007/s11024-024-09525-3.

Chagula:1968:EAA

[Cha68] W. K. Chagula. The East African Academy. Minerva, 6(3): 408–418, Spring 1968. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096426.

Chandrasekhar:1984:PS

[Cha84] S. Chandrasekhar. The pursuit of science. Minerva, 22(3–4):410– 420, Autumn–Winter 1984. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF02207374.

Chayut:1994:HSR

[Cha94] Michael Chayut. The hybridisation of scientific roles and ideas in the context of centres and peripheries. Minerva, 32(3): 297–308, Autumn 1994. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01098664.

Chankseliani:2024:BRD

[Cha24] Maia Chankseliani. Book review: David P. Baker and Justin J. W. Powell, Global Mega–Science: Universities, Research Collaborations and Knowledge Production. Minerva, 62(3):485– 489, September 2024. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL https://link.springer.

com/article/10.1007/s11024-024-09533-3.

Cherns:1971:SSR

[Che71] A. B. Cherns. Social science research policies in the United States. Minerva, 9(4):550–553, October 1971. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01558027.

Calderon-Hernandez:2020:BUS

[CHJZSG20] Gregorio Calderón-Hernández, Yudy Andrea Jiménez-Zapata, and Héctor Mauricio Serna-Gomez. Barriers to university spinoff creation in an emerging context: an institutional theory of organizations approach. Minerva, 58(4):625–650, December 2020. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s11024-020-09407-4.

Chlup:2023:CVD

[Chl23] Radek Chlup. Covid vaccination disputes in Czechia: Political myth-making and boundary work. Minerva, 61(3):383– 405, September 2023. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL https://link.springer.

com/article/10.1007/s11024-023-09505-z.

Christian:2008:ERA

[Chr08] Rochelle Christian. Essay review: Assessing and managing environmental risks. Minerva, 46(2):275–283, June 2008. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-008-9099-6.

Chubin:1985:MRI

[Chu85] Daryl E. Chubin. Misconduct in research: an issue of science policy and practice. Minerva, 23(2):175–202, Summer 1985. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01099941.

Cai:2022:THQ

[CL22] Yuzhuo Cai and Annina Lattu. Triple helix or quadruple helix: Which model of innovation to choose for empirical studies? Minerva, 60(2):257–280, June 2022. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL https://link.

springer.com/article/10.1007/s11024-021-09453-6.

Clark:1970:SCS

[Cla70a] Ronalde W. Clark. Scientists and civil servants: The struggle over the national physical laboratory in 1918. Minerva, 8(1–4): 148–150, January 1970. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01553207.

Clark:1970:RDF

[Cla70b] Terry N. Clark. The rise and decline of France as a scientific centre. Minerva, 8(1–4):599–601, January 1970. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01553243;   http://link.springer.com/article/
10.1007/BF01553243.

Clark:1974:EAC

[Cla74] Ian D. Clark. Expert advice in the controversy about supersonic transport in the United States. Minerva, 12(4):416–432, October 1974. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01557638.

Clark:1995:RCV

[Cla95] C. A. Clark. The response of the Committee of Vice-Chancellors and Principals to the Department for Education’s Review of Higher Education. Minerva, 33(3):279–289, Autumn 1995. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01117504.

Clarke:2008:MM

[Cla08] Steve Clarke. Moral minds. Minerva, 46(1):147–150, March 2008. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-007-9080-9.

Cournand:1976:SC

[CM76] André Cournand and Michael Meyer. The scientist’s code. Minerva, 14(1):79–96, Spring 1976. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01096215.

Carlson:1995:ISM

[CMR95] Timothy Carlson and Dominique Martin-Rovet. The implications of scientific mobility between France and the United States. Minerva, 33(3):211–250, Autumn 1995. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01117501.

Coben:1976:FOF

[Cob76] Stanley Coben. Foundation officials and fellowships: Innovation in the patronage of science. Minerva, 14(2):225–240, Summer 1976. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01107887.

Coccia:2009:BPR

[Coc09] Mario Coccia. Bureaucratization in public research institutions. Minerva, 47(1):31–50, March 2009. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/content/pdf/10.1007/s11024-
008-9113-z.pdf.

Coenen:2006:LSS

[Coe06] Reinhard Coenen. The legacy of the studiengruppe Für systemforschung. Minerva, 44(2):143–148, June 2006. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
s11024-006-0003-y.

Cohn:1968:UT

[Coh68] Edwin J. Cohn. Universities in Turkey. Minerva, 6(4):607– 609, Summer 1968. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096554.

Cohen:1998:BRO

[Coh98] Robin Cohen. Book review: Open the Social Sciences: Report of the Gulbenkian Commission on the Restructuring of the Social Sciences, edited by Immanuel Wallerstein. Minerva, 36 (2):198–200, June 1998. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1004352221670.

Coing:1981:RHF

[Coi81] Helmut Coing. Remarks on the history of foundations and their role in the promotion of learning. Minerva, 19(2):271–281, Summer 1981. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096568.

Coleman:1986:IDU

[Col86] James S. Coleman. The idea of the developmental university. Minerva, 24(4):476–494, Winter 1986. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01096709.

Cole:1998:HDP

[Col98] Stephen Cole. How does peer review work and can it be improved? Minerva, 36(2):179–189, June 1998. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1023/
A%3A1004348103015.

Collier:2000:DOB

[Col00] James H. Collier. Divining the oracle of big science: Steps on the path to a new republicanism. Minerva, 38(1):109–120, ???? 2000. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1026553810453.

Colyvas:2007:FHC

[Col07] Jeannette A. Colyvas. Factory, hazard, and contamination: The use of metaphor in the commercialization of recombinant DNA. Minerva, 45(2):143–159, June 2007. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
007-9036-0.

Conia:1976:CAA

[Con77] Jean-Marie Conia. Criteria of academic appointment. Minerva, 14(4):570–574, Winter 1976–1977. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01097639.

Connelly:1996:IBE

[Con96] John Connelly. Internal bolshevisation? Elite social science training in Stalinist Poland. Minerva, 34(4):323–346, ???? 1996. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00127070.

Cooper:1967:RUC

[Coo67a] C. M. Cooper. Research in an underdeveloped country. Minerva, 5(4):585–587, Summer 1967. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01096791.

Cooper:1967:BUP

[Coo67b] W. Mansfield Cooper. British universities and the public purse. Minerva, 6(1):99–102, Autumn 1967. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01101609.

Cooper:1972:UUU

[Coo72] William Mansfield Cooper. The urban university and its urban environment. Minerva, 10(1):158–160, January 1972. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01881393.

Cope:2002:PTA

[Cop02] David Cope. Parliaments and technology assessment. Minerva, 40(4):421–424, ???? 2002. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1020969503706.

Court:1979:ISS

[Cou79] David Court. The idea of social science in East Africa: an aspect of the development of higher education. Minerva, 17(2): 244–282, Summer 1979. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096283.

Cox:1994:PCF

[Cox94] Caroline Cox. “Political correctness” and freedom of speech in British universities. Minerva, 32(2):193–195, Summer 1994. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01103359.

Cole:1999:SPN

[CP99] Stephen Cole and Thomas J. Phelan. The scientific productivity of nations. Minerva, 37(1):1–23, March 1999. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1023/A%3A1004591413630.

Cramer:2025:ERI

[CR25] Katharina C. Cramer and Nicolas V. Rüffin. The EUropeanisation of research infrastructure policy. Minerva, 63(1):45– 68, March 2025. CODEN MINEFY. ISSN 0026-4695 (print),

1573-1871 (electronic). URL https://link.springer.com/
article/10.1007/s11024-024-09544-0.

Crawford:1966:TCR

[Cra66] Malcolm Crawford. Thoughts on chemical research and teaching in East Africa. Minerva, 4(2):170–185, Winter 1966. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01584853.

Crawford:1967:GUE

[Cra67a] Malcolm Crawford. Government and the universities in East Africa: I — on academic freedom in Africa. Minerva, 5(3): 376–381, Spring 1967. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF02388576.

Crawford:1967:AFA

[Cra67b] Malcolm Crawford. On academic freedom in Africa. Minerva, 6 (1):116–118, Autumn 1967. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01101613.

Craver:1986:PDR

[Cra86] Earlene Craver. Patronage and the directions of research in economics: The Rockefeller Foundation in Europe, 1924–1938. Minerva, 24(2–3):205–222, Summer–Autumn 1986. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01096611.

Crosland:1970:RDF

[Cro70] Maurice Crosland. The rise and decline of France as a scientific centre. Minerva, 8(1–4):453–454, January 1970. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01553231;   http://link.springer.com/article/
10.1007/BF01553231.

Crosland:1975:DPC

[Cro75] Maurice Crosland. The development of a professional career in science in France. Minerva, 13(1):38–57, Spring 1975. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096241.

Crosland:1978:FAS

[Cro78] Maurice Crosland. The French Academy of Sciences in the nineteenth century. Minerva, 16(1):73–102, Spring 1978. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01102182.

Crosland:1979:PGS

[Cro79] Maurice Crosland. From prizes to grants in the support of scientific research in France in the nineteenth century: The Montyon legacy. Minerva, 17(3):355–380, Autumn 1979. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01096220.

Crosland:1981:SCR

[Cro81] Maurice Crosland. Scientific credentials: Record of publications in the assessment of qualifications for election to the French Académie des sciences. Minerva, 19(4):605–631, Winter 1981. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096195.

Crosland:1986:APN

[Cro86] Maurice Crosland. Assessment by peers in nineteenth-century France: The manuscript reports on candidates for election to the Académie des sciences. Minerva, 24(4):413–432, Winter 1986. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096705.

Croham:1994:RTB

[Cro94] Douglas Croham. The restoration of trust between government and universities. Minerva, 32(2):196–200, Summer 1994. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01103360.

Carr-Saunders:1963:SAU

[CS63] A. M. Carr-Saunders. Staffing African universities. Minerva, 1 (3):302–318, Spring 1963. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF02251986.

Choi:2015:PMS

[CS15] Hyungsub Choi and Brit Shields. A place for materials science: Laboratory buildings and interdisciplinary research at the University of Pennsylvania. Minerva, 53(1):21–42, March 2015. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-015-9265-6.

Cooper:1974:BR

[CSGZ^{+}74] William Mansfield Cooper, Brigitte Schroeder-Gudehus, Harriet Zuckerman, Richard Mayne, F. R. Jevons, and Yehuda Elkana. Book reviews. Minerva, 12(1):124–149, January 1974. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01558545.

Caston:1985:AFP

[CSKF85] Geoffrey Caston, E. S., Keith, and S. G. Fleet. Academic freedom and permanent tenure in academic appointments. Minerva, 23(1):96–150, Spring 1985. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01097842.

Cantwell:2013:GSI

[CT13] Brendan Cantwell and Barrett J. Taylor. Global status, intrainstitutional stratification and organizational segmentation: a time-dynamic tobit analysis of ARWU position among U.S. universities. Minerva, 51(2):195–223, June 2013. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
s11024-013-9228-8.

Chang:2022:EOK

[CTLL22] Yuan-Chieh Chang, Tung-Fei Tsai-Lin, and Tian Liang. Entrepreneurial orientation and knowledge transfer effectiveness: The effect of organizational commercial slack. Minerva, 60(3): 441–462, September 2022. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL https://link.springer.

com/article/10.1007/s11024-022-09463-y.

Cueto:1997:SUA

[Cue97] Marcos Cueto. Science under adversity: Latin American medical research and American private philanthropy, 1920– 1960. Minerva, 35(3):233–245, September 1997. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1023/
A%3A1004230000979.

Cummings:1973:JPU

[Cum73] William K. Cummings. The Japanese private University. Minerva, 11(3):348–371, July 1973. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01557689.

Cummings:1990:CES

[Cum90] William K. Cummings. The culture of effective science: Japan and the United States. Minerva, 28(4):426–445, Winter 1990. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096650.

Carter:1964:GSP

[CW64] C. F. Carter and B. R. Williams. Government scientific policy and the growth of the British economy. Minerva, 3(1): 114–125, Autumn 1964. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01630154.

Dahrendorf:1993:TNH

[DA93] Ralf Dahrendorf and Stefan Amsterdamski. Transformation of the national higher education and research systems of central Europe. Minerva, 31(4):504–522, Winter 1993. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01096452.

Daalder:1974:DUB

[Daa74] Hans Daalder. The Dutch universities between the “new democracy” and the “new management”. Minerva, 12(2): 221–257, April 1974. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01553179.

Dannatt:1967:BIR

[Dan67] R. J. Dannatt. Books, information and research: Libraries for technological universities. Minerva, 5(2):209–226, Winter 1967. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02388525.

DasGupta:1964:RHE

[Das64] A. K. Das Gupta. Reflections on higher education in India in the light of the Robbins Report. Minerva, 2(2):160–168, Winter 1964. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096592.

Davies:2001:BHE

[Dav01] John L. Davies. Borderless higher education in continental Europe. Minerva, 39(1):27–48, ???? 2001. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1023/
A%3A1010370201943.

Davies:2020:ELS

[Dav20] Sarah R. Davies. Epistemic living spaces, international mobility, and local variation in scientific practice. Minerva, 58(1):97–114, March 2020. CODEN MINEFY. ISSN 0026-4695 (print), 1573-

1871 (electronic). URL http://link.springer.com/article/
10.1007/s11024-019-09387-0.

Dupree:1977:BR

[DBDM77] A. Hunter Dupree, Joseph Ben-Dravid, and P. B. Medawar. Book reviews. Minerva, 15(1):94–114, Spring 1977. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01096322.

Davis:1992:BR

[DBL^{+}92] Bernard D. Davis, Jack N. Barkenbus, Grahame Lock, Roy Porter, Malcolm Richardson, and Norriss S. Hetherington. Book reviews. Minerva, 30(1):101–126, Spring 1992. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01096398.

DeMoortel:2022:DIN

[DCJ22] Kevin De Moortel, Thomas Crispeels, and Qiaosong Jing. Do interpersonal networks mediate the relationship between international academic mobility and entrepreneurial knowledge? Minerva, 60(1):29–55, March 2022. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL https://link.

springer.com/article/10.1007/s11024-021-09448-3.

D:1987:IMS

[DDS^{+}87] Wolfgang Epstein M. D., Harry Fozzard M. D., Bernard Roizman Sc.D., E. H. Uhlenhuth M. D., and Christopher Zarins M. D. II the morality of scientists. Minerva, 25(3):358–361, Autumn 1987. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01097789.

DeSilva:1978:UGS

[De 78] Kingsley De Silva. The universities and the government in Sri Lanka. Minerva, 16(2):251–272, Summer 1978. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01096016.

DeRudder:1997:TEG

[De 97] Helmut De Rudder. The transformation of East German higher education: Renewal as adaptation, integration and innovation. Minerva, 35(2):99–125, June 1997. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1023/
A%3A1004228410448.

Debiter:1968:WWH

[Deb69] Stevan Debiter. Wanted: a world handbook of research and development statistics. Minerva, 7(1–2):88–91, Autumn– Winter 1968–1969. CODEN MINEFY. ISSN 0026-4695 (print),

1573-1871 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF01096813.

Dedijer:1963:USU

[Ded63] Stevan Dedijer. Underdeveloped science in underdeveloped countries. Minerva, 2(1):61–81, Autumn 1963. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01100781.

Dedijer:1966:SSP

[Ded66] Stevan Dedijer. The science of science: a programme and a plea. Minerva, 4(4):489–504, Summer 1966. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF02207980.

Deich:1979:RAN

[Dei79] Ingrid Deich. The redistribution of authority in national laboratories in Western Germany. Minerva, 17(3):413–444, Autumn 1979. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096222.

DeJager:1993:PSP

[DeJ93] Timothy DeJager. Pure science and practical interests: The origins of the agricultural research council, 1930–1937. Minerva, 31 (2):129–150, Summer 1993. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01125432.

Demortain:2017:ERS

[Dem17] David Demortain. Expertise, regulatory science and the evaluation of technology and risk: Introduction to the special issue. Minerva, 55(2):139–159, June 2017. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

Deshmukh:1965:DKI

[Des65] C. D. Deshmukh. Dr. karve and the Indian universities. Minerva, 3(3):398–401, Spring 1965. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01099964.

Desrosieres:2009:HRC

[Des09] Alain Desrosières. How to be real and conventional: a discussion of the quality criteria of official statistics. Minerva, 47 (3):??, September 2009. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-009-9125-3.

Deuraseh:2007:ITI

[Deu07] Nurdeen Deuraseh. Islamic–technoscientific identities. Minerva, 45(1):97–99, March 2007. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-006-9025-8.

Devorkin:1980:MSI

[Dev80] David H. Devorkin. The maintenance of a scientific institution: Otto Struve, the Yerkes Observatory, and its optical bureau during the Second World War. Minerva, 18(4):595–623, Winter 1980. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096126.

Domes:1975:TFU

[DF75] Jürgen Domes and Armin Paul Frank. The tribulations of the free University of Berlin. Minerva, 13(2):183–199, Summer 1975. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01097794.

David-Fox:1996:SPE

[DF96] Michael David-Fox. Science, political enlightenment and agitprop: On the typology of social knowledge in the early Soviet period. Minerva, 34(4):347–366, ???? 1996. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF00127071.

deGroot:2021:CSB

[dGBM^{+}21] Esther de Groot, Yvette Baggen, Nienke Moolenaar, Diede Stevens, Jan van Tartwijk, Roger Damoiseaux, and Manon Kluijtmans. Clinician–scientists in-and-between research and practice: How social identity shapes brokerage. Minerva, 59(1):123–137, March 2021. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http:/

/link.springer.com/article/10.1007/s11024-020-09420-
7;   http://link.springer.com/content/pdf/10.1007/s11024-
020-09420-7.pdf.

deGaulle:1969:IHE

[dGdMS^{+}69] Charles de Gaulle, Maurice Couve de Murville, Maurice Schumann, François Ortoli, Edgar Faure, and Robert Galley. II the higher education law of 12 November, 1968. Minerva, 7(4): 712–727, Summer 1969. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01099543.

Dabos:2019:PSS

[DGR19] Marcelo P. Dabós, Ernesto R. Gantman, and Carlos J. Fernández Rodríguez. The prestige of social scientists in Spain and France: an examination of their h-index values using Scopus and Google Scholar. Minerva, 57(1):47–66, March 2019. CODEN MINEFY.

ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/article/10.1007/s11024-018-9358-0.

Dedijer:1964:USP

[DH64] Stevan Dedijer and Guy Hunter. The unity of scientific policy: dvazhcy dva = two times two = 2 × 2. Minerva, 3(1): 126–130, Autumn 1964. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01630155.

Dhanagare:1984:EIU

[Dha84] D. N. Dhanagare. Equality in the Indian university. Minerva, 22 (3–4):388–403, Autumn–Winter 1984. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF02207371.

Dietrich:2005:TMW

[Die05] Michael R. Dietrich. Two men who invented genetics. Minerva, 43(4):429–433, December 2005. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
005-2468-5.

Diner:1975:DDD

[Din75] Steven J. Diner. Department and discipline: The department of sociology at the University of Chicago, 1892–1920. Minerva, 13 (4):514–553, Winter 1975. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096176.

Divall:1991:FSV

[Div91] Colin Divall. Fundamental science versus design: Employers and engineering studies in British universities, 1935–1976. Minerva, 29(2):167–194, Summer 1991. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01096407.

Divall:1994:POE

[Div94] Colin Divall. Professional organisation, employers and the education of engineers for management: a comparison of mechanical, electrical and chemical engineers in Britain, 1897–1977. Minerva, 32(3):241–266, Autumn 1994. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01098662.

Divall:1998:BRM

[Div98] Colin Divall. Book review: The Missing Stratum: Technical School Education in England, 1900–1990s, by Michael Sanderson. Minerva, 36(1):91–98, March 1998. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1023/
A%3A1004303108087.

Davis:1974:NRF

[DK74] Lance E. Davis and Daniel J. Kevles. The national research fund: a case study in the industrial support of academic science. Minerva, 12(2):207–220, April 1974. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01553178.

DiFiore:2023:CQI

[DKCL^{+}23] Monica Di Fiore, Marta Kuc-Czarnecka, Samuele Lo Piano, Arnald Puy, and Andrea Saltelli. The challenge of quantification: an interdisciplinary reading. Minerva, 61(1):53–70, March 2023. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL https://link.springer.com/article/10.
1007/s11024-022-09481-w.

Dresselhaus:1992:C

[DKM^{+}92] M. S. Dresselhaus, Clark Kerr, Walter E. Massey, John Roberts, and Charles H. Townes. Comments. Minerva, 30(2):148– 162, Summer 1992. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096879.

delaMothe:2003:RTP

[dlM03] John de la Mothe. Re-thinking policy in the New Republic of Knowledge. Minerva, 41(3):195–205, ???? 2003. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1023/A%3A1025561613229.

Dean:1973:SIC

[DM73] Genevieve Dean and Manfredo Macioti. Scientific institutions in China. Minerva, 11(3):318–334, July 1973. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01557687.

Dobrov:1966:PDS

[Dob66] G. M. Dobrov. Predicting the development of science. Minerva, 4(2):218–230, Winter 1966. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01584857.

Doern:1970:BSG

[Doe70] G. Bruce Doern. “Big science”, government and the scientific community in Canada: The ING affair. Minerva, 8(1–4): 357–375, January 1970. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01553223;   http://
link.springer.com/article/10.1007/BF01553223.

Dolan:2005:EV

[Dol05] Brian Dolan. Encyclopedic visions. Minerva, 43(1):87–98, ???? 2005. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-004-6631-1.

Dongerkery:1971:BC

[Don71] S. R. Dongerkery. Bombay colleges. Minerva, 9(2):294–297, April 1971. CODEN MINEFY. ISSN 0026-4695 (print),

1573-1871 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF01553163.

Donnison:1972:RP

[Don72] David Donnison. Research for policy. Minerva, 10(4):519– 536, October 1972. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01695904.

Doty:1972:CII

[Dot72] Paul Doty. Can investigations improve scientific advice? The case of the ABM. Minerva, 10(2):280–294, April 1972. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01682422.

Douglass:2003:SMH

[Dou03] John Aubrey Douglass. Scanning the market horizon: Educational futures in historical perspective. Minerva, 41(4):381– 395, ???? 2003. CODEN MINEFY. ISSN 0026-4695 (print),

1573-1871 (electronic). URL http://link.springer.com/
accesspage/article/10.1023/B%3AMINE.0000005160.65307.
59.

Downer:2017:APW

[Dow17] John Downer. The aviation paradox: Why we can ‘know’ jetliners but not reactors. Minerva, 55(2):229–248, June 2017. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). See erratum [Dow18].

Downer:2018:EAP

[Dow18] John Downer. Erratum to: The aviation paradox: Why we can ‘know’ jetliners but not reactors. Minerva, 56(2):259, June 2018. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s11024-017-9334-0;   http://link.springer.com/content/
pdf/10.1007/s11024-017-9334-0.pdf. See [Dow17].

Dusdal:2020:UVR

[DPB^{+}20] Jennifer Dusdal, Justin J. W. Powell, David P. Baker, Yuan Chih Fu, Yahya Shamekhi, and Manfred Stock. University vs. research institute? the dual pillars of German science production, 1950–2010. Minerva, 58(3):319–342, September 2020. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/article/10.1007/s11024-
019-09393-2;   http://link.springer.com/content/pdf/
10.1007/s11024-019-09393-2.pdf.

Dietz:2012:MPI

[DR12] James S. Dietz and Juan D. Rogers. Meanings and policy implications of “transformative research”: Frontiers, hot science, evolution, and investment risk. Minerva, 50(1):21–44, March 2012. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-012-9190-x.

Drake:1984:RUA

[Dra84] Keith Drake. The recovery of university autonomy in Great Britain. Minerva, 22(3–4):346–364, Autumn–Winter 1984. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF02207369.

Delicado:2014:WRS

[DRC^{+}14] Ana Delicado, Raquel Rego, Cristina Palma Conceição, Inês Pereira, and Luís Junqueira. What roles for scientific associations in contemporary science? Minerva, 52(4):439–465, December 2014. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-014-9260-3.

Drever:1971:RD

[Dre71] James Drever. Reflections on a debate. Minerva, 9(3): 406–410, July 1971. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01553364.

DSouza:2006:GL

[D’S06] Rohan D’Souza. Global learning. Minerva, 44(2):235–237, June 2006. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-006-0009-5.

Derrick:2016:ESE

[DS16] Gemma E. Derrick and Gabrielle N. Samuel. The evaluation scale: Exploring decisions about societal impact in peer review panels. Minerva, 54(1):75–97, March 2016. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s11024-016-9290-0.pdf.

Neto:2023:PST

[dSNC23] Victo José da Silva Neto and Tulio Chiarini. The platformization of science: Towards a scientific digital platform taxonomy. Minerva, 61(1):1–29, March 2023. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL https://link.

springer.com/article/10.1007/s11024-022-09477-6.

Delvenne:2012:BCC

[DT12] Pierre Delvenne and François Thoreau. Beyond the “charmed Circle” of OECD: New directions for studies of national innovation systems. Minerva, 50(2):205–219, June 2012. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/s11024-012-9195-5.

Dubarle:1963:PPS

[Dub63] D. Dubarle. The proper public of science: Reflections on a Cartesian theme concerning humanity and the state as audiences of the scientific community. Minerva, 1(4):405–427, Summer 1963. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01107185.

Dubarle:1965:SRC

[Dub65] D. Dubarle. Some remarks on “criteria for scientific choice: II”. Minerva, 3(2):261–264, Winter 1965. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01096339.

Duff:1966:EHE

[Duf66] James Duff. The ecology of higher education. Minerva, 5(1): 39–46, Autumn 1966. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF02389792.

Dupree:1963:CSO

[Dup63] A. Hunter Dupree. Central scientific organisation in the United States government. Minerva, 1(4):453–469, Summer 1963. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01107188.

Dupree:1984:BR

[Dup84] A. Hunter Dupree. Book reviews. Minerva, 22(1):116–127, Spring 1984. CODEN MINEFY. ISSN 0026-4695 (print),

1573-1871 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF02207559.

Dupree:1990:SPU

[Dup90] A. Hunter Dupree. Science policy in the United States: The legacy of John Quincy Adams. Minerva, 28(3):259–271, Autumn 1990. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096291.

Dagnino:1998:UIG

[DV98] Renato Dagnino and Léa Velho. University–industry–government relations on the periphery: The University of Campinas, Brazil. Minerva, 36(3):229–251, September 1998. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1023/
A%3A1004335804375.

Eamon:1985:SNP

[Eam85] William Eamon. From the secrets of nature to public knowledge: The origins of the concept of openness in science. Minerva, 23 (3):321–347, Autumn 1985. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096442.

Eriksson:2022:AUM

[EB22] Lena Eriksson and Helena Bergman. Acceptable use: Morality and credibility struggles in Swedish 1960s alcohol and illicit drug (ab)use research and policy. Minerva, 60(3):419–440, September 2022. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL https://link.springer.com/article/10.
1007/s11024-022-09462-z.

Entradas:2024:CFU

[EBMP24] Marta Entradas, Martin W. Bauer, Frank Marcinkowski, and Giuseppe Pellegrini. The communication function of universities: Is there a place for science communication? Minerva, 62 (1):25–47, March 2024. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL https://link.springer.

com/article/10.1007/s11024-023-09499-8.

Eckart:2004:STR

[Eck04] Wolfgang Uwe Eckart. Science in the Third Reich: Margit Szöllösi-Janze (ed.), Science in the Third Reich (German Historical Perspectives, XII) (Oxford/New York: Berg, 2001), 289 pp., ISBN: 1-85973-421-9. Minerva, 42(4):451–454, ???? 2004. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-004-3002-x.

Eisemon:1991:CQS

[ED91] Thomas O. Eisemon and Charles H. Davis. Can the quality of scientific training and research in Africa be improved? Minerva, 29(1):1–26, Spring 1991. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096296.

Edlund:2024:MTE

[Edl24] Peter Edlund. More than euros: Exploring the construction of project grants as prizes and consolations. Minerva, 62(1):1– 23, March 2024. CODEN MINEFY. ISSN 0026-4695 (print),

1573-1871 (electronic). URL https://link.springer.com/
article/10.1007/s11024-023-09501-3.

Edqvist:2003:LSS

[Edq03] Olle Edqvist. Layered science and science policies. Minerva, 41 (3):207–221, ???? 2003. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1025509629159.

Education:1966:IUR

[Edu66] Higher Education. Indian university reform. Minerva, 5(1): 47–81, Autumn 1966. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF02389793.

Etzkowitz:2006:WSF

[EG06] Henry Etzkowitz and Namrata Gupta. Women in science: a fair shake? Minerva, 44(2):185–199, June 2006. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
006-0007-7.

Eisemon:1979:ISN

[Eis79] Thomas Owen Eisemon. The implantation of science in Nigeria and Kenya. Minerva, 17(4):504–526, Winter 1979. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01101901.

Eisenstadt:1987:CSK

[Eis87] S. N. Eisenstadt. The classical sociology of knowledge and beyond. Minerva, 25(1–2):77–91, Spring–Summer 1987. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01096857.

Eisler:2013:EUS

[Eis13] Matthew N. Eisler. “The ennobling unity of science and technology”: Materials sciences and engineering, the Department of Energy, and the nanotechnology enigma. Minerva, 51(2): 225–251, June 2013. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-013-9224-z.

Etzkowitz:1998:RRC

[EK98] Henry Etzkowitz and Carol Kemelgor. The role of research centres in the collectivisation of academic science. Minerva, 36(3): 271–288, September 1998. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1004348123030.

Etzkowitz:2001:GIS

[EK01] Henry Etzkowitz and Carol Kemelgor. Gender inequality in science: a universal condition? Minerva, 39(2):239–257, ???? 2001. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1010337925943.

Edler:2022:UCI

[EKB22] Jakob Edler, Maria Karaulova, and Katharine Barker. Understanding conceptual impact of scientific knowledge on policy: The role of policymaking conditions. Minerva, 60(2):209– 233, June 2022. CODEN MINEFY. ISSN 0026-4695 (print),

1573-1871 (electronic). URL https://link.springer.com/
article/10.1007/s11024-022-09459-8.

Edlund:2022:SIA

[EL22] Peter Edlund and Inti Lammi. Stress-inducing and anxietyridden: a practice-based approach to the construction of statusbestowing evaluations in research funding. Minerva, 60(3):397– 418, September 2022. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL https://link.springer.

com/article/10.1007/s11024-022-09466-9.

Elkana:1977:DUS

[Elk77] Yehuda Elkana. The distinctiveness and universality of science: Reflections on the work of Professor Robin Horton. Minerva, 15 (2):155–173, Summer 1977. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096675.

Elmgren:2024:EPD

[ELSS24] Maja Elmgren, Åsa Lindberg-Sand, and Anders Sonesson. Evaluation practices of doctoral examination committees: Boundarywork under pressure. Minerva, 62(3):427–456, September 2024. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL https://link.springer.com/article/10.
1007/s11024-024-09523-5.

Elvin:1968:HU

[Elv68] Lionel Elvin. Hands off the universities? Minerva, 6(4): 603–607, Summer 1968. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096553.

Elzinga:2012:RDI

[Elz12] Aant Elzinga. The rise and demise of the International Council for Science Policy Studies (ICSPS) as a Cold War bridging organization. Minerva, 50(3):277–305, September 2012. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-012-9205-7.

Encel:1970:PSS

[Enc70] Sol Encel. The problems of a small scientific community: The Norwegian case. Minerva, 8(1–4):295–297, January 1970. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01553217;   http://link.springer.com/article/
10.1007/BF01553217.

Encel:1971:SSS

[Enc71] S. Encel. The support of science without science policy in Australia. Minerva, 9(3):349–360, July 1971. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01553360.

Encel:2000:UDM

[Enc00] Sol Encel. Universities do matter: Australian universities in crisis. Minerva, 38(2):241–251, ???? 2000. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1023/
A%3A1026509413392.

Encel:2003:SRU

[Enc03] Sol Encel. Student radicalism at the University of Sydney. Minerva, 41(4):415–419, ???? 2003. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1023/
B%3AMINE.0000005289.45571.4a.

Encel:2006:UC

[Enc06] Sol Encel. A University off course? Minerva, 44(2):201– 208, June 2006. CODEN MINEFY. ISSN 0026-4695 (print),

1573-1871 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11024-006-0005-9.

Encel:2007:VVC

[Enc07] Sol Encel. Vicissitudes of a Vice-Chancellor. Minerva, 45(4): 475–479, December 2007. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-007-9063-x.

Endersby:2006:WR

[End06] Jim Endersby. Wallace redux? Minerva, 44(2):209–218, June 2006. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-006-0006-8.

Enders:2009:BRR

[End09] Jürgen Enders. Book review: Richard Whitley, Jochen Gläser (eds.), The Changing Governance of the Sciences. The Advent of Research Evaluation Systems. Sociology of the Sciences Yearbook. Minerva, 47(4):465–468, December 2009. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s11024-009-9132-4.pdf.

Engerman:1999:NSN

[Eng99] David C. Engerman. New society, new scholarship: Soviet studies programmes in interwar America. Minerva, 37(1): 25–43, March 1999. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1004599118173.

Engerman:2024:BRM

[Eng24] David C. Engerman. Book review: Mark Solovey and Christian Dayé, eds., Cold War Social Science: Transnational Entanglements. Minerva, 62(3):481–483, September 2024. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL https://link.springer.com/article/10.
1007/s11024-024-09540-4.

Etzkowitz:1991:PKO

[EP91] Henry Etzkowitz and Lois S. Peters. Profiting from knowledge: Organisational innovations and the evolution of academic norms. Minerva, 29(2):133–166, Summer 1991. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01096406.

Epstein:1986:IMS

[Eps86] Richard A. Epstein. II the morality of scientists. Minerva, 24(2– 3):344–347, Summer–Autumn 1986. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01096615.

Es:1977:SSP

[Es77] Es. Social science as public opinion. Minerva, 15(3–4):273– 285, Autumn–Winter 1977. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096725.

Epstein:1988:AFA

[ESC^{+}88] Joseph Epstein, Professor Carol Simpson Stern, Professor Buckley Christ, Jr., Professor Richard Hughes, Professor Ennio Rossi, and Professor Addison Stone. Academic freedom and academic agitation at Northwestern University. Minerva, 26(2): 199–272, Summer 1988. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096696.

Esterle:2004:PDN

[Est04] Laurence Esterle. Participative democracy: a new form of rationality. Minerva, 42(1):101–104, ???? 2004. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1023/
B%3AMINE.0000017837.79006.68.

Elvin:1967:C

[ET67] Lionel Elvin and John Turkevich. Correspondence. Minerva, 5 (3):429–434, Spring 1967. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF02388581.

Eisaku:1970:CSD

[ETMT70] Satō Eisaku, Fukuda Takeo, Sakata Michita, and Noda Takeo. Coping with student disorder in Japan. Minerva, 8(1–4): 116–135, January 1970. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01553203.

Etzkowitz:1983:ESE

[Etz83] Henry Etzkowitz. Entrepreneurial scientists and entrepreneurial universities in American academic science. Minerva, 21(2–3): 198–233, Summer–Autumn 1983. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01097964.

Etzkowitz:1992:IIT

[Etz92] Henry Etzkowitz. Individual investigators and their research groups. Minerva, 30(1):28–50, Spring 1992. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01096395.

Etzkowitz:1993:ESO

[Etz93] Henry Etzkowitz. Enterprises from science: The origins of science-based regional economic development. Minerva, 31(3): 326–360, Autumn 1993. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01098626.

Etzkowitz:1994:KPM

[Etz94] Henry Etzkowitz. Knowledge as property: The Massachusetts Institute of Technology and the debate over academic patent policy. Minerva, 32(4):383–421, Winter 1994. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01098031.

Etzkowitz:1996:CIC

[Etz96] Henry Etzkowitz. Conflicts of interest and commitment in academic science in the United States. Minerva, 34(3):259–277, ???? 1996. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00120327.

Eustace:1995:CDC

[Eus95] Rowland Eustace. A comment on the discussion of Conrad Russell’s academic freedom. Minerva, 33(1):67–73, Spring 1995. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01098620.

Eustace:1997:BR

[Eus97] Rowland Eustace. Book review. Minerva, 35(2):195–201, June 1997. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1004258225939.

Evans:2012:ECT

[EV12] Samuel A. W. Evans and Walter D. Valdivia. Export controls and the tensions between academic freedom and national security. Minerva, 50(2):169–190, June 2012. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
012-9196-4.

Ezrahi:1980:UPR

[Ezr80] Yaron Ezrahi. Utopian and pragmatic rationalism: The political context of scientific advice. Minerva, 18(1):111–131, Spring 1980. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096662.

Fisher:2001:CAI

[FAGH01] Donald Fisher, Janet Atkinson-Grosjean, and Dawn House. Changes in academy/industry/state relations in Canada: The creation and development of the networks of Centres of Excellence. Minerva, 39(3):299–325, ???? 2001. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1023/
A%3A1017924027522.

Falkenberg:2021:RIY

[Fal21] Ruth I. Falkenberg. Re-invent yourself! How demands for innovativeness reshape epistemic practices. Minerva, 59(4):423– 444, December 2021. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL https://link.springer.

com/article/10.1007/s11024-021-09447-4.

Farrington:2001:BHE

[Far01] Dennis J. Farrington. Borderless higher education: Challenges to regulation, accreditation and intellectual property rights. Minerva, 39(1):63–84, ???? 2001. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1023/
A%3A1010322318781.

Fara:2005:NIE

[Far05] Patricia Fara. Newton, industry, and empire. Minerva, 43(4): 435–439, December 2005. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-005-2465-8.

Frodeman:2012:DPR

[FB12] Robert Frodeman and Adam Briggle. The dedisciplining of peer review. Minerva, 50(1):3–19, March 2012. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
012-9192-8.

Fernandez:2021:SAP

[FBF^{+}21] Frank Fernandez, David P. Baker, Yuan-Chih Fu, Ismael G. Muñoz, and Karly Sarita Ford. A symbiosis of access: Proliferating STEM PhD training in the U.S. from 1920–2010. Minerva, 59(1):79–98, March 2021. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/article/10.1007/s11024-020-09422-5.

Faria:2024:SIC

[FBGGH24] João Ricardo Faria, Christopher J. Boudreaux, Rajeev K. Goel, and Devrim Göktepe-Hultén. Science and innovation: a cyclical approach. Minerva, 62(4):593–610, December 2024. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL https://link.springer.com/article/10.
1007/s11024-024-09531-5.

Franssen:2023:SDR

[FBH23] Thomas Franssen, Siri Brorstad Borlaug, and Anders Hylmö. Steering the direction of research through organizational identity formation. Minerva, 61(4):495–519, December 2023. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (elec-

tronic).  URL https://link.springer.com/article/10.
1007/s11024-023-09494-z.

Feldman:2004:TOS

[FD04] Maryann P. Feldman and Pierre Desrochers. Truth for its own sake: Academic culture and technology transfer at Johns Hopkins University. Minerva, 42(2):105–126, ???? 2004. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/B%3AMINE.0000030019.99709.a0.

Feller:2009:PMG

[Fel09] Irwin Feller. Performance measurement and the governance of American academic science. Minerva, 47(3):??, September 2009. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-009-9129-z.

Feld:2019:SPP

[Fel19] Adriana Feld. Science, politics/policy and the Cold War in Argentina: From concepts to institutional models in the 1950s and ’60s. Minerva, 57(4):523–547, December 2019. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/article/10.1007/s11024-
019-09379-0.

Ferpozzi:2023:PPP

[Fer23] Hugo Ferpozzi. Public-private partnerships and the landscape of neglected tropical disease research: The shifting logic and spaces of knowledge production. Minerva, 61(4):607–629, December 2023. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL https://link.springer.com/article/10.
1007/s11024-023-09496-x.

Friedman:1990:CUP

[FF90] Robert S. Friedman and Renee C. Friedman. The Canadian universities and the promotion of economic development. Minerva, 28(3):272–293, Autumn 1990. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01096292.

Felt:2010:MMP

[FF10] Ulrike Felt and Maximilian Fochler. Machineries for making publics: Inscribing and De-scribing publics in public engagement. Minerva, 48(3):219–238, September 2010. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/s11024-010-9155-x.

Foladori:2015:NMK

[FFL^{+}15] Guillermo Foladori, Edgar Arteaga Figueroa, Edgar Záyago Lau, Richard Appelbaum, Eduardo Robles-Belmont, Liliana Villa, Rachel Parker, and Vanessa Leos. Nanotechnology in Mexico: Key findings based on OECD criteria. Minerva, 53(3): 279–301, September 2015. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-015-9281-6.

Fochler:2016:UGH

[FFM16] Maximilian Fochler, Ulrike Felt, and Ruth Müller. Unsustainable growth, hyper-competition, and worth in life science research: Narrowing evaluative repertoires in doctoral and postdoctoral scientists’ work and lives. Minerva, 54(2):175–200, June 2016. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/content/pdf/
10.1007/s11024-016-9292-y.pdf.

Freeman:1995:BR

[FFML95] Christopher Freeman, Kenneth R. Foster, and Heinz Maier-Leibnitz. Book reviews. Minerva, 33(4):395–405, Winter 1995. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096520.

Fach:1992:ERT

[FG92] Wolfgang Fach and Edgar Grande. Emergent rationality in technological policy: Nuclear energy in the Federal Republic of Germany. Minerva, 30(1):14–27, Spring 1992. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01096394.

Feller:2007:SIR

[FG07] Irwin Feller and George Gamota. Science indicators as reliable evidence. Minerva, 45(1):17–30, March 2007. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
s11024-006-9017-8.

Falade:2020:DPS

[FG20] Bankole A. Falade and Lars Guenther. Dissonance and polyphasia as strategies for resolving the potential conflict between science and religion among South Africans. Minerva, 58(3):459– 480, September 2020. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/article/10.1007/s11024-020-09403-8.

Fenkner:2025:SSC

[FGR25] Tim Fenkner and Jonas Gottschalk-Rayling. Surviving science – coping with exit-decisions in physics and history. Minerva, 63(3): 449–469, September 2025. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL https://link.springer.

com/article/10.1007/s11024-024-09548-w.

Fox:1980:BR

[FGT80] Robert Fox, Geoffrey J. Giles, and Ernst Topitsch. Book reviews. Minerva, 18(1):164–195, Spring 1980. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01096664.

Fumasoli:2013:SAS

[FH13] Tatiana Fumasoli and Jeroen Huisman. Strategic agency and system diversity: Conceptualizing institutional positioning in higher education. Minerva, 51(2):155–169, June 2013. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/s11024-013-9225-y.

Floud:1994:BR

[FHE94] Jean Floud, Lillian Hoddeson, and Henry Etzkowitz. Book reviews. Minerva, 32(2):220–237, Summer 1994. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01103362.

Folker:2009:WGW

[FHS09] Anna Paldam Folker, Lotte Holm, and Peter Sandøe. ‘We have to go where the money is’ — dilemmas in the role of nutrition scientists: an interview study. Minerva, 47(2):217–236, June 2009. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-009-9122-6.

Fielden:2001:MBE

[Fie01] John Fielden. Markets for ‘borderless education’. Minerva, 39(1):49–62, ???? 2001. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1010374319689.

Finn:1976:FPU

[Fin77] Chester E. Finn, Jr. Federal patronage of universities in the United States: a rose by many other names? Minerva, 14(4): 496–529, Winter 1976–1977. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01097637.

Finlay:2007:ADG

[Fin07] Mark R. Finlay. Academic drift in German agricultural education. Minerva, 45(3):349–352, September 2007. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
s11024-007-9052-0.

Fischer:1963:USS

[Fis63] Joseph Fischer. The university student in South and South-East Asia. Minerva, 2(1):39–53, Autumn 1963. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01100779.

Fisher:1978:RFD

[Fis78] Donald Fisher. The Rockefeller Foundation and the development of scientific medicine in Great Britain. Minerva, 16 (1):20–41, Spring 1978. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01102180.

Fisher:2007:BSE

[Fis07] Donald Fisher. British sociology: an eyewitness account. Minerva, 45(2):235–240, June 2007. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
007-9037-z.

Fleck:2020:BRC

[Fle20] Christian Fleck. Book review: Christopher Adair-Toteff and Stephen Turner (eds.), The Calling of Social Thought: Rediscovering the Work of Edward Shils. Manchester University Press, Manchester, 2019, 270 pp., ISBN 978-1-5261-2005-2. Minerva, 58(3):481–483, September 2020. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/article/10.1007/s11024-020-09410-9.

Feuer:2010:SAP

[FM10] Michael J. Feuer and Christina J. Maranto. Science advice as procedural rationality: Reflections on the national research council. Minerva, 48(3):259–275, September 2010. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
s11024-010-9152-0.

Forman:1974:FSP

[For74] Paul Forman. The financial support and political alignment of physicists in Weimar Germany. Minerva, 12(1):39–66, January 1974. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01558540.

Fortescue:1992:RAS

[For92] Stephen Fortescue. The Russian Academy of Sciences and the Soviet Academy of Sciences: Continuity or disjunction? Minerva, 30(4):459–478, Winter 1992. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01096573.

Forstner:2008:EHD

[For08] Christian Forstner. The early history of David Bohm’s quantum mechanics through the perspective of Ludwik Fleck’s thought-collectives. Minerva, 46(2):215–229, June 2008. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/s11024-008-9090-2.

Foster:1997:RST

[Fos97] Kennith R. Foster. Risk, scientific testimony and the burden of proof: Science on trial. Minerva, 35(1):73–81, ???? 1997. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/B%3AMINE.0000019492.15173.65.

Fox:1973:SEP

[Fox73] Robert Fox. Scientific enterprise and the patronage of research in France 1800–70. Minerva, 11(4):442–473, October 1973. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01557797.

Fox:1997:PVA

[Fox97] Daniel M. Fox. Policy and vulnerability: American foundations in Twentieth-Century health affairs. Minerva, 35(3):311– 319, September 1997. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1004282017817.

Fox:2006:FDH

[Fox06] Robert Fox. Fashioning the discipline: History of science in the European intellectual tradition. Minerva, 44(4):410– 432, December 2006. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-006-9015-x.

Faulkner:2017:SCB

[FP17] Alex Faulkner and Lonneke Poort. Stretching and challenging the boundaries of law: Varieties of knowledge in biotechnologies regulation. Minerva, 55(2):209–228, June 2017. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

Flink:2018:EFR

[FP18] Tim Flink and Tobias Peter. Excellence and frontier research as travelling concepts in science policymaking. Minerva, 56(4): 431–452, December 2018. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/article/10.1007/s11024-018-9351-7.

Frankel:1978:FC

[Fra78] Charles Frankel. A further comment. Minerva, 16(1):142– 148, Spring 1978. CODEN MINEFY. ISSN 0026-4695 (print),

1573-1871 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF01102185.

Franke:1991:PUA

[Fra91] Ann H. Franke. Private universities in Australia. Minerva, 29 (3):294–320, Autumn 1991. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01097880.

Franssen:2022:KSJ

[Fra22] Thomas Franssen. Book review: Katherine E. Smith, Justyna Bandola-Gill, Nasar Meer, Ellen Stewart and Richard Watermeyer, The Impact Agenda: Controversies, Consequences and Challenges. Minerva, 60(2):325–328, June 2022. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (elec-

tronic).  URL https://link.springer.com/article/10.
1007/s11024-022-09464-x.

Freeman:1967:CER

[Fre67] Christopher Freeman. A comment on “expenditures on research and development in Poland”. Minerva, 5(3):371–375, Spring 1967. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02388575.

Freeman:1968:RDC

[Fre68] Christopher Freeman. Research and development in Czechoslovakia. Minerva, 6(4):598–601, Summer 1968. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01096551.

Freudenthal:1987:JBD

[Fre87] Gad Freudenthal. Joseph Ben-David’s sociology of scientific knowledge. Minerva, 25(1–2):135–149, Spring–Summer 1987. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096861.

Freeden:2023:BRU

[Fre23a] Michael Freeden. Book review: Ursula van Beek (ed.), Democracy under Pressure. Resilience or Retreat? Minerva, 61(4): 631–633, December 2023. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL https://link.springer.

com/article/10.1007/s11024-023-09513-z.

Freiman:2023:VHC

[Fre23b] Ori Freiman. Vaccine hesitancy and the concept of trust: an analysis based on the Israeli COVID-19 vaccination campaign. Minerva, 61(3):357–381, September 2023. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL https://

link.springer.com/article/10.1007/s11024-023-09498-9.

Friborg:1969:RBD

[Fri69] Göran Friborg. The reduction of the brain drain: Problems and polices. Minerva, 7(4):760–761, Summer 1969. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01099548.

Fox:1981:BR

[FS81] Robert Fox and Terry Shinn. Book reviews. Minerva, 19(2): 329–339, Summer 1981. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096571.

Floud:1988:BR

[FSH88] Jean Floud, Allen Salzman, and Lillian Hoddeson. Book reviews. Minerva, 26(2):273–296, Summer 1988. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01096697.

Franssen:2018:DPF

[FSHdR18] Thomas Franssen, Wout Scholten, Laurens K. Hessels, and Sarah de Rijcke. The drawbacks of project funding for epistemic innovation: Comparing institutional affordances and constraints of different types of research funding. Minerva, 56(1):11–33, March 2018. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/article/10.1007/s11024-017-9338-9;

 http://link.springer.com/content/pdf/10.1007/s11024-
017-9338-9.pdf.

Foster:1975:BR

[FSMH75] Philip Foster, A. B. Shah, Kenneth Mellanby, and Richard G. Hewlett. Book reviews. Minerva, 13(3):466–496, Autumn 1975. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01097850.

Fuller:2000:RSP

[Ful00] Steve Fuller. The Republic of science: Its political and economic theory. Minerva, 38(1):26–32, ???? 2000. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1023/
A%3A1026548212754.

Fabian:1979:CCH

[FV79] Bernhard Fabian and Rudolf Vierhaus. The calling and condition of the humanistic disciplines. Minerva, 17(4):549–554, Winter 1979. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01101903.

Fuchs:2001:GSS

[FvSA01] Stefan Fuchs, Janina von Stebut, and Jutta Allmendinger. Gender, science, and scientific organizations in Germany. Minerva, 39(2):175–201, ???? 2001. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1010380510013.

Frumkin:1994:BR

[FW94] Peter Frumkin and George Weisz. Book reviews. Minerva, 32 (4):469–479, Winter 1994. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01098034.

Gajendragadkar:1968:MII

[Gaj68] P. B. Gajendragadkar. The medium of instruction in Indian higher education. Minerva, 6(2):257–262, Winter 1968. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01097261.

Glaser:2022:IRR

[GAL22] Jochen Gläser, Mitchell Ash, and Klaus Lieb. The independence of research — a review of disciplinary perspectives and outline of interdisciplinary prospects. Minerva, 60(1):105–138, March 2022. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL https://link.springer.com/article/10.
1007/s11024-021-09451-8.

Gamble:1997:BAV

[Gam97] Vanessa Northington Gamble. Black autonomy versus white control: Black hospitals and the dilemmas of white philanthropy, 1920–1940. Minerva, 35(3):247–267, September 1997. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1004234101887.

Garrett:2001:CCP

[Gar01] Richard Garrett. The coming challenge:private competition in English higher education. Minerva, 39(1):99–114, ???? 2001. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1010330504668.

Gaston:1971:SCP

[Gas71] Jerry Gaston. Secretiveness and competition for priority of discovery in physics. Minerva, 9(4):472–492, October 1971. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01558020.

Gascoigne:2007:WSW

[Gas07] John Gascoigne. Wider still and wider: science and the quest for exploration. Minerva, 45(4):481–483, December 2007. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-007-9060-0.

Gaussen:1973:HCF

[Gau73] Frédéric Gaussen. I: The human cost of French University expansion. Minerva, 11(3):372–386, July 1973. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01557690.

Grove:1991:BR

[GC91] J. W. Grove and R. C. Curtis. Book reviews. Minerva, 29 (1):109–130, Spring 1991. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096300.

Gruber:2023:WIK

[GCD23] Marie Gruber, Thomas Crispeels, and Pablo D’Este. Who am I? The influence of knowledge networks on PhD students’ formation of a researcher role identity. Minerva, 61(4):521–552, December 2023. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL https://link.springer.com/article/10.
1007/s11024-023-09492-1.

Genuth:2000:HEB

[GCS00] Joel Genuth, Ivan Chompalov, and Wesley Shrum. How experiments begin: The formation of scientific collaborations. Minerva, 38(3):311–348, ???? 2000. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1023/
A%3A1026573717027.

Greene:1985:PRA

[GDHH85] Penelope J. Greene, Jane S. Durch, Wendy Horwitz, and Valwyn S. Hooper. Policies for responding to allegations of fraud in research. Minerva, 23(2):203–215, Summer 1985. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01099942.

Gedikoglu:1995:CMU

[Ged95] Tokay Gedıkoğlu. Changing models of university government in Turkey. Minerva, 33(2):149–169, Summer 1995. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01097910.

Geertz:1974:SSP

[Gee74] Clifford Geertz. Social science policy in a new state. Minerva, 12(3):365–381, July 1974. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01102530.

Geiger:1988:AFA

[Gei88] Roger L. Geiger. American foundations and academic social science, 1945–1960. Minerva, 26(3):315–341, Autumn 1988. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096400.

Geiger:1997:WHA

[Gei97] Roger L. Geiger. What happened after Sputnik? Shaping university research in the United States. Minerva, 35(4):349– 367, December 1997. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1004324331995.

Geiger:2008:RV

[Gei08] Roger L. Geiger. The riddle of the valley. Minerva, 46(1): 127–132, March 2008. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-007-9074-7.

Gellert:2000:ABF

[Gel00] Claudius Gellert. Akademgorodok: Building for the future — and finding it does not work. Minerva, 37(4):405–415, December 2000. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1004702029154.

Gemelli:2003:LMF

[Gem03] Giuliana Gemelli. ‘Leadership and mind’: Frederic C. Lane as cultural entrepreneur and diplomat. Minerva, 41(2):115– 132, ???? 2003. CODEN MINEFY. ISSN 0026-4695 (print),

1573-1871 (electronic). URL http://link.springer.com/
accesspage/article/10.1023/A%3A1023654730130.

Genuth:1987:GTS

[Gen87] Joel Genuth. Groping towards science policy in the United States in the 1930s. Minerva, 25(3):238–268, Autumn 1987. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01097784.

Genov:2023:LLI

[Gen23] Nikolai Genov. ‘Lasting impact on the studies of knowledge and society’: a review of Nico Stehr, Knowledge Capitalism. Minerva, 61(3):461–465, September 2023. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL https://

link.springer.com/article/10.1007/s11024-023-09504-0.

Geuna:1998:IEU

[Geu98] Aldo Geuna. The internationalisation of European universities: A return to medieval roots. Minerva, 36(3):253–270, September 1998. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1004391922122.

Gould:2002:RPS

[GF02] Chandré Gould and Peter Folb. The role of professionals in the south African chemical and biological warfare programme. Minerva, 40(1):77–91, ???? 2002. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1023/
A%3A1015061614159.

Gonzalez:2021:SPC

[GFM21] Candelaria Barrios González, Esther Flores, and M. Ángeles Martínez. Scientific production convergence: an empirical analysis across nations. Minerva, 59(4):445–467, December 2021. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL https://link.springer.com/article/10.
1007/s11024-021-09443-8.

Gilbert:1973:BR

[GG73] Irene A. Gilbert and Herbert J. Grubel. Book reviews. Minerva, 11(1):140–152, January 1973. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01544986.

Ghosh:1988:GCU

[Gho88] Suresh Chandra Ghosh. The genesis of Curzon’s university reform: 1899–1905. Minerva, 26(4):463–492, Winter 1988. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096494.

Gibbons:1970:CGA

[Gib70] Michael Gibbons. The CERN 300 GeV accelerator: a case study in the application of the Weinberg criteria. Minerva, 8(1–4): 180–191, January 1970. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01553211.

Gibbons:2000:CPU

[Gib00] Michael Gibbons. Changing patterns of university– industry relations. Minerva, 38(3):352–361, ???? 2000. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1023/
A%3A1026559728608.

Gill:1967:PDM

[Gil67] Richard Rockingham Gill. Problems of decision-making in Soviet science policy. Minerva, 5(2):198–208, Winter 1967. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF02388524.

Gilbert:1969:API

[Gil69] Trene Gilbert. The academic profession in India. Minerva, 7 (4):755–758, Summer 1969. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01099546.

Gilbert:1972:IAP

[Gil72] Irene A. Gilbert. The Indian academic profession: The origins of a tradition of subordination. Minerva, 10(3):384–411, July 1972. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01556921.

Gillick:1977:CCM

[Gil77] Muriel Gillick. The criteria of choice in medical policy: Radiotherapy in Massachusetts. Minerva, 15(1):15–31, Spring 1977. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096318.

Giles:1978:UGN

[Gil78] Geoffrey J. Giles. University government in Nazi Germany: Hamburg. Minerva, 16(2):196–221, Summer 1978. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01096014.

GirodDeLain:1979:TVS

[Gir79] Bertrand Girod De L’ain. Two views of the state of the universities and of scientific research in France in the late 1970s. Minerva, 17(2):283–304, Summer 1979. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01096284.

Gish:1969:ESD

[Gis69] Oscar Gish. Emigration and the supply and demand for medical manpower: The Irish case. Minerva, 7(4):668–679, Summer 1969. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01099540.

Gish:1970:ESD

[Gis70] Oscar Gish. Emigration and the supply and demand for medical manpower: The Irish case. Minerva, 8(1–4):449– 451, January 1970. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01553229;   http://
link.springer.com/article/10.1007/BF01553229.

Gissing:2004:BHB

[Gis04] Philip Gissing. Britain and the h-Bomb. Minerva, 42(3):315– 319, ???? 2004. CODEN MINEFY. ISSN 0026-4695 (print),

1573-1871 (electronic). URL http://link.springer.com/
accesspage/article/10.1023/B%3AMINE.0000038302.42899.
9d.

Gornitzka:1998:BU

[GKL98] Åse Gornitzka, Svein Kyvik, and Ingvild Marheim Larsen. The bureaucratisation of universities. Minerva, 36(1):21–47, March 1998. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1004382403543.

Gulbrandsen:2004:SMN

[GL04] Magnus Gulbrandsen and Liv Langfeldt. In search of ‘mode 2’: the nature of knowledge production in Norway. Minerva, 42(3):237–250, ???? 2004. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1023/
B%3AMINE.0000038305.35138.31.

Grunert:2022:ARF

[GL22] Cathleen Grunert and Katja Ludwig. Academic reform in fractured disciplines — on the interaction of Bologna, new-public-management and the dynamics of disciplinary development. Minerva, 60(1):57–80, March 2022. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL https://link.

springer.com/article/10.1007/s11024-021-09452-7.

Glazer:1974:SMP

[Gla74] Nathan Glazer. The schools of the minor professions. Minerva, 12(3):346–364, July 1974. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01102529.

Glassheim:2008:CN

[Gla08] Eagle Glassheim. The conquest of nature. Minerva, 46(1): 165–167, March 2008. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-007-9078-3.

Glaser:2024:EGC

[Glä24] Jochen Gläser. From effects of governance to causes of epistemic change. Minerva, 62(3):309–337, September 2024. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (elec-

tronic).  URL https://link.springer.com/article/10.
1007/s11024-024-09526-2.

Glick:2005:DDA

[Gli05] Thomas F. Glick. Dictating to the dictator: Augustus Trowbridge, the Rockefeller Foundation, and the support of physics in Spain, 1923–1927. Minerva, 43(2):121–145, June 2005. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-005-2327-4.

Globerman:1976:CSP

[Glo76] Steven Globerman. Canadian science policy and economic nationalism. Minerva, 14(2):191–208, Summer 1976. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01107885.

Gemelli:2003:I

[GM03a] Giuliana Gemelli and Roy Macleod. Introduction. Minerva, 41(2):95–99, ???? 2003. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1023680029762.

Geuna:2003:URE

[GM03b] Aldo Geuna and Ben R. Martin. University research evaluation and funding: An international comparison. Minerva, 41(4):277– 304, ???? 2003. CODEN MINEFY. ISSN 0026-4695 (print),

1573-1871 (electronic). URL http://link.springer.com/
accesspage/article/10.1023/B%3AMINE.0000005155.70870.
bd.

Geuna:2009:GUK

[GM09] Aldo Geuna and Alessandro Muscio. The governance of University knowledge transfer: A critical review of the literature. Minerva, 47(1):93–114, March 2009. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
009-9118-2.

Gajary:2024:CRL

[GMD^{+}24] Lisa C. Gajary, Shalini Misra, Anand Desai, Dean M. Evasius, Joy Frechtling, David A. Pendlebury, Joshua D. Schnell, Gary Silverstein, and John Wells. Convergence research as a ‘systemof-systems’: a framework and research agenda. Minerva, 62 (2):253–286, June 2024. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL https://link.springer.

com/article/10.1007/s11024-023-09503-1.

Gvishiani:1973:ISP

[GMY73] D. M. Gvishiani, S. R. Mikulinsky, and M. G. Yaroshevsky. II the sociological and psychological study of scientific activity. Minerva, 11(1):121–129, January 1973. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01544982.

Grove:1984:BR

[GNNBD84] J. W. Grove, Thomas Nipperdey, Ernst Nolte, and Joseph Ben-David. Book reviews. Minerva, 22(3–4):437–462, Autumn– Winter 1984. CODEN MINEFY. ISSN 0026-4695 (print),

1573-1871 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF02207377.

Godin:2002:NMF

[God02] Benoit Godin. The numbers makers: Fifty years of science and technology official statistics. Minerva, 40(4):375–397, ???? 2002. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1020921620544.

Godin:2008:SSW

[God08] Benoît Godin. In the shadow of Schumpeter: W. Rupert Maclaurin and the study of technological innovation. Minerva, 46(3): 343–360, September 2008. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-008-9100-4.

Godin:2010:IWW

[God10] Benoît Godin. Innovation without the word: William F. Ogburn’s contribution to the study of technological innovation. Minerva, 48(3):277–307, September 2010. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
010-9151-1.

Godin:2012:ISI

[God12] Benoît Godin. “Innovation studies”: The invention of a specialty. Minerva, 50(4):397–421, December 2012. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
s11024-012-9212-8.

Godin:2014:ISS

[God14] Benoît Godin. “Innovation studies”: Staking the claim for a new disciplinary “tribe”. Minerva, 52(4):489–495, December 2014. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-014-9262-1.

Gold:1990:FED

[Gol90] Allan R. Gold. Freedom of expression and disruptions at meetings of student societies in university buildings. Minerva, 28 (1):96–97, Spring 1990. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096329.

Goni:2025:WDP

[Goñ25] Julian Iñaki Goñi. What is “Dialogue” in public engagement with science and technology? Bridging STS and deliberative democracy. Minerva, 63(3):607–631, September 2025. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL https://link.springer.com/article/10.
1007/s11024-024-09551-1.

Goodlad:1999:BRE

[Goo99] Sinclair Goodlad. Book review: Education in a Research University, edited by Kenneth J. Arrow, Richard W. Cottle, B. Curtis Eaves and Ingram Olkin. Minerva, 37(1):98–101, March 1999. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1004548526351.

Goodlad:2002:BUS

[Goo02] Sinclair Goodlad. The British universities– surviving change. Minerva, 40(4):399–406, ???? 2002. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1023/A%3A1020973604614.

Gavrila:2025:NDA

[GOR25] S. Gabriela Gavrila, Lisa Overbey, and Francisco O. Ramirez. In the name of diversity: Analyzing the adoptions of diversityrelated offices in U.S. higher education, 1968–2020. Minerva, 63(3):423–448, September 2025. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL https://link.

springer.com/article/10.1007/s11024-025-09576-0.

Gummett:1977:ACP

[GP77] Philip J. Gummett and Geoffrey L. Price. An approach to the central planning of British science: The formation of the advisory council on scientific policy. Minerva, 15(2):119–143, Summer 1977. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096673.

Goncalves:2004:IST

[GP04] Maria Eduarda Gonçalves and Pierre Papon. Introduction — scientific and technological institutions and the new knowledgebased society: A European perspective. Minerva, 42(1):3–9, ???? 2004. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/B%3AMINE.0000017696.44267.c5.

Gould:1973:BR

[GPM73] Julius Gould, Keith Pavitt, and Samuel McCracken. Book reviews. Minerva, 11(2):269–281, April 1973. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01557721.

Glick:2006:SFE

[GR06] Thomas F. Glick and José M. Sánchez Ron. Science frustrated: The ‘Einstein Institute’ in Madrid. Minerva, 44(4): 355–378, December 2006. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-006-9013-z.

Grant:1966:NBR

[Gra66] Robert P. Grant. National biomedical research agencies: a comparative study of fifteen countries. Minerva, 4(4):466–488, Summer 1966. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02207979.

Granick:1971:SRD

[Gra71] David Granick. The Soviet research and development system. Minerva, 9(1):129–132, January 1971. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01553191.

Gramelsberger:2012:BRP

[Gra12] Gabriele Gramelsberger. Book review: Paul N. Edwards, A Vast Machine: Computer Models, Climate Data, and the Politics of Global Warming. Minerva, 50(4):533–537, December 2012. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-012-9216-4.

Greenfeld:1987:SNG

[Gre87] Liah Greenfeld. Science and national greatness in seventeenthcentury England. Minerva, 25(1–2):107–122, Spring–Summer 1987. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096859.

Greenberg:2005:BRW

[Gre05] Daniel Greenberg. Book review: When economists go to Washington, DC: Randall Lutter and Jason F. Shogren (eds.), Painting the White House Green: Rationalizing Environmental Policy inside the Executive Office of the President (Washington, DC: RFF Press, 2004), 201 pp., ISBN 1-891853-73-2 (cloth) and ISBN 1-89153-72-4 (paper). Minerva, 43(1):109–112, ???? 2005. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-004-6644-9.

Greenberg:2007:SSS

[Gre07a] Daniel S. Greenberg. Saving the soul of science. Minerva, 45(4): 485–487, December 2007. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-007-9061-z.

Greene:2007:FO

[Gre07b] Mott Greene. Fathoming the ocean. Minerva, 45(1):101–103, March 2007. CODEN MINEFY. ISSN 0026-4695 (print),

1573-1871 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11024-006-9026-7.

Grosart:1971:CSP

[Gro71] Allister Grosart. Canadian science policy. Minerva, 9(4): 538–544, October 1971. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01558024.

Grove:1980:STA

[Gro80] J. W. Grove. Science as technology: Aspects of a potent myth. Minerva, 18(2):293–312, Summer 1980. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01100250.

Grossman:1982:AFS

[Gro82] David M. Grossman. American foundations and the support of economic research, 1913–29. Minerva, 20(1–2):59–82, Spring–Summer 1982. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01098190.

Grove:1985:RRS

[Gro85] J. W. Grove. Rationality at risk: Science against pseudoscience. Minerva, 23(2):216–240, Summer 1985. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01099943.

Grosby:1992:UTF

[Gro92] Steven Grosby. “The university of the twenty-first century”: Report on the discussions. Minerva, 30(2):269–295, Summer 1992. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096885.

Grove:1996:MSR

[Gro96a] J. W. Grove. The morality of scientists revisited. Minerva, 34(1):57–67, ???? 1996. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00124201.

Grove:1996:SBS

[Gro96b] J. W. Grove. Stalin’s bomb: Soviet physicists and the Cold War. Minerva, 34(4):381–392, ???? 1996. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00127073.

Grove:1999:BRR

[Gro99] J. W. Grove. Book review: The Racial Economy of Science: Toward a Democratic Future, edited by Sandra Harding; Feminism and Science, edited by Evelyn Fox Keller and Helen Longino. Minerva, 37(2):191–198, June 1999. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1023/A%3A1004612723645.

Grubel:1968:RBD

[Gru68] Herbert G. Grubel. The reduction of the brain drain: Problems and policies. Minerva, 6(4):541–558, Summer 1968. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01096547.

Grubel:1970:ESD

[Gru70] Herbert Grubel. Emigration and the supply and demand for medical manpower: The Irish case. Minerva, 8(1–4): 146–148, January 1970. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01553206.

Grubel:1976:RPS

[Gru76] Herbert G. Grubel. Reflections on the present state of the brain drain and a suggested remedy. Minerva, 14(2):209–224, Summer 1976. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01107886.

Grundmann:2017:PEK

[Gru17] Reiner Grundmann. The problem of expertise in knowledge societies. Minerva, 55(1):25–48, March 2017. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s11024-016-9308-7.pdf.

Geiger:2005:BTT

[GS05] Roger L. Geiger and Creso Sá. Beyond technology transfer: US state policies to harness University research for economic development. Minerva, 43(1):1–21, ???? 2005. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
004-6623-1.

Grove:1989:BR

[GT89] J. W. Grove and R. Steven Turner. Book reviews. Minerva, 27 (4):535–548, Winter 1989. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096543.

Guagnini:1988:HEE

[Gua88] Anna Guagnini. Higher education and the engineering profession in Italy: The Scuole of Milan and Turin, 1859–1914. Minerva, 26 (4):512–548, Winter 1988. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096496.

Guilhot:2006:NIF

[Gui06] Nicolas Guilhot. A network of influential friendships: The fondation pour une entraide intellectuelle Européenne and East–West cultural dialogue, 1957–1991. Minerva, 44(4):379–409, December 2006. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-006-9014-y.

Gunn:1967:OSB

[Gun67] Lewis A. Gunn. Organising for science in Britain: Some relevant questions. Minerva, 5(2):167–197, Winter 1967. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF02388523.

Guo:2013:CGW

[Guo13] Sun-Wei Guo. China’s “gene war of the Century” and its aftermath: The contest goes on. Minerva, 51(4):485–512, December 2013. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-013-9237-7.

Guston:1994:CSM

[Gus94] David H. Guston. Congressmen and scientists in the making of science policy: The Allison commission, 1884–1886. Minerva, 32(1):25–52, Spring 1994. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096655.

Guston:2012:PTM

[Gus12] David H. Guston. The pumpkin or the tiger? Michael Polanyi, Frederick Soddy, and anticipating emerging technologies. Minerva, 50(3):363–379, September 2012. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s11024-012-9204-8.pdf.

Glaser:2018:CFA

[GV18] Jochen Gläser and Kathia Serrano Velarde. Changing funding arrangements and the production of scientific knowledge: Introduction to the special issue. Minerva, 56(1):1–10, March 2018. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s11024-018-9344-6;   http://link.springer.com/content/
pdf/10.1007/s11024-018-9344-6.pdf.

Haber:1973:GIF

[Hab73] L. F. Haber. Government intervention at the frontiers of science: British dyestuffs and synthetic organic chemicals 1914–39. Minerva, 11(1):79–94, January 1973. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01544978.

Hafele:1974:HNC

[Häf74] Wolf Häfele. Hypotheticality and the new challenges: The pathfinder role of nuclear energy. Minerva, 12(3):303–322, July 1974. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01102526.

Hafele:1997:RIL

[Häf97] Wolf Häfele. Reshaping and integrating a large scientific institution of the former German democratic Republic after reunification. Minerva, 35(2):127–137, June 1997. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1023/
A%3A1004242827287.

Hagendijk:2004:PUS

[Hag04] R. P. Hagendijk. The public understanding of science and public participation in regulated worlds. Minerva, 42(1):41–59, ???? 2004. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/B%3AMINE.0000017699.19747.f0.

Hahn:1975:SRO

[Hah75] Roger Hahn. Scientific research as an occupation in eighteenthcentury Paris. Minerva, 13(4):501–513, Winter 1975. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01096175.

Halsey:1963:SGS

[Hal63] A. H. Halsey. Science and government in Sweden: Impressions from an OECD conference. Minerva, 2(1):54–60, Autumn 1963. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01100780.

Halliday:1975:PUP

[Hal75] Terence C. Halliday. The politics of “universal participatory democracy”: A Canadian case study. Minerva, 13(3): 404–427, Autumn 1975. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01097848.

Halsey:2000:ESS

[Hal00] A. H. Halsey. Edward Shils, sociology and universities. Minerva, 37(4):391–404, December 2000. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1023/A%3A1004782815042.

Hammerstein:1993:BR

[Ham93] Notker Hammerstein. Book review. Minerva, 31(4):523–528, Winter 1993. CODEN MINEFY. ISSN 0026-4695 (print),

1573-1871 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF01096453.

Hamlin:2005:VFS

[Ham05] Christopher Hamlin. A virtue-free science for public policy? Minerva, 43(4):397–418, December 2005. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
005-2499-y.

Hamlin:2008:JDC

[Ham08] Christopher Hamlin. Just don’t call it science. Minerva, 46 (1):99–116, March 2008. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-007-9071-x.

Haniotis:1965:SNS

[Han65] George V. Haniotis. The search for a national scientific policy in Greece. Minerva, 3(3):312–320, Spring 1965. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01099957.

Haniotis:1968:SUG

[Han68] George Haniotis. The situation of the universities in Greece. Minerva, 6(2):163–184, Winter 1968. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01097257.

Harding:2002:EPD

[Har02] Rebecca Harding. Evolution, path dependence, learning and innovation: A review of four recent books. Minerva, 40(3): 289–299, ???? 2002. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1019516424249.

Hardtwig:2004:PAS

[Har04] Wolfgang Hardtwig. The Prussian Academy of Sciences and humanities during the Weimar Republic. Minerva, 42(4): 333–357, ???? 2004. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-004-2112-9.

Harrison:2006:SD

[Har06] Peter Harrison. Science and dissent. Minerva, 44(2):223– 227, June 2006. CODEN MINEFY. ISSN 0026-4695 (print),

1573-1871 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11024-006-0011-y.

Harman:2007:PI

[Har07] Oren Harman. On the power of ideas. Minerva, 45(2): 175–189, June 2007. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-007-9029-z.

Harman:2008:WPS

[Har08] Oren Harman. When philosophy of science counts. Minerva, 46 (1):133–138, March 2008. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-007-9075-6.

Harwood:2010:UAD

[Har10] Jonathan Harwood. Understanding academic drift: On the institutional dynamics of higher technical and professional education. Minerva, 48(4):413–427, December 2010. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
010-9156-9.

Hauray:2017:RKR

[Hau17] Boris Hauray. From regulatory knowledge to regulatory decisions: The European evaluation of medicines. Minerva, 55(2): 187–208, June 2017. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

Havemann:1966:GDR

[Hav66] Robert Havemann. German Democratic Republic prof, Havemann and the East German Academy. Minerva, 4(3):419– 424, Spring 1966. CODEN MINEFY. ISSN 0026-4695 (print),

1573-1871 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF01584640.

Hayward:1988:RCU

[Hay88] John Hayward. Responses to contraction: The University of Hull, 1979–1984. Minerva, 24(1):74–97, Spring 1988. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01102554.

Hayhoe:1993:CUS

[Hay93] Ruth Hayhoe. Chinese universities and the social sciences. Minerva, 31(4):478–503, Winter 1993. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01096451.

Hayhoe:1994:IHL

[Hay94] Ruth Hayhoe. Ideas of higher learning, east and west: Conflicting values in the development of the Chinese university. Minerva, 32(4):361–382, Winter 1994. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01098030.

Hayhoe:1998:DJI

[Hay98] Ruth Hayhoe. Dilemmas in Japan’s intellectual culture. Minerva, 36(1):1–19, March 1998. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1023/A%3A1004376819473.

Harsh:2018:RCR

[HBW^{+}18] Matthew Harsh, Ravtosh Bal, Jameson Wetmore, G. Pascal Zachary, and Kerry Holden. The rise of computing research in East Africa: The relationship between funding, capacity and research community in a nascent field. Minerva, 56(1):35–58, March 2018. CODEN MINEFY. ISSN 0026-4695 (print), 1573-

1871 (electronic). URL http://link.springer.com/article/
10.1007/s11024-017-9341-1.

Holton:1971:BR

[HC71] Gerald Holton and Charles F. Carter. Book reviews. Minerva, 9 (4):568–577, October 1971. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01558034.

Hammerstein:1988:BR

[HCRG88] Nötker Hammerstein, Maurice Crosland, Marc Raeff, and J. W. Grove. Book reviews. Minerva, 26(3):429–459, Autumn 1988. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096405.

Hill:1993:BR

[HD93] Stephen Hill and David DeVorkin. Book reviews. Minerva, 31 (2):242–252, Summer 1993. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01125437.

Hoang:2022:SPS

[HD22] Cuong Huu Hoang and Trang Thi Doan Dang. A sociocultural perspective on scholars developing research skills via research communities in Vietnam. Minerva, 60(1):81–104, March 2022. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL https://link.springer.com/article/10.
1007/s11024-021-09454-5.

Hammarfelt:2025:AIA

[HD25] Björn Hammarfelt and Johanna Dahlin. Abstracting it all: The Soviet Institute of Scientific Information (VINITI) and the promise of centralisation, 1952–1977. Minerva, 63(1):115– 133, March 2025. CODEN MINEFY. ISSN 0026-4695 (print),

1573-1871 (electronic). URL https://link.springer.com/
article/10.1007/s11024-024-09545-z.

Hammarfelt:2017:EME

[HdRW17] Björn Hammarfelt, Sarah de Rijcke, and Paul Wouters. From eminent men to excellent universities: University rankings as calculative devices. Minerva, 55(4):391–411, December 2017. CODEN MINEFY. ISSN 0026-4695 (print),

1573-1871 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s11024-017-9329-x.pdf;   http://
link.springer.com/article/10.1007/s11024-017-9329-x.

Heim:2006:EPR

[Hei06] Susanne Heim. Expansion policy and the role of agricultural research in Nazi Germany. Minerva, 44(3):267–284, June 2006. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-006-9007-x.

Heidler:2011:CSS

[Hei11] Richard Heidler. Cognitive and social structure of the elite collaboration network of astrophysics: a case study on shifting network structures. Minerva, 49(4):461–488, December 2011. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-011-9184-0.

Heidler:2017:ECC

[Hei17] Richard Heidler. Epistemic cultures in conflict: The case of astronomy and high energy physics. Minerva, 55(3):249– 277, September 2017. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/article/10.1007/s11024-017-9315-3.

Hennis:1977:LGU

[Hen77] Wilhelm Hennis. The legislator and the German university. Minerva, 15(3–4):286–315, Autumn–Winter 1977. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01096726.

Hendriks:2021:RMD

[Hen21] Barbara Hendriks. A review of Mark Dennis Robinson, The Market in Mind — How Financialization is Shaping Neuroscience, Translational Medicine and Innovation in Biotechnology, The MIT Press, Cambridge, MA, 2019, 309 pp., ISBN: 978-0-262-53687-5. Minerva, 59(1):139–143, March 2021. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/article/10.1007/s11024-
021-09432-x;   http://link.springer.com/content/pdf/
10.1007/s11024-021-09432-x.pdf.

Hermanowicz:2013:CM

[Her13] Joseph C. Hermanowicz. The culture of mediocrity. Minerva, 51(3):363–387, September 2013. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
013-9231-0.

Hermanowicz:2021:HEA

[Her21] Joseph C. Hermanowicz. Honest evaluation in the academy. Minerva, 59(3):311–329, September 2021. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL https://link.

springer.com/article/10.1007/s11024-021-09434-9.

Hermanowicz:2024:TU

[Her24] Joseph C. Hermanowicz. The therapeutic university. Minerva, 62(4):505–526, December 2024. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL https://link.

springer.com/article/10.1007/s11024-024-09539-x.

Hess:2011:BSS

[Hes11] David J. Hess. Bourdieu and science studies: Toward a reflexive sociology. Minerva, 49(3):??, September 2011. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/s11024-011-9178-y.

Hessels:2013:CSS

[Hes13] Laurens K. Hessels. Coordination in the science system: Theoretical framework and a case study of an intermediary organization. Minerva, 51(3):317–339, September 2013. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s11024-013-9230-1.pdf.

Hesselmann:2023:TTA

[Hes23] Felicitas Hesselmann. A tale of two academic communities: Digital imaginaries of automatic screening tools in editorial practice. Minerva, 61(2):221–241, June 2023. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL https://link.

springer.com/article/10.1007/s11024-022-09484-7.

Hetherington:1990:NAC

[Het90] Norriss S. Hetherington. The National Advisory Committee for Aeronautics: A forerunner of federal governmental support for scientific research. Minerva, 28(1):59–80, Spring 1990. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096326.

Hetherington:1991:APG

[Het91] Norriss S. Hetherington. Air power and governmental support for scientific research: The approach to the Second World War. Minerva, 29(4):420–439, Winter 1991. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01113490.

Heterington:1997:HRN

[Het97] Norris Heterington. A handful of recent NASA history books i; More than Merely Men, Machinery, Missions and Political Machinations?; The Birth of NASA: The Diary of T. Keith Glennan, edited by J. D. Hunley, with an introduction by Roger D. Launius; The Problem of Specs Travel: The Rocket Motor, by Hermann Noordung, edited by Ernst Stuhlinger and J. D. Hunley with Jennifer Garland; Powering Apollo: James E. Webb of NASA, by W. Henry Lambright; Spaceflights Revolution: NASA Langley Research Center From Sputnik to Apollo, by James R. Hansen; Suddenly, Tomorrow Came … A History of the Johnson Space Center, by Henry C. Dethoff. Minerva, 35(4): 387–396, December 1997. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1004379303770.

Heyd:1987:NEP

[Hey87] Michael Heyd. The new experimental philosophy: a manifestation of “enthusiasm” or an antidote to it? Minerva, 25(4): 423–440, Winter 1987. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01099441.

Hessels:2019:VVH

[HFSdR19] Laurens K. Hessels, Thomas Franssen, Wout Scholten, and Sarah de Rijcke. Variation in valuation: How research groups accumulate credibility in four epistemic cultures. Minerva, 57 (2):127–149, June 2019. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/article/10.1007/s11024-018-09366-x.

Hellstrom:2020:CSM

[HH20a] Tomas Hellström and Christina Hellström. Cross-sectoral mobility funding and the challenge of immersion: The case of SSH. Minerva, 58(3):389–407, September 2020. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/article/10.1007/s11024-
020-09398-2;   http://link.springer.com/content/pdf/
10.1007/s11024-020-09398-2.pdf.

Hicks:2020:CPE

[HH20b] Diana Hicks and J. Britt Holbrook. A cartography of philosophy’s engagement with society. Minerva, 58(1):25–45, March 2020. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/article/10.
1007/s11024-019-09384-3.

Horbach:2020:JPR

[HH20c] Serge P. J. M. Horbach and Willem Halffman. Journal peer review and editorial evaluation: Cautious innovator or sleepy giant? Minerva, 58(2):139–161, June 2020. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/article/10.1007/s11024-
019-09388-z;   http://link.springer.com/content/pdf/
10.1007/s11024-019-09388-z.pdf.

Hagendijk:2006:PDG

[HI06] Rob Hagendijk and Alan Irwin. Public deliberation and governance: Engaging with science and technology in contemporary Europe. Minerva, 44(2):167–184, June 2006. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
006-0012-x.

Heffernan:2007:DIP

[HJ07] Michael Heffernan and Heike Jöns. Degrees of influence: the politics of honorary degrees in the universities of Oxford and Cambridge, 1900–2000. Minerva, 45(4):389–416, December 2007. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-007-9065-8.

Heffernan:2008:DIP

[HJ08] Michael Heffernan and Heike Jöns. Degrees of influence: The politics of honorary degrees in the universities of Oxford and Cambridge, 1900–2000. Minerva, 46(1):169, March 2008. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s11024-008-9088-9.pdf.

Hellstrom:2012:RWC

[HJ12] Tomas Hellström and Merle Jacob. Revisiting ‘Weinberg’s choice’: Classic tensions in the concept of scientific merit. Minerva, 50(3):381–396, September 2012. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
012-9203-9.

Haakansta:2016:MTB

[HJ16] Carin Håkansta and Merle Jacob. Mode 2 and the tension between excellence and utility: The case of a policyrelevant research field in Sweden. Minerva, 54(1):1–20, March 2016. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-015-9288-z.

Hellstrom:2025:CRS

[HJ25] Tomas Hellström and Merle Jacob. Claiming relevance for social science and the humanities: Relevance expressions across methodological divides. Minerva, 63(2):257–279, June 2025. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL https://link.springer.com/article/10.
1007/s11024-024-09561-z.

Heilbron:1988:FPM

[HK88] J. L. Heilbron and Daniel J. Kevles. Finding a policy for mapping and sequencing the human genome: Lessons from the history of particle physics. Minerva, 26(3):299–314, Autumn 1988. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096399.

Hoddeson:2000:SSC

[HK00] Lillian Hoddeson and Adrienne W. Kolb. The Superconducting Super Collider’s frontier outpost, 1983–1988. Minerva, 38 (3):271–310, ???? 2000. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1026569616118.

Hicks:2011:EER

[HK11] Diana Hicks and J. Sylvan Katz. Equity and excellence in research funding. Minerva, 49(2):137–151, June 2011. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/s11024-011-9170-6.

Halffman:2010:IAU

[HL10] Willem Halffman and Loet Leydesdorff. Is inequality among universities increasing? Gini coefficients and the elusive rise of elite universities. Minerva, 48(1):55–72, March 2010. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s11024-010-9141-3.pdf.

Herzog:2022:CSD

[HL22] Lisa Herzog and Robert Lepenies. Citizen science in deliberative systems: Participation, epistemic injustice, and civic empowerment. Minerva, 60(4):489–508, December 2022. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (elec-

tronic).  URL https://link.springer.com/article/10.
1007/s11024-022-09467-8.

Hacking:2024:MAL

[HLE24] Nick Hacking, Jamie Lewis, and Robert Evans. Mapping approaches to ‘citizen science’ and ‘community science’ and everything in-between: The evolution of new epistemic territory?

Minerva, 62(4):549–572, December 2024. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL https://

link.springer.com/article/10.1007/s11024-024-09529-z.

Hofmanner:2021:SPS

[HM21] Alexandra Hofmänner and Elisio Macamo. The science policy script, revised. Minerva, 59(3):331–354, September 2021. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL https://link.springer.com/article/10.
1007/s11024-020-09427-0.

Hirsch:1977:IOA

[HN77] Helmut Hirsch and Helga Nowotny. Information and opposition in Austrian nuclear energy policy. Minerva, 15(3–4):316– 334, Autumn–Winter 1977. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096727.

Hohlfeld:2004:I

[HNW04] Rainer Hohlfeld, Peter Nötzoldt, and Peter Th. Walther. Introduction. Minerva, 42(4):323–331, ???? 2004. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
004-2110-y.

Hoch:1987:MGN

[Hoc87] Paul K. Hoch. Migration and the generation of new scientific ideas. Minerva, 25(3):209–237, Autumn 1987. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01097783.

Hoddeson:1980:EQT

[Hod80] Lillian Hartmann Hoddeson. The entry of the quantum theory of solids into the Bell telephone laboratories, 1925–40: a case-study of the industrial application of fundamental science. Minerva, 18 (3):422–447, Autumn 1980. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096951.

Holloway:1968:SUG

[Hol68] John Holloway. The situation of the universities in Greece. Minerva, 6(3):427–431, Spring 1968. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01096429.

Holton:1974:SGS

[Hol74] Gerald Holton. Striking gold in science: Fermi’s group and the recapture of Italy’s place in physics. Minerva, 12(2): 159–198, April 1974. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01553176.

Holton:1981:FAP

[Hol81] Gerald Holton. The formation of the American physics community in the 1920s and the coming of Albert Einstein. Minerva, 19 (4):569–581, Winter 1981. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096193.

Hook:1966:TAK

[Hoo66] Sidney Hook. Thoughts after Knopfelmacher: Appointments on university. Minerva, 4(2):279–285, Winter 1966. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01584864.

Hook:1984:AEA

[Hoo84] Sidney Hook. The academic ethic in abeyance: Recollections of Walpurgisnacht at New York University. Minerva, 22(3–4):297– 315, Autumn–Winter 1984. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF02207366.

Hook:1987:CMA

[Hoo87] Sidney Hook. Communists, McCarthy and American universities. Minerva, 25(3):331–348, Autumn 1987. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01097787.

Hook:1989:ROH

[Hoo89] Sidney Hook. Reflections on the obligation of honesty in the university. Minerva, 27(4):505–534, Winter 1989. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01096542.

Hooker:2003:SLA

[Hoo03] Cliff Hooker. Science: Legendary, academic- and postacademic? Minerva, 41(1):71–81, ???? 2003. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1023/
A%3A1022253704644.

Hooker:2008:CBT

[Hoo08] C. A. Hooker. Constructivism between transcendentalism and convention. Minerva, 46(1):87–97, March 2008. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
s11024-007-9070-y.

Hoppe:2011:BRM

[Hop11] Robert Hoppe. Book review: Mark B. Brown, Science in Democracy. Expertise, Institutions, and Representation. Minerva, 49 (3):??, September 2011. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s11024-011-9179-x.pdf.

Horta:2022:AIA

[Hor22] Hugo Horta. Academic inbreeding: Academic oligarchy, effects, and barriers to change. Minerva, 60(4):593–613, December 2022. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL https://link.springer.com/article/10.
1007/s11024-022-09469-6.

Hoydal:2024:CWL

[Høy24] Øyunn Syrstad Høydal. Could I write like Carol Weiss? Minerva, 62(4):491–504, December 2024. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL https://link.

springer.com/article/10.1007/s11024-024-09534-2.

Hladchenko:2019:ITH

[HP19] Myroslava Hladchenko and Romulo Pinheiro. Implementing the triple helix model: Means–ends decoupling at the state level? Minerva, 57(1):1–22, March 2019. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/article/10.1007/s11024-018-9355-3;
 http://link.springer.com/content/pdf/10.1007/s11024-
018-9355-3.pdf.

Hoddeson:1990:BR

[HR90] Lillian Hoddeson and Nathan Reingold. Book reviews. Minerva, 28(4):540–545, Winter 1990. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096653.

Hackett:2009:SCI

[HR09] Edward J. Hackett and Diana R. Rhoten. The Snowbird Charrette: Integrative interdisciplinary collaboration in environmental research design. Minerva, 47(4):407–440, December 2009. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-009-9136-0.

Halffman:2015:AMO

[HR15] Willem Halffman and Hans Radder. The academic manifesto: From an occupied to a public University. Minerva, 53 (2):165–187, June 2015. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s11024-015-9270-9.pdf.

Habermas:1971:BR

[HRKZ71] Jürgen Habermas, Fritz K. Ringer, Norman Kaplan, and John Ziman. Book reviews. Minerva, 9(3):422–437, July 1971. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01553369.

Hrotic:2013:SPD

[Hro13] Steven Hrotic. Survey of the philosophic discipline. Minerva, 51(1):93–122, March 2013. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-013-9218-x.

Hunt:1999:RSP

[HS99] Jane Hunt and Simon Shackley. Reconceiving science and policy: Academic, fiducial and bureaucratic knowledge. Minerva, 37 (2):141–164, June 1999. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1004696104081.

Hendriks:2019:WCS

[HSR19] Barbara Hendriks, Arno Simons, and Martin Reinhart. What are clinician scientists expected to do? The undefined space for professionalizable work in translational biomedicine. Minerva, 57(2):219–237, June 2019. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/article/10.1007/s11024-019-09367-4.

Hughes:2006:ASC

[Hug06] Thomas P. Hughes. America as second creation. Minerva, 44 (2):219–222, June 2006. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-006-0010-z.

Hunter:1980:UNL

[Hun80] Howard O. Hunter. Universities and the needs of local and regional communities comments on the outlook of the centre for educational research and innovation of the organisation for economic co-operation and development. Minerva, 18(4): 624–643, Winter 1980. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096127.

Hunter:1981:CSA

[Hun81] Howard O. Hunter. The constitutional status of academic freedom in the United States. Minerva, 19(4):519–568, Winter 1981. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096192.

Hunter:1985:ASG

[Hun85] Howard O. Hunter. Academic self-government in the United States. Minerva, 23(1):1–28, Spring 1985. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/BF01097838.

Husen:1976:SUR

[Hus77] Torsten Husén. Swedish university research at the crossroads. Minerva, 14(4):419–446, Winter 1976–1977. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01097634.

Husen:1983:ERM

[Hus83] Torsten Husén. Educational research and the making of policy in education: an international perspective. Minerva, 21 (1):81–100, Spring 1983. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01099504.

Hutchinson:1969:SCS

[Hut69] Eric Hutchinson. Scientists and civil servants: The struggle over the national physical laboratory in 1918. Minerva, 7(3): 373–398, Spring 1969. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01103267.

Hutchinson:1970:SIC

[Hut70a] Eric Hutchinson. Scientists as an inferior class: The early years of the DSIR. Minerva, 8(1–4):396–411, January 1970. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01553225;   http://link.springer.com/article/
10.1007/BF01553225.

Hutchinson:1970:SGO

[Hut70b] Eric Hutchinson. Scientists, government and organised research in Great Britain, 1914–16. Minerva, 8(1–4):594– 597, January 1970. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01553241;   http://
link.springer.com/article/10.1007/BF01553241.

Hutchinson:1972:FCB

[Hut72] Eric Hutchinson. A fruitful cooperation between government and academic science: Food research in the United Kingdom. Minerva, 10(1):19–50, January 1972. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01881389.

Hutchinson:1975:OUG

[Hut75] Eric Hutchinson. The origins of the University grants committee. Minerva, 13(4):583–620, Winter 1975. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01096178.

Harambam:2023:CTW

[HV23] Jaron Harambam and Ehler Voss. The corona truth wars: Epistemic disputes and societal conflicts around a pandemic — an introduction to the special issue. Minerva, 61(3):299–313, September 2023. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL https://link.springer.com/article/10.
1007/s11024-023-09511-1.

Hessels:2011:PAS

[HvL11] Laurens K. Hessels and Harro van Lente. Practical applications as a source of credibility: A comparison of three fields of Dutch academic chemistry. Minerva, 49(2):??, June 2011. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s11024-011-9167-1.pdf.

Hyslop:1964:UEA

[Hys64] J. M. Hyslop. The University of East Africa. Minerva, 2(3): 286–302, Spring 1964. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01097318.

Ibrahim:2023:SPC

[Ibr23] Youssef Ibrahim. The science and politics of climate engineering — social science perspectives. Minerva, 61(2):291–297, June 2023. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL https://link.springer.com/article/10.
1007/s11024-023-09488-x.

Irwin:2025:ARF

[IH25] Alan Irwin and Maja Horst. Always relevant? finding a place for the social sciences in the technical University and the business school. Minerva, 63(2):281–300, June 2025. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL https://

link.springer.com/article/10.1007/s11024-024-09553-z.

Innis:1992:LCL

[Inn92] Nancy K. Innis. Lessons from the controversy over the loyalty oath at the University of California. Minerva, 30(3): 337–365, Autumn 1992. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01097643.

Israel:2008:PIP

[Isr08] Paul Israel. Pathways in intellectual property. Minerva, 46 (1):151–154, March 2008. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-007-9081-8.

Jacobsen:2006:WC

[Jac06] Anja Skaar Jacobsen. The water controversy. Minerva, 44(4): 459–462, December 2006. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-006-9011-1.

Jacobsen:2008:CBC

[Jac08] Anja Skaar Jacobsen. The complementarity between the collective and the individual. Minerva, 46(2):195–214, June 2008. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-008-9091-1.

Jacob:2009:CGA

[Jac09] Merle Jacob. On commodification and the governance of academic research. Minerva, 47(4):391–405, December 2009. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/s11024-009-9134-2.

James:1968:UGP

[Jam68] Walter James. The universities, the government and the public accounts committee. Minerva, 6(2):265–267, Winter 1968. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01097264.

Jasanoff:2003:THC

[Jas03] Sheila Jasanoff. Technologies of humility: Citizen participation in governing science. Minerva, 41(3):223–244, ???? 2003. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1025557512320.

Josephson:1997:DPR

[JE97] Paul R. Josephson and Igor Egorov. The deceptive promise of reform: Ukrainian science in crisis. Minerva, 35(4):321– 347, December 1997. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1004319001044.

Jenkins:1997:BRN

[Jen97] Edgar W. Jenkins. Book review: a new social contract for science? Minerva, 35(1):83–90, ???? 1997. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1023/
B%3AMINE.0000019493.52810.15.

Jenkin:2001:UPW

[Jen01] John Jenkin. A unique partnership: William and Lawrence Bragg and the 1915 Nobel Prize in Physics. Minerva, 39(4): 373–392, ???? 2001. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1012783802528.

Jasanoff:2009:CAS

[JK09] Sheila Jasanoff and Sang-Hyun Kim. Containing the atom: Sociotechnical imaginaries and nuclear power in the United States and South Korea. Minerva, 47(2):??, June 2009. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/s11024-009-9124-4.

Johann:2019:PSA

[JM19] David Johann and Sabrina Jasmin Mayer. The perception of scientific authorship across domains. Minerva, 57(2):175–196, June 2019. CODEN MINEFY. ISSN 0026-4695 (print), 1573-

1871 (electronic). URL http://link.springer.com/article/
10.1007/s11024-018-9363-3.

Johnson:1965:EBD

[Joh65a] Harry G. Johnson. The economics of the “brain drain”: The Canadian case. Minerva, 3(3):299–311, Spring 1965. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01099956.

Johnson:1965:FSB

[Joh65b] Harry G. Johnson. Federal support of basic research: Some economic issues. Minerva, 3(4):500–514, Summer 1965. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01099414.

Johnson:1966:EBD

[Joh66] Harry G. Johnson. The economics of the “brain drain”. Minerva, 4(2):273–274, Winter 1966. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01584861.

Johnson:1967:ICH

[Joh67] Harry G. Johnson. The international circulation of human capital. Minerva, 6(1):105–112, Autumn 1967. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01101611.

Johnson:1970:ESD

[Joh70] Harry G. Johnson. Emigration and the supply and demand for medical manpower: The Irish case. Minerva, 8(1–4): 144–146, January 1970. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01553205.

Johnson:1971:CSG

[Joh71a] Harry G. Johnson. Comments on Senator Grosart’s paper. Minerva, 9(4):544–547, October 1971. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01558025.

Johnson:1971:EBB

[Joh71b] Harry G. Johnson. The economic benefits of basic research. Minerva, 9(2):291–293, April 1971. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01553161.

Johnson:1972:DAP

[Joh72a] Harry G. Johnson. The dilution of academic power in Canada. Minerva, 10(3):486–490, July 1972. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01556928.

Johnson:1972:STS

[Joh72b] Harry G. Johnson. Science and trans-science. Minerva, 10 (3):484–486, July 1972. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01556927.

Johnson:1972:SEA

[Joh72c] Harry G. Johnson. Some economic aspects of science. Minerva, 10(1):10–18, January 1972. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01881388.

Johnson:1973:UCU

[Joh73a] Harry G. Johnson. The uneasy case for universal graduate programmes in economics. Minerva, 11(2):263–268, April 1973. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01557720.

Johnson:1973:USW

[Joh73b] Harry G. Johnson. The university and social welfare. Minerva, 11(1):30–52, January 1973. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01544976.

Johnson:1974:EGW

[Joh74a] Harry G. Johnson. Economic growth and welfare. Minerva, 12 (1):115–116, January 1974. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01558543.

Johnson:1974:ORU

[Joh74b] Harry G. Johnson. Observations on the role of the University in development planning. Minerva, 12(1):32–38, January 1974. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01558539.

Johnson:1975:LL

[Joh75] Harry G. Johnson. Learning and libraries. Minerva, 13(4): 621–632, Winter 1975. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096179.

Johnson:1985:ASR

[Joh85] Jeffrey A. Johnson. Academic self-regulation and the chemical profession in imperial Germany. Minerva, 23(2):241–271, Summer 1985. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01099944.

Johnson:2002:CWG

[Joh02] Jeffrey Allan Johnson. Chemical warfare in the great war. Minerva, 40(1):93–106, ???? 2002. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1023/A%3A1015065730997.

Johnston:2009:IDA

[Joh09] Sean F. Johnston. Implanting a discipline: The academic trajectory of nuclear engineering in the USA and UK. Minerva, 47(1):51–73, March 2009. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-009-9114-6.

Johnston:2015:BRA

[Joh15] Sean F. Johnston. Book review: Anne Marcovich and Terry Shinn, Toward a New Dimension: Exploring the Nanoscale. Minerva, 53(4):431–434, December 2015. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
015-9282-5.

Johnes:2019:IRL

[Joh19] Geraint Johnes. The incidence of and returns to ‘overeducation’: PIAAC evidence on the G7. Minerva, 57(1):85–107, March 2019. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s11024-018-9357-1;   http://link.springer.com/content/
pdf/10.1007/s11024-018-9357-1.pdf.

Jones:1966:SSE

[Jon66] R. V. Jones. Scientists and statesmen: The example of Henry tizard. Minerva, 4(2):202–214, Winter 1966. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01584855.

Jones:1971:SIC

[Jon71] R. V. Jones. Scientists as an inferior class. Minerva, 9(1): 136–140, January 1971. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01553193.

Jones:1972:ITR

[Jon72] R. V. Jones. II temptations and risks of the scientific adviser. Minerva, 10(3):441–451, July 1972. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01556925.

Jones:1976:KPT

[Jon76] R. V. Jones. Knowledge and power: Thoughts on intelligence in a democracy. Minerva, 14(2):241–250, Summer 1976. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01107888.

Jordan:1970:STR

[Jor70] Lloyd Jordan. Scientific and technical relations among Eastern European communist countries. Minerva, 8(1–4):376– 395, January 1970. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01553224;   http://
link.springer.com/article/10.1007/BF01553224.

Josephson:1988:SPS

[Jos88] Paul R. Josephson. Science policy in the Soviet Union, 1917–1927. Minerva, 26(3):342–369, Autumn 1988. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01096401.

Josephson:1990:BR

[Jos90] Paul R. Josephson. Book reviews. Minerva, 28(2):248–257, Summer 1990. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02219661.

Josephson:1994:RSI

[Jos94] Paul R. Josephson. Russian scientific institutions: Internationalisation, democracy and dispersion. Minerva, 32(1):1–24, Spring 1994. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096654.

Johnson:1965:PTB

[JS65] Harry G. Johnson and Richard Stone. Productivity trends in British university education. Minerva, 4(1):95–105, Autumn 1965. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01585987.

Judson:1980:RHM

[Jud80] Horace Freeland Judson. Reflections on the historiography of molecular biology. Minerva, 18(3):369–421, Autumn 1980. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096950.

Julius:1967:SPN

[Jul67] H. W. Julius. Scientific policy in the Netherlands. Minerva, 5 (4):507–519, Summer 1967. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096780.

Jung:2022:IDE

[Jun22] Jisun Jung. Imagining doctoral education in the fourth industrial revolution: Driving technology or being driven by technology. Minerva, 60(4):615–632, December 2022. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL https://

link.springer.com/article/10.1007/s11024-022-09473-w.

Kaataja:2011:URC

[Kaa11] Sampsa Kaataja. University researchers contributing to technology markets 1900–85. A long-term analysis of academic patenting in Finland. Minerva, 49(4):447–460, December 2011. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-011-9185-z.

Kahn:2016:MSA

[Kah16] Michael Kahn. Mitigating south Africa’s HIV epidemic: The interplay of social entrepreneurship and the innovation system. Minerva, 54(2):129–150, June 2016. CODEN MINEFY.

ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
016-9293-x.

Kaiserfeld:2013:WNH

[Kai13] Thomas Kaiserfeld. Why new hybrid organizations are formed: Historical perspectives on epistemic and academic drift. Minerva, 51(2):171–194, June 2013. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s11024-013-9226-x.pdf.

Kakiuchi:1969:PPJ

[Kak69] Yoshinobu Kakiuchi. Policy and politics in Japanese science: The persistence of a tradition. Minerva, 7(4):758–760, Summer 1969. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01099547.

Kaldewey:2018:GCD

[Kal18a] David Kaldewey. The grand challenges discourse: Transforming identity work in science and science policy. Minerva, 56(2):161–182, June 2018. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/article/10.1007/s11024-017-9332-2;
 http://link.springer.com/content/pdf/10.1007/s11024-
017-9332-2.pdf.

Kaltenbrunner:2018:SKP

[Kal18b] Wolfgang Kaltenbrunner. Situated knowledge production, international impact: Changing publishing practices in a German engineering department. Minerva, 56(3):283–303, September 2018. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s11024-017-9337-x.

Kamara:2009:TGA

[Kam09] Mercy W. Kamara. The typology of the game that American, British, and Danish crop and plant scientists play. Minerva, 47(4):441–463, December 2009. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
009-9135-1.

Kapitza:1966:SPUa

[Kap66a] Peter L. Kapitza. Scientific policy in the U.S.S.R. Minerva, 4 (3):391–397, Spring 1966. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01584634.

Kapitza:1966:SPUb

[Kap66b] Peter L. Kapitza. Scientific policy in the U.S.S.R. Minerva, 4 (4):555–560, Summer 1966. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF02207983.

Karve:1963:UPI

[Kar63] D. D. Karve. The universities and the public in India. Minerva, 1(3):263–284, Spring 1963. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF02251984.

Karve:1965:IIU

[Kar65] D. D. Karve. On the improvement of the Indian universities. Minerva, 3(2):159–171, Winter 1965. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01096332.

Karl:1997:PMD

[Kar97] Barry D. Karl. Philanthropy and the maintenance of democratic elites. Minerva, 35(3):207–220, September 1997. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1023/A%3A1004286118726.

Kargon:2006:WMH

[Kar06] Robert Kargon. Why Mit’s history matters. Minerva, 44(4): 468–471, December 2006. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-006-9008-9.

Karakas:2024:PGI

[Kar24] Öznur Karakaş. The persistence of gender inequality in escience: The case of eSec. Minerva, 62(4):611–634, December 2024. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL https://link.springer.com/article/10.
1007/s11024-024-09530-6.

Kaysen:1966:BRN

[Kay66a] Carl Kaysen. Basic research and national goals. Minerva, 4 (2):254–272, Winter 1966. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01584860.

Kaysen:1966:CSC

[Kay66b] Carl Kaysen. The complexity of scientific choice. Minerva, 4 (3):398–400, Spring 1966. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01584635.

Kaynard:1984:UTT

[Kay84] Samuel M. Kaynard. University teachers and trade unions. Minerva, 22(1):96–115, Spring 1984. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF02207558.

Kay:1997:RIP

[Kay97] Lily E. Kay. Rethinking institutions: Philanthropy as an historiographic problem of knowledge and power. Minerva, 35(3): 283–293, September 1997. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1004238202796.

Kay:2012:OCU

[Kay12] Luciano Kay. Opportunities and challenges in the use of innovation prizes as a government policy instrument. Minerva, 50 (2):191–196, June 2012. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-012-9198-2.

Krueck:1999:SPC

[KB99] Carsten P. Krueck and Jutta Borchers. Science in politics: a comparison of climate modelling centres. Minerva, 37(2): 105–123, June 1999. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1004692000447.

Kastenhofer:2023:YPT

[KB23] Karen Kastenhofer and Anja Bauer. “Are you a TA practitioner, then?” — identity constructions in post-normal science. Minerva, 61(1):93–115, March 2023. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL https://link.

springer.com/article/10.1007/s11024-022-09480-x.

Kuhn:1972:BR

[KBD72] Thomas S. Kuhn and Joseph Ben-David. Book reviews. Minerva, 10(1):166–187, January 1972. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01881397.

Katz:1975:SRA

[KBD75] Shaul Katz and Joseph Ben-David. Scientific research and agricultural innovation in Israel. Minerva, 13(2):152–182, Summer 1975. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01097793.

Krucken:2013:MTH

[KBK13] Georg Krücken, Albrecht Blümel, and Katharina Kloke. The managerial turn in higher education? On the interplay of organizational and occupational change in German academia. Minerva, 51(4):417–442, December 2013. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
013-9240-z.

Koch:2021:SCC

[KBWA21] Tomás Koch, Gustavo Blanco-Wells, and Ricardo A. Ayala. Scholarly communities at the crossroads: Internationalizing sociological networks in Valparaíso, Chile (2003–2019). Minerva, 59(1):99–122, March 2021. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/article/10.1007/s11024-020-09421-6.

Kemelgor:2001:OIW

[KE01] Carol Kemelgor and Henry Etzkowitz. Overcoming isolation: Women’s dilemmas in American academic science. Minerva, 39 (2):153–174, ???? 2001. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1010344929577.

Kearnes:2013:GLN

[Kea13] Matthew Kearnes. On guidebooks, lists and nanotechnology. Minerva, 51(4):513–519, December 2013. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
013-9241-y.

Kedourie:1993:BUUa

[Ked93a] Elie Kedourie. The British universities under duress: Two essays by Professor Elie Kedourie. Minerva, 31(1):56–75, Spring 1993. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096172.

Kedourie:1993:BUUb

[Ked93b] Elie Kedourie. The British universities under duress: Two essays by Professor Elie Kedourie. Minerva, 31(1):76–105, Spring 1993. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096173.

Keith:1981:IPC

[Kei81] S. T. Keith. Inventions, patents and commercial development from governmentally financed research in Great Britain: The origins of the national research development corporation. Minerva, 19(1):92–122, Spring 1981. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF02192550.

Keller:1980:BR

[Kel80] A. G. Keller. Book reviews. Minerva, 18(4):652–671, Winter 1980. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096129.

Keller:1984:SCT

[Kel84] Alexander Keller. Has science created technology? Minerva, 22 (2):160–182, Summer 1984. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF02207947.

Keller:1985:MMO

[Kel85] Alexander Keller. Mathematics, mechanics and the origins of the culture of mechanical invention. Minerva, 23(3):348– 361, Autumn 1985. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096443.

Kemple:2006:UOU

[Kem06] Thomas M. Kemple. ’unfashionable observations’ on the use and abuse of Weber. Minerva, 44(3):325–337, June 2006. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-006-9004-0.

Kemple:2008:RRR

[Kem08] Thomas M. Kemple. Review: Re-reading Max Weber. Minerva, 46(3):385–389, September 2008. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
008-9104-0.

Kerr:1989:AEU

[Ker89] Clark Kerr. The academic ethic and university teachers: A “disintegrating profession”? Minerva, 27(2–3):139–156, Summer–Autumn 1989. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01102565.

Kennedy:1972:DAB

[KHB^{+}72] Donald Kennedy, David A. Hamburg, G. L. Bach, Robert McAfee Brown, Sanford M. Dornbusch, David M. Mason, and Wolfgang K. H. Panofsky. Decision of the advisory board of Stanford University in the matter of Professor H. Bruce Franklin, 5 January, 1972. Minerva, 10(3):452–483, July 1972. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01556926.

Kim:2019:RKU

[KHCJ19] Bo Kyoung Kim, Hokyu Hwang, Hee Jin Cho, and Yong Suk Jang. The rationalization of Korean universities. Minerva, 57(4): 501–521, December 2019. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/article/10.1007/s11024-019-09377-2.

Kohoutek:2024:AIU

[KHS24] Jan Kohoutek, Karel Hanuš, and Marián Sekerák. Academic inbreeding at universities in the Czech Republic: Beyond immobile inbred employees? Minerva, 62(2):287–304, June 2024. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL https://link.springer.com/article/10.
1007/s11024-023-09515-x.

Kidd:1965:EBD

[Kid65] Charles V. Kidd. The economics of the “brain drain”. Minerva, 4 (1):105–107, Autumn 1965. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01585988.

Kidwai:1966:CBU

[Kid66] A. J. Kidwai. Collaboration between universities and government laboratories. Minerva, 4(3):402–406, Spring 1966. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01584637.

Kidd:1983:MYS

[Kid83] Charles V. Kidd. The movement of younger scientists into and out of the United States from 1967 to 1980: Some aspects of the international movement of scientific knowledge. Minerva, 21 (4):387–409, Winter 1983. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096524.

Kidwell:2007:WM

[Kid07] Peggy Aldrich Kidwell. Women in mathematics. Minerva, 45(3): 353–356, September 2007. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-007-9053-z.

Kielmansegg:1985:TAE

[Kie85] Peter Graf Kielmansegg. Is there an academic ethic? Minerva, 23(4):504–507, Winter 1985. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01098703.

Kim:1982:NKT

[Kim82] Yung Sik Kim. Natural knowledge in a traditional culture: Problems in the study of the history of Chinese science. Minerva, 20 (1–2):83–104, Spring–Summer 1982. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01098191.

Kim:2008:DEH

[Kim08] Eun-Sung Kim. Directed evolution: a historical exploration into an evolutionary experimental system of nanobiotechnology, 1965–2006. Minerva, 46(4):??, December 2008. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
s11024-008-9108-9.

King:1963:HEP

[Kin63] Alexander King. Higher education, professional manpower and the state: Reflections on education and professional employment in the U.S.S.R. Minerva, 1(2):182–190, Winter 1963. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096250.

King:1997:CFR

[Kin97] Desmond King. Creating a funding regime for social research in Britain: The Heyworth Committee on Social Studies and the Founding of the Social Science Research Council. Minerva, 35(1):1–26, ???? 1997. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1004276108684.

Kissi:2024:BRC

[Kis24] Foster Oduro Kissi. Book review: Clapperton Chakanetsa Mavhunga, What Do Science, Technology and Innovation Mean from Africa? The MIT Press, 2017, 241pp., ISBN 978-0-262-34232-2. Minerva, 62(4):635–638, December 2024. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (elec-

tronic).  URL https://link.springer.com/article/10.
1007/s11024-024-09549-9.

Kitagawa:2010:PRE

[Kit10] Fumi Kitagawa. Pooling resources for excellence and relevance: An evolution of universities as multi-scalar network organisations. Minerva, 48(2):169–187, June 2010. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
010-9147-x.

Karl:1981:APP

[KK81] Barry D. Karl and Stanley N. Katz. The American private philanthropic foundation and the public sphere 1890–1930. Minerva, 19(2):236–270, Summer 1981. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01096567.

Kauffman:1995:BR

[KKF95] George B. Kauffman, Laurie M. Kauffman, and Stephen Fortescue. Book reviews. Minerva, 33(1):99–105, Spring 1995. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01098622.

Kodate:2010:MCC

[KKK10] Naonori Kodate, Kashiko Kodate, and Takako Kodate. Mission completed? Changing visibility of Women’s colleges in England and Japan and their roles in promoting gender equality in science. Minerva, 48(3):309–330, September 2010. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-010-9150-2.

Kargon:1994:IGP

[KL94] Robert Kargon and Stuart Leslie. Imagined geographies: Princeton, Stanford and the boundaries of useful knowledge in postwar America. Minerva, 32(2):121–143, Summer 1994. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01103354.

Kreimer:2003:PVB

[KL03] Pablo Kreimer and Manuel Lugones. Pioneers and victims: The birth and death of Argentina’s first molecular biology laboratory. Minerva, 41(1):47–69, ???? 2003. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1023/
A%3A1022209922391.

Klein:2023:BDC

[Kle23] Julie Thompson Klein. Boundary discourse of crossdisciplinary and cross-sector research: Refiguring the landscape of science. Minerva, 61(1):31–52, March 2023. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL https://link.

springer.com/article/10.1007/s11024-022-09474-9.

Klenk:2025:PSM

[Kle25] Nicole Klenk. From problem-solving to meaning-making: Reframing the societal relevance of research. Minerva, 63(2):409– 422, June 2025. CODEN MINEFY. ISSN 0026-4695 (print),

1573-1871 (electronic). URL https://link.springer.com/
article/10.1007/s11024-025-09583-1.

Kloss:1968:URW

[Klo68] G. Kloss. University reform in West Germany. Minerva, 6 (3):323–353, Spring 1968. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096421.

Kloss:1971:GFP

[Klo71] G. Kloss. The growth of federal power in the West German University system. Minerva, 9(4):510–527, October 1971. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01558022.

Kladakis:2024:CEP

[KMB24] Alexander Kladakis, Philippe Mongeon, and Carter W. Bloch. Citation elites in polytheistic and umbrella disciplines: Patterns of stratification and concentration in Danish and British science. Minerva, 62(3):397–426, September 2024. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL https://

link.springer.com/article/10.1007/s11024-024-09521-7.

Kuzhabekova:2024:RRS

[KMBC24] Aliya Kuzhabekova, Kairat Moldashev, Altyn Baigazina, and Vichny Chanchem. Readjustment of returning scholars: Experiences of Cambodian researchers. Minerva, 62(3):457–479, September 2024. CODEN MINEFY. ISSN 0026-4695 (print),

1573-1871 (electronic). URL https://link.springer.com/
article/10.1007/s11024-024-09522-6.

Koertge:2007:SW

[Koe07] Noretta Koertge. The science wars. Minerva, 45(1):105–111, March 2007. CODEN MINEFY. ISSN 0026-4695 (print),

1573-1871 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11024-006-9020-0.

Kogan:1997:DHE

[Kog97] Maurice Kogan. Diversification in higher education: Differences and commonalities. Minerva, 35(1):47–62, ???? 1997. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1004285524614.

Kogan:1999:CA

[Kog99] Maurice Kogan. The culture of academe. Minerva, 37(1): 63–74, March 1999. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1004532121808.

Kohler:1976:MSE

[Koh76] Robert E. Kohler. The management of science: The experience of Warren Weaver and the Rockefeller Foundation programme in molecular biology. Minerva, 14(3):279–306, Autumn 1976. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096274.

Kohler:1978:PAS

[Koh78] Robert E. Kohler. A policy for the advancement of science: The Rockefeller Foundation, 1924–29. Minerva, 16(4):480–515, Winter 1978. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01100330.

Kohler:1985:SPW

[Koh85] Robert E. Kohler. Science and philanthropy: Wickliffe rose and the international education board. Minerva, 23(1):75–95, Spring 1985. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01097841.

Kohler:1993:SEP

[Koh93] Robert E. Kohler. Sanitarians, engineers and public science in the gilded age. Minerva, 31(2):184–210, Summer 1993. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01125434.

Koizumi:1995:DIT

[Koi95] Kenkichiro Koizumi. The development of industrial technology in Japan: Will versus absorptive capacity. Minerva, 33 (1):19–35, Spring 1995. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01098617.

Kolman:1978:LTS

[Kol78] Arnost Kolman. A life-time in Soviet science reconsidered: The adventure of cybernetics in the Soviet Union. Minerva, 16(3): 416–424, Autumn 1978. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096669.

Kolinsky:1983:DIU

[Kol83] Martin Kolinsky. The demise of the inter–University council for higher education overseas: a chapter in the history of the idea of the university. Minerva, 21(1):37–80, Spring 1983. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01099503.

Kolinsky:1985:GNU

[Kol85] Martin Kolinsky. The growth of Nigerian universities 1948– 1980: The British share. Minerva, 23(1):29–61, Spring 1985. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01097839.

Koloski:2004:MVT

[Kol04] Laurie Koloski. More variations than theme: The Sovietization of Eastern European universities. Minerva, 42(3):309–314, ???? 2004. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/B%3AMINE.0000038301.39533.11.

Konig:2004:AES

[Kön04] Wolfgang König. The Academy and the engineering sciences: An unwelcome royal gift. Minerva, 42(4):359–377, ???? 2004. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-004-2111-x.

Konig:2007:DDM

[Kön07] Ariane König. Democratizing decision-making on food safety in the EU: Closing gaps between principles of governance and practice. Minerva, 45(3):275–294, September 2007. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
s11024-007-9044-0.

Kornhauser:1962:SAI

[Kor62] William Kornhauser. Strains and accommodations in industrial research organisations in the United States. Minerva, 1 (1):30–42, Autumn 1962. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01101451.

Kordon:1979:TVS

[Kor79] Claude Kordon. Two views of the state of the universities and of scientific research in France in the late 1970s. Minerva, 17 (2):305–312, Summer 1979. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096285.

Korsmo:2006:MST

[Kor06] Fae L. Korsmo. Marine science and technology: History, politics, and people. Minerva, 44(1):119–123, March 2006. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/s11024-005-5396-5.

Korom:2020:HDA

[Kor20] Philipp Korom. How do academic elites march through departments? A comparison of the most eminent economists and sociologists’ career trajectories. Minerva, 58(3):343–365, September 2020. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/article/10.
1007/s11024-020-09399-1;   http://link.springer.com/
content/pdf/10.1007/s11024-020-09399-1.pdf.

Kleinman:2014:UCC

[KOT14] Daniel Lee Kleinman and Robert Osley-Thomas. Uneven commercialization: Contradiction and conflict in the identity and practices of American universities. Minerva, 52(1):1–26, March 2014. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-014-9248-z.

Kayrooz:2002:AFI

[KP02] Carole Kayrooz and Paul Preston. Academic freedom: Impressions of Australian social scientists. Minerva, 40(4):341– 358, ???? 2002. CODEN MINEFY. ISSN 0026-4695 (print),

1573-1871 (electronic). URL http://link.springer.com/
accesspage/article/10.1023/A%3A1020925721453.

Kennedy:1989:DAP

[KPL^{+}89] Donald Kennedy, John Perky, Carolyn Lougee, Marsh McCall, Paul Robinson, James Gibb, Clara N. Bush, Judith Brown, George Dekker, Bill King, William Chace, Carlos Camargo, J. Martin Evans, Ronald Rebholz, Carl Degler, and et al. The discussion about proposals to change the Western Culture program at Stanford University. Minerva, 27(2–3):223–411, Summer–Autumn 1989. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01102569.

Krauch:2006:BSP

[Kra06] Helmut Krauch. Beginning science policy research in Europe: The Studiengruppe für Systemforschung, 1957–1973. Minerva, 44(2):131–142, June 2006. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-006-0002-z.

Krasnodebski:2018:BPP

[Kra18] Marcin Krasnodebski. Beyond private and public research: The legal and organizational reality behind industrial research institutes in interwar France. Minerva, 56(3):333–355, September 2018. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/article/10.
1007/s11024-018-9345-5.

Kristeller:1978:H

[Kri78] Paul Oskar Kristeller. Humanism. Minerva, 16(4):586–595, Winter 1978. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01100334.

Kristeller:1980:HSP

[Kri80] Paul Oskar Kristeller. I historical scholarship and philosophical thought. Minerva, 18(2):313–323, Summer 1980. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01100251.

Krige:1997:BRI

[Kri97] John Krige. Book review: Inside NASA: High Technology and Organizational Change in the U.S. Space Program, by H. E. McCurdy. Minerva, 35(4):397–399, December 1997. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1004371002862.

Krige:2000:NSW

[Kri00] John Krige. NATO and the strengthening of Western science in the post-Sputnik era. Minerva, 38(1):81–108, ???? 2000. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1026549723657.

Krige:2001:NPP

[Kri01] John Krige. The 1984 Nobel Physics Prize for heterogeneous engineering. Minerva, 39(4):425–443, ???? 2001. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1023/
A%3A1012735919366.

Krige:2007:DN

[Kri07] John Krige. Defending the nation. Minerva, 45(4):489– 493, December 2007. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-007-9062-y.

Krick:2022:PGP

[Kri22] Eva Krick. Participatory governance practices at the democracy– knowledge-nexus. Minerva, 60(4):467–487, December 2022. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (elec-

tronic).  URL https://link.springer.com/article/10.
1007/s11024-022-09470-z.

Kroll:2003:PSB

[Kro03] Gary Kroll. The Pacific science board in Micronesia: Science, government, and conservation on the post-war Pacific frontier. Minerva, 41(1):25–46, ???? 2003. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1023/
A%3A1022205821483.

Krohn:2023:BRH

[Kro23] Wolfgang Krohn. Book review: Hans-Jörg Rheinberger, Split and Splice. A Phenomenology of Experimentation. Minerva, 61(4):635–637, December 2023. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL https://link.

springer.com/article/10.1007/s11024-023-09509-9.

Krull:2004:TRP

[Kru04] Wilhelm Krull. Towards a research policy for the new Europe: Changes and challenges for public and private funders. Minerva, 42(1):29–39, ???? 2004. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1023/
B%3AMINE.0000017697.58827.c8.

Krull:2005:EHU

[Kru05] Wilhelm Krull. Exporting the Humboldtian University. Minerva, 43(1):99–102, ???? 2005. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-004-6633-z.

Krull:2008:RPP

[Kru08] Wilhelm Krull. Review: Past–present–future: The ETH Zurich. Minerva, 46(2):271–274, June 2008. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
008-9098-7.

Krementsov:2001:GTA

[KS01] Nikolai Krementsov and Susan Gross Solomon. Giving and taking across borders: The Rockefeller Foundation and Russia, 1919–1928. Minerva, 39(3):265–298, ???? 2001. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1023/A%3A1017909113410.

Karve:1964:HEI

[KSC^{+}64] D. D. Karve, A. B. Shah, C. F. Carter, Alvin M. Weinberg, E. Barton Worthington, and D. Odhiambo. Higher education in India. Minerva, 2(3):379–388, Spring 1964. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01097325.

Kuhner:1984:BAP

[Küh84] Hans Kühner. Between autonomy and planning: The Chinese Academy of Sciences in transition. Minerva, 22(1):13–44, Spring 1984. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02207555.

Kuklick:2011:SMP

[Kuk11] Henrika Kuklick. Stuart Macintyre, the poor relation. A history of social sciences in Australia. Minerva, 49(3):??, September 2011. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-011-9173-3.

Kuwahara:2001:JWS

[Kuw01] Motoko Kuwahara. Japanese women in science and technology. Minerva, 39(2):203–216, ???? 2001. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1023/
A%3A1010332611830.

Koch:2019:LLA

[KV19] Natalie Koch and Neha Vora. Laboratories of liberalism: American higher education in the Arabian Peninsula and the discursive production of authoritarianism. Minerva, 57(4):549–564, December 2019. CODEN MINEFY. ISSN 0026-4695 (print), 1573-

1871 (electronic). URL http://link.springer.com/article/
10.1007/s11024-019-09382-5.

Koch:2025:AMF

[KV25] Susanne Koch and Judit Varga. All that matters are forests and seas? practising relevance in interdisciplinary environmentfocused social science fields. Minerva, 63(2):301–324, June 2025. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL https://link.springer.com/article/10.
1007/s11024-024-09556-w.

Kearnes:2011:TTU

[KW11] Matthew Kearnes and Matthias Wienroth. Tools of the trade: UK research intermediaries and the politics of impacts. Minerva, 49(2):153–174, June 2011. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-011-9172-4.

Kirkman:1974:BR

[KZP^{+}74] W. P. Kirkman, John Ziman, Keith Pavitt, Dietrich Goldschmidt, Ulrich Teichler, E. Barton Worthington, and P. B. Medawar. Book reviews. Minerva, 12(3):382–400, July 1974. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01102531.

Laboulais:2008:SSS

[Lab08] Isabelle Laboulais. Serving science and the state: Mining science in France, 1794–1810. Minerva, 46(1):17–36, March 2008. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-008-9085-z.

Ladd:1970:AUT

[Lad70] Everett Carll Ladd. American university teachers and opposition to the Vietnam war. Minerva, 8(1–4):542– 556, January 1970. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01553238;   http://
link.springer.com/article/10.1007/BF01553238.

Lagemann:1987:PFW

[Lag87] Ellen Condliffe Lagemann. A philanthropic foundation at work: Gunnar Myrdal’s American dilemma and the Carnegie corporation. Minerva, 25(4):441–470, Winter 1987. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01099442.

Laidlaw:2019:PSR

[Lai19] Tess Laidlaw. Pandemic stories: Rhetorical motifs in journalists’ coverage of biomedical risk. Minerva, 57(4):433–451, December 2019. CODEN MINEFY. ISSN 0026-4695 (print), 1573-

1871 (electronic). URL http://link.springer.com/article/
10.1007/s11024-019-09383-4.

Laird:2020:SPD

[Lai20] Frank N. Laird. Sticky policies, dysfunctional systems: Path dependency and the problems of government funding for science in the United States. Minerva, 58(4):513–533, December 2020. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/article/10.
1007/s11024-020-09409-2;   http://link.springer.com/
content/pdf/10.1007/s11024-020-09409-2.pdf.

Lakhtin:1968:ORM

[Lak68] George A. Lakhtin. Operational research methods in the management of scientific research. Minerva, 6(4):524–540, Summer 1968. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096546.

Lakoff:1973:VAS

[Lak73] Sanford A. Lakoff. The vicissitudes of American science policy at home and abroad. Minerva, 11(2):175–190, April 1973. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01557715.

Lankford:1987:PPG

[Lan87] John Lankford. Private patronage and the growth of knowledge: The J. Lawrence Smith fund of the National Academy of Sciences, 1884–1940. Minerva, 25(3):269–281, Autumn 1987. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01097785.

Lane:2005:LDP

[Lan05] Jonathon Lane. Loyalty, democracy and the public intellectual. Minerva, 43(1):73–85, ???? 2005. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
004-6627-x.

Logar:2014:SRC

[LAN14] Nathaniel Logar, Laura Diaz Anadon, and Venkatesh Narayanamurti. Semiconductor research corporation: a case study in cooperative innovation partnerships. Minerva, 52(2):237–261, June 2014. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-014-9253-2.

Laudel:2005:MCA

[Lau05] Grit Laudel. Migration currents among the scientific elite. Minerva, 43(4):377–395, December 2005. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
005-2474-7.

Laudel:2017:HDN

[Lau17] Grit Laudel. How do national career systems promote or hinder the emergence of new research lines? Minerva, 55(3): 341–369, September 2017. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/article/10.1007/s11024-017-9314-4.

Lawless:1981:CUU

[Law81] David J. Lawless. The Canadian university under the impact of academic trade unions. Minerva, 19(3):464–479, Autumn 1981. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02192825.

Layard:1973:NMH

[Lay73] Richard Layard. The new media and higher education. Minerva, 11(2):211–227, April 1973. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01557717.

Lepori:2013:SPC

[LB13] Benedetto Lepori and Andrea Bonaccorsi. The socio-political construction of a European census of higher education institutions: Design, methodological and comparability issues. Minerva, 51(3):271–293, September 2013. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
013-9235-9.

Lewis:2016:HMC

[LBA16] Jamie Lewis, Andrew Bartlett, and Paul Atkinson. Hidden in the middle: Culture, value and reward in bioinformatics. Minerva, 54(4):471–490, December 2016. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s11024-016-9304-y.pdf.

Lepori:2010:PSM

[LBP10] Benedetto Lepori, Lukas Baschung, and Carole Probst. Patterns of subject mix in higher education institutions: a first empirical analysis using the AQUAMETH database. Minerva, 48(1):73–99, March 2010. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-010-9143-1.

Linz:1972:BR

[LC72] Juan Linz and William K. Cummings. Book reviews. Minerva, 10(4):627–638, October 1972. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01695909.

Leach:1998:HMC

[Lea98] Colin Leach. A history of Merton College, Oxford. Minerva, 36(4):392–395, December 1998. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1023/A%3A1004439300206.

Lecuyer:1987:SRS

[Léc87] Bernard-Pierre Lécuyer. The statistician’s role in society: The institutional establishment of statistics in France. Minerva, 25 (1–2):35–55, Spring–Summer 1987. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01096855.

Lecuyer:2005:WDU

[Léc05] Christophe Lécuyer. What do universities really owe industry? The case of solid state electronics at Stanford. Minerva, 43 (1):51–71, ???? 2005. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-004-6618-y.

Lee:1998:OSB

[Lee98] J. M. Lee. Overseas students in Britain: How their presence was politicised in 1966–1967. Minerva, 36(4):305–321, December 1998. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1004455800689.

Lee:2006:CSU

[Lee06] J. M. Lee. Commonwealth students in the United Kingdom, 1940–1960: Student welfare and world status. Minerva, 44 (1):1–24, March 2006. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-005-5411-x.

Lee:2007:LAP

[Lee07] J. M. Lee. Leaders for Africa: Postcolonial University reform. Minerva, 45(1):85–92, March 2007. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
006-9023-x.

Lemay:2020:RES

[Lem20] Margaret A. Lemay. The role of expectations of science in shaping research policy: a discursive analysis of the creation of Genome Canada. Minerva, 58(2):235–260, June 2020. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/article/10.1007/s11024-
020-09395-5.

Levi:1973:ETE

[Lev73] Edward H. Levi. Equality through education. Minerva, 11 (2):157–161, April 1973. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01557713.

Levy:1998:IRO

[Lev98] Daniel C. Levy. Internationalised reform: Overlapping agendas in East Asian and Latin American higher education. Minerva, 36(4):367–379, December 1998. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1023/
A%3A1004406409775.

Lewis:1977:GTS

[Lew77] Robert A. Lewis. Government and the technological sciences in the Soviet Union: The rise of the Academy of Sciences. Minerva, 15(2):174–199, Summer 1977. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01096676.

Lewy:1980:PHW

[Lew80] Guenter Lewy. The persisting heritage of the 1960s in West German higher education. Minerva, 18(1):1–28, Spring 1980. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096659.

Lewis:1984:HTI

[Lew84] Robert Lewis. Hierarchy and technological innovation in Soviet industry: The science-production associations. Minerva, 22(2): 129–159, Summer 1984. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF02207946.

Lindvig:2019:SUO

[LH19] Katrine Lindvig and Line Hillersdal. Strategically unclear? Organising interdisciplinarity in an excellence programme of interdisciplinary research in Denmark. Minerva, 57(1):23–46, March 2019. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/article/10.
1007/s11024-018-9361-5.

Lim:1974:RUD

[Lim74] David Lim. The role of the university in development planning in Malaysia. Minerva, 12(1):18–32, January 1974. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01558538.

Limerick:1995:GMG

[Lim95] Lord Limerick. Guide for members of governing bodies of universities and colleges in England and Wales. Minerva, 33(4): 373–394, Winter 1995. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096519.

Lindsay:1981:DPA

[Lin81] Alan Lindsay. The developing pattern of Australian tertiary education: an analysis and critique of three reports. Minerva, 19(1):165–184, Spring 1981. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF02192552.

Lindqvist:2001:NEC

[Lin01] Svante Lindqvist. The Nobel exhibition cultures of creativity: the centennial exhibition of the Nobel Prize, 1901– 2001. Minerva, 39(4):461–465, ???? 2001. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1023/
A%3A1012733001619.

Lipset:1964:USP

[Lip64] Seymour Martin Lipset. University students and politics in underdeveloped countries. Minerva, 3(1):15–56, Autumn 1964. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01630148.

Lisichkin:1967:SPU

[Lis67] Professor S. Lisichkin. Scientific policy in the USSR. Minerva, 5(3):387–390, Spring 1967. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF02388578.

Littlewood:2007:BWM

[Lit07] Jez Littlewood. Biological weapons: Much ado and little action. Minerva, 45(2):191–203, June 2007. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
007-9030-6.

Lloyd-Jones:1984:MN

[LJ84] Hugh Lloyd-Jones. Malign neglect. Minerva, 22(3–4):405–409, Autumn–Winter 1984. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF02207373.

Lloyd-Jones:1993:BR

[LJ93] Hugh Lloyd-Jones. Book review. Minerva, 31(3):375–378, Autumn 1993. CODEN MINEFY. ISSN 0026-4695 (print),

1573-1871 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF01098628.

Lekka-Kowalik:2022:AGW

[LK22] Agnieszka Lekka-Kowalik. Academia in the grip of the wolf and its utopia. Minerva, 60(1):139–158, March 2022. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (elec-

tronic).  URL https://link.springer.com/article/10.
1007/s11024-021-09449-2.

Lomnitz:1994:VMA

[LM94] Larissa Adler Lomnitz and Leticia Mayer. Veterinary medicine and animal husbandry in Mexico: From empiricism to science and technology. Minerva, 32(2):144–157, Summer 1994. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01103355.

Laredo:2004:PSR

[LM04] Philippe Larédo and Philippe Mustar. Public sector research: a growing role in innovation systems. Minerva, 42(1):11–27, ???? 2004. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/content/pdf/
10.1023/B%3AMINE.0000017698.38154.a6.pdf.

Langfeldt:2020:CEN

[LNST20] Liv Langfeldt, Maria Nedeva, Sverker Sörlin, and Duncan A. Thomas. Co-existing notions of research quality: a framework to study context-specific understandings of good research. Minerva, 58(1):115–137, March 2020. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/article/10.1007/s11024-
019-09385-2;   http://link.springer.com/content/pdf/
10.1007/s11024-019-09385-2.pdf.

LoPiano:2025:RIS

[Lo 25] Samuele Lo Piano. A review of Ian Scoones, Navigating Uncertainty: Radical Rethinking for a Turbulent World. Minerva, 63(3):633–636, September 2025. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL https://link.

springer.com/article/10.1007/s11024-025-09579-x.

Lobkowicz:1984:AEP

[Lob84a] Nikolaus Lobkowicz. The academic ethic: Politics and academic citizenship. Minerva, 22(2):236–244, Summer 1984. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF02207950.

Lobkowicz:1984:REY

[Lob84b] Nikolaus Lobkowicz. Reflections on eleven years as president of a German university. Minerva, 22(3–4):365–387, Autumn–Winter 1984. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02207370.

Lock:1989:CDU

[Loc89] Grahame Lock. The collectivisation of the Dutch universities. Minerva, 27(2–3):157–176, Summer–Autumn 1989. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01102566.

Logar:2009:TCA

[Log09] Nathaniel Logar. Towards a culture of application: Science and decision making at the National Institute of Standards & Technology. Minerva, 47(4):345–366, December 2009. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-009-9137-z.

Logar:2011:CGC

[Log11] Nathaniel Logar. Chemistry, green chemistry, and the instrumental valuation of sustainability. Minerva, 49(1):113–136, March 2011. CODEN MINEFY. ISSN 0026-4695 (print),

1573-1871 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11024-011-9165-3.

Lomnitz:1979:HPO

[Lom79] Larissa Lomnitz. Hierarchy and peripherality: The organisation of a Mexican research institute. Minerva, 17(4):527–548, Winter 1979. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01101902.

Long:1969:PPJ

[Lon69] T. Dixon Long. Policy and politics in Japanese science. Minerva, 7(3):426–453, Spring 1969. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01103269.

Lowy:1987:IMP

[Löw87] Ilana Löwy. The impact of medical practice on biomedical research: The case of human leucocyte antigens studies. Minerva, 25(1–2):171–200, Spring–Summer 1987. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01096863.

Lowen:1992:EWO

[Low92] Rebecca S. Lowen. “Exploiting a wonderful opportunity”: The patronage of scientific research at Stanford University, 1937–1965. Minerva, 30(3):391–421, Autumn 1992. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01097645.

Low:2001:ESN

[Low01] Morris Low. From Einstein to Shirakawa: the Nobel Prize in Japan. Minerva, 39(4):445–460, ???? 2001. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1023/
A%3A1012740020275.

Low:2005:ESJ

[Low05] Morris Low. Environmental science in Japan. Minerva, 43(4): 441–444, December 2005. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-005-2469-4.

Leprieur:1979:SDR

[LP79] François Leprieur and Pierre Papon. Synthetic dyestuffs: The relations between academic chemistry and the chemical industry in nineteenth-century France. Minerva, 17(2):197–224, Summer 1979. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096281.

Law:2021:PEB

[LS21] Angus Law and Graham Spinardi. Performing expertise in building regulation: ‘Codespeak’ and fire safety experts. Minerva, 59(4):515–538, December 2021. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL https://link.

springer.com/article/10.1007/s11024-021-09446-5.

Luukkonen:2016:NSU

[LT16] Terttu Luukkonen and Duncan A. Thomas. The ‘negotiated space’ of University researchers’ pursuit of a research agenda. Minerva, 54(1):99–127, March 2016. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
016-9291-z.

Lubenow:2002:MWF

[Lub02] William C. Lubenow. Making words flesh: Changing roles of University learning and the professions in 19th century England. Minerva, 40(3):217–234, ???? 2002. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1023/
A%3A1019568408319.

Lubenow:2004:RUA

[Lub04] W. C. Lubenow. Religion in the University: Authority, faith, and learning. Minerva, 42(3):269–283, ???? 2004. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1023/
B%3AMINE.0000038306.21128.e4.

Lubenow:2007:WSC

[Lub07] W. C. Lubenow. ‘we have such a class and they are known as p-p-p-prigs’: The anxieties of intellectuals. Minerva, 45(4): 459–474, December 2007. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-007-9055-x.

Lakoff:1983:BR

[LWZB83] Sanford Lakoff, Alvin M. Weinberg, Michael Zöller, and Martin Bulmer. Book reviews. Minerva, 21(2–3):320–347, Summer– Autumn 1983. CODEN MINEFY. ISSN 0026-4695 (print),

1573-1871 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF01097970.

Lynch:2003:BCW

[Lyn03] William T. Lynch. Beyond Cold War paradigms for science and democracy. Minerva, 41(4):365–379, ???? 2003. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1023/B%3AMINE.0000005159.50541.9b.

Lewis:1987:BR

[LZ87] Robert Lewis and J. M. Ziman. Book reviews. Minerva, 25 (4):513–522, Winter 1987. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01099447.

MacLeod:1969:CCR

[MA69] Roy M. MacLeod and E. Kay Andrews. The committee of civil research: Scientific advice for economic development 1925–30. Minerva, 7(4):680–705, Summer 1969. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01099541.

Minogue:1975:BR

[MA75] K. R. Minogue and Austen Albu. Book reviews. Minerva, 13 (2):322–336, Summer 1975. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01097801.

MacLeod:1970:SGO

[Mac70] Roy M. MacLeod. Scientists, government and organised research in Great Britain 1914–16. Minerva, 8(1–4):454– 457, January 1970. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01553232;   http://
link.springer.com/article/10.1007/BF01553232.

Macleod:1971:SVS

[Mac71] Roy M. Macleod. The support of Victorian science: The endowment of research movement in Great Britain, 1868–1900. Minerva, 9(2):197–230, April 1971. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01553156.

Macrae:1973:SFP

[Mac73] Duncan Macrae, Jr. I science and the formation of policy in a democracy. Minerva, 11(2):228–242, April 1973. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01557718.

Macrae:1976:TCP

[Mac76] Duncan Macrae, Jr. Technical communities and political choice. Minerva, 14(2):169–190, Summer 1976. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01107884.

Macdonald:1988:CFH

[Mac88] Stuart Macdonald. Controlling the flow of high-technology information from the United States to the Soviet Union: A labour of Sisyphus? Minerva, 24(1):39–73, Spring 1988. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01102553.

Macrakis:1989:RFG

[Mac89] Kristie Macrakis. The Rockefeller Foundation and German physics under national socialism. Minerva, 27(1):33–57, Spring 1989. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096714.

Macleod:1994:IMM

[Mac94] Roy Macleod. “Instructed men” and mining engineers: The associates of the Royal School of Mines and British Imperial science, 1851–1920. Minerva, 32(4):422–439, Winter 1994. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01098032.

Macleod:1997:SDH

[Mac97] Roy Macleod. Science and democracy: Historical reflections on present discontents. Minerva, 35(4):369–384, December 1997. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1004362816974.

MacLeod:1999:EII

[Mac99a] Roy MacLeod. Empire and Information: Intelligence Gathering and Social Communication in India 1780–1870, by C. A. Bayley. Minerva, 37(2):183–185, June 1999. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1023/
A%3A1004677030006.

MacLeod:1999:SAF

[Mac99b] Roy MacLeod. Secrets among friends: The research information service and the “special relationship” in Allied scientific information and intelligence, 1916–1918. Minerva, 37(3):201– 233, September 1999. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1004734120928.

Macilwain:2000:RPA

[Mac00a] Colin Macilwain. Revisiting the politics of American science. Minerva, 38(3):349–352, ???? 2000. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1023/
A%3A1026577817936.

Mackie:2000:WCE

[Mac00b] Robin Mackie. ‘but what is a chemical engineer?’: Profiling the membership of the British institution of chemical engineers, 1922–1956. Minerva, 38(2):171–199, ???? 2000. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1026549011575.

MacLeod:2001:P

[Mac01] Roy MacLeod. Preface. Minerva, 39(4):363, ???? 2001. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1023/A%3A1012792008117.

MacLeod:2002:P

[Mac02] Roy MacLeod. Preface. Minerva, 40(1):1–2, ???? 2002. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1023/A%3A1015060525546.

MacLeod:2003:P

[Mac03] Roy MacLeod. Preface. Minerva, 41(3):177–178, ???? 2003. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1025525018703.

MacLeod:2004:Pa

[Mac04a] Roy MacLeod. Preface. Minerva, 42(1):1, ???? 2004. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1023/B%3AMINE.0000017835.57509.ee.

MacLeod:2004:Pb

[Mac04b] Roy MacLeod. Preface. Minerva, 42(4):321, ???? 2004. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/s11024-004-2109-4.

MacLeod:2005:P

[Mac05] Roy MacLeod. Preface. Minerva, 43(2):119–120, June 2005. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-005-2325-6.

MacLeod:2006:Pa

[Mac06a] Roy MacLeod. Preface. Minerva, 44(2):129, June 2006. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/s11024-006-0001-0.

MacLeod:2006:Pb

[Mac06b] Roy MacLeod. Preface. Minerva, 44(3):239–240, June 2006. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-006-9006-y.

MacLeod:2007:AAV

[Mac07a] Roy MacLeod. Ave atque vale. Minerva, 45(4):361–363, December 2007. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-007-9068-5.

MacLeod:2007:Pa

[Mac07b] Roy MacLeod. Preface. Minerva, 45(1):1–2, March 2007. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/s11024-006-9027-6.

MacLeod:2007:Pb

[Mac07c] Roy MacLeod. Preface. Minerva, 45(2):113–114, June 2007. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-007-9038-y.

Macleod:2008:BMP

[Mac08] Roy Macleod. Balfour’s mission to Palestine: Science, strategy, and the inauguration of the Hebrew University in Jerusalem. Minerva, 46(1):53–76, March 2008. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
008-9087-x.

MacLeod:2016:CCC

[Mac16] Roy MacLeod. Consensus, civility, and community: The origins of Minerva and the vision of Edward Shils. Minerva, 54(3): 255–292, September 2016. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-016-9305-x.

Maddox:1964:CSC

[Mad64] John Maddox. Choice and the scientific community. Minerva, 2 (2):141–159, Winter 1964. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096591.

Madsen:2018:EPU

[Mad18] Dorte Madsen. Epistemological or political? unpacking ambiguities in the field of interdisciplinarity studies. Minerva, 56(4):

453–477, December 2018. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/article/10.1007/s11024-018-9353-5.

Maheshwari:1964:ISP

[Mah64] P. Maheshwari. Indian scientific policy. Minerva, 3(1):99– 113, Autumn 1964. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01630153.

Mahroum:2000:GMS

[Mah00] Sami Mahroum. Global magnets: Science and technology disciplines and departments in the United Kingdom. Minerva, 37(4): 379–390, December 2000. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1004749905956.

Mardin:1968:UT

[Mar69] Serif Mardin. Universities in Turkey. Minerva, 7(1–2):91– 92, Autumn–Winter 1968–1969. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01096814.

Marcus:1988:RUS

[Mar88] Alfred A. Marcus. Risk, uncertainty, and scientific judgement. Minerva, 26(2):138–152, Summer 1988. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01096693.

Marginson:2006:AAU

[Mar06] Simon Marginson. The Anglo–American University at its global high tide. Minerva, 44(1):65–87, March 2006. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/s11024-005-5399-2.

Maricle:2011:PIP

[Mar11] Genevieve E. Maricle. Prediction as an impediment to preparedness: Lessons from the US hurricane and earthquake research enterprises. Minerva, 49(1):87–111, March 2011. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-011-9166-2.

Martin:2020:BRT

[Mar20] Joseph D. Martin. Book review: Thomas C. Lassman, Edward Condon’s Cooperative Vision: Science, Industry, and Innovation in Modern America. University of Pittsburgh Press, Pittsburgh, 2018, 320 pp., $49.00 (hardcover). Minerva, 58(2):315–317, June 2020. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/article/10.
1007/s11024-020-09406-5.

Martin-Alcazar:2020:DCM

[MARMSG20] Fernando Martín-Alcázar, Marta Ruiz-Martínez, and Gonzalo Sánchez-Gardey. Deepening the consequences of multidisciplinarity on research: The moderating role of social capital. Minerva, 58(4):559–583, December 2020. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/article/10.1007/s11024-020-09404-7.

Matejko:1969:PTP

[Mat69] Aleksander Matejko. Planning and tradition in Polish higher education. Minerva, 7(4):621–648, Summer 1969. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01099538.

Matsumura:2008:AKJ

[Mat08] Janice Matsumura. The “all-knowing’ Japanese state? New scholarship on medicine, science, technology, and industrial relations. Minerva, 46(1):77–86, March 2008. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
007-9069-4.

Mazur:1973:IDB

[Maz73] Allan Mazur. II disputes between experts. Minerva, 11(2): 243–262, April 1973. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01557719.

Mazur:1975:OTI

[Maz75] Allan Mazur. Opposition to technological innovation. Minerva, 13(1):58–81, Spring 1975. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096242.

Mazur:1977:SC

[Maz77] Allan Mazur. Science courts. Minerva, 15(1):1–14, Spring 1977. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096317.

Mazur:1984:JTR

[Maz84] Allan Mazur. The journalists and technology: Reporting about love canal and three mile Island. Minerva, 22(1):45–66, Spring 1984. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02207556.

Mazur:1989:ADR

[Maz89] Allan Mazur. Allegations of dishonesty in research and their treatment by American universities. Minerva, 27(2–3):177–194, Summer–Autumn 1989. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01102567.

Mazur:1990:NPC

[Maz90] Allan Mazur. Nuclear power, chemical hazards, and the quantity of reporting. Minerva, 28(3):294–323, Autumn 1990. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096293.

Mazur:1997:REI

[Maz97] Allan Mazur. The reporting of environmental issues in Britain and the United States. Minerva, 35(1):94–97, ???? 1997. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/B%3AMINE.0000019489.41989.47.

Mazawi:2005:APR

[Maz05] André Elias Mazawi. The academic profession in a rentier state: The professoriate in Saudi Arabia. Minerva, 43(3):221–244, ???? 2005. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-005-6473-5.

Merz:2010:HTP

[MB10] Martina Merz and Peter Biniok. How technological platforms reconfigure science–industry relations: The case of micro- and nanotechnology. Minerva, 48(2):105–124, June 2010. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-010-9146-y.

Munch:2012:ISR

[MB12] Richard Münch and Christian Baier. Institutional struggles for recognition in the academic field: The case of University departments in German chemistry. Minerva, 50(1):97–126, March 2012. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-012-9189-3.

Murphy:1992:PAA

[MBG^{+}92] Joseph B. Murphy, Sarah R. Blanshei, James F. Guyot, Howard L. Simmons Ph.D., Joel Segall, Robert H. Chambers III, Baruch College, and Jim Sleeper. The progress of affirmative action: Accreditation and diversity. Minerva, 30(4):531–552, Winter 1992. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096577.

Mott:1967:C

[MBJ67] N. F. Mott, George Z. F. Bereday, and Harry G. Johnson. Correspondence. Minerva, 5(2):265–273, Winter 1967. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF02388528.

Mars:2014:SED

[MBS14] Matthew M. Mars, Kate Bresonis, and Katalin Szelényi. Science and engineering doctoral student socialization, logics, and the national economic agenda: Alignment or disconnect? Minerva, 52(3):351–379, September 2014. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
014-9256-z.

McCaughey:1976:AUT

[McC76] Robert A. McCaughey. American university teachers and opposition to the Vietnam war: a reconsideration. Minerva, 14(3): 307–329, Autumn 1976. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096275.

McGill:1972:UDU

[McG72] William J. McGill. Universities in danger: The United States Office for Civil Rights contra Columbia University. Minerva, 10 (2):319–322, April 1972. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01682424.

McGucken:1978:FPS

[McG78] William McGucken. On freedom and planning in science: The society for freedom in science, 1940–46. Minerva, 16 (1):42–72, Spring 1978. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01102181.

McGucken:1979:COS

[McG79] William McGucken. The central organisation of scientific and technical advice in the United Kingdom during the Second World War. Minerva, 17(1):33–69, Spring 1979. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01096260.

McGuire:2016:CFE

[McG16] Wendy McGuire. Cross-field effects of science policy on the biosciences: Using Bourdieu’s relational methodology to understand change. Minerva, 54(3):325–351, September 2016. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/s11024-016-9300-2.

McLay:1999:LME

[McL99] David B. McLay. Lise Meitner and Erwin Schrödinger: Biographies of two Austrian physicists of Nobel stature. Minerva, 37 (1):75–94, March 1999. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1004536223625.

McLeod:2003:P

[McL03] Roy McLeod. Preface. Minerva, 41(2):93–94, ???? 2003. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1023/A%3A1023632413833.

Mazur:1995:BR

[MDG95] Allan Mazur, Charles R. Day, and J. W. Grove. Book reviews. Minerva, 33(3):291–302, Autumn 1995. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01117505.

Monahan:1994:BR

[MDH94] Edward J. Monahan, Lorraine Daston, and V. A. Hughes. Book reviews. Minerva, 32(1):99–117, Spring 1994. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01096658.

Mellanby:1963:ENU

[Mel63] Kenneth Mellanby. Establishing a new university in Africa. Minerva, 1(2):149–158, Winter 1963. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01096247.

Mellanby:1966:CAF

[Mel66] Kenneth Mellanby. Colonialism and academic freedom. Minerva, 4(4):562–563, Summer 1966. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF02207985.

Mellanby:1971:C

[Mel71] K. Mellanby. Correspondence. Minerva, 9(4):562–565, October 1971. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01558032.

Mellanby:1972:ESS

[Mel72] Kenneth Mellanby. Environmental studies: The search for an institutional form. Minerva, 10(1):161–163, January 1972. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01881394.

Mellanby:1974:DSR

[Mel74] Kenneth Mellanby. The disorganisation of scientific research. Minerva, 12(1):67–82, January 1974. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01558541.

Meltsner:1979:CSI

[Mel79] Arnold J. Meltsner. The communication of scientific information to the wider public: The case of seismology in California. Minerva, 17(3):331–354, Autumn 1979. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01096219.

Melsen:2024:PBO

[Mel24] Lieke A. Melsen. The politics behind overinterpreted and underexplored models: a review of Andrea Saltelli and Monica Di Fiore (eds.), The Politics of Modelling — Numbers between Science and Policy. Minerva, 62(1):141–144, March 2024. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL https://link.springer.com/article/10.
1007/s11024-024-09524-4.

Makinen:2020:PCB

[MEM20] Elina I. Mäkinen, Eliza D. Evans, and Daniel A. McFarland. The patterning of collaborative behavior and knowledge culminations in interdisciplinary research centers. Minerva, 58(1):71–95, March 2020. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http:/

/link.springer.com/article/10.1007/s11024-019-09381-
6;   http://link.springer.com/content/pdf/10.1007/s11024-
019-09381-6.pdf.

Merritt:1979:CUR

[Mer79] Richard L. Merritt. The courts, the universities and the right of admission in the Federal German Republic. Minerva, 17 (1):1–32, Spring 1979. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096259.

Mercelis:2024:BRE

[Mer24] Joris Mercelis. Book review: Eric S. Hintz, American Independent Inventors in an Era of Corporate R&D. Minerva, 62 (2):305–308, June 2024. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL https://link.springer.

com/article/10.1007/s11024-024-09532-4.

Metcalfe:2008:PNT

[Met08] Amy Scott Metcalfe. For-profit and non-traditional higher education in the Wild, Wild West. Minerva, 46(1):155–158, March 2008. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-007-9082-7.

Metcalfe:2010:UBR

[Met10] J. Stanley Metcalfe. University and business relations: Connecting the knowledge economy. Minerva, 48(1):5–33, March 2010. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-010-9140-4.

Meyer:2011:PVF

[Mey11] Ryan Meyer. The public values failures of climate science in the US. Minerva, 49(1):47–70, March 2011. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
011-9164-4.

Marques:2024:PBE

[MG24] Marcelo Marques and Lukas Graf. Pushing boundaries: The European universities initiative as a case of transnational institution building. Minerva, 62(1):93–112, March 2024. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL https://link.springer.com/article/10.
1007/s11024-023-09516-w.

Mahony:2016:MNI

[MH16] Martin Mahony and Mike Hulme. Modelling and the nation: Institutionalising climate prediction in the UK, 1988– 92. Minerva, 54(4):445–470, December 2016. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/content/pdf/10.1007/s11024-
016-9302-0.pdf.

Martin:1981:ICS

[MI81] Benjamin R. Martin and John Irvine. Internal criteria for scientific choice: an evaluation of research in high-energy physics using electron accelerators. Minerva, 19(3):408–432, Autumn 1981. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02192823.

Michael:1978:AAG

[Mic78] Ian Michael. Academic autonomy and governmental demands: The case of Malawi. Minerva, 16(4):465–479, Winter 1978. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01100329.

Middlehurst:2001:UCB

[Mid01] Robin Middlehurst. University challenges: Borderless higher education, today and tomorrow. Minerva, 39(1):3–26, ???? 2001. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1010343517872.

Miller:1972:BRT

[Mil72] Henry Miller. The balance of research, teaching and service in medical education. Minerva, 10(1):163–164, January 1972. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01881395.

Milanovic:2007:WSM

[Mil07] Fabien Milanovic. What ‘signature’ means in the sciences. Minerva, 45(3):341–343, September 2007. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
007-9054-y.

Mirowski:2008:LM

[Mir08] Philip Mirowski. Livin’ with the MTA. Minerva, 46(3):317– 342, September 2008. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-008-9102-2.

Miser:1973:SAR

[Mis73] Hugh J. Miser. I the scientist as adviser: The relevance of the early operations research experience. Minerva, 11(1): 95–108, January 1973. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01544979.

Mishan:1974:EGW

[Mis74] E. J. Mishan. Economic growth and welfare. Minerva, 12(1): 117–123, January 1974. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01558544.

Morris-Jones:1978:LTS

[MJ78] W. H. Morris-Jones. A life-time in Soviet science reconsidered: The study of political science in the universities of Bangladesh. Minerva, 16(3):425–444, Autumn 1978. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01096670.

Morris-Jones:1985:BR

[MJSB85] W. H. Morris-Jones, Terry Shinn, and Martin Bulmer. Book reviews. Minerva, 23(1):151–172, Spring 1985. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01097843.

Mossner:2017:KDI

[MK17] Nicola Mößner and Philip Kitcher. Knowledge, democracy, and the Internet. Minerva, 55(1):1–24, March 2017. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/s11024-016-9310-0.

Muller:2019:RDA

[MK19] Ruth Müller and Wolfgang Kaltenbrunner. Re-disciplining academic careers? interdisciplinary practice and career development in a Swedish environmental sciences research center. Minerva, 57(4):479–499, December 2019. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/article/10.1007/s11024-
019-09373-6;   http://link.springer.com/content/pdf/
10.1007/s11024-019-09373-6.pdf.

Miller:2022:RHF

[MKK22] Georgia Miller, Declan Kuch, and Matthew Kearnes. Reimagining health as a ‘flow on effect’ of biomedical innovation: Research policy as a site of state activism. Minerva, 60(2):235–256, June 2022. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL https://link.springer.com/article/10.
1007/s11024-021-09456-3.

Maier-Leibnitz:1981:RRW

[ML81] Heinz Maier-Leibnitz. Remarks on the recommendations of the Wissenschaftsrat on the promotion of especially talented persons. Minerva, 19(3):498–501, Autumn 1981. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF02192827.

Maier-Leibnitz:1983:IAP

[ML83] Heinz Maier-Leibnitz. The influence and the accountability of the physicist. Minerva, 21(4):353–369, Winter 1983. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01096522.

Maier-Leibnitz:1986:SEP

[ML86] Heinz Maier-Leibnitz. Should everything be published in English? Minerva, 24(2–3):242–245, Summer–Autumn 1986. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01096613.

Maier-Leibnitz:1989:MQR

[ML89] Heinz Maier-Leibnitz. The measurement of quality and reputation in the world of learning. Minerva, 27(4):483–504, Winter 1989. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096541.

Maier-Leibnitz:1991:SAR

[MLS91] Heinz Maier-Leibnitz and Christoph Schneider. The status of academic research in the Federal German Republic: a report on two surveys and the testimony of individual scientists. Minerva, 29(1):27–60, Spring 1991. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096297.

Meltsner:1988:BR

[MM88] Arnold J. Meltsner and Michael J. Moravcsik. Book reviews. Minerva, 24(1):134–142, Spring 1988. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01102558.

McEntee:2023:PRS

[MMMS^{+}23] Marie McEntee, Fabien Medvecky, Sara MacBride-Stewart, Vicki Macknight, and Michael Martin. Park rangers and sciencepublic expertise: Science as care in biosecurity for kauri trees in Aotearoa/ new Zealand. Minerva, 61(1):117–140, March 2023. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL https://link.springer.com/article/10.
1007/s11024-022-09482-9.

Miller:2013:FSP

[MN13] Thaddeus R. Miller and Mark W. Neff. De-facto science policy in the making: How scientists shape science policy and why it matters (or, why STS and STP scholars should socialize). Minerva, 51(3):295–315, September 2013. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-013-9234-x.

Musson:1975:BR

[MNS75] A. E. Musson, Richard Nelson, and Steven Shapin. Book reviews. Minerva, 13(4):633–644, Winter 1975. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01096180.

Montgomery:2017:CFS

[MO17] Kathleen Montgomery and Amalya L. Oliver. Conceptualizing fraudulent studies as viruses: New models for handling retractions. Minerva, 55(1):49–64, March 2017. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
016-9311-z.

Molella:2005:HTT

[Mol05] Arthur Molella. Hughes on technology; Thomas P. Hughes, Human–Built World: How to Think About Technology and Culture (Chicago: University of Chicago Press, 2004), 235 pp., ISBN 0-226-35933-6, $22.50. Minerva, 43(1):113–117, ???? 2005. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-004-6645-8.

Momigliano:1983:ITH

[Mom83] Arnaldo Momigliano. The introduction of the teaching of history as an academic subject and its implications. Minerva, 21(1):1–15, Spring 1983. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01099501.

Monahan:1988:AFS

[Mon88] Edward J. Monahan. The allocation of financial support by formula: The experience of Ontario. Minerva, 26(3):370– 391, Autumn 1988. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096402.

Monahan:1995:FCC

[Mon95] Edward J. Monahan. The fabrikant case at concordia University: Some lessons for the better management of universities and improved academic ethics. Minerva, 33(2):129–148, Summer 1995. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01097909.

Monahan:1998:UGR

[Mon98] Edward J. Monahan. University-government relations in Ontario: The history of a buffer body, 1958–1996. Minerva, 36(4): 347–366, December 1998. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1004446124796.

Moodie:1996:JDC

[Moo96] Graeme C. Moodie. On justifying the different claims to academic freedom. Minerva, 34(2):129–150, ???? 1996. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF00122897.

Moodie:1999:SPU

[Moo99] Graeme C. Moodie. Student politics in the United States and Britain. Minerva, 37(3):295–299, September 1999. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1023/A%3A1004722428197.

Moodie:2007:CHE

[Moo07] Gavin Moodie. Changing higher education. Minerva, 45(1): 73–84, March 2007. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-006-9022-y.

Moravcsik:1964:TAF

[Mor64] Michael J. Moravcsik. Technical assistance and fundamental research in underdeveloped countries. Minerva, 2(2):197–209, Winter 1964. CODEN MINEFY. ISSN 0026-4695 (print),

1573-1871 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF01096594.

Moravcsik:1966:SPS

[Mor66a] Michael J. Moravcsik. Some practical suggestions for the improvement of science in developing countries. Minerva, 4(3): 381–390, Spring 1966. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01584633.

Morrow:1966:TAK

[Mor66b] Glenn R. Morrow. Thoughts after Knopfelmacher: Appointments on university. Minerva, 4(2):276–279, Winter 1966. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01584863.

Moravcsik:1971:PES

[Mor71a] Michael Moravcsik. Patterns of evaluation in science. Minerva, 9(3):419–421, July 1971. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01553368.

Moravcsik:1971:SMP

[Mor71b] Michael Moravcsik. Some modest proposals. Minerva, 9(1): 55–65, January 1971. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01553187.

Moravcsik:1973:IUI

[Mor73] Michael J. Moravcsik. II the universal intellectual versus the expert. Minerva, 11(1):109–112, January 1973. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01544980.

Morgan:1982:PAE

[Mor82] J. Graham Morgan. Preparation for the advent: The establishment of sociology as a discipline in American universities in the late nineteenth century. Minerva, 20(1–2):25–58, Spring–Summer 1982. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01098189.

Moran:1985:RAH

[Mor85] Neil C. Moran. I. report of ad hoc committee to evaluate research of Dr John R. Darsee at Emory University. Minerva, 23(2): 276–305, Summer 1985. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01099946.

Moravcsik:1989:UB

[Mor89] Michael J. Moravcsik. The ultimate bottleneck. Minerva, 27 (1):21–32, Spring 1989. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096713.

Morgan:1992:IEA

[Mor92] Harriet P. Morgan. Intellectuals as expert advisers: The case of the mandatory retirement of teachers on indefinite tenure. Minerva, 30(4):479–496, Winter 1992. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01096574.

Moseley:1978:OEY

[Mos78] Russell Moseley. The origins and early years of the national physical laboratory: a chapter in the pre-history of British science policy. Minerva, 16(2):222–250, Summer 1978. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096015.

Moss:2008:NOA

[Mos08] Michael Moss. ‘nine o’clock and All’s well’, or ‘fire, fire, the Library’s burning’: The future of the academic library. Minerva, 46(1):117–125, March 2008. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-007-9077-4.

Mothes:1964:EVS

[Mot64] Kurt Mothes. The educational value of science. Minerva, 2 (3):370–378, Spring 1964. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01097324.

McCaughey:1978:BR

[MP78] Robert A. McCaughey and Keith Pavitt. Book reviews. Minerva, 16(1):149–155, Spring 1978. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01102186.

Medawar:1976:BR

[MPM^{+}76] P. B. Medawar, Don K. Price, Allan Mazur, Pearl Kibre, and Norman W. Storer. Book reviews. Minerva, 14(3):389– 403, Autumn 1976. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096279.

Murphy:2016:EPR

[MPW16] Joseph Murphy, Sarah Parry, and John Walls. The EP-SRC’s policy of responsible innovation from a trading zones perspective. Minerva, 54(2):151–174, June 2016. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/content/pdf/10.1007/s11024-
016-9294-9.pdf.

MacDonald:1974:IOU

[MR74] John B. MacDonald and Murray G. Ross. II: Ontario and its universities. Minerva, 12(4):515–521, October 1974. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01557643.

McCaughey:1979:BR

[MR79] Robert A. McCaughey and Yakov M. Rabkin. Book reviews. Minerva, 17(3):469–479, Autumn 1979. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01096225.

Martin-Rovet:1995:IESa

[MR95] Dominique Martin-Rovet. The international exchange of scholars: The training of young scientists through research abroad. Minerva, 33(1):75–98, Spring 1995. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01098621.

Martin-Rovet:1995:IESb

[MRC95] Dominique Martin-Rovet and Timothy Carlson. The international exchange of scholars: The training of young scientists through research abroad. Minerva, 33(2):171–191, Summer 1995. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01097911.

Marcovich:2011:THQ

[MS11] Anne Marcovich and Terry Shinn. From the triple helix to a quadruple helix? The case of dip-pen nanolithography. Minerva, 49(2):175–190, June 2011. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-011-9169-z.

Mata:2012:NSF

[MS12] Tiago Mata and Tom Scheiding. National science foundation patronage of social science, 1970s and 1980s: Congressional scrutiny, advocacy network, and the prestige of economics. Minerva, 50(4):423–449, December 2012. CODEN MINEFY.

ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
012-9213-7.

Marcovich:2013:RCS

[MS13] Anne Marcovich and Terry Shinn. Respiration and cognitive synergy: Circulation in and between scientific research spheres. Minerva, 51(1):1–23, March 2013. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
013-9220-3.

Makinen:2023:MSS

[MS23] Elina I. Mäkinen and Adi Sapir. Making sense of science, university, and industry: Sensemaking narratives of Finnish and Israeli scientists. Minerva, 61(2):175–198, June 2023. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL https://link.springer.com/article/10.
1007/s11024-022-09485-6.

Mizrahi-Shtelman:2021:WRA

[MSD21] Ravit Mizrahi-Shtelman and Gili S. Drori. World-rank and/or locally relevant? Organizational identity in the mission statements of higher education organizations in Israel, 2008–2018. Minerva, 59(1):1–25, March 2021. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/article/10.1007/s11024-020-09414-5.

Mazrui:1967:GUE

[MT67] Ali A. Mazrui and Yash Tandon. Government and the universities in East Africa: II — the university of East Africa as a political institution. Minerva, 5(3):381–386, Spring 1967. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02388577.

Matthies:2019:AHI

[MT19] Hildegard Matthies and Marc Torka. Academic habitus and institutional change: Comparing two generations of German scholars. Minerva, 57(3):345–371, September 2019. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/article/10.1007/s11024-
019-09370-9.

Miettinen:2015:EAI

[MTE15] Reijo Miettinen, Juha Tuunainen, and Terhi Esko. Epistemological, artefactual and interactional-institutional foundations of social impact of academic research. Minerva, 53(3):257– 277, September 2015. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-015-9278-1.

Macq:2020:DPP

[MTS20] Hadrien Macq, Élise Tancoigne, and Bruno J. Strasser. From deliberation to production: Public participation in science and technology policies of the European Commission (1998–2019). Minerva, 58(4):489–512, December 2020. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/article/10.1007/s11024-020-09405-6.

Mucha:1985:ULD

[Muc85] Janusz Mucha. University legislation and the decline of academic autonomy in Poland. Minerva, 23(3):362–382, Autumn 1985. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096444.

Muir:1969:RBD

[Mui69] J. Douglas Muir. The reduction of the brain drain: Problems and policies. Minerva, 7(3):494–498, Spring 1969. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01103273.

Mukerjee:1964:ISP

[Muk64] Dilip Mukerjee. Indian science: Policy, organisation and application. Minerva, 2(3):360–369, Spring 1964. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01097323.

Mullins:1972:DSS

[Mul72] Nicholas C. Mullins. The development of a scientific specialty: The phage group and the origins of molecular biology. Minerva, 10(1):51–82, January 1972. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01881390.

Muller:2014:PLS

[Mül14] Ruth Müller. Postdoctoral life scientists and supervision work in the contemporary University: a case study of changes in the cultural norms of science. Minerva, 52(3):329–349, September 2014. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-014-9257-y.

Muller:2020:ERR

[Mul20] Johan Muller. ‘From excellence to relevance?’ A review of Erika Kraemer-Mbula, Robert Tijssen, Matthew L. Wallace and Robert McLean (eds.), Transforming Research Excellence: New Ideas from the Global South. Minerva, 58(3):485–488, September 2020. CODEN MINEFY. ISSN 0026-4695 (print), 1573-

1871 (electronic). URL http://link.springer.com/article/
10.1007/s11024-020-09415-4.

Musselin:2022:PSE

[Mus22] Christine Musselin. Peter Scott, Retreat or Resolution? Tackling the Crisis of Mass Higher Education. Minerva, 60(3):463– 466, September 2022. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL https://link.springer.

com/article/10.1007/s11024-022-09475-8.

Miller:1975:BR

[MZP75] Henry Miller, John Ziman, and Keith Pavitt. Book reviews. Minerva, 13(1):121–130, Spring 1975. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01096245.

Nastesjo:2021:NUE

[Näs21] Jonatan Nästesjö. Navigating uncertainty: Early career academics and practices of appraisal devices. Minerva, 59(2):237– 259, June 2021. CODEN MINEFY. ISSN 0026-4695 (print),

1573-1871 (electronic). URL https://link.springer.com/
article/10.1007/s11024-020-09425-2.

Nastesjo:2025:BDL

[Näs25] Jonatan Nästesjö. Between delivery and luck: Projectification of academic careers and conflicting notions of worth at the postdoc level. Minerva, 63(1):69–92, March 2025. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL https://

link.springer.com/article/10.1007/s11024-024-09541-3.

Nayudamma:1967:PIA

[Nay67] Y. Nayudamma. Promoting the industrial application of research in an underdeveloped country. Minerva, 5(3):323–339, Spring 1967. CODEN MINEFY. ISSN 0026-4695 (print),

1573-1871 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF02388572.

Nayudamma:1973:DMR

[Nay73] Y. Nayudamma. Decentralised management of R&D in a developing country. Minerva, 11(4):516–536, October 1973. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01557800.

Nixdorff:2002:EUR

[NB02] Kathryn Nixdorff and Wolfgang Bender. Ethics of University research, biotechnology and potential military spin-off. Minerva, 40(1):15–35, ???? 2002. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1015045127363.

Navarro-Brotons:2005:BPP

[NBGT05] Víctor Navarro-Brotóns, Jorge Velasco González, and José Doménech Torres. The birth of particle physics in Spain. Minerva, 43 (2):183–196, June 2005. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-005-2331-8.

Neff:2014:RPP

[Nef14] Mark William Neff. Research prioritization and the potential pitfall of path dependencies in coral reef science. Minerva, 52 (2):213–235, June 2014. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-014-9250-5.

Nehring:2006:PET

[Neh06] Holger Nehring. Politics and the ‘environment’ in Twentieth-Century Germany. Minerva, 44(3):338–354, June 2006. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/s11024-006-9005-z.

Nelson:1971:WLT

[Nel71] Richard R. Nelson. “World leadership”, the “technological gap” and national science policy. Minerva, 9(3):386–399, July 1971. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01553362.

Nelson:2021:PVP

[Nel21] John P. Nelson. Public value promises and outcome reporting in advanced research projects agency-energy. Minerva, 59(4): 493–513, December 2021. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL https://link.springer.

com/article/10.1007/s11024-021-09444-7.

Nelson:1974:BR

[NGZG74] Richard R. Nelson, Loren R. Graham, John Ziman, and Herbert G. Grubel. Book reviews. Minerva, 12(2):277–290, April 1974. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01553181.

Nicholls:1965:SPY

[Nic65] Frank G. Nicholls. Scientific policy in Yugoslavia. Minerva, 3 (3):405–407, Spring 1965. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01099966.

Nichols:1972:SPP

[Nic72] Rodney W. Nichols. Some practical problems of scientistadvisers. Minerva, 10(4):603–613, October 1972. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01695907.

Nicholson:1977:AAB

[Nic77] Heather Johnston Nicholson. Autonomy and accountability of basic research. Minerva, 15(1):32–61, Spring 1977. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01096319.

Nguyen:2016:KPO

[NM16] Huong Thi Lan Nguyen and Vincent Lynn Meek. Key problems in organizing and structuring University research in Vietnam: The lack of an effective research “behaviour Formalization” system. Minerva, 54(1):45–73, March 2016. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
016-9289-6.

Nolden:1982:PUP

[Nol82] Hans-Willi Nolden. I the private university: Its possibilities and limitations in the German Federal Republic. Minerva, 20(1– 2):213–225, Spring–Summer 1982. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01098195.

Nolte:1983:TSP

[Nol83] Ernst Nolte. Thoughts on the state and prospects of the academic ethic in the universities of the Federal Republic of Germany. Minerva, 21(2–3):161–171, Summer–Autumn 1983. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01097962.

Nolin:1999:GPN

[Nol99a] Jan Nolin. Global policy and national research: The international shaping of climate research in four European union countries. Minerva, 37(2):125–140, June 1999. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1023/
A%3A1004641211833.

Nolin:1999:TSR

[Nol99b] Jan Nolin. Timing and sponsorship: The research to policy process and the European Union’s Kyoto proposal. Minerva, 37 (2):165–181, June 1999. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1004623529580.

Nour:2011:NRG

[Nou11] Samia Satti Osman Mohamed Nour. National, regional and global perspectives of higher education and science policies in the Arab region. Minerva, 49(4):387–423, December 2011. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-011-9183-1.

Novotny:2017:HAP

[Nov17] Adam Novotny. The heterogeneity of the academic profession: The effect of occupational variables on university scientists’ participation in research commercialization. Minerva, 55(4):485– 508, December 2017. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/article/10.1007/s11024-017-9321-5.

Niembro:2025:TIC

[NS25] Andrés Niembro and Fernando Svampa. Territorial inequalities and (de)Concentration of public investment in science: a study on CONICET (Argentina) and the tensions between academic excellence and equity. Minerva, 63(3):499–533, September 2025. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL https://link.springer.com/article/10.
1007/s11024-024-09550-2.

Neudert:2025:LRR

[NSB25] Philipp Neudert, Mareike Smolka, and Stefan Böschen. The limits of responsibilization? Responsibility boundary-work through visions in the case of neuromorphic computing. Minerva, 63(3): 579–606, September 2025. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL https://link.springer.

com/article/10.1007/s11024-024-09562-y.

Nowotny:2003:IMR

[NSG03] Helga Nowotny, Peter Scott, and Michael Gibbons. Introduction: ‘mode 2’ revisited: The new production of knowledge. Minerva, 41(3):179–194, ???? 2003. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1025505528250.

Notzoldt:2004:PAS

[NW04] Peter Nötzoldt and Peter Th. Walther. The Prussian Academy of Sciences during the third Reich. Minerva, 42(4):421–444, ???? 2004. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-004-2115-6.

Nye:1975:SPF

[Nye75] Mary Jo Nye. The scientific periphery in France: The faculty of sciences at Toulouse (1880–1930). Minerva, 13(3): 374–403, Autumn 1975. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01097847.

Nye:2007:PE

[Nye07] Mary Jo Nye. Passionate empiricism. Minerva, 45(4):495– 498, December 2007. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-007-9056-9.

Nekola:1968:RDC

[NZ68] J. Nekola and J. Zelinka. Research and development in Czechoslovakia. Minerva, 6(3):388–397, Spring 1968. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01096424.

OBrien:1966:AAF

[O’B66] Conor Cruise O’Brien. Autonomy and academic freedom in Britain and Africa. Minerva, 5(1):89–92, Autumn 1966. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF02389795.

Olson:2017:EIE

[OB17] Rebecca E. Olson and Caragh Brosnan. Examining interprofessional education through the lens of interdisciplinarity: Power, knowledge and new ontological subjects. Minerva, 55(3):299– 319, September 2017. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/article/10.1007/s11024-017-9316-2.

ofChesterfield:1968:UGP

[oC68] Bowden of Chesterfield. The universities, the government and the public accounts committee. Minerva, 6(4):612–614, Summer 1968. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096556.

Ortiga:2020:CAL

[OCW20] Yasmin Y. Ortiga, Meng-Hsuan Chou, and Jue Wang. Competing for academic labor: Research and recruitment outside the academic center. Minerva, 58(4):607–624, December 2020. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/article/10.1007/s11024-
020-09412-7.

Okyar:1968:UT

[Oky68] Osman Okyar. Universities in Turkey. Minerva, 6(2):213– 243, Winter 1968. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01097259.

Okyar:1970:SAM

[Oky70] Osman Okyar. Science in the Arab Middle East. Minerva, 8(1– 4):451–453, January 1970. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01553230;   http://
link.springer.com/article/10.1007/BF01553230.

Oleksiyenko:2013:OLI

[Ole13] Anatoly Oleksiyenko. Organizational legitimacy of international research collaborations: Crossing boundaries in the middle East. Minerva, 51(1):49–69, March 2013. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
013-9221-2.

Ossowska:1964:SS

[OO64] María Ossowska and Stanislaw Ossowski. The science of science. Minerva, 3(1):72–82, Autumn 1964. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01630150.

Olmos-Penuela:2015:WSR

[OPBCM15] Julia Olmos-Peñuela, Paul Benneworth, and Elena Castro-Martínez. What stimulates researchers to make their research usable? Towards an ‘openness’ approach. Minerva, 53(4): 381–410, December 2015. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-015-9283-4.

Opie:2004:TER

[Opi04] Brian Opie. Tertiary education and research in New Zealand. Minerva, 42(3):299–307, ???? 2004. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1023/B%3AMINE.0000038303.
80536.3a.

Orlans:1971:SSR

[Orl71a] Harold Orlans. Social science research policies in the United States. Minerva, 9(1):7–31, January 1971. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01553183.

Orlans:1971:SRU

[Orl71b] Harold Orlans. Sponsored research and university budgets. Minerva, 9(3):411–414, July 1971. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01553365.

Orlans:1972:CCS

[Orl72] Harold Orlans. Criteria of choice in social science research. Minerva, 10(4):571–602, October 1972. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01695906.

Orlans:1976:ASS

[Orl76] Harold Orlans. The advocacy of social science in Europe and America. Minerva, 14(1):6–32, Spring 1976. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01096212.

Orlans:1980:RS

[Orl80] Harold Orlans. On the responsibility of scientists. Minerva, 18 (3):521–528, Autumn 1980. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096955.

Orlans:1986:ASS

[Orl86] Harold Orlans. Academic social scientists and the presidency: From Wilson to Nixon. Minerva, 24(2–3):172–204, Summer–Autumn 1986. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096610.

Orlans:1991:BR

[Orl91] Harold Orlans. Book review. Minerva, 29(3):386–390, Autumn 1991. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01097882.

Orlans:1992:AAH

[Orl92] Harold Orlans. Accreditation in American higher education: The issue of “diversity”. Minerva, 30(4):513–530, Winter 1992. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096576.

Orlans:1996:ESB

[Orl96] Harold Orlans. Edward Shils’ beliefs about society and sociology. Minerva, 34(1):23–37, ???? 1996. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00124198.

Orlans:1999:BRS

[Orl99] Harold Orlans. Book review: The Shape of the River: Long-Term Consequences of Considering Race in College and University Admissions, by William G. Bowen and Derek Bok. Minerva, 37(2):185–190, June 1999. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1004605914076.

Orthia:2016:WWT

[Ort16] Lindy A. Orthia. What’s wrong with talking about the scientific revolution? Applying lessons from history of science to applied fields of science studies. Minerva, 54(3):353–373, September 2016. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-016-9299-4.

OSullivan:2001:HDN

[O’S01] Abigail O’Sullivan. Henry Dale’s Nobel Prize winning ‘discovery’. Minerva, 39(4):409–424, ???? 2001. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1023/
A%3A1012787903437.

Ohly:2021:GPM

[OS21] Sandra Ohly and Christian Schneijderberg. German professors’ motivation to act as peer reviewers in accreditation and evaluation procedures. Minerva, 59(2):217–236, June 2021. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL https://link.springer.com/article/10.
1007/s11024-020-09430-5.

Osborne:1971:UPS

[Osb71] Denis Osborne. The use and promotion of science in developing countries. Minerva, 9(1):45–55, January 1971. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01553186.

Osley-Thomas:2020:CAD

[OT20] Robert Osley-Thomas. The closing of academic departments and programs: a core and periphery approach to the liberal arts and practical arts. Minerva, 58(2):211–233, June 2020. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/article/10.1007/s11024-
019-09389-y.

Ozdas:1968:UT

[Ozd68] M. Nimet Ozdas. Universities in Turkey. Minerva, 6(4): 609–612, Summer 1968. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096555.

Page:2002:RFC

[Pag02] Benhamin B. Page. The Rockefeller Foundation and central Europe: a reconsideration. Minerva, 40(3):265–287, ???? 2002. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1019520525157.

Palladino:1990:PEA

[Pal90] Paolo Palladino. The political economy of applied research: Plant breeding in Great Britain, 1910–1940. Minerva, 28(4): 446–468, Winter 1990. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096651.

Pancaldi:1993:VVC

[Pan93] Giuliano Pancaldi. Vito Volterra: Cosmopolitan ideals and nationality in the Italian scientific community between the Belle époque and the First World War. Minerva, 31(1):21–37, Spring 1993. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096170.

Panofsky:2011:FAI

[Pan11] Aaron L. Panofsky. Field analysis and interdisciplinary science: Scientific capital exchange in behavior genetics. Minerva, 49 (3):??, September 2011. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-011-9175-1.

Papon:1973:TIS

[Pap73] P. Papon. The training of industrial scientists in France. Minerva, 11(2):191–210, April 1973. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01557716.

Papon:1988:STP

[Pap88] Pierre Papon. Science and technology policy in France: 1981–1986. Minerva, 26(4):493–511, Winter 1988. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01096495.

Papon:1996:NCS

[Pap96] Piere Papon. A new context for scientific expertise? Some lessons from the French experience. Minerva, 34(2):151–160, ???? 1996. CODEN MINEFY. ISSN 0026-4695 (print),

1573-1871 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF00122898.

Papon:2004:ESC

[Pap04] Pierre Papon. European scientific cooperation and research infrastructures: Past tendencies and future prospects. Minerva, 42(1):61–76, ???? 2004. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/B%3AMINE.0000017700.63978.
4a.

Parikh:1965:DKI

[Par65] G. D. Parikh. Dr. Karve and the Indian universities. Minerva, 3(3):401–404, Spring 1965. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01099965.

Parsons:1968:FNC

[Par68] Talcott Parsons. The future of the nineteenth century idea of a university. Minerva, 6(2):267–271, Winter 1968. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01097265.

Parikh:1971:BC

[Par71a] G. D. Parikh. Bombay College. Minerva, 9(3):417–419, July 1971. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01553367.

Parsons:1971:C

[Par71b] Talcott Parsons. Correspondence. Minerva, 9(4):553–556, October 1971. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01558028.

Parmar:2000:CCM

[Par00] Inderjeet Parmar. The Carnegie corporation and the mobilisation of opinion in the United States’ rise to globalism, 1939–1945. Minerva, 37(4):355–378, December 2000. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1023/A%3A1004797821885.

Parmar:2002:RKA

[Par02] Inderjeet Parmar. ‘to relate knowledge and action’: the impact of the Rockefeller Foundation on foreign policy thinking during America’s rise to globalism 1939–1945. Minerva, 40 (3):235–263, ???? 2002. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1019572526066.

Passmore:1978:CCG

[Pas78] John Passmore. A comment by the Chairman of the German Universities Commission. Minerva, 16(1):139–142, Spring 1978. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01102184.

Pattullo:1985:GRI

[Pat85] Edward L. Pattullo. Governmental regulation of the investigation of human subjects in social research. Minerva, 23(4): 521–533, Winter 1985. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01098705.

Paulley:1967:BUP

[Pau67] J. W. Paulley. British universities and the public purse. Minerva, 6(1):102–105, Autumn 1967. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01101610.

Pavitt:1976:GSI

[Pav76] Keith Pavitt. Governmental support for industrial research and development in France: Theory and practice. Minerva, 14(3): 330–354, Autumn 1976. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096276.

Payne:1980:OUM

[Pay80] Anthony Payne. One university, many governments: Regional integration, politics and the university of the West Indies. Minerva, 18(3):474–498, Autumn 1980. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01096953.

Pollock:2002:TPV

[PC02] Neil Pollock and James Cornford. The theory and practice of the virtual University: Working through the work of making work mobile. Minerva, 40(4):359–373, ???? 2002. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1023/A%3A1020977705523.

Pavitt:1979:BR

[PCW79] KEith Pavitt, Cmarles Carter, and John T. Wilson. Book reviews. Minerva, 17(1):178–192, Spring 1979. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01096264.

Parsons:1971:UUU

[PD71] Kermit C. Parsons and Georgia K. Davis. The urban University and its urban environment. Minerva, 9(3):361–385, July 1971. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01553361.

Powell:2017:SPG

[PD17] Justin J. W. Powell and Jennifer Dusdal. Science production in Germany, France, Belgium, and Luxembourg: Comparing the contributions of research universities and institutes to science, technology, engineering, mathematics, and health. Minerva, 55(4):413–434, December 2017. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/content/pdf/10.1007/s11024-
017-9327-z.pdf;   http://link.springer.com/article/10.
1007/s11024-017-9327-z.

Pedersen:1977:DUB

[Ped77] Mogens N. Pedersen. The Danish university between the millstones. Minerva, 15(3–4):335–376, Autumn–Winter 1977. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01096728.

Peierls:1984:BRF

[Pei84] Rudolf Peierls. Book reviews: Forty Years into the Atomic Age (pp. 285–291). The Nuclear Chain Reaction Forty Years Later by Robert G. Sachs. Minerva, 22(2):281–291, Summer 1984. CODEN MINEFY. ISSN 0026-4695 (print),

1573-1871 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF02207952;   http://www.
jstor.org/stable/41820567.

Peine:2011:CIW

[Pei11] Alexander Peine. Challenging incommensurability: What we can learn from Ludwik Fleck for the analysis of configurational innovation. Minerva, 49(4):489–508, December 2011. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s11024-011-9180-4.pdf.

Perkins:1968:RS

[Per68] James A. Perkins. The restless student. Minerva, 6(4):487– 496, Summer 1968. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096544.

Pereira:1971:IRA

[Per71] H. C. Pereira. The integration of research agencies for African agricultural development. Minerva, 9(1):38–45, January 1971. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01553185.

Pestre:2003:RKP

[Pes03] Dominique Pestre. Regimes of knowledge production in society: Towards a more political and social reading. Minerva, 41 (3):245–261, ???? 2003. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1025553311412.

Pestre:2009:UFG

[Pes09] Dominique Pestre. Understanding the forms of government in Today’s liberal and democratic societies: an introduction. Minerva, 47(3):??, September 2009. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
009-9126-2.

Peteri:1993:SVS

[Pét93] György Péteri. “Scientists versus scholars”: The prelude to communist takeover in Hungarian science, 1945–1947. Minerva, 31 (3):291–325, Autumn 1993. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01098625.

Peteri:1995:LSS

[Pét95] György Péteri. On the legacy of state socialism in academia. Minerva, 33(4):305–324, Winter 1995. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01096515.

Peteri:1996:CFA

[Pét96a] György Péteri. Controlling the field of academic economics in Hungary, 1953–1976. Minerva, 34(4):367–380, ???? 1996. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF00127072.

Peteri:1996:I

[Pét96b] György Péteri. Introduction. Minerva, 34(4):321–322, ???? 1996. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00127069.

Petitjean:2008:ISP

[Pet08a] Patrick Petitjean. Introduction: Science, politics, philosophy and history. Minerva, 46(2):175–180, June 2008. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
s11024-008-9095-x.

Petitjean:2008:JEW

[Pet08b] Patrick Petitjean. The joint establishment of the world federation of scientific workers and of UNESCO after world war II. Minerva, 46(2):247–270, June 2008. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
008-9092-0.

Pettersson:2018:NUI

[Pet18] Ingemar Pettersson. The nomos of the university: Introducing the professor’s privilege in 1940s Sweden. Minerva, 56(3):381–403, September 2018. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/article/10.1007/s11024-018-9348-2;
 http://link.springer.com/content/pdf/10.1007/s11024-
018-9348-2.pdf.

Paradeise:2021:SIM

[PF21] Catherine Paradeise and Ghislaine Filliatreau. Scientific integrity matters. Minerva, 59(3):289–309, September 2021. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (elec-

tronic).  URL https://link.springer.com/article/10.
1007/s11024-021-09440-x.

Pereira:2018:CAH

[PFC18] Tiago Santos Pereira, Paulo F. C. Fonseca, and António Carvalho. Carnation atoms? A history of nuclear energy in Portugal. Minerva, 56(4):505–528, December 2018. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/article/10.1007/s11024-018-9354-4.

Pfetsch:1970:SOS

[Pfe70] Frank Pfetsch. Scientific organisation and science policy in imperial Germany, 1871–1914: The foundation of the Imperial Institute of Physics and Technology. Minerva, 8(1–4): 557–580, January 1970. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01553239;   http://
link.springer.com/article/10.1007/BF01553239.

Pfister:2015:CEI

[Pfi15] Thomas Pfister. Coproducing European integration studies: Infrastructures and epistemic movements in an interdisciplinary field. Minerva, 53(3):235–255, September 2015. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
s11024-015-9275-4.

Philippczyck:2024:FDS

[PGO24] Nicole Philippczyck, Jan Grundmann, and Simon Oertel. The framing of diversity statements in European universities: The role of imprinting and institutional legacy. Minerva, 62(1):69– 92, March 2024. CODEN MINEFY. ISSN 0026-4695 (print),

1573-1871 (electronic). URL https://link.springer.com/
article/10.1007/s11024-023-09507-x.

Pedersen:1980:IRR

[PH80] Mogens-N. Pedersen and Howard O. Hunter. II recent reforms in Swedish higher education. Minerva, 18(2):324–351, Summer 1980. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01100252.

Porter:1983:BR

[PHA83] Roy Porter, A. H. Halsey, and Andrew Abbott. Book reviews. Minerva, 21(1):144–154, Spring 1983. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01099507.

Philip:1978:TDA

[Phi78] John R. Philip. Towards diversity and adaptability: an Australian view of governmentally supported science. Minerva, 16 (3):397–415, Autumn 1978. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096668.

pHughes:1979:BR

[pHPP^{+}79] Thomas p. Hughes, Pierre Papon, Keith Pavitt, John Ziman, and Michael J. Moravcsik. Book reviews. Minerva, 17(4): 555–573, Winter 1979. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01101904.

Pickstone:2007:SSM

[Pic07] John V. Pickstone. The social study of medicine. Minerva, 45 (2):231–233, June 2007. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-007-9041-3.

Pieris:1964:UPP

[Pie64] Ralph Pieris. Universities, politics and public opinion in Ceylon. Minerva, 2(4):435–454, Summer 1964. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01097537.

Pielke:2012:BRP

[Pie12] Roger Pielke, Jr. Basic research as a political symbol. Minerva, 50(3):339–361, September 2012. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
012-9207-5.

Piganiol:1968:SPE

[Pig68] Pierre Piganiol. Scientific policy and the European community. Minerva, 6(3):354–365, Spring 1968. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01096422.

Prieur:2016:LSS

[PJLP16] Annick Prieur, Sune Qvotrup Jensen, Julie Laursen, and Oline Pedersen. “Social skills”: Following a travelling concept from American academic discourse to contemporary Danish welfare institutions. Minerva, 54(4):423–443, December 2016. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-016-9307-8.

Priebe:1980:MGP

[PK80] Paul M. Priebe and George B. Kauffman. Making governmental policy under conditions of scientific uncertainty: a century of controversy about saccharin in congress and the laboratory. Minerva, 18(4):556–574, Winter 1980. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01096124.

Porter:1985:BR

[PK85] Roy Porter and Peter Kneen. Book reviews. Minerva, 23(4): 534–541, Winter 1985. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01098706.

Pielke:2009:RFS

[PK09] Roger Pielke, Jr. and Roberta Klein. The rise and fall of the science advisor to the President of the United States. Minerva, 47(1):7–29, March 2009. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-009-9117-3.

Peterson:2023:BTS

[PKvdW23] Jesse D. Peterson, Dick Kasperowski, and René van der Wal. Bringing together species observations: a case story of Sweden’s biodiversity informatics infrastructures. Minerva, 61(2): 265–289, June 2023. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL https://link.springer.

com/article/10.1007/s11024-023-09491-2.

Pielke:2019:GRG

[PL19] Roger Pielke, Jr. and Björn-Ola Linnér. From Green Revolution to Green Evolution: a critique of the political myth of averted famine. Minerva, 57(3):265–291, September 2019. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/article/10.1007/s11024-
019-09372-7.

Plane:1999:SMD

[Pla99] Brian Plane. The “Sputnik myth” and dissent over scientific policies under the new economic system in East Berlin, 1961–1964. Minerva, 37(1):45–62, March 1999. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1023/
A%3A1004519720899.

Plommer:1975:ICR

[Plo75] W. H. Plommer. II a comment on the report of the general board. Minerva, 13(2):297–302, Summer 1975. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01097798.

Polanyi:1962:RS

[Pol62] Michael Polanyi. The Republic of science. Minerva, 1(1): 54–73, Autumn 1962. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01101453.

Polanyi:1967:GSS

[Pol67] Michael Polanyi. The growth of science in society. Minerva, 5 (4):533–545, Summer 1967. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096782.

Polkinghorne:1972:SCP

[Pol72] J. C. Polkinghorne. Secretiveness and competition for priority of discovery in physics. Minerva, 10(2):329–331, April 1972. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01682426.

Polanyi:2000:RSP

[Pol00] Michael Polanyi. The Republic of science: Its political and economic theory. Minerva, 38(1):1–21, ???? 2000. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1023/
A%3A1026591624255.

Pols:2005:BPT

[Pol05] Hans Pols. Biomedical platforms: Technology in modern science. Minerva, 43(3):331–334, ???? 2005. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
005-6514-0.

Porter:1970:DCU

[Por70] John Porter. The democratisation of the Canadian universities and the need for a national system. Minerva, 8(1–4): 325–356, January 1970. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01553222;   http://
link.springer.com/article/10.1007/BF01553222.

Poser:2004:SWL

[Pos04] Hans Poser. Sisyphus at work: The Leibniz edition, the Kaiserreich and divided Germany. Minerva, 42(4):379–392, ???? 2004. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-004-2113-8.

Powell:2007:IEL

[POSC07] Walter W. Powell, Jason Owen-Smith, and Jeannette A. Colyvas. Innovaton and emulation: Lessons from American universities in selling private rights to public knowledge. Minerva, 45 (2):121–142, June 2007. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-007-9034-2.

Pouris:2007:ERE

[Pou07] Anastassios Pouris. Estimating R&D expenditures in the higher education sector. Minerva, 45(1):3–16, March 2007. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
s11024-006-9018-7.

Parsons:1968:CAA

[PP68] Talcott Parsons and Gerald M. Platt. Considerations on the American academic system. Minerva, 6(4):497–523, Summer 1968. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096545.

Peterson:2023:MSS

[PP23] David Peterson and Aaron Panofsky. Metascience as a scientific social movement. Minerva, 61(2):147–174, June 2023. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL https://link.springer.com/article/10.
1007/s11024-023-09490-3.

Petruhina:2025:CSH

[PP25] Polina S. Petruhina and Vitaly Pronskikh. Cultural shifts in high energy physics collaboration from the Cold War to the present: a historical and philosophical perspective. Minerva, 63 (1):135–154, March 2025. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL https://link.springer.

com/article/10.1007/s11024-024-09543-1.

Pippard:1989:BR

[PPF89] Brian Pippard, Roy Porter, and A. H. Franke. Book reviews. Minerva, 27(1):126–137, Spring 1989. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01096717.

Pratt:1967:AFA

[Pra67] R. Cranford Pratt. On academic freedom in Africa. Minerva, 5 (4):574–575, Summer 1967. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096787.

Prainsack:2012:BRE

[Pra12] Barbara Prainsack. Book review: Elias G. Carayannis and David F. J. Campbell, Mode 3 Knowledge Production in Quadruple Helix Innovation Systems: 21st-Century Democracy, Innovation, and Entrepreneurship for Development. Minerva, 50(1): 139–142, March 2012. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-012-9194-6.

Prest:2023:BRM

[Pre23] Wilfrid Prest. Book review: Maxwell Bennett, The Search for Truth: History and Future of Universities, (Sydney, Sydney University Press, 2022). xxi + 333 pp. ISBN 978-1-74210-520-8 (paperback), 978-1-74210-521-5 (e-book). Minerva, 61 (1):141–146, March 2023. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL https://link.springer.

com/article/10.1007/s11024-022-09483-8.

Price:1978:EBU

[Pri78] Geoffrey L. Price. The expansion of British universities and their struggle to maintain autonomy: 1943–46. Minerva, 16(3): 357–381, Autumn 1978. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096666.

Price:1988:NSS

[Pri88] Don K. Price. The natural sciences, the social sciences and politics. Minerva, 26(3):416–428, Autumn 1988. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01096404.

Pritchard:1998:AFA

[Pri98] Rosalind M. O. Pritchard. Academic freedom and autonomy in the United Kingdom and Germany. Minerva, 36(2): 101–124, June 1998. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1004301400289.

Prywes:1971:BRT

[Pry71] Moshe Prywes. The balance of research, teaching and service in medical education. Minerva, 9(4):451–471, October 1971. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01558019.

Pye:1965:HEPb

[PS65a] Lucian Pye and Arthur L. Singer. Higher education and politics in Singapore. Minerva, 4(1):107–110, Autumn 1965. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01585989.

Pye:1965:HEPa

[PS65b] Lucian W. Pye and Arthur L. Singer, Jr. Higher education and politics in Singapore. Minerva, 3(3):321–335, Spring 1965. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01099958.

Plonski:2001:GST

[PS01] Guilherme Ary Plonski and Rochelle G. Saidel. Gender, science and technology in Brazil. Minerva, 39(2):217–238, ???? 2001. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1010340812739.

Psacharopoulos:1971:SPA

[Psa71] George Psacharopoulos. On some positive aspects of the economics of the brain drain. Minerva, 9(2):231–242, April 1971. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01553157.

Psacharopoulos:1972:SPA

[Psa72] George Psacharopoulos. On some positive aspects of the economics of the brain drain. Minerva, 10(1):165, January 1972. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01881396.

Psacharopoulos:1988:EEG

[Psa88] George Psacharopoulos. Efficiency and equity in Greek higher education. Minerva, 26(2):119–137, Summer 1988. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01096692.

Pavitt:1980:BR

[PSNB80] Keith Pavitt, Edward Shils, Joseph Needham, and Gregory Blue. Book reviews. Minerva, 18(2):352–363, Summer 1980. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01100253.

Puccetti:1972:AGA

[Puc72] Roland Puccetti. Authoritarian government and academic subservience: The University of Singapore. Minerva, 10(2): 223–241, April 1972. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01682419.

Puig:2005:SPS

[Pui05] Albert Presas I Puig. Science on the periphery. The Spanish reception of nuclear energy: an attempt at modernity? Minerva, 43(2):197–218, June 2005. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-005-2332-7.

Pansters:2016:EIT

[PvR16] Wil G. Pansters and Henk J. van Rinsum. Enacting identity and transition: Public events and rituals in the university (Mexico and South Africa). Minerva, 54(1):21–43, March 2016. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/content/pdf/
10.1007/s11024-015-9287-0.pdf.

Pestre:2009:GTS

[PW09] Dominique Pestre and Peter Weingart. Governance of and through science and numbers: Categories, tools and technologies — preface. Minerva, 47(3):??, September 2009. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/s11024-009-9130-6.

Quartey:1971:SDC

[Qua71] J. A. K. Quartey. Science in developing countries. Minerva, 9 (4):548–550, October 1971. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01558026.

Qian:2025:ECI

[QZL25] Yuchen Qian, Liyin Zhang, and Jiang Li. Enhancing China’s incentive system for scientific innovation: a review and recommendations. Minerva, 63(3):557–578, September 2025. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (elec-

tronic).  URL https://link.springer.com/article/10.
1007/s11024-024-09554-y.

Rabkin:1976:NSS

[Rab76] Yakov M. Rabkin. “Naukovedenie”: The study of scientific research in the Soviet Union. Minerva, 14(1):61–78, Spring 1976. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096214.

Rabi:2025:RGH

[Rab25] Michael Rabi. The rise of global health emergency governance. Minerva, 63(3):471–497, September 2025. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL https://

link.springer.com/article/10.1007/s11024-024-09555-x.

Radnitzky:1983:STP

[Rad83] Gerard Radnitzky. Science, technology, and political responsibility. Minerva, 21(2–3):234–264, Summer–Autumn 1983. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01097965.

Radin:2019:AFS

[Rad19] Joanna Radin. Alternative facts and states of fear: Reality and STS in an age of climate fictions. Minerva, 57(4):411–431, December 2019. CODEN MINEFY. ISSN 0026-4695 (print), 1573-

1871 (electronic). URL http://link.springer.com/article/
10.1007/s11024-019-09374-5.

Rae:2006:AEP

[Rae06] Ian D. Rae. All environmental politics is local? Minerva, 44(4): 447–458, December 2006. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-006-9016-9.

Rae:2008:FT

[Rae08] Ian Rae. Flash trash. Minerva, 46(1):139–141, March 2008. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-007-9072-9.

Rahman:1965:SPY

[Rah65] A. Rahman. Scientific policy in Yugoslavia. Minerva, 3(4): 524–526, Summer 1965. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01099416.

Raiser:1964:GVR

[Rai64] Ludwig Raiser. A German view of the Robbins Committee’s report. Minerva, 2(3):336–342, Spring 1964. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01097321.

Raina:1996:RCS

[Rai96] Dhruv Raina. Reconfiguring the centre: The structure of scientific exchanges between colonial India and Europe. Minerva, 34(2):161–176, ???? 1996. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00122899.

Rainger:2001:CLP

[Rai01] Ronald Rainger. Constructing a landscape for postwar science: Roger Revelle, the Scripps Institution and the University of California, San Diego. Minerva, 39(3):327–352, ???? 2001. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1017952811593.

Rakovic:1965:SPY

[Rak65] Branko Raković. Scientific policy in Yugoslavia. Minerva, 3 (2):187–209, Winter 1965. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096334.

Rasmussen:2002:SMB

[Ras02] Nicolas Rasmussen. Of ‘small men’, big science and bigger business: The Second World War and biomedical research in the United States. Minerva, 40(2):115–146, ???? 2002. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1015760920428.

Rasmussen:2003:PS

[Ras03] Nicolas Rasmussen. Politics by science. Minerva, 41(1): 83–87, ???? 2003. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1022201720574.

Rasmussen:2006:LBD

[Ras06] Nicolas Rasmussen. Lives of a blockbuster drug. Minerva, 44 (2):229–234, June 2006. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-006-0008-6.

Ratnam:1965:HEP

[Rat65] K. J. Ratnam. Higher education and policies in Singapore. Minerva, 3(4):526–529, Summer 1965. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01099417.

Ravetz:1974:SS

[Rav74] Jerome R. Ravetz. The safety of safeguards. Minerva, 12 (3):323–325, July 1974. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01102527.

Rothblatt:1976:BR

[RAV76] Sheldon Rothblatt, Noel Annan, and Rainald Von Gizycki. Book reviews. Minerva, 14(2):251–274, Summer 1976. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01107889.

Ravetz:1978:SKE

[Rav78] Jerome Ravetz. Scientific knowledge and expert advice in debates about large technological innovations. Minerva, 16(2): 273–282, Summer 1978. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096017.

Ray:2009:RPC

[Ray09] Tim Ray. Rethinking Polanyi’s concept of tacit knowledge: From personal knowing to imagined institutions. Minerva, 47 (1):75–92, March 2009. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-009-9119-1.

Razem:1994:RPF

[Raž94] Dušan Ražem. Radiation processing in the former Yugoslavia, 1947–1966: From “big science” to nullity. Minerva, 32(3):

309–326, Autumn 1994. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01098665.

Rodman:2013:IQH

[RBB13] Karmen Rodman, Roberto Biloslavo, and Silva Bratož. Institutional quality of a higher education institution from the perspective of employers. Minerva, 51(1):71–92, March 2013. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-013-9219-9.

Ravetz:1973:BR

[RBM73] J. R. Ravetz, Michael J. Beloff, and Henry Miller. Book reviews. Minerva, 11(4):633–647, October 1973. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01557803.

Ritchie-Calder:1967:RWR

[RC67] Ritchie-Calder. The rejection of the who research centre. Minerva, 5(4):571–573, Summer 1967. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01096786.

Ronan:1988:BR

[RC88] Colin Ronan and Elisabeth Crawford. Book reviews. Minerva, 26(4):611–626, Winter 1988. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096503.

Rappa:1993:YSI

[RD93] Michael Rappa and Koenraad Debackere. Youth and scientific innovation: The role of young scientists in the development of a new field. Minerva, 31(1):1–20, Spring 1993. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096169.

Rushforth:2015:AIJ

[RdR15] Alexander Rushforth and Sarah de Rijcke. Accounting for impact? The journal impact factor and the making of biomedical research in the Netherlands. Minerva, 53(2):117–139, June 2015. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/content/pdf/
10.1007/s11024-015-9274-5.pdf.

Reed:1975:HME

[Ree75] Howard A. Reed. Hacettepe and Middle East technical universities: New universities in Turkey. Minerva, 13(2):200–235, Summer 1975. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01097795.

Reif:1980:HBS

[Rei80] Arnold E. Reif. Hippocrates and the biomedical scientist: The problems of controlling the dangers of science. Minerva, 18 (1):29–50, Spring 1980. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096660.

Reingold:1997:FFF

[Rei97] Nathan Reingold. Form, function and fecundity in American institutions. Minerva, 35(3):221–232, September 1997. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1023/A%3A1004290219634.

Reinisch:2007:NBG

[Rei07] Jessica Reinisch. A new beginning? German medical and political traditions in the aftermath of the Second World War. Minerva, 45(3):241–257, September 2007. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
007-9045-z.

Rentetzi:2008:URI

[Ren08] Maria Rentetzi. The U.S. radium industry: Industrial in-house research and the commercialization of science. Minerva, 46(4): 437–462, December 2008. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-008-9111-1.

Reynolds:1987:DMC

[Rey87] Philip Reynolds. Decision-making in a contracting university: The case of the University of Lancaster, 1979–1985. Minerva, 25 (4):397–422, Winter 1987. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01099440.

Reymert:2021:BAR

[Rey21] Ingvild Reymert. Bibliometrics in academic recruitment: a screening tool rather than a game changer. Minerva, 59(1): 53–78, March 2021. CODEN MINEFY. ISSN 0026-4695 (print),

1573-1871 (electronic). URL http://link.springer.com/
article/10.1007/s11024-020-09419-0;   http://link.springer.
com/content/pdf/10.1007/s11024-020-09419-0.pdf.

Riesman:1973:EEE

[RG73] David Riesman and Gerald Grant. Evangelism, egalitarianism, and educational reform. Minerva, 11(3):296–317, July 1973. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01557686.

Risien:2021:BCB

[RG21] Julie Risien and Bruce Evan Goldstein. Boundaries crossed and boundaries made: The productive tension between learning and influence in transformative networks. Minerva, 59(4): 539–563, December 2021. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL https://link.springer.

com/article/10.1007/s11024-021-09442-9.

Reingold:1989:BR

[RH89] Nathan Reingold and A. H. Halsey. Book reviews. Minerva, 27(2–3):412–423, Summer–Autumn 1989. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01102570.

Rudwick:1990:BR

[RH90] Martin Rudwick and Notker Hammerstein. Book reviews. Minerva, 28(3):386–399, Autumn 1990. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01096295.

Richards:1965:AUA

[Ric65] Audrey I. Richards. The adaptation of universities to the African situation: Review article. Minerva, 3(3):336–342, Spring 1965. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01099959.

Richards:1988:GBA

[Ric88] Pamela Spence Richards. Great Britain and Allied scientific information: 1939–1945. Minerva, 26(2):177–198, Summer 1988. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096695.

Richards:1990:MSK

[Ric90a] Pamela Spence Richards. The movement of scientific knowledge from and to Germany under national socialism. Minerva, 28 (4):401–425, Winter 1990. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096649.

Richardson:1990:PIL

[Ric90b] Malcolm Richardson. Philanthropy and the internationality of learning: The Rockefeller Foundation and national socialist Germany. Minerva, 28(1):21–58, Spring 1990. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01096325.

Richardson:2007:LEP

[Ric07] Alan Richardson. ‘logical empiricism’ and the philosophy of science. Minerva, 45(3):357–360, September 2007. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
s11024-007-9048-9.

Riesman:1969:UGS

[Rie69] David Riesman. Universities and the growth of science in Germany and the United States. Minerva, 7(4):751–755, Summer 1969. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01099545.

Riesman:1983:SPT

[Rie83] David Riesman. Some personal thoughts on the academic ethic. Minerva, 21(2–3):265–284, Summer–Autumn 1983. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01097966.

Rimmington:1998:BRC

[Rim98] Don Rimmington. Book review: The Cold War and Academic Governance: The Lattimore Case at Johns Hopkins, by Lionel S. Lewis. Minerva, 36(1):81–84, March 1998. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1023/A%3A1004338622199.

Ritterbush:1966:RTG

[Rit66] Philip C. Ritterbush. Research training in governmental laboratories in the United States. Minerva, 4(2):186–201, Winter 1966. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01584854.

Ritterbush:1971:ESS

[Rit71] Philip C. Ritterbush. Environmental studies: The search for an institutional form. Minerva, 9(4):493–509, October 1971. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01558021.

Rose:1991:IPC

[RLH91] Kenneth W. Rose, Erwin Levold, and Lee R. Hiltzik. Ivan Pavlov on communist dogmatism and the autonomy of science in the Soviet Union in the early 1920s. Minerva, 29(4):463–475, Winter 1991. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01113492.

Rabkin:1984:TTC

[RLL84] Yakov M. Rabkin and J. Ann Lévi-Lloyd. Technology and two cultures: One hundred years of engineering education in Montreal. Minerva, 22(1):67–95, Spring 1984. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF02207557.

Roberts:1994:UBT

[Rob94] John Roberts. The universities between their internal and external enemies: II. Minerva, 32(3):327–333, Autumn 1994. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01098666.

Roberts:1998:BRS

[Rob98] Lissa Roberts. Book review: Studies in the Culture of Science in France and Britain Since the Enlightenment, by Maurice Crosland. Minerva, 36(1):86–89, March 1998. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1023/
A%3A1004398807178.

Robbins:2003:UPP

[Rob03] Keith Robbins. Universities: Past, present, and future. Minerva, 41(4):397–406, ???? 2003. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/B%3AMINE.0000005158.32821.
aa.

Roberson:2021:SSQ

[Rob21] Tara M. Roberson. On the social shaping of quantum technologies: an analysis of emerging expectations through grant proposals from 2002–2020. Minerva, 59(3):379–397, September 2021. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL https://link.springer.com/article/10.
1007/s11024-021-09438-5.

Rodgers:2012:BBU

[Rod12] Diane M. Rodgers. Busy as a bee or unemployed?: Shifting scientific discourse on work. Minerva, 50(1):45–64, March 2012. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-012-9187-5.

Rodder:2017:CSA

[Röd17] Simone Rödder. The climate of science–art and the art–science of the climate: Meeting points, boundary objects and boundary work. Minerva, 55(1):93–116, March 2017. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
016-9312-y.

Rogers:2013:BRR

[Rog13] Juan D. Rogers. Book review: Richard Whitley, Jochen Gläser and Lars Engwall (eds.), Reconfiguring Knowledge Production: Changing Authority Relationships in the Sciences and Their Consequences for Intellectual Innovation. Minerva, 51 (1):127–129, March 2013. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-013-9222-1.

Roland:2008:HB

[Rol08] Alex Roland. Heyday of the boffins. Minerva, 46(1):159– 163, March 2008. CODEN MINEFY. ISSN 0026-4695 (print),

1573-1871 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11024-007-9083-6.

Ronayne:1979:FTD

[Ron79] Jarlath Ronayne. Further thoughts on diversity and adaptability in Australian science policy. Minerva, 17(3):445–458, Autumn 1979. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096223.

Rose:1967:RWR

[Ros67a] Hilary Rose. The rejection of the who research centre: a case study of decision-making in international scientific collaboration. Minerva, 5(3):340–356, Spring 1967. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF02388573.

Rose:1967:OSB

[Ros67b] Steven P. R. Rose. Organising for science in Britain: How far is there still to go? Minerva, 5(4):578–581, Summer 1967. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096789.

Rose:1968:RWR

[Ros68] Hilary Rose. The rejection of the WHO research centre. Minerva, 6(2):263–264, Winter 1968. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01097262.

Ross:1972:DAP

[Ros72] Murray G. Ross. The dilution of academic power in Canada: The University of Toronto act. Minerva, 10(2):242–258, April 1972. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01682420.

Rose:2002:CPH

[Ros02] Steven Rose. Can philosophy help biology, or philosophers understand biologists? Minerva, 40(2):181–187, ???? 2002. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1023/A%3A1015756819520.

Rottenberg:1966:WBR

[Rot66] Simon Rottenberg. The warrants for basic research. Minerva, 5 (1):30–38, Autumn 1966. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF02389791.

Rottenberg:1981:ESP

[Rot81] Simon Rottenberg. The economy of science: The proper role of government in the growth of science. Minerva, 19 (1):43–71, Spring 1981. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF02192548.

Rottenberg:1986:USA

[Rot86] Simon Rottenberg. The universities and south Africa: The campaign for divestment. Minerva, 24(2–3):223–241, Summer– Autumn 1986. CODEN MINEFY. ISSN 0026-4695 (print),

1573-1871 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF01096612.

Rothman:1990:JBS

[Rot90] Stanley Rothman. Journalists, broadcasters, scientific experts and public opinion. Minerva, 28(2):117–133, Summer 1990. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02219656.

Rothblatt:1995:CKP

[Rot95] Sheldon Rothblatt. Clark Kerr and the pursuit of excellence in the modern university. Minerva, 33(3):265–277, Autumn 1995. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01117503.

Rothblatt:1999:HMR

[Rot99] Sheldon Rothblatt. Historical methods of reshaping the map of learning. Minerva, 37(3):281–293, September 1999. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1023/A%3A1004762011359.

Rothblatt:2004:TTB

[Rot04] Sheldon Rothblatt. A tale of two Berkeleys. Minerva, 42(2):173– 189, ???? 2004. CODEN MINEFY. ISSN 0026-4695 (print),

1573-1871 (electronic). URL http://link.springer.com/
accesspage/article/10.1023/B%3AMINE.0000030021.29930.
5b.

Rothblatt:2007:TPP

[Rot07] Sheldon Rothblatt. Thinking in perfect paragraphs. Minerva, 45(4):445–458, December 2007. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
007-9059-6.

Rousseau:2007:TKK

[Rou07] George Rousseau. Two kinds of knowledge at the crossroads: Literature and science, literature and medicine, as types of cultural understanding. Minerva, 45(1):63–71, March 2007. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-006-9021-z.

Roumbanis:2019:SVA

[Rou19] Lambros Roumbanis. Symbolic violence in academic life: a study on how junior scholars are educated in the art of getting funded. Minerva, 57(2):197–218, June 2019. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/article/10.1007/s11024-018-9364-2;

 http://link.springer.com/content/pdf/10.1007/s11024-
018-9364-2.pdf.

Roumbanis:2024:NAE

[Rou24] Lambros Roumbanis. New arguments for a pure lottery in research funding: a sketch for a future science policy without timeconsuming grant competitions. Minerva, 62(2):145–165, June 2024. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL https://link.springer.com/article/10.
1007/s11024-023-09514-y.

Rowe:1964:SIP

[Row64] A. P. Rowe. From scientific idea to practical use. Minerva, 2 (3):303–319, Spring 1964. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01097319.

Roy:1984:ARP

[Roy84] Rustum Roy. Alternatives to review by peers: a contribution to the theory of scientific choice. Minerva, 22(3–4):316–328, Autumn–Winter 1984. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF02207367.

Rozwadowski:2004:IEN

[Roz04] Helen M. Rozwadowski. Internationalism, environmental necessity, and national interest: Marine science and other sciences. Minerva, 42(2):127–149, ???? 2004. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1023/
B%3AMINE.0000030023.04586.45.

Rozwadowski:2005:OAS

[Roz05] Helen M. Rozwadowski. Oceans apart? STS and international relations. Minerva, 43(4):445–448, December 2005. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/s11024-005-2470-y.

Rose:1971:LFA

[RR71] Hilary Rose and Steven Rose. Legal framework for the assessment and control of technology. Minerva, 9(4):560–562, October 1971. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01558031.

Roca-Rosell:2005:PTE

[RR05] Antoni Roca-Rosell. Professionalism and technocracy: Esteve Terradas and science policy in the early years of the Franco regime. Minerva, 43(2):147–162, June 2005. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
005-2329-2.

Russell:1976:LGU

[RS77] Beverly Russell and Arnold Shore. Limitations on the governmental use of social science in the United States. Minerva, 14(4): 475–495, Winter 1976–1977. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01097636.

Ritchie:1970:BR

[RSR^{+}70] E. Ritchie, M. G. Simpson, Ray Richardson, Sylvia G. Haim, Julius Gould, Cyril Smith, and Sheldon Rothblatt. Book reviews. Minerva, 8(1–4):458–473, January 1970. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01553233;   http://link.springer.com/article/
10.1007/BF01553233.

Rose:1997:PIB

[RSS97] Kenneth W. Rose, Benjamin R. Shute, and Darwin H. Stapleton. Philanthropy and institution-building in the twentieth century. Minerva, 35(3):203–205, September 1997. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1023/A%3A1004287300070.

Rudd:1968:REG

[Rud68] Ernest Rudd. The rate of economic growth, technology and the Ph.D. Minerva, 6(3):366–387, Spring 1968. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01096423.

Ruegg:1975:ISG

[Rüe75] Walter Rüegg. The intellectual situation in German higher education. Minerva, 13(1):103–120, Spring 1975. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01096244.

Ruegg:1983:AAC

[Rue83] Walter Ruegg. I: an achievement of academic citizenship: Professors, government and the people of the Canton of Berne. Minerva, 21(1):101–140, Spring 1983. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01099505.

Ruegg:1986:AE

[Rüe86] Walter Rüegg. The academic ethos. Minerva, 24(4):393–412, Winter 1986. CODEN MINEFY. ISSN 0026-4695 (print),

1573-1871 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF01096704.

Ruegg:1992:TUF

[Rüe92] Walter Rüegg. The traditions of the university in the face of the demands of the twenty-first century. Minerva, 30(2): 189–205, Summer 1992. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096882.

Ruland:2023:BSB

[Rül23] Anna-Lena Rüland. Big science, big trouble? Understanding conflict in and around big science projects and networks. Minerva, 61(4):553–580, December 2023. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL https://link.

springer.com/article/10.1007/s11024-023-09497-w.

Russell:1994:AFR

[Rus94] Conrad Russell. Academic freedom: a reply. Minerva, 32(3): 334–343, Autumn 1994. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01098667.

Rommetveit:2020:MWR

[RvDG20] Kjetil Rommetveit, Niels van Dijk, and Kristrún Gunnarsdóttir. Make way for the robots! Human- and machinecentricity in constituting a European public–private partnership. Minerva, 58(1):47–69, March 2020. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/article/10.1007/s11024-
019-09386-1;   http://link.springer.com/content/pdf/
10.1007/s11024-019-09386-1.pdf.

Ryan:2001:HEB

[Rya01] Yoni Ryan. Higher education as a business:lessons from the corporate world. Minerva, 39(1):115–135, ???? 2001. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1023/A%3A1010334721507.

Ramayah:2013:EIK

[RYI13] T. Ramayah, Jasmine A. L. Yeap, and Joshua Ignatius. An empirical inquiry on knowledge sharing among academicians in higher learning institutions. Minerva, 51(2):131–154, June 2013. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-013-9229-7.

S:1970:HMS

[S.70a] E. S. How many scientists and technologists? Minerva, 8(1– 4):155–159, January 1970. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01553209.

S:1970:NPS

[S.70b] E. S. A neglected problem of science policy. Minerva, 8(1–4): 321–324, January 1970. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01553221;   http://
link.springer.com/article/10.1007/BF01553221.

S:1970:PUA

[S.70c] E. S. The political university and academic freedom. Minerva, 8(1–4):479–491, January 1970. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01553235;
http://link.springer.com/article/10.1007/BF01553235.

S:1971:AAU

[S.71a] E. S. Academic appointment, university autonomy and the federal government. Minerva, 9(2):161–170, April 1971. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01553154.

S:1971:AS

[S.71b] E. S. Anti-science. Minerva, 9(4):441–450, October 1971. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01558018.

S:1971:NSO

[S.71c] E. S. No salvation outside higher education. Minerva, 9 (3):313–321, July 1971. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01553358.

S:1972:MPD

[S.72a] E. S. Minerva: The past decade and the next. Minerva, 10 (1):1–9, January 1972. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01881387.

S:1972:IC

[S.72b] E. S. The invitation to Caesar. Minerva, 10(4):513–518, October 1972. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01695903.

S:1972:SBS

[S.72c] E. S. Stanford and Berlin: The spheres of politics and intellect. Minerva, 10(3):351–361, July 1972. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01556919.

S:1973:FTR

[S.73a] E. S. The freedom of teaching and research. Minerva, 11(4): 433–441, October 1973. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01557796.

S:1973:MSS

[S.73b] E. S. Muting the social sciences at Berkeley. Minerva, 11 (3):290–295, July 1973. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01557685.

S:1973:TH

[S.73c] E. S. Trojan horses. Minerva, 11(3):285–289, July 1973. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01557684.

S:1974:E

[S74a] E S. Elitism. Minerva, 12(1):1–7, January 1974. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01558536.

S:1974:EAF

[S.74b] E. S. The enemies of academic freedom. Minerva, 12(4): 405–415, October 1974. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01557637.

S:1974:IOH

[S.74c] E. S. Identity and openness in higher education. Minerva, 12 (2):258–276, April 1974. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01553180.

S:1974:PUS

[S.74d] E. S. The public understanding of science. Minerva, 12(2): 153–158, April 1974. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01553175.

S:1974:RHE

[S74e] E S. The reorganisation of higher education in Sweden. Minerva, 12(1):83–114, January 1974. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01558542.

S:1974:UD

[S.74f] E. S. An unresolved dilemma. Minerva, 12(3):295–302, July 1974. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01102525.

S:1975:AJP

[S.75a] E. S. Alternatives to judgement by peers. Minerva, 13(3): 341–348, Autumn 1975. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01097845.

S:1975:RSA

[S.75b] E. S. The right to speak in American universities: Yale University. Minerva, 13(2):305–321, Summer 1975. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01097800.

S:1976:GCR

[S.76] E. S. A great citizen of the republic of science: Michael Polanyi, 1892–1976. Minerva, 14(1):1–5, Spring 1976. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01096211.

S:1976:CAAa

[S.77] E. S. The criteria of academic appointment. Minerva, 14(4): 407–418, Winter 1976–1977. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01097633.

S:1982:IIG

[S.82] E. S. II an innovation in German higher education: The private university. Minerva, 20(1–2):213, Spring–Summer 1982. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01098194.

S:1983:CAS

[S.83] E. S. The constitution, academic self-government and academic trade unions in American state universities and colleges: a decision of the United States supreme court. Minerva, 21(2–3):

296–319, Summer–Autumn 1983. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01097969.

S:1984:CHE

[S.84a] E. S. The condition of humanistic education in the United States. Minerva, 22(3–4):404, Autumn–Winter 1984. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF02207372.

S:1984:SFC

[S.84b] E. S. Secrecy and freedom of communication in American science. Minerva, 22(3–4):421–423, Autumn–Winter 1984. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02207375.

S:1985:MS

[S.85] E. S. The morality of scientists. Minerva, 23(2):272–275, Summer 1985. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01099945.

S:1986:UNS

[S.86] E. S. Universities in the new states of Africa and Asia. Minerva, 24(4):476, Winter 1986. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096708.

S:1987:JBD

[S.87] E. S. Joseph Ben-David, 1920–1986. Minerva, 25(1–2):1–2, Spring–Summer 1987. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096852.

S:1988:APC

[S.88a] E. S. The academic profession and contemporary politics. Minerva, 26(4):575–579, Winter 1988. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01096498.

S:1988:AAR

[S.88b] E. S. Affirmative action reaffirmed. Minerva, 26(4):598–599, Winter 1988. CODEN MINEFY. ISSN 0026-4695 (print),

1573-1871 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF01096501.

S:1988:SRU

[S.88c] E. S. Some reflections on universities after the disturbances at the end of the 1960s. Minerva, 24(1):98–99, Spring 1988. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01102555.

S:1990:RCR

[S.90] E. S. I report of the committee on recruitment and retention of minority group members on the faculty at Yale, 16 May, 1989. Minerva, 28(2):221–242, Summer 1990. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF02219659.

S:1992:UTF

[S.92] E. S. The university of the twenty-first century: A symposium to celebrate the centenary of the University of Chicago. Minerva, 30(2):129, Summer 1992. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096877.

S:1993:TU

[S.93] E. S. The teaching of undergraduates. Minerva, 31(2):211– 227, Summer 1993. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01125435.

Sa:2005:RPE

[Sá05] Creso Sá. Research policy in emerging economies: Brazil’s sector funds. Minerva, 43(3):245–263, ???? 2005. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
005-6474-4.

Sadlak:1991:UAU

[Sad91] Jan Sadlak. The use and abuse of the university: Higher education in Romania, 1860–1990. Minerva, 29(2):195–225, Summer 1991. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096408.

Simmons:1992:PAA

[SAJ92] Howard Simmons, Lamar Alexander, and Scott Jaschik. The progress of affirmative action: Accreditation and diversity. Minerva, 30(4):552–569, Winter 1992. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01096578.

Salomon:1964:ISP

[Sal64] Jean-Jacques Salomon. International scientific policy. Minerva, 2(4):411–434, Summer 1964. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01097536.

Salam:1966:ISD

[Sal66] Abdus Salam. The isolation of the scientist in developing countries. Minerva, 4(4):461–465, Summer 1966. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF02207978.

Salam:1991:NST

[Sal91] Abdus Salam. Notes on science, technology and science education in the development of the south. Minerva, 29(1): 90–108, Spring 1991. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096299.

Salomon:2000:STD

[Sal00] Jean-Jacques Salomon. Science, technology and democracy. Minerva, 38(1):33–51, ???? 2000. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1023/A%3A1026552331409.

Salom:2005:NPC

[Sal05] Francesc X. Barca Salom. Nuclear power for Catalonia: The role of the Official Chamber of Industry of Barcelona, 1953– 1962. Minerva, 43(2):163–181, June 2005. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
005-2330-9.

Sapolsky:1971:SPA

[Sap71] Harvey M. Sapolsky. Science policy in American state government. Minerva, 9(3):322–348, July 1971. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01553359.

Sapir:2017:PPP

[Sap17] Adi Sapir. Protecting the purity of pure research: Organizational boundary-work at an institute of basic research. Minerva, 55 (1):65–91, March 2017. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-016-9309-6.

Shiva:1980:LFC

[SB80] V. Shiva and J. Bandyopadhyay. The large and fragile community of scientists in India. Minerva, 18(4):575–594, Winter 1980. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096125.

Schleusser:1992:OSN

[SB92] Heinz Schleusser and Diethwr Breitenbach. Old strains and new initiatives in the universities of The Federal German Republic. Minerva, 30(3):422–429, Autumn 1992. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01097646.

Stensaker:2013:DEI

[SB13] Bjørn Stensaker and Mats Benner. Doomed to be entrepreneurial: Institutional transformation or institutional lockins of ‘new’ universities? Minerva, 51(4):399–416, December 2013. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-013-9238-6.

Smit:2025:BID

[SBvE25] Jorrit P. Smit, Lisa Burghardt, and Lucy van Eck. Becoming an impact-driven University: A socio-technical analysis of the reconfiguration of relevance during institutional transformation. Minerva, 63(2):381–407, June 2025. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL https://link.

springer.com/article/10.1007/s11024-024-09557-9.

Soo:2004:MRU

[SC04] Mary Soo and Cathryn Carson. Managing the research university: Clark Kerr and the University of California. Minerva, 42 (3):215–236, ???? 2004. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/B%3AMINE.0000038304.39259.
94.

Santiago:2012:MRP

[SC12] Rui Santiago and Teresa Carvalho. Managerialism rhetorics in Portuguese higher education. Minerva, 50(4):511–532, December 2012. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-012-9211-9.

Scazzieri:2003:URK

[Sca03] Roberto Scazzieri. University reform and the knowledge economy. Minerva, 41(4):407–413, ???? 2003. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1023/
B%3AMINE.0000005290.46876.ff.

Schelsky:1972:WSD

[Sch72] Helmut Schelsky. The wider setting of disorder in the German universities. Minerva, 10(4):614–626, October 1972. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01695908.

Schwartzman:1978:SBS

[Sch78] Simon Schwartzman. Struggling to be born: The scientific community in Brazil. Minerva, 16(4):545–580, Winter 1978. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01100332.

Schoijet:1979:CMS

[Sch79] Mauricio Schoijet. The condition of Mexican science. Minerva, 17(3):381–412, Autumn 1979. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01096221.

SchOtt:1980:FRS

[Sch80a] Thomas SchØtt. Fundamental research in a small country: Mathematics in Denmark 1928–1977. Minerva, 18(2):243– 283, Summer 1980. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01100248.

Schultz:1980:PRP

[Sch80b] Theodore W. Schultz. The productivity of research the politics and economics of research. Minerva, 18(4):644–651, Winter 1980. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096128.

Schultz:1983:IUP

[Sch83] Theodore W. Schultz. II an unpersuasive plea for centralised control of agricultural research: On a report of the Rockefeller Foundation. Minerva, 21(1):141–143, Spring 1983. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01099506.

Schwartzman:1986:CFC

[Sch86] Simon Schwartzman. Coming full circle: a reappraisal of university research in Latin America. Minerva, 24(4):456–475, Winter 1986. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096707.

Schott:1987:SPI

[Sch87a] Thomas Schøtt. Scientific productivity and international integration of small countries: Mathematics in Denmark and Israel. Minerva, 25(1–2):3–20, Spring–Summer 1987. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01096853.

Schultz:1987:USS

[Sch87b] Theodore W. Schultz. I are university scholars and scientists free agents? Minerva, 25(3):349–357, Autumn 1987. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01097788.

Schott:1991:WSC

[Sch91] Thomas Schott. The world scientific community: Globality and globalisation. Minerva, 29(4):440–462, Winter 1991. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01113491.

Schott:1993:MSS

[Sch93] Thomas Schott. The movement of science and of scientific knowledge: Joseph Ben-David’s contribution to its understanding. Minerva, 31(4):455–477, Winter 1993. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01096450.

Schwartzman:1994:STB

[Sch94] Simon Schwartzman. Science and technology in Brazil: a new policy for a global world. Minerva, 32(4):440–468, Winter 1994. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01098033.

Schneider:2003:MAF

[Sch03a] William H. Schneider. The model American foundation officer: Alan Gregg and the Rockefeller Foundation medical divisions. Minerva, 41(2):155–166, ???? 2003. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1023/
A%3A1023658831039.

Schneider:2003:WPN

[Sch03b] William H. Schneider. War, philanthropy, and the national institute of hygiene in France. Minerva, 41(1):1–23, ???? 2003. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1022257805553.

Schimank:2005:NPM

[Sch05] Uwe Schimank. ‘new public management’ and the academic profession: Reflections on the German situation. Minerva, 43(4): 361–376, December 2005. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-005-2472-9.

Schuring:2006:ECL

[Sch06] Michael Schüring. Expulsion, compensation, and the legacy of the Kaiser Wilhelm Society. Minerva, 44(3):307–324, June 2006. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-006-9003-1.

Schuetze:2007:RUS

[Sch07] Hans Georg Schuetze. Research universities and the spectre of academic capitalism. Minerva, 45(4):435–443, December 2007. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-007-9058-7.

Schauz:2014:WBR

[Sch14] Désirée Schauz. What is basic research? Insights from historical semantics. Minerva, 52(3):273–328, September 2014. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s11024-014-9255-0.pdf.

Schonbauer:2025:ECH

[Sch25] Sarah Maria Schönbauer. Environmental care: How marine scientists relate to environmental changes. Minerva, 63(1):93– 113, March 2025. CODEN MINEFY. ISSN 0026-4695 (print),

1573-1871 (electronic). URL https://link.springer.com/
article/10.1007/s11024-024-09538-y.

Scott:1971:BR

[SCJP71] Anthony Scott, Charles F. Carter, R. G. Jobling, and Moshe Prywes. Book reviews. Minerva, 9(2):298–311, April 1971. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01553164.

Scott:1967:ICH

[Sco67] Anthony Scott. The international circulation of human capital. Minerva, 6(1):112–116, Autumn 1967. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01101612.

Scott:1971:SPA

[Sco71] Anthony Scott. On some positive aspects of the economics of the brain drain. Minerva, 9(4):558–560, October 1971. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01558030.

Scott:2001:HES

[Sco01] Peter Scott. Higher education sans Frontières. (French) []. Minerva, 39(1):137–141, ???? 2001. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1023/A%3A1010386722415.

Scott:2005:UKE

[Sco05] Peter Scott. Universities and the knowledge economy. Minerva, 43(3):297–309, ???? 2005. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-005-6510-4.

Scott:2007:PKW

[Sco07] Peter Scott. From Professor to ‘knowledge worker’: Profiles of the academic profession. Minerva, 45(2):205–215, June 2007. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-007-9031-5.

Smith-Doerr:2008:DPP

[SD08] Laurel Smith-Doerr. Decoupling policy and practice: How life scientists respond to ethics education. Minerva, 46(1): 1–16, March 2008. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-007-9084-5.

Samuel:2019:EEP

[SDvL19] Gabrielle Samuel, Gemma E. Derrick, and Thed van Leeuwen. The ethics ecosystem: Personal ethics, network governance and regulating actors governing the use of social media research data. Minerva, 57(3):317–343, September 2019. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/article/10.1007/s11024-
019-09368-3;   http://link.springer.com/content/pdf/
10.1007/s11024-019-09368-3.pdf.

Segerstedt:1985:CPE

[Seg85] Torgny Segerstedt. On certain problems in the ethical obligations of University teachers. Minerva, 23(1):62–74, Spring 1985. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01097840.

Seifart:1981:SRG

[Sei81] Werner Seifart. The support of research by German foundations: Functional and legal aspects. Minerva, 19(1):72–91, Spring 1981. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02192549.

Seitz:1988:ISP

[Sei88] Frederick Seitz. II some personal observations on the aftermath of the disturbances. Minerva, 24(1):130–133, Spring 1988. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01102557.

Selgelid:2008:MGB

[Sel08] Michael J. Selgelid. Michael L. Gross, bioethics and armed conflict: Moral dilemmas of medicine and war. Minerva, 46(3): 381–384, September 2008. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-008-9103-1.

Sigl:2025:THD

[SF25] Lisa Sigl and Maximilian Fochler. Towards a “Hinterland” for doing relevance. a typology of practices and competencies to guide the development of more relevant research and career paths. Minerva, 63(2):351–380, June 2025. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL https://

link.springer.com/article/10.1007/s11024-024-09569-5.

Schroeder-Gudehus:1972:ASG

[SG72] Brigitte Schroeder-Gudehus. The argument for the selfgovernment and public support of science in Weimar Germany. Minerva, 10(4):537–570, October 1972. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01695905.

Simon:1992:BR

[SGM92] Herbert A. Simon, J. W. Grove, and Allan Mazur. Book reviews. Minerva, 30(4):570–584, Winter 1992. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01096579.

Schmitt:2023:CII

[SGSB23] Josephine B. Schmitt, Anne Goldmann, Samuel T. Simon, and Christoph Bieber. Conception and interpretation of interdisciplinarity in research practice: Findings from group discussions in the emerging field of digital transformation. Minerva, 61 (2):199–220, June 2023. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL https://link.springer.

com/article/10.1007/s11024-023-09489-w.

Stigler:1978:BR

[SH78] George J. Stigler and R. A. Hodgkin. Book reviews. Minerva, 16 (4):596–605, Winter 1978. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01100335.

Simon:2002:EIR

[SH02] Jacqueline Simon and Melissa Hersh. An educational imperative: The role of ethical codes and normative prohibitions in CBW-applicable research. Minerva, 40(1):37–55, ???? 2002. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1015001429180.

Sousa:2007:OOD

[SH07] Celio A. A. Sousa and Paul H. J. Hendriks. That obscure object of desire: the management of academic knowledge. Minerva, 45(3):259–274, September 2007. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/content/pdf/10.1007/s11024-
007-9046-y.pdf.

Sejersen:2018:MEP

[SH18] Nadja Sejersen and Janus Hansen. From a means to an end: Patenting in the 1999 Danish ‘Act on Inventions’ and its effect on research practice. Minerva, 56(3):261–281, September 2018. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s11024-017-9336-y.

Shah:1973:IAP

[Sha73] A. B. Shah. The Indian academic profession. Minerva, 11(1): 137–139, January 1973. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01544985.

Shattuck:1984:HUB

[Sha84] John Shattuck. Harvard University basic science, secrecy and national security. Minerva, 22(3–4):424–436, Autumn–Winter 1984. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02207376.

Shattock:1987:LDU

[Sha87] Michael Shattock. The last days of the University grants committee. Minerva, 25(4):471–485, Winter 1987. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01099443.

Shattock:1989:HER

[Sha89a] Michael Shattock. Higher education and the research councils. Minerva, 27(2–3):195–222, Summer–Autumn 1989. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01102568.

Shaw:1989:APC

[Sha89b] Thurstan Shaw. The academic profession and contemporary politics. Minerva, 27(1):58–86, Spring 1989. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01096715.

Shapiro:1992:FRA

[Sha92a] Harold T. Shapiro. The functions and resources of the American University of the twenty-first century. Minerva, 30(2): 163–174, Summer 1992. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096880.

Shattock:1992:IET

[Sha92b] Michael Shattock. The internal and external threats to the university of the twenty-first century. Minerva, 30(2):130–147, Summer 1992. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096878.

Shattock:1995:ES

[Sha95] Michael Shattock. Edward Shils. Minerva, 33(1):1–3, Spring 1995. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01098615.

Shattock:2000:E

[Sha00] Michael Shattock. Editorial. Minerva, 37(4):301–302, December 2000. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1004729825211.

Shah:2017:WSS

[Sha17] Esha Shah. Who is the scientist–subject? A critique of the neo-Kantian scientist–subject in Lorraine Daston and Peter Galison’s objectivity. Minerva, 55(1):117–138, March 2017. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s11024-017-9313-5.pdf.

Shimshoni:1965:ISP

[Shi65] Daniel Shimshoni. Israeli scientific policy. Minerva, 3(4): 441–456, Summer 1965. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01099410.

Shils:1969:API

[Shi69] Edward Shils. The academic profession in India. Minerva, 7 (3):345–372, Spring 1969. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01103266.

Shimshoni:1970:MSA

[Shi70] Daniel Shimshoni. The mobile scientist in the American instrument industry. Minerva, 8(1–4):59–89, January 1970. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01553199.

Shils:1973:APU

[Shi73] Edward Shils. The American private university. Minerva, 11 (1):6–29, January 1973. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01544975.

Shils:1975:AEU

[Shi75] Edward Shils. The academic ethos under strain. Minerva, 13 (1):1–37, Spring 1975. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096240.

Shils:1978:OLU

[Shi78] Edward Shils. The order of learning in the United States from 1865 to 1920: The ascendancy of the universities. Minerva, 16 (2):159–195, Summer 1978. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096013.

Shils:1979:GUU

[Shi79] Edward Shils. Government and universities in the United States. Minerva, 17(1):129–177, Spring 1979. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01096263.

Shinn:1980:OIS

[Shi80] Terry Shinn. Orthodoxy and innovation in science: The atomist controversy in French chemistry. Minerva, 18(4):539–555, Winter 1980. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096123.

Shinn:1984:RTS

[Shi84] Terry Shinn. Reactionary technologists: The struggle over the école polytechnique, 1880–1914. Minerva, 22(3–4):329–345, Autumn–Winter 1984. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF02207368.

Shils:1987:JBD

[Shi87] Edward Shils. Joseph Ben-David: a memoir. Minerva, 25(1– 2):201–205, Spring–Summer 1987. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01096864.

Shils:1989:MUL

[Shi89] Edward Shils. The modern university and liberal democracy. Minerva, 27(4):425–460, Winter 1989. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01096539.

Shils:1991:RTC

[Shi91] Edward Shils. Reflections on tradition, centre and periphery and the universal validity of science: The significance of the life of S. Ramanujan. Minerva, 29(4):393–419, Winter 1991. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01113489.

Shils:1992:PDI

[Shi92a] Edward Shils. Points of departure: The idea of the university: Obstacles and opportunities in contemporary societies. Minerva, 30(2):301–313, Summer 1992. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01096887.

Shils:1992:PDS

[Shi92b] Edward Shils. Points of departure: The situation of the universities in the twenty-first century. Minerva, 30(2):296–301, Summer 1992. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096886.

Shils:1992:SSA

[Shi92c] Edward Shils. The service of society and the advancement of learning in the twenty-first century. Minerva, 30(2):242– 268, Summer 1992. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096884.

Shils:1994:BUT

[Shi94a] Edward Shils. The British universities in tribulation. Minerva, 32(2):200–219, Summer 1994. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01103361.

Shils:1994:DWS

[Shi94b] Edward Shils. Do we still need academic freedom? Minerva, 32(1):79–98, Spring 1994. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096657.

Shils:1995:AFP

[Shi95] Edward Shils. Academic freedom and permanent tenure. Minerva, 33(1):5–17, Spring 1995. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01098616.

Shinn:2016:TRS

[Shi16] Terry Shinn. Toward a reactionary science? Minerva, 54 (2):241–253, June 2016. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-016-9296-7.

Shin:2018:PEI

[Shi18] Youjung Shin. A policy entrepreneur in the information society: Shaping the interdisciplinarity of brain research in Korea. Minerva, 56(2):231–257, June 2018. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/article/10.1007/s11024-017-9328-y.

Shonfield:1972:SSG

[Sho72] Andrew Shonfield. The social sciences in the great debate on science policy. Minerva, 10(3):426–438, July 1972. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01556923.

Silcock:1964:DUS

[Sil64] T. H. Silcock. The development of universities in South-East Asia to 1960. Minerva, 2(2):169–196, Winter 1964. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01096593.

Singh:1962:IFS

[Sin62] Amar Kumar Singh. The impact of foreign study: The Indian experience. Minerva, 1(1):43–53, Autumn 1962. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01101452.

Singh:1992:SIU

[Sin92] Amrik Singh. Science in the Indian universities. Minerva, 30 (1):51–61, Spring 1992. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096396.

Siotis:1969:SUG

[Sio69] Jean Siotis. The situation of the universities in Greece. Minerva, 7(3):490–493, Spring 1969. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01103272.

S:1977:SAA

[SJ77] E. S and Harry G. Johnson. The soil and air of academic life. Minerva, 15(2):200–213, Summer 1977. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01096677.

Shinn:1990:FPR

[SJ90] Terry Shinn and Ivor Jennings. “Formation par la recherche”: Technological training through research in France. Minerva, 28 (2):134–146, Summer 1990. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF02219657.

Szollosi-Janze:2005:SSS

[SJ05] Margit Szöllösi-Janze. Science and social space: Transformations in the institutions of Wissenschaft from the Wilhelmine Empire to the Weimar Republic. Minerva, 43(4):339–360, December 2005. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-005-2473-8.

Sjogren:2007:DMP

[Sjö07] Ebba Sjögren. Defining ‘markets’ for pharmaceuticals in Sweden: Public policy and commercialization. Minerva, 45(2): 161–173, June 2007. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-007-9040-4.

Shin:2014:RPR

[SJPA14] Jung Cheol Shin, Jisun Jung, Gerard A. Postiglione, and Norzaini Azman. Research productivity of returnees from study abroad in Korea, Hong Kong, and Malaysia. Minerva, 52(4): 467–487, December 2014. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-014-9259-9.

Smith:2000:TCA

[SK00] Bruce L. R. Smith and David Korn. Is there a crisis of accountability in the American research University? Minerva, 38 (2):129–145, ???? 2000. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1026544926596.

Skilbeck:2006:EKS

[Ski06] Malcolm Skilbeck. Educating the knowledge society. Minerva, 44(1):89–101, March 2006. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-005-5403-x.

Skoie:1969:PSS

[Sko69] Hans Skoie. The problems of a small scientific community: The Norwegian case. Minerva, 7(3):399–425, Spring 1969. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01103268.

Sa:2013:TNC

[SKS13] Creso Sá, Andrew Kretz, and Kristjan Sigurdson. Technonationalism and the construction of University technology transfer. Minerva, 51(4):443–464, December 2013. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
s11024-013-9242-x.

Skupien:2019:SMM

[Sku19] Stefan Skupien. Searching for macro-meso-micro-level links in studies of North–South research collaborations. Minerva, 57(3): 391–410, September 2019. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/article/10.1007/s11024-019-09371-8.

Smith:1989:CRS

[SL89] Cyril S. Smith and Otto N. Larsen. The criterion of “relevance” in the support of research in the social sciences: 1965–1985. Minerva, 27(4):461–482, Winter 1989. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01096540.

Sigl:2018:IIM

[SL18] Lisa Sigl and Liudvika Leišytė. Imaginaries of invention management: Comparing path dependencies in East and West Germany. Minerva, 56(3):357–380, September 2018. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/article/10.1007/s11024-018-9347-3;
 http://link.springer.com/content/pdf/10.1007/s11024-
018-9347-3.pdf.

Skrentny:2022:BSP

[SL22] John D. Skrentny and Kevin Lewis. Beyond the “STEM Pipeline”: Expertise, careers, and lifelong learning. Minerva, 60(1):1–28, March 2022. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL https://link.springer.

com/article/10.1007/s11024-021-09445-6.

Sundqvist:2024:MSR

[SL24] Göran Sundqvist and Sebastian Linke. Making science relevant: Comparing two science advisory organizations beyond the linear knowledge model. Minerva, 62(4):527–547, December 2024. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL https://link.springer.com/article/10.
1007/s11024-024-09528-0.

Slade:2011:PVM

[Sla11] Catherine P. Slade. Public value mapping of equity in emerging nanomedicine. Minerva, 49(1):71–86, March 2011. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/s11024-011-9163-5.

Sarac-Lesavre:2019:STE

[SLL19] Başak Saraç-Lesavre and Brice Laurent. Stress-testing Europe: Normalizing the post-Fukushima crisis. Minerva, 57(2):239–260, June 2019. CODEN MINEFY. ISSN 0026-4695 (print), 1573-

1871 (electronic). URL http://link.springer.com/article/
10.1007/s11024-018-9362-4.

Shu:2022:CRE

[SLL22] Fei Shu, Sichen Liu, and Vincent Larivière. China’s research evaluation reform: What are the consequences for global science? Minerva, 60(3):329–347, September 2022. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL https://

link.springer.com/article/10.1007/s11024-022-09468-7.

Schofer:2022:IRH

[SLM22] Evan Schofer, Julia C. Lerch, and John W. Meyer. Illiberal reactions to higher education. Minerva, 60(4):509–534, December 2022. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL https://link.springer.com/article/10.
1007/s11024-022-09472-x.

S:1972:OSC

[SM72] E. S. and Robert E. Machol. The obligations of scientists as counsellors. Minerva, 10(1):107–157, January 1972. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01881392.

Santesmases:1997:SPS

[SM97] María Jesús Santesmases and Emilio Muñoz. The scientific periphery in Spain: The establishment of a biomedical discipline at the Centro de Investigaciones Biológicas, 1956–1967. Minerva, 35(1):27–45, ???? 1997. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1004267304624.

Samaras:2025:WIG

[SM25] Georgia Samaras and Ruth Müller. What’s it got to do with the brain? mobilising and doing clinical relevance in epigenetic psychiatric research. Minerva, 63(2):325–349, June 2025. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL https://link.springer.com/article/10.
1007/s11024-025-09575-1.

Salmela:2021:III

[SMaR21] Mikko Salmela, Miles MacLeod, and Johan Munck af Rosenschöld. Internally incentivized interdisciplinarity: Organizational restructuring of research and emerging tensions. Minerva, 59(3):355–377, September 2021. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL https://link.

springer.com/article/10.1007/s11024-020-09431-4.

Smith:1973:NSP

[Smi73] Bruce L. R. Smith. A new science policy in the United States. Minerva, 11(2):162–174, April 1973. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01557714.

Smith:1979:BDR

[Smi79] Bruce L. R. Smith. The brain drain re-emergent: Foreign medical graduates in American medical schools. Minerva, 17(4): 483–503, Winter 1979. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01101900.

Smith:1996:AS

[Smi96] Bruce L. R. Smith. The accountability of science. Minerva, 34(1):45–56, ???? 1996. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF00124200.

Smith:1999:UTW

[Smi99] David Smith. The use of “team work” in the practical management of research in the inter-war period: John Boyd Orr at the Rowett Research Institute. Minerva, 37(3):259–280, September 1999. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1004753825471.

Smith:2007:AFU

[Smi07] Bruce L. R. Smith. Academic freedom in the United States. Minerva, 45(3):321–329, September 2007. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
007-9042-2.

Smith:2008:BRM

[Smi08] Robyn Smith. Book review: Maasen, Sabine and Sutter, Barbara (eds): On Willing Selves: Neoliberal Politics vis-à-vis the Neuroscientific Challenge. Minerva, 46(4):491–494, December 2008. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/content/pdf/
10.1007/s11024-008-9107-x.pdf.

Snellen:1983:SMC

[Sne83] I. Th. M. Snellen. Social merit as a criterion of scientific choice: Its application in Dutch science policy. Minerva, 21 (1):16–36, Spring 1983. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01099502.

S:1980:ZAF

[SNSB^{+}80] E. S., Ernst Nolte, Dr. Hans Schulze-Berndt, Burkhard Zipfel, Prof. Dr. Bernd Riithers, Prof. Dr. Reinhard Mussgnug, and Dr. Peter Glotz. The Zipfel affair at the Free University of Berlin. Minerva, 18(1):132–163, Spring 1980. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01096663.

Solomon:1975:CSS

[Sol75] Susan Gross Solomon. Controversy in social science: Soviet rural studies in the 1920s. Minerva, 13(4):554–582, Winter 1975. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096177.

Solomon:2003:BBA

[Sol03] Susan Gross Solomon. Building bridges: Alan Gregg and Soviet Russia, 1925–1928. Minerva, 41(2):167–176, ????

2003. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1023636529221.

Sommer:1992:AAC

[Som92] Robert Sommer. Administrative advisory committees at the University of California at Davis. Minerva, 30(1):1–13, Spring 1992. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096393.

Somsen:2008:HUC

[Som08a] Geert J. Somsen. A history of universalism: Conceptions of the internationality of science from the enlightenment to the Cold War. Minerva, 46(3):??, September 2008. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s11024-008-9105-z.pdf.

Somsen:2008:VLS

[Som08b] Geert J. Somsen. Value-laden science: Jan Burgers and scientific politics in the Netherlands. Minerva, 46(2):231–245, June 2008. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-008-9093-z.

Song:2021:BRD

[Son21] Jieun Song. Book review: David John Frank & John W. Meyer, University and the Global Knowledge Society, Princeton University Press, 2020. 182 pp. paperback; $29.95; ISBN 978-0-691-20205-1. Minerva, 59(3):419–422, September 2021. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL https://link.springer.com/article/10.
1007/s11024-021-09450-9.

Sorlin:2006:EAA

[Sör06] Sverker Sörlin. European and American approaches to the environment. Minerva, 44(4):433–438, December 2006. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/s11024-006-9012-0.

Spann:1966:TAK

[Spa66] R. N. Spann. Thoughts after Knopfelmacher: Appointments on university. Minerva, 4(2):274–276, Winter 1966. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01584862.

Speiser:1985:ETB

[Spe85] A. P. Speiser. European technology between two poles: The United States and the Far East. Minerva, 23(4):508–520, Winter 1985. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01098704.

Sarewitz:2012:FL

[SR12] Daniel Sarewitz and Arie Rip. A forward look. Minerva, 50 (2):143–148, June 2012. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-012-9202-x.

Shir-Raz:2023:CSC

[SREM^{+}23] Yaffa Shir-Raz, Ety Elisha, Brian Martin, Natti Ronel, and Josh Guetzkow. Censorship and suppression of Covid-19 heterodoxy: Tactics and counter-tactics. Minerva, 61(3):407–433, September 2023. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL https://link.springer.com/article/10.
1007/s11024-022-09479-4.

Sivin:1992:BR

[SRM92] Nathan Sivin, George Rosen, and Edward J. Monahan. Book reviews. Minerva, 30(3):432–456, Autumn 1992. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01097648.

Sarrico:2010:AQE

[SRTC10] Cláudia S. Sarrico, Maria J. Rosa, Pedro N. Teixeira, and Margarida F. Cardoso. Assessing quality and evaluating performance in higher education: Worlds apart or complementary views? Minerva, 48(1):35–54, March 2010. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
010-9142-2.

Steiner:1977:BR

[SS77] Arthur Steiner and Bruce L. R. Smith. Book reviews. Minerva, 15(2):247–269, Summer 1977. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01096679.

S:1979:GFB

[SS79] E. S. and Theodore W. Schultz. Governments, foundations and the bias of research. Minerva, 17(3):459–468, Autumn 1979. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096224.

Satalkina:2022:SIR

[SS22] Liliya Satalkina and Gerald Steiner. Social innovation: a retrospective perspective. Minerva, 60(4):567–591, December 2022. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL https://link.springer.com/article/10.
1007/s11024-022-09471-y.

Sui:2015:CGT

[SSF15] Suli Sui and Margaret Sleeboom-Faulkner. Commercial genetic testing and its governance in Chinese society. Minerva, 53(3): 215–234, September 2015. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-015-9279-0.

Stone:1968:EDA

[SSG68] Richard Stone, Giovanna Stone, and Jane Gunton. An example of demographic accounting: The school ages. Minerva, 6 (2):185–212, Winter 1968. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01097258.

Sabzalieva:2021:SDP

[SSK21] Emma Sabzalieva, Creso M. Sá, and Nadiia Kachynska. Science diplomacy policy processes in comparative perspective: The use of scientific cooperation agreements in Canada, India, Norway, and the UK. Minerva, 59(2):149–172, June 2021. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL https://link.springer.com/article/10.
1007/s11024-020-09429-y.

Sutherland:1971:BR

[SSS^{+}71] Gordon Sutherland, Helmut Schelsky, Kalman H. Silvert, Donald G. MacRae, and A. Hunter Dupree. Book reviews. Minerva, 9(1):141–158, January 1971. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01553194.

Salter:1993:ASS

[ST93] Brian Salter and Ted Tapper. The application of science and scientific autonomy in Great Britain: a case study of the science and engineering research council. Minerva, 31(1):38–55, Spring 1993. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096171.

Schmidtlein:1996:RAR

[ST96] Frank A. Schmidtlein and Alton L. Taylor. Responses of American research universities to issues posed by the changing environment of higher education. Minerva, 34(3):291–308, ???? 1996. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00120329.

Staley:1995:RFP

[Sta95] David J. Staley. The Rockefeller Foundation and the patronage of German sociology, 1946–1955. Minerva, 33(3): 251–264, Autumn 1995. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01117502.

Stamhuis:2002:RDS

[Sta02a] Ida H. Stamhuis. Recapturing Dutch science. Minerva, 40 (4):407–415, ???? 2002. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1020981806432.

Stamhuis:2002:WAS

[Sta02b] Ida H. Stamhuis. Women, actors and subjects in science. Minerva, 40(2):211–213, ???? 2002. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1023/
A%3A1015713518611.

Stapleton:2003:JWR

[Sta03] Darwin H. Stapleton. Joseph willits and the Rockefeller’s European programme in the social sciences. Minerva, 41(2): 101–114, ???? 2003. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1023602713292.

Stanley:2008:MME

[Sta08] Matthew Stanley. Mysticism and Marxism: A. S. Eddington, Chapman Cohen, and political engagement through science popularization. Minerva, 46(2):181–194, June 2008. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/s11024-008-9089-8.

Steiner:1974:SSP

[Ste74] Arthur Steiner. Scientists, statesmen, and politicians: The competing influences on American atomic energy policy 1945–46. Minerva, 12(4):469–509, October 1974. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01557641.

S:1976:CAAb

[STH^{+}77] E. S., Paul Trappe, Jeanne Hersch, Bernard Ducret, and Herbert Ltithy. Criteria of academic appointment. Minerva, 14(4):530– 569, Winter 1976–1977. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01097638.

Stilgoe:2012:ESP

[Sti12] Jack Stilgoe. Experiments in science policy: an autobiographical note. Minerva, 50(2):197–204, June 2012. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
012-9199-1.

S:1991:AFU

[STK^{+}91] E. S., Tidningarnas Telegrambyrå, Mats Knutson, Jacob Sundberg, Anki Gundhäll, Professor Lars Gustafsson, Alan Dershowitz, Svante Nycander, Bengt Johansson, Magnus Eriksson, Lotta Gustavson, Marianne Gunnarsson, Kristina Vallström, Monique Wadsted, Mary Ann Glendon, and et al. Academic freedom at the University of Stockholm. Minerva, 29(3): 321–385, Autumn 1991. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01097881.

Stone:1965:MES

[Sto65a] Richard Stone. A model of the educational system. Minerva, 3 (2):172–186, Winter 1965. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096333.

Stout:1965:UAT

[Sto65b] A. K. Stout. On university appointments: Thoughts after Knopfelmacher. Minerva, 4(1):55–72, Autumn 1965. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01585984.

Stone:1966:IOD

[Sto66] Richard Stone. Input-output and demographic accounting: a tool for educational planning. Minerva, 4(3):365–380, Spring 1966. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01584632.

Stone:1972:EPB

[Sto72] Richard Stone. The evaluation of pollution: Balancing gains and losses. Minerva, 10(3):412–425, July 1972. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01556922.

Stone:2011:AIN

[Sto11] Diane Stone. The ASEAN–ISIS network: Interpretive communities, informal diplomacy and discourses of region. Minerva, 49(2):??, June 2011. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-011-9171-5.

Stretch:1964:AEL

[Str64] K. L. Stretch. Academic ecology: On the location of institutions of higher education. Minerva, 2(3):320–335, Spring 1964. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01097320.

Strathern:2003:RDS

[Str03] Marilyn Strathern. Re-describing society. Minerva, 41(3): 263–276, ???? 2003. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1025586327342.

Sturdy:1998:BRL

[Stu98] David J. Sturdy. Book review: Labour, Science and Technology in France, 1500–1620, by Henry Heller. Minerva, 36 (2):191–200, June 1998. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1004300219853.

Subarsky:1966:TCR

[Sub66] Zachariah Subarsky. Thoughts on chemical research and teaching in East Africa. Minerva, 4(4):561–562, Summer 1966. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02207984.

Sullivan:1991:PUW

[Sul91] Antony T. Sullivan. Palestinian universities in the West Bank and Gaza Strip. Minerva, 29(3):249–268, Autumn 1991. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01097878.

Sullivan:1995:RDP

[Sul95] Antony T. Sullivan. Recent developments in Palestinian higher education and the CEEPAT programme. Minerva, 33(1): 37–43, Spring 1995. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01098618.

Sutherland:1973:TOS

[Sut73] Gordon Sutherland. Is there an optimum size for a university? Minerva, 11(1):53–78, January 1973. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01544977.

Suttmeier:1997:PLS

[Sut97] Richard P. Suttmeier. The political life of science in post-Mao China. Minerva, 35(1):90–94, ???? 1997. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1023/
B%3AMINE.0000019490.45560.a8.

Suttmeier:1998:HET

[Sut98] Richard P. Suttmeier. Higher education in Twentieth-Century China: Promise, turmoil, transformation? Minerva, 36(3): 293–298, September 1998. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1004368727574.

Suzuki:2001:MST

[Suz01] Kenji Suzuki. Managing science and technology in Japan. Minerva, 39(3):358–362, ???? 2001. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1023/A%3A1017956912501.

Serrano-Velarde:2010:FWM

[SV10] Kathia Serrano-Velarde. A fish out of water? Management consultants in academia. Minerva, 48(2):125–144, June 2010. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-010-9148-9.

Serrano-Velarde:2014:RAI

[SV14] Kathia Serrano-Velarde. Rising above institutional constraints? The quest of German accreditation agencies for autonomy and professional legitimacy. Minerva, 52(1):97–118, March 2014. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-014-9245-2.

Sanchez-Vazquez:2015:NEP

[SVMN15] Luis Sánchez-Vázquez and Alfredo Menéndez-Navarro. Nuclear energy in the public sphere: Anti-nuclear movements vs. industrial lobbies in Spain (1962–1979). Minerva, 53(1):69–88, March 2015. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-014-9263-0.

Swatez:1970:SOU

[Swa70] Gerald M. Swatez. The social organisation of a university laboratory. Minerva, 8(1–4):36–58, January 1970. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01553198.

Szczepanski:1968:FNC

[Szc68] Jan Szczepanski. The future of the nineteenth century idea of a University. Minerva, 6(3):419–423, Spring 1968. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01096427.

Schultz:1981:BR

[SZL81] Theodore W. Schultz, J. M. Ziman, and Saunders Mac Lane. Book reviews. Minerva, 19(3):502–517, Autumn 1981. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02192828.

Tallacchini:2009:GVE

[Tal09] Mariachiara Tallacchini. Governing by values. EU ethics: Soft tool, hard effects. Minerva, 47(3):??, September 2009. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/s11024-009-9127-1.

Taubert:2012:MDS

[Tau12] Niels C. Taubert. Minerva and the development of science (policy) studies. Minerva, 50(3):261–275, September 2012. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-012-9208-4.

Taubert:2021:GOA

[Tau21] Niels Taubert. Green open access in astronomy and mathematics: The complementarity of routines among authors and readers. Minerva, 59(2):173–194, June 2021. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL https://

link.springer.com/article/10.1007/s11024-020-09424-3.

Turpin:2005:TAP

[TC05] Tim Turpin and Russel Cooper. Technology, adaptation, and public policy in developing countries: The ‘ins and outs’ of the digital divide. Minerva, 43(4):419–427, December 2005. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-005-2471-x.

Tlili:2010:MSS

[TD10] Anwar Tlili and Emily Dawson. Mediating science and society in the EU and UK: From information-transmission to deliberative democracy? Minerva, 48(4):429–461, December 2010. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-010-9160-0.

Tedeschi:1995:FTT

[Ted95] Christopher M. Tedeschi. Foetal tissue transplantation research: Scientific progress and the role of special interest groups. Minerva, 33(1):45–66, Spring 1995. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01098619.

Tellmann:2022:STA

[Tel22] Silje Maria Tellmann. The societal territory of academic disciplines: How disciplines matter to society. Minerva, 60(2):159– 179, June 2022. CODEN MINEFY. ISSN 0026-4695 (print),

1573-1871 (electronic). URL https://link.springer.com/
article/10.1007/s11024-022-09460-1.

Teubal:1983:NSP

[Teu83] Morris Teubal. Neutrality in science policy: The promotion of sophisticated industrial technology in Israel. Minerva, 21(2– 3):172–197, Summer–Autumn 1983. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01097963.

Treue:1983:BR

[TGGS83] Wilhelm Treue, Penelope M. Gouk, J. W. Grove, and Harold Stone. Book reviews. Minerva, 21(4):456–476, Winter 1983. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096528.

Tang:2024:SAW

[TH24] Li Tang and Hugo Horta. Supporting academic women’s careers: Male and female academics’ perspectives at a Chinese research university. Minerva, 62(1):113–139, March 2024. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL https://link.springer.com/article/10.
1007/s11024-023-09506-y.

Thackray:1980:PHA

[Tha80] Arnold Thackray. The pre-history of an academic discipline: The study of the history of science in the United States, 1891–1941. Minerva, 18(3):448–473, Autumn 1980. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01096952.

Taylor:2008:GSR

[THE08] Peter J. Taylor, Michael Hoyler, and David M. Evans. A geohistorical study of ‘the rise of modern science’: Mapping scientific practice through urban networks, 1500–1900. Minerva, 46(4):??, December 2008. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-008-9109-8.

Thiel:2004:PAF

[Thi04] Jens Thiel. Paul Abraham: a forgotten scholar of the Prussian Academy of Sciences and Humanities. Minerva, 42(4): 393–420, ???? 2004. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-004-2114-7.

Tao:2017:GDP

[THM17] Yu Tao, Wei Hong, and Ying Ma. Gender differences in publication productivity among academic scientists and engineers in the U.S. and China: Similarities and differences. Minerva, 55(4): 459–484, December 2017. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/article/10.1007/s11024-017-9320-6.

Thomas:1967:ICH

[Tho67] Brinley Thomas. The international circulation of human capital. Minerva, 5(4):479–506, Summer 1967. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01096779.

Thomas:1968:ICH

[Tho68] Brinley Thomas. The international circulation of human capital. Minerva, 6(3):423–427, Spring 1968. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01096428.

Thomas:1968:RBD

[Tho69] Brinley Thomas. The reduction of the brain drain: Problems and policies. Minerva, 7(1–2):92–94, Autumn–Winter 1968– 1969. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096815.

Thorpe:2010:PPF

[Tho10] Charles Thorpe. Participation as post-Fordist politics: Demos, new labour, and science policy. Minerva, 48(4):389–411, December 2010. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/content/pdf/
10.1007/s11024-010-9157-8.pdf.

Thune:2010:TTH

[Thu10] Taran Thune. The training of “triple helix workers”? Doctoral students in university–industry–government collaborations. Minerva, 48(4):463–483, December 2010. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
s11024-010-9158-7.

Turpin:2002:IHE

[TIC02] Tim Turpin, Robyn Iredale, and Paola Crinnion. The internationalization of higher education: Implications for Australia and its education ‘clients’. Minerva, 40(4):327–340, ???? 2002. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1020933923270.

Tilak:2000:HED

[Til00] Jandhyala B. G. Tilak. Higher education in developing countries. Minerva, 38(2):233–240, ???? 2000. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1023/
A%3A1026557329322.

Tilak:2024:BRM

[Til24] Jandhyala B. G. Tilak. Book review: Michael D. Smith, The Abundant University: Remaking Higher Education for a Digital World, The MIT Press, Cambridge, MA, 2023, pp. xxviii + 258 + Index ISBN: 978-0-262-04855-2 (hardcover) Price: US$35.00. Minerva, 62(4):643–646, December 2024. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL https://

link.springer.com/article/10.1007/s11024-024-09536-0.

Turner:1986:BR

[TK86] R. Steven Turner and Yung Sik Kim. Book reviews. Minerva, 24 (4):495–509, Winter 1986. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096710.

Thackray:1972:BR

[TKMB72] Arnold Thackray, G. Kloss, Kenneth Mellanby, and K. M. Baker. Book reviews. Minerva, 10(3):491–508, July 1972. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01556929.

Teich:1976:RLN

[TL77] Albert H. Teich and W. Henry Lambright. The redirection of a large national laboratory. Minerva, 14(4):447–474, Winter 1976– 1977. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01097635.

Tchalakov:2010:ASO

[TMP10] Ivan Tchalakov, Tihomir Mitev, and Venelin Petrov. The academic spin-offs as an engine of economic transition in Eastern Europe. A path-dependent approach. Minerva, 48(2): 189–217, June 2010. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-010-9149-8.

Tobias:1988:PPD

[Tob88] Phillip V. Tobias. Prehistory and politicl discrimination. Minerva, 26(4):588–597, Winter 1988. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01096500.

Tobiason:2019:CEU

[Tob19] Glory Tobiason. Countering expert uncertainty: Rhetorical strategies from the case of value–added modeling in teacher evaluation. Minerva, 57(1):109–126, March 2019. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/article/10.1007/s11024-
018-9359-z.

Tomusk:1996:RTE

[Tom96] Voldemar Tomusk. Recent trends in Estonian higher education: Emergence of the binary division from the point of view of staff development. Minerva, 34(3):279–289, ???? 1996. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF00120328.

Tomusk:1998:DRH

[Tom98] Voldemar Tomusk. Developments in Russian higher education: Legislative and policy reform within a Central and East European context. Minerva, 36(2):125–146, June 1998. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1004387802106.

Topham:2002:TAS

[Top02] Jonathan R. Topham. Not thinking about science and religion. Minerva, 40(2):203–209, ???? 2002. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1023/
A%3A1015704702681.

Torka:2018:PDT

[Tor18] Marc Torka. Projectification of doctoral training? How research fields respond to a new funding regime. Minerva, 56(1):59–83, March 2018. CODEN MINEFY. ISSN 0026-4695 (print), 1573-

1871 (electronic). URL http://link.springer.com/article/
10.1007/s11024-018-9342-8.

Toulmin:1964:CSC

[Tou64] Stephen Toulmin. The complexity of scientific choice: a stocktaking. Minerva, 2(3):343–359, Spring 1964. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01097322.

Toulmin:1966:CSCb

[Tou66a] Stephen Toulmin. The complexity of scientific choice. Minerva, 4 (4):564–565, Summer 1966. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF02207986.

Toulmin:1966:CSCa

[Tou66b] Stephen Toulmin. The complexity of scientific choice II: Culture, overheads or tertiary industry? Minerva, 4(2):155–169, Winter 1966. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01584852.

Toulmin:1967:WBR

[Tou67] Stephen Toulmin. The warrants for basic research. Minerva, 5 (4):576–578, Summer 1967. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096788.

Thoenig:2014:OGP

[TP14] Jean-Claude Thoenig and Catherine Paradeise. Organizational governance and the production of academic quality: Lessons from two top U.S. research universities. Minerva, 52(4):381– 417, December 2014. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-014-9261-2.

Thoenig:2016:SCO

[TP16] Jean-Claude Thoenig and Catherine Paradeise. Strategic capacity and organisational capabilities: A challenge for universities. Minerva, 54(3):293–324, September 2016. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
016-9297-6.

Trapeznikov:1967:SPU

[Tra67] Academician V. Trapeznikov. Scientific policy in the USSR. Minerva, 5(4):546–552, Summer 1967. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01096783.

Travis:2005:CA

[Tra05] Anthony Travis. Chemical affinities. Minerva, 43(3):335–338, ???? 2005. CODEN MINEFY. ISSN 0026-4695 (print),

1573-1871 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11024-005-6515-z.

Toulmin:1974:BR

[TRGN74] Stephen Toulmin, Martin J. S. Rudwick, Irene A. Gilbert, and Richard R. Nelson. Book reviews. Minerva, 12(4):522–539, October 1974. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01557644.

Tribe:1971:LFA

[Tri71a] Laurence H. Tribe. Legal frameworks for the assessment and control of technology. Minerva, 9(2):243–255, April 1971. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01553158.

Trist:1971:C

[Tri71b] Eric Trist. Correspondence. Minerva, 9(4):556–558, October 1971. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01558029.

Trilling:1979:TEF

[Tri79] Leon Trilling. Technological elites in France and the United States. Minerva, 17(2):225–243, Summer 1979. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01096282.

Tribe:1994:BEM

[Tri94] Keith Tribe. Business education at the Mannheim Handelshochschule, 1907–1933. Minerva, 32(2):158–185, Summer 1994. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01103356.

Trow:1976:EHE

[Tro76] Martin Trow. “Elite higher education”: an endangered species? Minerva, 14(3):355–376, Autumn 1976. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01096277.

Trow:2000:MHE

[Tro00] Martin Trow. From mass higher education to universal access: The American advantage. Minerva, 37(4):303–328, December 2000. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1004708520977.

Trunk:2006:BWL

[Tru06] Achim Trunk. Biochemistry in wartime: The life and lessons of Adolf Butenandt, 1936–1946. Minerva, 44(3):285–306, June 2006. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-006-9002-2.

Taubert:2024:MGD

[TSB24] Niels Taubert, Linda Sterzik, and Andre Bruns. Mapping the German diamond open access journal landscape. Minerva, 62 (2):193–227, June 2024. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL https://link.springer.

com/article/10.1007/s11024-023-09519-7.

Treue:1986:BR

[TSH86] Wilhelm Treue, George J. Stigler, and L. F. Haber. Book reviews. Minerva, 24(2–3):359–389, Summer–Autumn 1986. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01096617.

Tavares:2019:IRP

[TSL19] Orlanda Tavares, Cristina Sin, and Vasco Lança. Inbreeding and research productivity among sociology PhD holders in Portugal. Minerva, 57(3):373–390, September 2019. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/article/10.1007/s11024-019-09378-1.

Tsvetkova:2014:MNP

[Tsv14] Natalia Tsvetkova. Making a new and pliable Professor: American and Soviet transformations in German universities, 1945– 1990. Minerva, 52(2):161–185, June 2014. CODEN MINEFY.

ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
014-9252-3.

Trubnikov:2024:BJP

[TT24] Dmitrii Trubnikov and Ekaterina Trubnikova. From bogus journals to predatory universities: The evolution of the Russian academic sphere within the predatory settings of the state. Minerva, 62(1):49–68, March 2024. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL https://link.springer.

com/article/10.1007/s11024-023-09502-2.

Tucker:2000:RD

[Tuc00] Nicholas Tucker. Those regal dons. Minerva, 38(2):221–231, ???? 2000. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1026505212484.

Turner:1987:GTS

[Tur87a] R. Steven Turner. The great transition and the social patterns of German science. Minerva, 25(1–2):56–76, Spring–Summer 1987. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096856.

Turner:1987:SNC

[Tur87b] Stephen P. Turner. The survey in nineteenth-century American geology: The evolution of a form of patronage. Minerva, 25(3): 282–330, Autumn 1987. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01097786.

Turner:1998:DFM

[Tur98] Stephen P. Turner. Did funding matter to the development of research methods in sociology? Minerva, 36(1):69–79, March 1998. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1004386504452.

Turtle:2006:JJ

[Tur06] Alison M. Turtle. Jensen on Jensen. Minerva, 44(1):125– 128, March 2006. CODEN MINEFY. ISSN 0026-4695 (print),

1573-1871 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11024-005-5397-4.

Turner:2010:NAE

[Tur10] Stephen P. Turner. Normal accidents of expertise. Minerva, 48(3):239–258, September 2010. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
010-9153-z.

Twomey:1999:MLT

[Two99] Christopher P. Twomey. The McNamara line and the turning point for civilian scientist–advisers in American defence policy, 1966–1968. Minerva, 37(3):235–258, September 1999. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1023/A%3A1004741523654.

Tyfield:2012:CPE

[Tyf12] David Tyfield. A cultural political economy of research and innovation in an age of crisis. Minerva, 50(2):149–167, June 2012. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-012-9201-y.

Ulam:1983:AEI

[Ula83] Adam Ulam. The academic ethic: II the danger lie within the universities themselves. Minerva, 21(2–3):292–295, Summer– Autumn 1983. CODEN MINEFY. ISSN 0026-4695 (print),

1573-1871 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF01097968.

Undurraga:2023:TDU

[UMC^{+}23] Tomas Undurraga, Sasha Mudd, Dusan Cotoras, Gonzalo Aguirre, and Tamara Orellana. “They don’t understand us, but we have to understand them”: Interrogating the making of interdisciplinary research in Chilean climate science. Minerva, 61(4): 581–606, December 2023. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL https://link.springer.

com/article/10.1007/s11024-023-09495-y.

Umunc:1986:SIR

[Umu86] Himmet Umunc. In search of improvement: The reorganisation of higher education in Turkey. Minerva, 24(4):433–455, Winter 1986. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096706.

Valdivia:2011:SBD

[Val11] Walter D. Valdivia. The stakes in Bayh–Dole: Public values beyond the pace of innovation. Minerva, 49(1):25–46, March 2011. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/content/pdf/
10.1007/s11024-011-9162-6.pdf.

Vanderstraeten:2015:MPA

[Van15] Raf Vanderstraeten. The making of Parsons’s The American University. Minerva, 53(4):307–325, December 2015. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/s11024-015-9285-2.

Varcoe:1970:SGO

[Var70] Ian Varcoe. Scientists, government and organised research in Great Britain 1914–16: The early history of the DSIR. Minerva, 8(1–4):192–216, January 1970. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01553212.

Varcoe:1971:SIC

[Var71] Ian Varcoe. Scientists as an inferior class. Minerva, 9(1): 132–135, January 1971. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01553192.

Varcoe:1981:COR

[Var81] Ivan Varcoe. Co-operative research associations in British industry, 1918–34. Minerva, 19(3):433–463, Autumn 1981. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF02192824.

Varughese:2012:WMM

[Var12] Shiju Sam Varughese. Where are the missing masses? The quasi-publics and non-publics of technoscience. Minerva, 50 (2):239–254, June 2012. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-012-9197-3.

Vavakova:1998:NSC

[Vav98] Blanka Vavakova. The new social contract between governments, universities and society: Has the old one failed? Minerva, 36(3): 209–228, September 1998. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1004384519396.

vandenBelt:2019:ESW

[vdB19] Henk van den Belt. Esha Shah, who is the scientist–subject? Affective history of the gene. Minerva, 57(2):261–264, June 2019. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s11024-019-09376-3.

vanderWaals:2001:FWS

[vdW01] J. H. van der Waals. The fate of women in the science pipeline. Minerva, 39(3):353–358, ???? 2001. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1023/
A%3A1017904928431.

Velarde:2018:WWA

[Vel18] Kathia Serrano Velarde. The way we ask for money… the emergence and institutionalization of grant writing practices in academia. Minerva, 56(1):85–107, March 2018. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/article/10.1007/s11024-
018-9346-4.

Verhoog:1981:RS

[Ver81] Henk Verhoog. The responsibilities of scientists. Minerva, 19 (4):582–604, Winter 1981. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096194.

Vessuri:1984:SSC

[Ves84] Hebe M. C. Vessuri. The search for a scientific community in Venezuela: From isolation to applied research. Minerva, 22(2): 196–235, Summer 1984. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF02207949.

Vessuri:1988:USR

[Ves88] Hebe M. C. Vessuri. The universities, scientific research and the national interest in Latin America. Minerva, 24(1):1–38, Spring 1988. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01102552.

Vessuri:1994:FSR

[Ves94] Hebe M. C. Vessuri. Foreign scientists, the Rockefeller Foundation and the origins of agricultural science in Venezuela. Minerva, 32(3):267–296, Autumn 1994. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01098663.

Villanueva-Felez:2013:MPN

[VFMGEE13] Africa Villanueva-Felez, Jordi Molas-Gallart, and Alejandro Escribá-Esteve. Measuring personal networks and their relationship with scientific production. Minerva, 51(4):465–483, December 2013. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-013-9239-5.

Velho:2024:WSP

[VGM^{+}24] Raquel Velho, Michael Gastrow, Caroline Mason, Marina Ulguim, and Yoliswa Sikhosana. What is the space for “Place” in social studies of astronomy? Minerva, 62(4):573–591, December 2024. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL https://link.springer.com/article/10.
1007/s11024-024-09527-1.

Vickery:1963:SIP

[Vic63] B. C. Vickery. Scientific information: Problems and prospects. Minerva, 2(1):21–38, Autumn 1963. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01100778.

Vuolanto:2017:GPN

[VL17] Pia Vuolanto and Anne Laiho. The gender perspective in nursing research: A theoretical treasure chest or a ‘thorn’ in the side? Minerva, 55(3):371–390, September 2017. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/article/10.1007/s11024-017-9318-0.

Vucinich:1993:BR

[VLJG93] Alexander Vucinich, Hugh Lloyd-Jones, and J. W. Grove. Book reviews. Minerva, 31(1):106–124, Spring 1993. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01096174.

Vucinich:1991:BR

[VMD^{+}91] Alexander Vucinich, Edward J. Monahan, Bernard D. Davis, Arthur Hearnden, Paolo Palladino, and Bruce L. R. Smith. Book reviews. Minerva, 29(4):476–516, Winter 1991. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01113493.

Volf:2021:BRR

[Vol21a] Darina Volf. Book review: Roger D. Launius, Reaching for the Moon: a Short History of the Space Race. Yale University Press, New Haven and London, 2019, 247 pp. Minerva, 59(1):145–147, March 2021. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http:/

/link.springer.com/article/10.1007/s11024-021-09437-
6;   http://link.springer.com/content/pdf/10.1007/s11024-
021-09437-6.pdf.

Volf:2021:EAS

[Vol21b] Darina Volf. Evolution of the Apollo–Soyuz Test Project: The effects of the “Third” on the interplay between cooperation and competition. Minerva, 59(3):399–418, September 2021. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL https://link.springer.com/article/10.
1007/s11024-021-09435-8.

VonGizycki:1973:CPI

[Von73] Rainald Von Gizycki. Centre and periphery in the international scientific community: Germany, France and Great Britain in the 19th century. Minerva, 11(4):474–494, October 1973. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01557798.

Voss:2023:CCCa

[Vos23a] Ehler Voss. Conjuration and conspiracy. the controversy over the German Covid policy as a mediumistic trial, or: The medium is the mess. Minerva, 61(3):435–458, September 2023. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL https://link.springer.com/article/10.
1007/s11024-023-09508-w. See correction [Vos23b].

Voss:2023:CCCb

[Vos23b] Ehler Voss. Correction to: Conjuration and conspiracy. the controversy over the German Covid policy as a mediumistic trial, or: The medium is the mess. Minerva, 61(3):459–460, September 2023. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL https://link.springer.com/article/10.
1007/s11024-023-09512-0. See [Vos23a].

Vasen:2022:HTP

[VP22] Federico Vasen and Miguel Sierra Pereiro. “The hardest task” — peer review and the evaluation of technological activities. Minerva, 60(3):375–395, September 2022. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL https://

link.springer.com/article/10.1007/s11024-022-09461-0.

vanRijswoud:2010:VEB

[vR10] Erwin van Rijswoud. Virology experts in the boundary zone between science, policy and the public: a biographical analysis. Minerva, 48(2):145–167, June 2010. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s11024-010-9145-z.pdf.

vanRooij:2013:GPT

[vR13] Arjan van Rooij. Gaps and plugs: TNO, and the problems of getting knowledge out of laboratories. Minerva, 51(1):25–48, March 2013. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-013-9217-y.

vanRooij:2014:UKP

[vR14] Arjan van Rooij. University knowledge production and innovation: Getting a grip. Minerva, 52(2):263–272, June 2014. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-014-9254-1.

Veletanlic:2020:IIA

[VS20] Emina Veletanlić and Creso Sá. Implementing the innovation agenda: a study of change at a research funding agency. Minerva, 58(2):261–283, June 2020. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/article/10.1007/s11024-020-09396-4.

Volkmann:2014:TWA

[VSR14] Ute Volkmann, Uwe Schimank, and Markus Rost. Two worlds of academic publishing: Chemistry and German sociology in comparison. Minerva, 52(2):187–212, June 2014. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
014-9251-4.

Vuolanto:2015:BWP

[Vuo15] Pia Vuolanto. Boundary work and power in the controversy over therapeutic touch in Finnish nursing science. Minerva, 53(4): 359–380, December 2015. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-015-9284-3.

Vandermoere:2012:DNB

[VV12] Frédéric Vandermoere and Raf Vanderstraeten. Disciplinary networks and bounding: Scientific communication between science and technology studies and the history of science. Minerva, 50(4):451–470, December 2012. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
012-9210-x.

Varma:1995:IIS

[VW95] Roli Varma and Richard Worthington. Immiseration of industrial scientists in corporate laboratories in the United States. Minerva, 33(4):325–338, Winter 1995. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01096516.

Waardenburg:1970:SAM

[Waa70] Jacques Waardenburg. Science in the Arab Middle East. Minerva, 8(1–4):597–599, January 1970. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01553242;
http://link.springer.com/article/10.1007/BF01553242.

Waaijer:2015:CAA

[Waa15] Cathelijn J. F. Waaijer. The coming of age of the academic career: Differentiation and professionalization of German academic positions from the 19th century to the present. Minerva, 53(1):43–67, March 2015. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-014-9264-z.

Wallis:1988:UG

[Wal88] W. Allen Wallis. I University government. Minerva, 24(1): 100–129, Spring 1988. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01102556.

Wallace:1992:SDG

[Wal92] Robert W. Wallace. Starting a department and getting it under way: Sociology at Columbia University, 1891–1914. Minerva, 30 (4):497–512, Winter 1992. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096575.

Walsh:1996:HTE

[Wal96] James Jackson Walsh. Higher technological education in Britain: The case of the Manchester municipal College of technology. Minerva, 34(3):219–257, ???? 1996. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF00120326.

Walsh:1998:PTE

[Wal98] James Jackson Walsh. Postgraduate technological education in Britain: Events leading to the establishment of Churchill College, Cambridge, 1950–1958. Minerva, 36(2):147–177, June 1998. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1004396018945.

Walker:2006:MW

[Wal06] Mark Walker. MARK WALKER. Minerva, 44(3):241–250, June 2006. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-006-9000-4.

Walker:2010:BRR

[Wal10] Mark Walker. Book review: Roberto Scazzieri and Raffaella Simili (eds.): The Migration of Ideas. Minerva, 48(1): 101–104, March 2010. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-010-9139-x.

Wang:1995:FWW

[Wan95] Zuoyue Wang. The first world war, academic science, and the “two cultures”: Educational reforms at the University of Cambridge. Minerva, 33(2):107–127, Summer 1995. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01097908.

Washburn:1988:APC

[Was88] Wilcomb E. Washburn. The academic profession and contemporary politics. Minerva, 26(3):392–415, Autumn 1988. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01096403.

Woodhall:1965:PTB

[WB65] Maureen Woodhall and Mark Blaug. Productivity trends in British university education, 1938–62. Minerva, 3(4):483– 498, Summer 1965. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01099412.

Wohlstetter:1977:MPC

[WBJ^{+}77] Albert Wohlstetter, Thomas A. Brown, Gregory Jones, David McGarvey, Henry Rowen, Vincent Taylor, and Roberta Wohlstetter. The military potential of civilian nuclear energy. Minerva, 15(3–4):387–538, Autumn–Winter 1977. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01096730.

Wagner:2015:RDC

[WBL15] Caroline S. Wagner, Lutz Bornmann, and Loet Leydesdorff. Recent developments in China–U.S. cooperation in science.

Minerva, 53(3):199–214, September 2015. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
015-9273-6.

Weinryb:2018:RSC

[WBW18] Noomi Weinryb, Maria Blomgren, and Linda Wedlin. Rationalizing science: a comparative study of public, industry, and nonprofit research funders. Minerva, 56(4):405–429, December 2018. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s11024-018-9352-6;   http://link.springer.com/content/
pdf/10.1007/s11024-018-9352-6.pdf.

Woelert:2018:MDI

[WC18] Peter Woelert and Gwilym Croucher. The multiple dynamics of isomorphic change: Australian law schools 1987–1996. Minerva, 56(4):479–503, December 2018. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/article/10.1007/s11024-018-9350-8.

Wynne-Edwards:1972:CFR

[WE72] V. C. Wynne-Edwards. The choice and formulation of research problems: Four comments on the Rothschild report. Minerva, 10(2):191–208, April 1972. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01682417.

Weber:1977:SAA

[Web77] Marianne Weber. The soil and air of academic life. Minerva, 15 (2):214–246, Summer 1977. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096678.

Wehling:2015:BRL

[Weh15] Peter Wehling. Book review: Luigi Pellizzoni and Marja Ylönen (eds.), Neoliberalism and Technoscience: Critical Assessments. Minerva, 53(2):193–198, June 2015. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
015-9268-3.

Weinberg:1963:CSC

[Wei63] Alvin M. Weinberg. Criteria for scientific choice. Minerva, 1 (2):159–171, Winter 1963. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096248.

Weightman:1964:HRP

[Wei64a] J. G. Weightman. Heretical remarks on Professor Aron and university expansion. Minerva, 2(4):477–483, Summer 1964. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01097539.

Weinberg:1964:CSC

[Wei64b] Alvin M. Weinberg. Criteria for scientific choice II: The two cultures. Minerva, 3(1):3–14, Autumn 1964. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01630147.

Weinberg:1965:SCBb

[Wei65a] Alvin M. Weinberg. Scientific choice and biomedical science. Minerva, 4(1):3–14, Autumn 1965. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01585982.

Weinberg:1965:SCBa

[Wei65b] Alvin M. Weinberg. Scientific choice, basic science and applied missions. Minerva, 3(4):515–523, Summer 1965. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01099415.

Weinberg:1966:CSC

[Wei66] Alvin M. Weinberg. The complexity of scientific choice. Minerva, 4(3):400–402, Spring 1966. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01584636.

Weinberg:1967:GSS

[Wei67] Alvin M. Weinberg. The growth of science in society. Minerva, 6 (1):121–122, Autumn 1967. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01101615.

Weinberg:1968:SCS

[Wei69] Alvin M. Weinberg. Scientific choice and the scientific muckrakers: Review article. Minerva, 7(1–2):52–63, Autumn– Winter 1968–1969. CODEN MINEFY. ISSN 0026-4695 (print),

1573-1871 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/BF01096808.

Weinberg:1971:EBB

[Wei71] Alvin M. Weinberg. The economic benefits of basic research. Minerva, 9(2):293–294, April 1971. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01553162.

Weinberg:1972:UIR

[Wei72a] Alvin M. Weinberg. I a useful institution of the republic of science. Minerva, 10(3):439–440, July 1972. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01556924.

Weinberg:1972:STS

[Wei72b] Alvin M. Weinberg. Science and trans-science. Minerva, 10 (2):209–222, April 1972. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01682418.

Weinberg:1974:ISP

[Wei74] Alvin M. Weinberg. Institutions and strategies in the planning of research. Minerva, 12(1):8–17, January 1974. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01558537.

Weisz:1976:EDF

[Wei76] George Weisz. Émile Durkheim on the French universities. Minerva, 14(3):377–388, Autumn 1976. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01096278.

Weinberg:1978:OCR

[Wei78a] Alvin M. Weinberg. The obligations of citizenship in the republic of science. Minerva, 16(1):1–3, Spring 1978. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01102178.

Weinberg:1978:UBR

[Wei78b] Alvin M. Weinberg. The use of the breeder reactor. Minerva, 16 (4):581–585, Winter 1978. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01100333.

Weisz:1978:BR

[Wei78c] George Weisz. Book reviews. Minerva, 16(3):445–460, Autumn 1978. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096671.

Weisz:1979:FUE

[Wei79] George Weisz. The French universities and education for the new professions, 1885–1914: an episode in French university reform. Minerva, 17(1):98–128, Spring 1979. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01096262.

Weinberg:1984:VSU

[Wei84] Alvin M. Weinberg. Values in science: Unity as a criterion of scientific choice. Minerva, 22(1):1–12, Spring 1984. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF02207554.

Weinberg:1987:SDI

[Wei87a] Alvin M. Weinberg. The strategic defense initiative, arms control, and the ethos of the University. Minerva, 25(4): 486–501, Winter 1987. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01099444.

Weisz:1987:MEF

[Wei87b] George Weisz. The medical elite in France in the early nineteenth century. Minerva, 25(1–2):150–170, Spring–Summer 1987. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096862.

Weinberg:1990:TD

[Wei90] Alvin M. Weinberg. Technology and democracy. Minerva, 28 (1):81–90, Spring 1990. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096327.

Weinberg:1993:CLT

[Wei93a] Alvin M. Weinberg. Chapters from the life of a technological fixer. Minerva, 31(4):379–454, Winter 1993. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01096449.

Weindling:1993:PHP

[Wei93b] Paul Weindling. Public health and political stabilisation: The Rockefeller Foundation in Central and Eastern Europe between the two world wars. Minerva, 31(3):253–267, Autumn 1993. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01098623.

Weinberg:1996:ESG

[Wei96] Alvin M. Weinberg. Edward Shils and the “Governmentalisation” of science. Minerva, 34(1):39–43, ???? 1996. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF00124199.

Weindling:1997:PWH

[Wei97] Paul Weindling. Philanthropy and world health: The Rockefeller Foundation and the league of nations health organisation. Minerva, 35(3):269–281, September 1997. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1023/
A%3A1004242303705.

Weingart:1999:CCS

[Wei99] Peter Weingart. Climate coalitions: The science and politics of climate change. Minerva, 37(2):103–104, June 1999. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1023/A%3A1004679732306.

Weinberg:2000:CSC

[Wei00] Alvin M. Weinberg. Criteria for scientific choice (Minerva, i (2), (1962), 158–171). Minerva, 38(3):253–266, ???? 2000. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1026512029530.

Weingart:2008:E

[Wei08] Peter Weingart. Editorial. Minerva, 46(2):171–174, June 2008. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-008-9094-y.

Weingart:2009:EIa

[Wei09a] Peter Weingart. Editorial for issue 47/1. Minerva, 47(1):1–5, March 2009. CODEN MINEFY. ISSN 0026-4695 (print), 1573-

1871 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s11024-009-9115-5.pdf.

Weingart:2009:EIb

[Wei09b] Peter Weingart. Editorial for issue 47/3. Minerva, 47(3): ??, September 2009. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-009-9131-5.

Weingart:2012:MYR

[Wei12a] Peter Weingart. Minerva — 50 years reflecting on science in society. Minerva, 50(3):255–259, September 2012. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s11024-012-9209-3.pdf.

Weiss:2012:TST

[Wei12b] Charles Weiss. On the teaching of science, technology and international affairs. Minerva, 50(1):127–137, March 2012. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s11024-012-9191-9.pdf.

Weiss:2015:HDS

[Wei15] Charles Weiss. How do science and technology affect international affairs? Minerva, 53(4):411–430, December 2015. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s11024-015-9286-1.pdf.

Werskey:2007:VCR

[Wer07] Gary Werskey. The visible College revisited: Second opinions on the red scientists of the 1930s. Minerva, 45(3):305–319, September 2007. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-007-9047-x.

Wettersten:1985:CAS

[Wet85] John Wettersten. On conservative and adventurous styles of scientific research. Minerva, 23(4):443–463, Winter 1985. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01098700.

Wilson:1974:RBD

[WG74] James A. Wilson and Jerry Gaston. Reflux from the “brain drain”. Minerva, 12(4):459–468, October 1974. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01557640.

Weisz:1985:BR

[WG85] George Weisz and Michael Gibbons. Book reviews. Minerva, 23 (2):309–317, Summer 1985. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01099948.

Whyte:1997:FBB

[WG97] Neil Whyte and Philip Gummett. Far beyond the bounds of science: The making of the United Kingdom’s first space policy. Minerva, 35(2):139–169, June 1997. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1023/
A%3A1004294811357.

Whitley:2018:ICF

[WGL18] Richard Whitley, Jochen Gläser, and Grit Laudel. The impact of changing funding and authority relationships on scientific innovations. Minerva, 56(1):109–134, March 2018. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/article/10.1007/s11024-
018-9343-7.

Weinberg:1980:BR

[WGN80] Alvin M. Weinberg, Anthony Grafton, and Richard R. Nelson. Book reviews. Minerva, 18(3):529–537, Autumn 1980. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/BF01096956.

Woodson:2018:NAC

[WHF18] Thomas S. Woodson, Matthew Harsh, and Rider Foley. Nonacademic careers for STS graduate students: Hopping off the tenure track. Minerva, 56(4):529–535, December 2018. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/article/10.1007/s11024-
018-9360-6.

White:1998:ABN

[Whi98] Paul S. White. All bureaucrats now? Minerva, 36(3):299– 303, September 1998. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1004372829391.

Whitley:2009:BRM

[Whi09] Richard Whitley. Book review: Michele Lamont, How Professors Think: Inside the Curious World of Academic Judgement. Minerva, 47(4):469–472, December 2009. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
009-9133-3.

Whitley:2011:CGA

[Whi11] Richard Whitley. Changing governance and authority relations in the public sciences. Minerva, 49(4):359–385, December 2011. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-011-9182-2.

Whitley:2012:TUN

[Whi12] Richard Whitley. Transforming universities: National conditions of their varied organisational actorhood. Minerva, 50(4): 493–510, December 2012. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-012-9215-5.

Widmalm:1995:SNN

[Wid95] Sven Widmalm. Science and neutrality: The Nobel prizes of 1919 and scientific internationalism in Sweden. Minerva, 33 (4):339–360, Winter 1995. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096517.

Widmalm:2001:I

[Wid01] Sven Widmalm. Introduction. Minerva, 39(4):365–372, ???? 2001. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1012731818458.

Widmalm:2007:ISC

[Wid07] Sven Widmalm. Introduction: Science and the creation of value. Minerva, 45(2):115–120, June 2007. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
007-9039-x.

Williams:1964:REG

[Wil64] B. R. Williams. Research and economic growth — what should we expect? Minerva, 3(1):57–71, Autumn 1964. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01630149.

Wilson:1966:EBS

[Wil66] James A. Wilson. The emigration of British scientists. Minerva, 5(1):20–29, Autumn 1966. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF02389790.

Wilson:1971:DAS

[Wil71] John T. Wilson. A dilemma of American science and higher educational policy: The support of individuals and fields versus the support of universities. Minerva, 9(2):171–196, April 1971. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01553155.

Williams:1972:UVU

[Wil72] Bruce R. Williams. University values and university organisation. Minerva, 10(2):259–279, April 1972. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01682421.

Williams:1974:LLE

[Wil74] Peter Williams. Lending for learning: an experiment in Ghana. Minerva, 12(3):326–345, July 1974. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01102528.

Williams:1983:LBT

[Wil83a] Bruce Williams. Living better with technology. Minerva, 21 (4):370–386, Winter 1983. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096523.

Wilson:1983:AEP

[Wil83b] James Q. Wilson. The academic ethic: I. Partisanship, judgement and the academic ethic. Minerva, 21(2–3):285–291, Summer–Autumn 1983. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01097967.

Willison:1985:BR

[Wil85] I. R. Willison. Book review. Minerva, 23(3):433–440, Autumn 1985. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096447.

Williams:1986:DIR

[Wil86] Bruce Williams. The direct and indirect role of higher education in industrial innovation: What should we expect? Minerva, 24(2–3):145–171, Summer–Autumn 1986. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01096609.

Williams:1997:RCC

[Wil97] Roger Williams. Reflections on Cornford’s Cambridge and the present dilemmas of British higher education. Minerva, 35(1): 63–72, ???? 1997. CODEN MINEFY. ISSN 0026-4695 (print),

1573-1871 (electronic). URL http://link.springer.com/
accesspage/article/10.1023/B%3AMINE.0000019491.38349.
c1.

Williams:2007:HBC

[Wil07] Robyn Williams. Halfway between chutzpah and megalomania. Minerva, 45(3):345–348, September 2007. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
007-9049-8.

Williams:2020:BRL

[Wil20] Logan D. A. Williams. Book review: ‘From blindness to light’: a review of Benoît Godin and Dominique Vinck (eds.), Critical Studies of Innovation: Alternative Approaches to the Pro– Innovation Bias. Minerva, 58(2):309–314, June 2020. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/article/10.1007/s11024-
020-09402-9.

Winters:1994:JHZ

[Win94] Stanley B. Winters. Josef Hlávka, Zdeněk Nejedlý, and the Czech Academy of Sciences and Arts, 1891–1952. Minerva, 32 (1):53–78, Spring 1994. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096656.

Winter:2020:ILI

[Win20] Katarina Winter. “I’ll look into it!” lubricants in conversational coproduction. Minerva, 58(2):285–307, June 2020. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/article/10.1007/s11024-
020-09394-6;   http://link.springer.com/content/pdf/
10.1007/s11024-020-09394-6.pdf.

Weinberg:1970:BR

[WKC^{+}70] Alvin M. Weinberg, Dharma Kumar, D. S. L. Cardwell, Edward Shils, and Eric Ashby. Book reviews. Minerva, 8(1–4): 300–315, January 1970. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01553219;   http://
link.springer.com/article/10.1007/BF01553219.

Wright:2020:ECR

[WM20] Jack Wright and Tiago Mata. Epistemic consultants and the regulation of policy knowledge in the Obama administration. Minerva, 58(4):535–558, December 2020. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/article/10.1007/s11024-
020-09411-8;   http://link.springer.com/content/pdf/
10.1007/s11024-020-09411-8.pdf.

Watermeyer:2016:EEI

[WO16] Richard Watermeyer and Mark Olssen. ‘Excellence’ and exclusion: The individual costs of institutional competitiveness. Minerva, 54(2):201–218, June 2016. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/content/pdf/10.1007/s11024-
016-9298-5.pdf.

Woelert:2013:EMP

[Woe13] Peter Woelert. The ‘economy of memory’: Publications, citations, and the paradox of effective research governance. Minerva, 51(3):341–362, September 2013. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
013-9232-z.

Woelert:2015:GKF

[Woe15] Peter Woelert. Governing knowledge: The formalization dilemma in the governance of the public sciences. Minerva, 53(1):1–19, March 2015. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-015-9266-5.

Wood:1972:NPP

[Woo72] Glynn Wood. National planning and public demand in Indian higher education: The case of Mysore. Minerva, 10(1): 83–106, January 1972. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01881391.

Woolf:1975:SMI

[Woo75] Patricia K. Woolf. The second messenger: Informal communication in cyclic AMP research. Minerva, 13(3):349–373, Autumn 1975. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01097846.

Woolf:1984:BRI

[Woo84] Harry Woolf. Basic research and industrial enterprise. Minerva, 22(2):183–195, Summer 1984. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF02207948.

Woolf:2003:URI

[Woo03] Stuart Woolf. On University reform in Italy: Contradictions and power relations in structure and function. Minerva, 41(4):347– 363, ???? 2003. CODEN MINEFY. ISSN 0026-4695 (print),

1573-1871 (electronic). URL http://link.springer.com/
accesspage/article/10.1023/B%3AMINE.0000005156.56860.
06.

Woods:2009:PAC

[Woo09] Abigail Woods. ‘partnership’ in action: Contagious abortion and the governance of livestock disease in Britain, 1885– 1921. Minerva, 47(2):195–216, June 2009. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
009-9123-5.

Wallace:2015:RPA

[WR15] Matthew L. Wallace and Ismael Rafols. Research portfolio analysis in science policy: Moving from financial returns to societal benefits. Minerva, 53(2):89–115, June 2015. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s11024-015-9271-8.

Wildavsky:1991:BR

[WS91] Aaron Wildavsky and Allen R. Salzman. Book reviews. Minerva, 29(2):231–246, Summer 1991. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01096410.

Woelert:2025:SBC

[WS25] Peter Woelert and Bjørn Stensaker. Strategic bureaucracy: The convergence of bureaucratic and strategic management logics in the organizational restructuring of universities. Minerva, 63 (1):1–21, March 2025. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL https://link.springer.

com/article/10.1007/s11024-024-09535-1.

Weingart:2012:E

[WT12] Peter Weingart and Niels C. Taubert. Editorial. Minerva, 50(1):1–2, March 2012. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s11024-012-9193-7.pdf.

Wright:2017:VIR

[WV17] Claire Wright and Simon Ville. Visualising the interdisciplinary research field: The life cycle of economic history in Australia. Minerva, 55(3):321–340, September 2017. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/article/10.1007/s11024-017-9319-z.

Williams:2012:FIS

[WW12] Logan D. A. Williams and Thomas S. Woodson. The future of innovation studies in less economically developed countries. Minerva, 50(2):221–237, June 2012. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
012-9200-z.

Williams:2019:EST

[WW19] Logan D. A. Williams and Thomas S. Woodson. Enhancing socio-technical governance: Targeting inequality in innovation through inclusivity mainstreaming. Minerva, 57(4):453–477, December 2019. CODEN MINEFY. ISSN 0026-4695 (print), 1573-

1871 (electronic). URL http://link.springer.com/article/
10.1007/s11024-019-09375-4.

Xu:2022:RSC

[XH22] Shaoxiong Brian Xu and Guangwei Hu. Retraction stigma and its communication via retraction notices. Minerva, 60(3):349– 374, September 2022. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL https://link.springer.

com/article/10.1007/s11024-022-09465-w.

Xie:2021:LPC

[Xie21] Jialu Xie. ‘Political controversy and social science public funding’: a review of Mark Solovey, Social Science for What? Battles over Public Funding for the “Other Sciences” at the National Science Foundation. Minerva, 59(2):285–288, June 2021. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL https://link.springer.com/article/10.
1007/s11024-021-09436-7.

Xu:2021:III

[XOR21] Xin Xu, Alis Oancea, and Heath Rose. The impacts of incentives for international publications on research cultures in Chinese humanities and social sciences. Minerva, 59(4):469–492, December 2021. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL https://link.springer.com/article/10.
1007/s11024-021-09441-w.

Xu:2008:CLA

[Xu08] Ang Xu. China looks abroad: Changing directions in international science. Minerva, 46(1):37–51, March 2008. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/s11024-008-9086-y.

Yang:2020:PPG

[Yan20] Chih-Yuan Yang. Performing public good: The statecraft of objective and optimal FIT. Minerva, 58(3):435–458, September 2020. CODEN MINEFY. ISSN 0026-4695 (print), 1573-

1871 (electronic). URL http://link.springer.com/article/
10.1007/s11024-020-09400-x.

Youtie:2016:DCA

[YB16] Jan Youtie and Barry Bozeman. Dueling co-authors: How collaborators create and sometimes solve contributorship conflicts. Minerva, 54(4):375–397, December 2016. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
016-9303-z.

Yeo:2002:MKE

[Yeo02] Richard Yeo. Managing knowledge in early modern Europe. Minerva, 40(3):301–314, ???? 2002. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1023/
A%3A1019576610136.

Ylijoki:2014:UUS

[Yli14] Oili-Helena Ylijoki. University under structural reform: a microlevel perspective. Minerva, 52(1):55–75, March 2014. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic).

URL http://link.springer.com/accesspage/article/10.
1007/s11024-014-9246-1.

Zahlan:1970:SAM

[Zah70] A. B. Zahlan. Science in the Arab Middle East. Minerva, 8 (1–4):8–35, January 1970. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01553197.

Zahlan:2006:ASK

[Zah06] A. B. Zahlan. Arab societies as knowledge societies. Minerva, 44(1):103–112, March 2006. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-005-5404-9.

Zanders:2002:I

[Zan02] Jean-Pascal Zanders. Introduction. Minerva, 40(1):3–13, ???? 2002. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1023/A%3A1015093009616.

Zapp:2022:RGK

[Zap22] Mike Zapp. Revisiting the global knowledge economy: The worldwide expansion of research and development personnel, 1980–2015. Minerva, 60(2):181–208, June 2022. CO-DEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (elec-

tronic).  URL https://link.springer.com/article/10.
1007/s11024-021-09455-4.

Zuckerman:1975:WAS

[ZC75] Harriet Zuckerman and Jonathan R. Cole. Women in American science. Minerva, 13(1):82–102, Spring 1975. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/
BF01096243.

Ziman:1963:CSO

[Zim63] John Ziman. The college system at Oxford and Cambridge. Minerva, 1(2):191–208, Winter 1963. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01096251.

Ziman:1971:TPR

[Zim71] John Ziman. Three patterns of research in developing countries. Minerva, 9(1):32–37, January 1971. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01553184.

Ziman:1973:PSP

[Zim73] John Ziman. The profession of science and its powers. Minerva, 11(1):133–137, January 1973. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01544984.

Ziman:1977:ISC

[Zim77] John M. Ziman. The international scientific community. Minerva, 15(1):83–93, Spring 1977. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01096321.

Ziman:1978:SWR

[Zim78] John Ziman. Solidarity within the republic of science. Minerva, 16(1):4–19, Spring 1978. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01102179.

Ziman:1981:WOS

[Zim81] John Ziman. What are the options? social determinants of personal research plants. Minerva, 19(1):1–42, Spring 1981. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF02192547.

Ziman:1987:PPC

[Zim87] J. M. Ziman. The problem of “problem choice”. Minerva, 25 (1–2):92–106, Spring–Summer 1987. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/BF01096858.

Ziman:1998:BRI

[Zim98] John Ziman. Book review: Image and Logic: a Material Culture of Microphysics, by Peter Galison. Minerva, 36(3):289– 293, September 1998. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1023/A%3A1004352207101.

Ziman:2000:CSC

[Zim00a] John Ziman. Criteria for scientific choice — commentary. Minerva, 38(3):253–269, ???? 2000. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1023/
A%3A1026565515210.

Ziman:2000:RSP

[Zim00b] John Ziman. The Republic of science: Its political and economic theory. Minerva, 38(1):21–25, ???? 2000. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL

http://link.springer.com/accesspage/article/10.1023/
A%3A1026552331313.

Ziman:2005:KA

[Zim05] John Ziman. Knowledge by agreement. Minerva, 43(3): 289–295, ???? 2005. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11024-005-6509-x.

Zimmerman:2006:SPS

[Zim06] David Zimmerman. The society for the protection of science and learning and the politicization of British science in the 1930s. Minerva, 44(1):25–45, March 2006. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://

link.springer.com/accesspage/article/10.1007/s11024-
005-5405-8.

Zuckerman:1981:BR

[ZJ81] Lord Zuckerman and R. V. Jones. Book reviews. Minerva, 19 (1):185–200, Spring 1981. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF02192553.

Zomer:2010:DSO

[ZJE10] Arend H. Zomer, Ben W. A. Jongbloed, and Jürgen Enders. Do spin-offs make the academics’ heads spin? Minerva, 48(3): 331–353, September 2010. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s11024-010-9154-y.pdf.

Zukerman:1976:BR

[ZLD^{+}76] Lord Zukerman, Imre Lakatos, A. Hunter Dupree, Charles E. McClelland, and Alvin M. Weinberg. Book reviews. Minerva, 14(1):118–165, Spring 1976. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/BF01096217.

Zuckerman:1971:PES

[ZM71] Harriet Zuckerman and Robert K. Merton. Patterns of evaluation in science: Institutionalisation, structure and functions of the referee system. Minerva, 9(1):66–100, January 1971. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01553188.

Zielenska:2025:BTA

[ZW25] Marianna Zieleńska and Magdalena Wnuk. Benchmarking and the technicization of academic discourse: The case of the EU atrisk of poverty or social exclusion composite indicator. Minerva, 63(1):23–43, March 2025. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871 (electronic). URL https://link.springer.

com/article/10.1007/s11024-024-09537-z.

Ziman:1978:BR

[ZWJK78] John Ziman, Alvin M. Wzinberg, R. V. Jones, and Robert E. Klitgaard. Book reviews. Minerva, 16(2):327–353, Summer 1978. CODEN MINEFY. ISSN 0026-4695 (print), 1573-1871

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/BF01096019.